A Complete Bibliography of The Journal of the American Statistical Association: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.08

Title word cross-reference

.471 [CB94]. .632+ [ET97]. 1 − α [MP99]. 2 × 2 [Agr98a, Agr98b, Ano99h, BS97c, BB98b, BB98c, DD90, Gui93, MPS98a, MPS98b, SW98b, SW98c, SW98d, VHE91]. α/2 [MP99]. AR[k] [Dha93]. C_{p} [RS94b]. D [AC95c]. F [LRR91, RSY93]. ∫_{−∞}^{+∞} f(t) exp(−t^{2}) dt [CS90b]. K [Wan95, EL90, GF96a, Pen97]. K2 × J [YF95]. k × l [Qui98a]. L [WM90]. L2 × K [YF90]. L^{1} [AHY93, dHY93]. L_{1} [EM92, WS94]. M [Akr91b, Ber91a, KP90, MY91]. N [Hal94, KT94a, jW97]. N, p [Hal94]. P [BB94, Hol95, LS97, Ano91i, Daw95, SMP95b, Wee91]. S [CRH94, RV90c]. T [Ede90, Ver93, Hjo94]. v [FOS95]. X [Bre92, Lit92].

-Dimensional [Daw95]. -estimators [CRH94]. -Fixed [Bre92]. -Level [MP99]. -ness [Hjo94]. -Optimum [AC95c]. -Ratio [Pen97]. -Sample [EL90]. -Spherical [FOS95]. -System [RV90c]. -Values [Ano91i].

1 [GHMW94]. 1800- [Sol95]. 1990s [Ano91y, Pea91, Roc91].

21st [Ket97b, Ket97a].

83 [LB94d]. 84 [Fab94, Whi92b]. 86 [MS95a]. 88 [Wee94]. 89 [Sta94a].

90 [SG95a]. 92 [Wes98]. 92e [Fab94, Whi92b]. 92i [MS95a]. 93 [MS98a]. 94e [Wee94]. 96e [SG95a]. 98c [Wes98].

A-Bomb [PSVS92]. Aalen [HM91]. Aboriginal [Giv99]. Absolute [RC93]. Accelerated [CM97a, Yam92]. Accelerated-Time [CM97a]. Acceleration [JJ93]. Acceptability [BS97c]. Acceptance [LS91b]. Accidents [Bro92b, Wei92]. Account [BMR99]. Account-Level [BMR99]. Accounting [MR94, Roc95]. Accuracy [Che90, CT90, Haw93a, MS93, SW93, SM92]. Accurate [FWW99, SW97a]. Achieve [CN94]. Achieving [Ano98h, HOLDG98, Min98]. Acid [Haa90]. Acknowledgement [Ano99a, MS95a]. across [CJ99]. Action [APR92, Gai96, LS91b]. Activity [GJDT98a, GJDT98b]. Adapting [DJ95, HT97a, HT97b, HT97c, Por97a, Por97b]. Adaptive [AMCS98, Ano97j, BF93a, BWC98a, BWC98b, BGH93, CKM97, GRS98, GR96b, Gu90, HOS94b, KP92, LT94a, LS91a, LW97, Par93a, Pri94, Ruk93, SS96a, Sch95b, TFAH94, Tho90c, Fan92]. Addendum [Ano96a, Ano98a, Wes98]. Additive [BR94a, BRW99, CT93b, Cho92, DS97c, GM99, Gu92, HBH93, HM91, KO95, KPR91, LRM96, LCWH97, MS99a, Mul99, Niu96, OR98, SKW99a, SKW99b]. Address [Ket97a, Moo98a]. Adequacy [CW97, HW92a]. Adjoint [Dua90]. Adjust [IRC98a, IRC98b]. Adjusted [CW94a, NC99, ZD97]. Adjusting [Dig99, FZ99, Fre99a, LP99, LR99, RG94, SRR99a, SRR99b, vdL99]. Adjustment [HB99, Sch91b, SK91, SR95]. Adjustments [FMSP90, MS93]. Admissible [Sob91]. Adolescence [MSMP92]. Adrian [DE96]. Adult [HHH92]. Advantages [MH99]. Adversaries [LS91b]. Adverse [GLB98, Kad90b, YKB97]. Advertising [MSL93]. Affine [HMO97a, HMO97b]. Affine-Invariant [HMO97a, HMO97b]. Africa [GJDT98a, GJDT98b]. After [OMF^{+}93, Rub96a, TMP93, XXP93]. Against [AP96, Ano91j, CS92a, CF95, Hay90, LGWC93, Shi91, SS95a, Wan96, dCP99a]. Age [AK92, Cha91a, HR90, NSS97, RS90a, XP92]. Ages [Par93b]. Aggregated [MST97]. Aggregates [PH90a]. Agreed [LS91b]. Agricultural [CD92]. AHEAD [JS97b]. AIDS [Bac90, Bro92d, DXP92, DA96c, EFS92a, EFS92b, GJDT98a, GJDT98b, Gre92, Har92b, Har90b, HMC^{+}93, JMV94, Lou92b, OBW92, PS93, RG94, TCS94, TDW95, TMP93, XXP93, ZDL96]. al [Kin99]. Algorithm [Cho93, Gu93, GR92a, JJ93, KT95, LK92b, Luc93b, MR91, Sch96c, Sho93, WT90a, WZ97a, XLR91]. Algorithms [Joh96b, Joh98b, LB94c, McC97, WT90a, LB94d, LB94e]. Aligning [RN97b].

Alignment [LNL95]. Alignments [LNL99]. All-or-None [Bak98b, Bak98a]. Allocation [KM91a, Ode91, Pag90, SN90, WSLP90]. Almost [Eas94]. Also [Goo96]. Alternative [EH91, Fay96a, Hay90, MSMP92, Mun96, NG94, SW95a]. Alternatives [AC95b, Ano91j, CF95, Fol96, GR92b, JSKK98, LH91, LGWC93, MP99, RC93, SS95a, dCP99a]. Always [RRZ94]. Ambiguous [SM90b]. Amer [Fab94, LB94d, MS98a, MS95a, SG95a, Sta94a, Wee94, Wes98, Whi92b]. American [BB90, Pea91]. Among [Moo98a, SMF94, Moo98b]. Amount [Nyc91, Hjo94]. Analgesic [DS97a, DS97b, GB97a, GB97b, LL97a, LL97b, SBD97a, SBD97b, SBD97c]. Analog [BW94]. Analyses [HAS97, PC93, PSVS92, WC92]. Analysis [Ada98, AC93, Alb93a, AASM93, Alb93c, AC95a, And91, ASL91, AL95, AD90, Ano96g, BRR92, BCdF91, BPS96, Bak98b, Bak98a, BH93a, Bea91, BHB96, Bel93a, BC96, Ben91b, BS93, BY95, BS91a, Bri91b, Bri92, BR98b, BR98c, BR98a, BD91b, BH97, Car94b, CWW95, CH91b, CP99, CW99, CJ96, CM97b, CM91a, CR91, CH93a, CNWW98a, CNWW98b, Con92, Cox91, CH92, CMMY91, DwSH94, DG99, DAO94, DR91, DLB91, DS97a, DS97b, DH99, DM96, DPWH99, DD90, dE97, EP94, EK93, EGGS97, Fab91a, FZ98a, FZ98b, FSE98, Fay96a, FR99, FP95, For92, FWW99, Fun95, GSL92, GBJ96, GB97a, GB97b, GWG98, GBA^{+}99, GH90, Goo91a, Goo96, GRB^{+}95, Gro90, GL96, GS93b, GY93, Hab91, HW95a, HTB94, HM97, Hei95, Hil93, HT97d, HH91, IL91, IL94, JT97]. Analysis [Joh90, Joh94, Joh96a, Kad90b, Kne98a, Kne98b, KC96, Lan98a, Lan98b, Lav91a, LWY93, LB94b, Lin91, LW92, LL97a, LL97b, LCG91, LCWH97, Man90, MM95, Mor91a, MSK96, Mur95, MSL93, Nai91, Niu96, OMF^{+}93, OV95, Pas93, PB98, PM97, PS96, Pri96, PHL98, QTC96, Qui98a, RS98a, RS98b, RSY93, RG93, RRZ95, RAGW93a, RAGW93b, Roc95, RT96b, Ryu94, STR97, SG96a, Sha91, SBD97a, SBD97b, SBD97c, SJ92, SW90, Sil97a, Sin93, SD97, Sol91, Sol95, SEH95, SRSD94, SL98b, SM93b, Sto91a, Sto91b, SK93b, SY95, Tar99, TCS94, Ver93, Wak96, Wan92, WSJ92, WWZtO97, WLGC98, WW98a, WW98b, Weg90b, Wes97a, WM91, Yam90, Yam92, YK96, YJW95, YB97, Zuc98, dCP99b, Fab91b, Goo91b, Gla91]. Analytical [AHY93, Hod94, dHY93]. Analyzing [Boc99, BO95, DB97, GH94, Gra92, Mar92b, WKL95]. Ancestral [GT95]. Ancillarity [He90]. Annealing [GT95]. ANOVA [Ano91e]. Answer [BZ99]. Answered [Bin98b, Bin98a, Bre98b, Bre98a, GKL98c, GKL98d, GKL98a, GKL98b, Jud98b, Jud98a, Kad98b, Kad98a, San98b, San98a]. Anti [MSL93]. Anti-Drinking [MSL93]. Any [Eas94]. Application [AME96, ALJ99, AK92, BSB98, BF92, BMR99, BPS95, CH95a, CT98, CG97, CS96b, Cra94, CS90b, DwSH94, DDV99, Fin90, GT98, GRB^{+}95, GS93b, GR92a, Haa95b, HW92a, HR90, HDG97, HMC^{+}93, JMV94, Joh96a, KS96, LR96, Li92, LTW96, LS91b, MW98a, MW98b, MW98c, MSK96, MST97, O’N94, O’S90, OO97, PJT96, PH90a, RW98a, RW98b, SC96a, Sas97, SPM99, SG96b, SR95, SH90, SEH95, The93a, WM90, WHG98, Yam92, YK96, Ye98b, ZDL96].

Applications [AML95, AHB97, AGM98a, BPS96, Bak98b, Bea91, Bin98a, BS97b, BO95, BWC98a, Bre98a, Bri91b, CNWW98a, CMMY91, DN98a, DG92, FG94, Gal94, GJ94, GKL98a, GKL98b, GJDT98a, GT95, Gho92, GL92, Gra92, Hab95a, HMN94, Hig98a, HLC92, Hou95, IRC98a, Jud98a, Kad98a, Las94a, LZ98a, LZ98b, Liu94, LWT94, LKC97, Mac94, ME96, Mor91a, MLRR92, MP99, NA99, Niu96, NGG97, QH98a, Riv98, SW95a, San98a, Sil97a, Sto91a, Sto91b, TDW95, WEST91, DH93, Las94b, O’C99]. Applied [CP99, MR97a, MR97b, dCP99b]. Apportionment [Sch91b]. Approach [Amm93, BR94a, Bro97, CP91, CPS92, CKLH92, CT98, CM97b, Cre94, DF98, DFGJ93, DG92, EH91, Fon92, FK94, Fyg97, GMP97, GNK96, GJP96, GWB99, Gut94b, HW95a, Han94, HL98, HZ98, HL94, HDG97, Hon99, IL94, KDG94, KA91, Kol96, KS97, LaR91a, LK97, LMS99, Lav91a, LT98, Lju93a, Lju93b, MM93, MV91, NCDF96, NGH^{+}95, O’S90, Pai97, Pre96b, Raz90a, Raz90b, RCL96, SH92, Shu97, SK97, SG97, Tho91b, WW91a, WLGC98, WZ97b, WK98, YL97, ZK91, HW94, KT94a]. Approaches [CG90, DN98a, DN98b, GS90a, SR94]. Appropriateness [Ueb93]. Approximate [Ano91e, BC93, CS90a, Che90, Hsi97, KT94b, KT94c, TL97, KT94d]. Approximately [Agr98a, Agr98b, Ano99h, BB98b, BB98c, EMM95, MPS98a, MPS98b, Sev94, SW98b, SW98c, SW98d, Wie98]. Approximating [GW91, HL97]. Approximation [BS98, GJ99b, Jen91, Jen95, Lie94, Pes93, Shu97]. Approximations [AHY93, BDM97, CT90, DKRW97, Erk94, FG94, GR96a, Haw93a, JS94, Joe95, PWK99, SGP98, WBB93, dHY93]. Archimedean [GR93]. Area [FG91, GNSC98, HH91, LR95a, MSML97, PR90b, Rag93]. Areas [WC91]. ARIMA [CNS90, TR97]. Arising [Bro92d, EFS92a, EFS92b, GMD95, Gre92, Har92b, Lou92b, OBW92, ZDL96]. ARMA [AP96, GL92, JV93, Mau95, NP95, Roc92]. Arms [RG94]. Army [Hod94]. Array [Tan93]. Array-Based [Tan93]. Arrays [WW91a]. Arrest [CNWW98a, CNWW98b]. Artificial [DM90, HD97]. Arts [Moo98a, Moo98b]. ARX [CT93b]. Ascertained [Wan92]. Ascertainment [SPM99]. Asked [Bin98b, Bin98a, Bre98b, Bre98a, GKL98c, GKL98d, GKL98a, GKL98b, Jud98b, Jud98a, Kad98b, Kad98a, San98b, San98a]. Aspects [HH91]. Assays [HZS^{+}90, Min93, ZDL96]. Assessed [CG92]. Assessing [CT90, CW97, Gle92, GCW96, HW90a, HL94, HS93b, HS93c, Kra90, LA93, QN98, Qui98a, Sch93a, SS91b, Vel98a, Ylv93]. Assessment [GBA^{+}99, GJR98, KM91b, MO98, MP99, ZT97]. Assignment [CH91a]. Assoc [Fab94, LB94d, MS98a, MS95a, SG95a, Sta94a, Wee94, Wes98, Whi92b]. Association [Alb97, ALHH92, Goo96, Hab95b, KMT99, KA95, LMS99, Lau92b, Yam90]. Association-Marginal [LMS99]. Associations [SK93a]. Assumption [CB97, PHS94, Rom90]. Assumptions [RT96a]. Assurance [Riv98]. Astronomical [VB97]. Astronomy [AML95, WM90]. Asymmetric [Efr92].

Asymmetrical [Sut93, WW91a]. Asymmetry [KP97]. Asymptotic [Cha93a, CD96, DKRW97, MR91, Whi89, Whi92a, WCH98, XLR91, Fab94, Whi92b]. Asymptotics [Ano97i, CKLS96, Ste95, SCW96, SCW97]. Athletes [MH94b]. Atlantic [NSS97]. ATS [CMM93]. Attainment [AK92]. Attainments [WHGA96]. Attenuation [SKC94]. Attraction [HW92b, Wan95]. Auditing [CS96b]. Augmentation [HBH93, LW99, WT90a]. author [Efr94b, Efr96c, Fab91b, Goo91b, Las94b, Li91b]. authors [DG93b, RHS93b, SCM91b, WMR91b]. Autocorrelated [BD91b, HZ90]. Autocorrelations [CH95b, RT96a]. Automated [EL97, Joh96a, OR98]. Automatic [Cho93, GKK91, Gu93]. Autoregression [DM95, TS90]. Autoregressions [SG97]. Autoregressive [AR90, Ahn97, Boc99, BDR92, CT93a, GH99, LRM96, LL97c, MT93, Qui98b, SL93a, Sch96c, SS90b, Ter94, YR95]. Auxiliary [Hig98a, PS99, WPD98]. Average [AI95, BDR92, HMT92, LL97c, SL93a, TR97, YR95, hHMT93]. Averaged [CHR94]. Averages [RP96]. Averaging [RMH97, RB90]. AZT [RG94].

B [Efr94b, HS98c, HS98d]. B-Spline [HS98c, HS98d]. Back [Ano90a, Ano90b, Ano90c, Ano90d, Ano91a, Ano91b, Ano91c, Ano91d, Ano92a, Ano92b, Ano92c, Ano92d, Ano93a, Ano93b, Ano93c, Ano93d, Ano94a, Ano94b, Ano94c, Ano94d, Ano95a, Ano95b, Ano95c, Ano95d, Ano96b, Ano96c, Ano96d, Ano96e, Ano97b, Ano97c, Ano97d, Ano97e, Ano98b, Ano98c, Ano98d, Ano98e, Ano99b, Ano99c, Ano99d, Ano99e]. Bad [GS97]. Bahadur [Ano97j, Ruk93]. Balance [GJDT98a, GJDT98b]. Balanced [CD91, KW96b, Pen97, RS96a, SCC98]. Ban [PB92]. Bands [BD93, ES93, FS95a, HMY97, Owe95]. Bandwidth [Ano96f, BGH93, FJ90, HMT92, JMS96, OP96, OR98, PM90, RSW95, Sch95b, WW93, jW97, hHMT93]. Bandwidths [Rup97]. Base [Pea91, SS90a, WW91b]. Baseball [Alb93a, Alb94, AASM93, Alb93c, BH93b, SM93b]. Based [Ano91f, BAP95, BPL97, BY95, BW94, BO97, BF93a, BRW99, CM93, CS93, CH95b, CH93a, CD96, DM90, DM93, DM96, Efr92, Fan96, GH99, GS90a, GM99, GW99, HYP97, HMY97, Hol99a, Hös94a, Hua97, Hun94a, Joe95, Joh90, Kai94, KBB93, KS94, LRR91, LQT97, Lin91, LS93, LT96b, LS97, MS99a, MSH90, MSH93, McW90, Mee90, Mul99, Mur95, OP96, Pap96, Pas93, Pes93, PMW93, RAGW93a, RAGW93b, SDW98, SJ92, SKW99a, SKW99b, Sta90, SG97, Tab90, Tan93, VS93, WH90, WZ97b, Yam90, DR98]. Baseline [GWB99, Peñ98]. Basis [Kra90]. Bates [LB94c]. Batting [CB94]. Bayes [AC95a, ALM97, BP96, BM99, BPR95, BD91b, CG90, CS96a, CLWZ95, DLML99, DKRW97, Efr96a, Efr96b, Efr96c, Gel96a, Gel96b, Geo96a, Geo96b, Gho92, Goo92, Hsi97, KR95, LCG^{+}92b, LCG92a, LR97c, Lou92a, Mor96a, Mor96b, NS93, NT92, PWK99, Pen97, Pet92, Rag93, Rei96a, Rei96b, ST96a, ST96b, Sob91, Sob93, TS92, Tom94, VS93, VW95, Vid98]. Bayes/Non [Goo92]. Bayes/Non-Bayes [Goo92]. Bayesian [AC93, Alb97, AC97b, AG96, BH93a, BS93, Ber91b, BWC98a, BWC98b, Buc95c, CO90, CS91, CMW97, CKLH92, Che94a, CIS99, CKM97, CGM98a, CGM98b, CGM98c, CRS^{+}91, CNWW98a, CNWW98b, CV95, CCC96, CMMY91, DK99, DKW97, DKS97, DSM98a, DSM98b, EGMP93, EGG95, EW95, EGGS97, FOS97, FS98b, Fon92, FR99, FS94, GHRPS90, GSL92, GNK96, GB98, HW95a, Hay94, Hei95, HH91, IL91, Joh96a, KR96, KW95, KP92, KKT98a, KKT98b, Kol99, KLW94, KY96, Lav91a, LWW91, Lav91c, LRM96, LB94a, Liu94, LNL95, Liu96, MT93, McC98b, McC98c, MV91, MR97a, MG97a, MR97b, NSS97, NCC96, OKS97a, PJT96, PSS93, PS96, Per97, PS94b, Qui98a, RMH97, Rig97, RW97, Rue95, SR94, SN96, Sev94, Sin93, SD97, SK97, SM92, SKF98, SY95, SHGL96, SCF93, TW98a, TW98b]. Bayesian [TW98c, Tom96, TS96, Var98a, Var98b, VK92, Vil90, Wak96, WB96, Weg91, Wei97, Wei95b, Wes95, WK98, Zha98a, Zha98c, DE96]. Bayesians [BD91a]. Be [PH90b, SN96]. Bearings [ASL96]. Bearings-Only [ASL96]. Behavior [AL94, Cha93a, GS93a, Hal94, Man90, Rom90, MY95]. Behavioral [EGG95]. Beliefs [SHBF94]. Believe [CB94]. Benchmarking [CCD97]. Benefits [GM93a]. Bernardo [DE96]. Best [Man90, Rol94, SC96a, dCP99a]. Best-Case [Man90]. Best-Choice [SC96a]. Beta [CW91a]. Better [Cha91a, RS90a, SN96]. Between [BDR92, GR97a, GR97b, HH90, HHS97, HPS98, HH94, HS93b, HS93c, HHH92, LS91b, Llo90, Rig97, Sch93a, SS90a, Ylv93, BJB95]. Between-Block [HS93b, HS93c, Sch93a, Ylv93]. Beyond [LT94b]. Bias [CDW93, JS97a, LT97, LB96, MW98a, MW98b, MW98c, O’N94, PMW93, RW98a, RW98b, Rup97, SS90b, Sto93, WLGC98]. Biased [BBH96, GS91b]. Bifurcation [KR93]. Bilinear [Yam90]. Binary [Agr97, AC93, BO95, CLWZ95, Coo96, CL99b, FSE98, FLZ96, GK96, GM96, HL98, IRC98a, IRC98b, LLR95, LR97a, MSML97, McC94, Neu99, NCC96, OK96, PHL98, QN98, RGH99b, RGH99a, RF99, SW93, SB95, Was99a, WK98, ZDL96, ZL99, Zha98b, Fyg97]. Binned [Min98]. Binomial [CB94, Che90, CV95, Dea92, Hal94, Pes93, SK90, VM96, KT94a]. Bioequivalence [MP99, Wan99, BCH95]. Biological [GMP97, LNL99]. Biomedical [GS90b]. Birth [Hof92]. Bivariate [ALHH92, BHNO92, CR92, CM91a, ES99, FL95, GR93, Gür96, Hay94, LK95, Pru93, Ryu93, SV93, SK97, WJ93, WY96, WKL95]. Blind [LC95b]. Block [HS93b, HS93c, JMS97, Sch93a, Ylv93]. Blocking [BS91a]. Blocks [Bre91, Har91, Hoc91, SCM91a, Spe91, SCM91b]. Blood [HL94, MR97a, MR97b, OO97]. Bodies [PC96, Riv98]. Body [DMLC97]. Bomb [PSVS92]. Bone [DwSH94]. Book [AA94a, AA94b, AA94c, AA94d, AA94e, AA94f, AA94g, AA94h, AA94i, AA94j, AA94k, AA95a, AA95b, AA95c, AA95d, AA95e, AA96, AB97a, AJM99, Ack91, Ack93, Ack98, Ahn99, Ahs90, Alb93b, Alb95, Ale98, Alt94, Ame99, Ame92, And95, Ang93, Ang90, Ano90i, Ano90j, Ano90k, Ano90l, Ano91q, Ano91r, Ano91s, Ano91t, Ano92n, Ano92o, Ano92p, Ano92q, Ano93k, Ano93l, Ano93m, Ano93n, Ano94j, Ano94k, Ano94l, Ano94m, Ano95l, Ano95m, Ano95n, Ano95o, Ano96n, Ano96o, Ano96p, Ano96q, Ano97f, Ano97g, Ano97h, Ano97q, Ano97r, Ano97s, Ano97t, Ano98f, Ano98g, Ano99f, Ano99g, Anr96, App97, Asm93, AS93, Ath99, Atr94, BB98a, BBB98, Bag90, Bag91, Bal94, Bal96, Bal97, Ban90, Ban93, Bar93a, Bar97, Bar92, Bar98, Bec91, Bec98, Bec99, Bel93b]. Book [Ben97, Ben91a, Ber94a, Ber94b, Ber93, Ber95, Ber92a, Ber92b, Bha94, Bha90, Bie96, Big98, Bil91, Bil96, Bla92, Blo98, Boi90, Bol91, Bol94, Boo94, Boo99, Bov98, Bra98, Bre96b, Bri98, Bri99, BGH94, Bro99, Bro90, Bro91, Bro92c, Bry97, Bud96, Bun96, Bur92, BS91b, Buz97, Car90a, Car94a, Car90b, Cha96a, Cha92, Cha91b, Cha93b, Cha91c, Che91, Che92b, Che92a, Che95b, Che98, Chi91, Chi90, CS95b, Cla90b, Cla90a, Cla92, Cla98, Cla93, Cla99, Cle99, Cli99, Clo94, Col91, Col93, Col97, Con97, Coo91b, Coo93, Coo95, Cor99, Cor92, Cow91, Cow98, Cox99, Cra95, CL94, Cry99, Cwi94, D’A93, D’A94a, D’A94b, D’A94c, D’A94d, D’A95a, D’A95b, D’A95c, D’A96a, D’A96b, DSS97a, DSS97b, DSS97c, DSS98, DSS99, Dab92, Das95, Dav92, Dav99, Dav91]. Book [Dav98, DeG95, Dev95, DiC95, Dia99, DE96, Dic96, Die93, Die97, Dog95, Dra94a, Dra94b, EJD97a, EJD97b, EJD98a, EJD98b, Eas92, Edd90, Ela91, Eld96, Eli98, Emb95, Emb96a, Emb96b, Emb97a, Emb97b, Emb98a, Emb98b, Emb98c, Emb98d, Eub93, FH95, Far97, Fau90, Fau94, Fer98a, Fie93b, Fis91, Fla90, Fla91, Fle90, Fle91, Fou92, Fou93, Fra99, FM93, FKOR98, Fre95, Fre98, Fre99b, Fri98, Fun90, Fun92, Gas91, Gas95, Gas96, Gav91, Gen93, Gen99, Gho90, Gho94, Gho96, Gho98, Ghy98, GH91, Gil94, Gin95, Gla91, Gly98, Gol95, Gol93, Goo94a, Gra91, Gre91, Gro92, Gru93, Gru98, Gue96, Guo93, Gut90a, Gut90b, Gut91a, Gut91b, Gut92a, Gut93a, Gut99, Gut90c, Gut92b, Gut92c, Gut93b, Haa95a, Haa96, Hab95a, Hag97, HRF96, Hal97, Ham90a]. Book [Ham93, Ham90b, Har96, Hau97a, Hau97b, Hau90a, Hau90b, Hav95, Haw93b, Haw96, Haz95, Hea90, Hei90, Her98, Her93, Her97, Her91, Hig98b, Hil92, Hil91, Hil95, Hip95, Hod91, Hol98, Hol99b, Hol99c, Hos97, Hsi90, Igl91, Igl92, Iye97a, Iye97b, JEB98, JFM98, JGB99, Jam98, Jan95, Jas92a, Jas92b, Joh95, Jon90, Jor96, KB97a, KB97b, KK94a, KK94b, KK94c, KK95a, KK95b, KK95c, KK95d, KK95e, KK96a, KK96b, KK96c, KK97a, KK97b, KK97c, KK97d, KK97e, KK97f, KK98a, KK98b, KK98c, KK98d, KK98e, KK98f, KK98g, KK98h, KNB98, Kab98, Kad90a, Kad91, Kaf94, Kaf95, Kaf97a, Kaf97b, Kaf97c, Kaf97d, Kaf97e, Kaf97f, Kaf97g, Kaf97h, Kaf99a, Kaf99b, Kan90, Kar90, Kas97a, Kas90, Kas97b, Kat97a, Kat97b, Kaw99, Kep97, Kfa97, Khu94, Kia90, Kir98, Klu94, Koe97, Koo91, Kot93]. Book [Koy91, Kri94, Kro90, Kul95a, Kup90, LD96a, LD96b, LD97a, LD97b, LD98a, LD98b, LDH99, LH96, LH97a, LH97b, LH97c, LH97d, LH97e, LaV90, Lac90a, Lac90b, Lac90c, Lac90d, Lac91, Lac92, Lac93a, Lac93b, Lac93c, Lam97, Lan93, Lan95a, Lan95b, Lan97, Lar91b, Lau92a, Lee93, Lee96b, Leh97, Leh98, Lel99, Len98, Len99a, Len99b, Len99c, Len99d, Leo92a, Leo92b, Lev94, Li98d, Li99, Lig98, Lig94, Lin98, Lin94, Lin99a, Lin99b, Lip95, Lit94, Lit96, Liu97, Lju95, Loc90, Loh96, Lon91, Lor92, Lou99, Lou96b, LuV91, Luc98, Lun97a, Lun97b, Lun98b, Lun98a, Lun98c, Lun98d, Lun99, Lye97, Lyn98, MTW96a, MTW96b, MTW96c, MTW96d, MTW97a, MTW97b, MTW97c, MTW98a, MTW98b, MTW98c, MTW98d, MTW98e, MTW98f, MTW99a, MTW99b, MTW99c, MTW99d, MTW99e, MTW99f, MTW99g, MTW99h, MTW99i, MTW99j, MTW99k, MTW99l, MTW99m]. Book [MTW99n, MW95a, MW95b, MW95c, MW96a, MW96b, MW96c, MW96d, MW97a, MW97b, MW97c, MW97d, MW97e, MW98d, MW98e, Mad93, Mal95, Mar94, Mar90a, Mar92a, Mar98a, Mar95a, MA95, Mar90b, Mar95b, Mar98b, Mar98c, Mar92c, Mar98d, Mas92, Mas98, Mat99, Mat94, Mat93, McC98a, McD90, McD92, McD97, McG90, McK92, McK95, McN97, McN98a, McN98b, Mee95, MS90, Mel90, Meu95, Mey99, Mic97, Mic98b, Mic98a, Mik93, Mik94, Mik95a, Mik95b, Mik97, Mik98, Mit94, Mod95, Moo91, Moo99, Mor90, Mor91b, Mor93a, Mor93b, Mou95, Muk90, Muk92, Mul91, Mul93, Mul98, Mur97, Mur90, Mur91, NL94a, NL94b, NL94c, NL95, Nac91, Nak92, Nar93, Nei97, Nel92, Nel97, Nel98a, Nel98b, Neu92, New92, New90, Nol92, Not90, Not94, O’C99, O’R91, Oeh93a, Ogu99, Olk91, Olk92a, Olk92b, Olk93]. Book [Olk97, Olk99, Ols90, Omo95, Ove91, Ove94, Owe92, PAL96a, PAL96b, PD99, PE94a, PE94b, PE94c, PE94d, PE95a, PE95b, PE95c, PE95d, PE95e, PE95f, PE95g, PE95h, PE96a, PE96b, PE97a, PE97b, PE98a, PE98b, PE98c, PE98d, PE99, Pad93, Par91, Pau96, Ped94, Per96, dBPI97, Per90, Pfe99, Pic97, Pie93, Pin97, Pit99a, Pit99b, Ple91, Pod94, Pol90, Pot96, Pre94, Pre96a, Pre92, Puh98, Qua90, RAJ96, RBD96, RBL97a, RBL97b, RBL99a, RBL99b, RBL99c, RC99, RJ95, RJ96, RL96, RL98a, RL98b, RL98c, RLS99a, RLS99b, RS95a, RS95b, RS95c, RS95d, RS95e, RS95f, RS96b, RVL97a, RVL97b, RVL97c, RVL97d, RVL97e, RVL97f, RVL99a, RVL99b, RVL99c, Rad95, Ram93, Rat98, Rat94, Rau91, Rav97, Ree98, RN97a, Rei90, Ric90, Ric97b, Ric98, Rob90, Rob91, Roc93, Rod90, Roe92, Ros94, Ros99, Ros97a, Ros97b]. Book [Ruk98, Run99, Rut91, SRJ99a, SRJ99b, SRJ99c, SRJ99d, SRJ99e, SRJ99f, SRJ99g, SW94a, SW94b, SW95b, SW95c, SW95d, SW96, SW97b, SW97c, SW98e, SW98f, SW99, Sal93, Sam90a, Sam97, San91, San90, Sap91, Sar91a, Sar94, Sar91b, Sat90, Sca92, Sch96a, Sch91a, Sch94a, Sch90b, Sch96b, Sch95a, Sch94b, Sch90c, Sch94c, Sch98a, Sch98b, Sch99c, Sch97, Sch91d, Sch95e, Sco99, Sco90, Sea98, See95, See94, SM93a, SM99, Sel95, Sel96, Sen92, Sen93, Sen94, Ser93, Ser91, Set95, Sey97, Sey99, Sha92a, Sha99, Sha98, Sha92b, Sha93a, She95, Shl90, Sho92, Shu90, Shu93, Sim94, Sim93b, Sin97, Ski96, Sla93, Smi94, Smi96b, Smi96a, Smi98, Sob96, Sob98, Sol94, Spe93, Spe98, Sri93, Sta98a, Sta99, Sta98b, Sta93, Ste90a, Ste98, Ste90b, Ste92, Sti95, Sti92]. Book [Sto94, Sto97, Sto90, Str98a, Str99a, Str99b, Str99c, Str99d, Str95, SZ99, Str93b, Str94, Str98b, Stu92, Sud90, Sun99, Swa99, Swi90, Sys96, Tah92, Tak92, Tan90, Tas98, Tay90, Tei90, The93b, Thi94, Tho90b, Tiw99, Tro95b, Tsa90, Tsu94, Tur98, Tur90, Utt93, Utt94, Vau90, Vel98b, Vel99, Vin93, Vog92, Vog99, WP92, Wak91, Wal91, Wal95, Wal96, Wal98, War91, War97, Was97, Was98a, Was98b, Was90, Was91, Was93, Was94, Was99b, Wat90, Weg90a, Weh94, Wei90a, Wei91a, Wei98, Wei91b, Wei90b, Wei94, Wei95a, Wel99, Wel94, Wil91, Win96, Wol90, Wol93, Wol91, Won93, Won96, Woo90, Woo91, Woo93, Wu98, Wue96, Yan90, Yan98, You97, Zaj93, Zeh95, Zer96, Zer99, Zim92, dL93]. Bootstrap [AHY93, Ber90b, BBH94, Bre92, Bur94, CHP96, DM95, Efr90, Efr94a, ET97, FJ90, FH89, FH90, GMCM96, HH90, JB92, LT96b, LS97, MBL98, Mar90c, NG94, Pap96, PC96, PR94, Rub94, SS96b, Sha96, SW97a, SGP98, TS90, Zel93, ZD97, ZB92, dHY93, Efr94b]. Bootstrap-Adjusted [ZD97]. Bootstrap-Based [Pap96]. Bootstrapping [DC94, SW91a]. Borg [O’C99]. Born [HHH92]. Borrowed [Pri96]. Both [WR95, KT94a]. Bought [GS93a]. Bound [BF90, Sat95]. Boundary [CK92, Erk94, HW92a, Hsi97, ZKJ99]. Bounded [CH93b, HR94, MH90, MH94a, Ric97a, Tab90, DH93]. Bounded-Influence [HR94, MH90, Tab90]. Bounds [BP97, BHL93, CS96b, Spu99, WK92, DH93]. Bowhead [Buc95a, Buc95b, RGZ95a, RGZ95b, RGZ95c, RZ98, Sch95c, Sch95d, TL95a, TL95b, Wol95a, Wol95b]. ’Box [Jen95, Jen91, BDM97, ST99a, ST99b]. Box-Type [BDM97]. Brain [RTF92]. Brazil [Sas97]. Breakdown [Alb94, BG99, CMNS99, CH93b, EM92, HM97, HSP90, Hös94a, MH94a, MY91, SW95a, SR92, Str93a]. Breast [GWG90, Gra92, GS93b, LZ98a, LZ98b]. Breast-Cancer [GWG90]. Breast-Feeding [GS93b]. Bridge [MS96]. Bridging [Alb99, BRL99a, BRL99b, Kad99, Sch99a]. Brief [Goo92, JMS96]. Broken [GMP97]. Brownian [Ste94]. Burden [ZDL96]. Burst [EP94]. Bursts [LMR96]. Business [The93a].

C [Eld96]. Calculating [GS90a]. Calculation [SP91]. Calculations [MLRR92]. Calibrated [PB92]. Calibrating [BR95, Ber90a]. Calibration [CDW93, DS92, JK96, MZ97, Thé99, ZD97]. California [BdL99]. Callbacks [PMW93]. Camouflaged [OO97]. Campaign [MSL93]. Can [PH90b, Hjo94]. Canadian [SR95]. Cancer [Ano92e, BC90, Bak98b, Bak98a, BHB96, BSB98, GWG90, Gra92, LZ98a, LZ98b, MPCB95]. Canonical [BY95, GF96a]. Capital [Gro90]. Capture [KMT99, Pol91]. Capture-Recapture [KMT99]. Capturing [Soo94]. Carcinogenicity [HZS^{+}90]. Card [RCH94]. Care [CS91, DG99, NGG97]. Carlo [Kas97b, Ros95b, BBGL97, BS96b, CPS92, CL95, CC96b, GT95, GRS98, GJR98, Hig98a, Joh96b, Joh98b, LC98a, LC98b, MP95, NG94, Ode91, OB93, Ros95c, SMP95a, SMP95b, WT90a]. CART [CGM98a, CGM98b, CGM98c, DSM98a, DSM98b, KKT98a, KKT98b, Zha98a, Zha98c]. Case [AGM98a, Bak98b, BHB96, Bin98a, BWC98a, Bre98a, Bre96a, CGL93, CWW95, CNWW98a, DN98a, DC94, DMLC97, Efr97, FS95b, Fyg97, GKL98a, GKL98b, GJDT98a, GWG90, Hei91, Hod94, IRC98a, Jud98a, Kad98a, LZ98a, Man90, Mei94, Per97, QH98a, RCL96, RLN99, RSV95, San98a, Sev91, SHBF94, TFAH94]. Case- [FS95b]. Case-Control [BHB96, Bre96a, CGL93, CWW95, RCL96]. Case-Deletion [Per97]. Catch [Mil92]. Catchability [DFGJ93]. Categorical [BZ92, BCG99, Con92, Cro95, DS97a, DS97b, dE97, EGGS97, Fyg97, GB97a, GB97b, GH94, GH95, HZ98, LA94, Lan96, LMS99, LL97a, LL97b, PB94, SBD97a, SBD97b, SBD97c]. Categories [AMP90]. Categorization [Roj96]. Category [GS93a]. Causal [AI95, AIR96a, AIR96b, AIR96c, DW99, GM96, Hec96a, Hec96b, Mof96a, Mof96b, Neu99, RG96a, RG96b, RGH99b, RGH99a, Ros96a, Ros96b, RF99, SG97, Was99a, ZL99]. Causality [BDR92]. Cause [AKM94, GFT^{+}93]. Cause-Specific [AKM94, GFT^{+}93]. Caused [BL93]. CD4 [LCG^{+}92b, LCG92a, Lou92a, TS92, TDW95]. CDF [WBB93]. Cell [Ano96a, Bri92, Cox95, HS94, LCG^{+}92b, LCG92a, Lou92a, PB98, TS92]. Censored [AML95, AB97b, CW94a, Dab90, DS97a, DS97b, DG92, FG94, FSE98, GB97a, GB97b, GC91, GWG90, GL96, GY93, HW95a, Har92a, HP92, JMV94, KS97, LLE93, LL97a, LL97b, LYY99, LGWC93, LWT94, Pru93, SDW98, SBD97a, SBD97b, SBD97c, WH90, WT90b, WZ97a, WY96, Yan97, Yan99]. Censoring [GW99, Har90a, HR97, SW97a]. Census [CT98, CRS^{+}91, CH93a, DFGJ93, GBN96, Joh90, MS93, Pas93, RZ98, RAGW93a, RAGW93b, Sch93b, Sch91b, SS95b, Zas93]. Central [Ano91j, WK94]. Central-Mixture [Ano91j]. Century [Ket97b, Ket97a]. Certain [BHNO92, Bur94, Dha93]. Chain [BBGL97, BM95, CC96b, FK94, GT95, GRS98, GJR98, Hig98a, Joh96b, Joh98b, Kas97b, Lou96a, MTY95, NA99, Ros95c, RP96, Sam90b, Ros95b]. Chains [GJR98]. Change [AL95, BH93a, For94, KC93, Wan98a]. Changepoint [Bro97, CH97b]. Changepoints [CG97]. Changes [BP99, DDV99, IT94, JD91, KM96a, LD91]. Changing [GWB99, Nor90]. Chaotic [Ber91b, Lel94, MEGN92]. Characteristic [Hon99, Llo98]. Characteristics [GBN96, SLS97, SRR95]. Characterization [MLMT94, Vos93]. Characterizations [Chr91]. Characterizing [Hof92]. Charles [Eld96]. Chart [RP96]. Charts [Liu95, LT96b]. Checking [Pap96]. Checks [SGP98]. Chemicals [HZS^{+}90]. Chemotherapy [Min93]. Chi [HP92, Ryu94]. Chi-Squared [HP92, Ryu94]. Child [GR92a, Sas97]. Children [SB90, WHGA96]. Chilean [EV92]. Choice [BH93b, BD97, BGH93, FJ90, HMT92, Par93a, SC96a, Sch95b, Sev96, Yam90, hHMT93]. Choosing [DC94, KP91, Nyc91]. Circadian [PF97]. City [DXP92]. Claims [Coo91a, GHCF94]. Clas [Jen95]. Class [ES99, FKYZ99, For92, Jen91, Joh90, Lin91, LCG91, OKS97b, RLN99, Xio95]. Classes [BL92, CL92b, CH99, GMWZ99, HS98a]. Classical [Bra90, FOS97, GK93, LB94a]. Classification [BdL99, FS94, KP92, LT96a, Pre96b, SR95, Sta91, Zha98b, Eld96]. Classified [And91, BCdF91, Ben91b, CR91, Cox91, Goo91a, Goo96, Hab91, Goo91b]. Classifiers [Moj99]. Clinical [BO95, Bro92d, CCC96, DLB91, EF91a, EF91b, EFS92a, EFS92b, GFT^{+}93, GM96, Gre92, Har92b, LD92, LB94a, Lou92b, Mei91, OBW92, Roc95, Rub91, TFAH94, ZL91]. Clonogenic [Min93]. Closeness [Bly93a, CW93a, GKS93, MKSB90, Ped93, Rao93, RHS93a, RHS93b]. Cloud [CS93]. Clumps [HL97]. Cluster [Ano91h, DD90, NM98, RSY93, Tam95, Tho90c]. Clustered [CR92, GR92a, HS98b, HZ96]. Clustering [BR92, DR98, FL95, KST98, LJK93, MNS99, RMS90]. Clusters [NW97]. Clutter [BR98d, DR98]. Coale [XP92]. Coarse [HR90]. Coastal [PM97]. Cod [NSS97]. Codes [CRS^{+}91]. Coefficient [CT93a, CTC96, SW98a, WCH98, XL99]. Coefficients [CW94a, Pag90, RRZ94, Tho90a]. Coherent [Mei94]. Cointegration [Ahn97, BY95]. Collaborators [Ano95f, Ano97a, Ano98p]. Collapsed [Liu94]. College [MH94b]. Collinearity [FM92, SB97]. Column [Ye98b]. Combination [SGC95]. Combinations [SL93b, DH93]. Combining [CV95, DKRW97, DPWH99, Efr96a, Efr96b, Efr96c, FK99, Gel96a, Gel96b, Geo96a, Geo96b, GS91b, JK96, LT96a, LB94b, MSZ93, MH99, Moj99, Mor96a, Mor96b, NA99, Rei96a, Rei96b, ST96a, ST96b, Zas93, ZM93]. Comment [Agr98a, Agr98b, Alb93a, Alb99, Alh92, And91, Are92, Atk91, BCdF91, Bar93b, Bea91, Ben91b, Bin96a, Bin96b, Bin98b, Bin98a, Bly93a, BB98b, BB98c, Bre98b, Bre98a, Bre91, Bri91a, Bri91b, Bro92a, Bro92b, Bro92d, BRW99, Buc95a, Buc95b, BK99a, BK99b, CRS99b, CH91b, CW93a, Chr99a, Chr99b, CR91, CH93a, CH90, CW91b, Cox90, Cox91, Cre94, Cre97a, Cre97b, DH91, DSM98a, DSM98b, DS97a, DS97b, Dig99, Dor90, Dua91, Elt96a, Elt96b, EKWS91, Fab94, FZ98a, FZ98b, FZ99, Fie93a, Fie90, Fre99a, FORK99, Gel96a, Gel96b, GB97a, GB97b, Geo96a, Geo96b, GM99, GKS93, Gre92, Gut94b, GMS97a, GMS97b, Hab91, Hah90, HMN94, Han99a, Han99b, HT91, Har92b, Har91, He99a, He99b, Hec96a, Hec96b, Hoc91, Joh98a, Joh98c, Jud96a, Jud96b, Jud98b, Jud98a, Kad98b, Kad98a]. Comment [Kad99, KM90, Ken91, KDG94, Kin99, Kne98a, Kne98b, KKT98a, KKT98b, Koe99b, KR90b, LP99, Lan98a, Lan98b, Li98b, Li98c, LL97a, LL97b, Lit93b, LR99, LS99b, Lou92a, Lou92b, MS99a, Mac94, McC98b, McC98c, MS99c, McN92, MT98a, MT98b, MPS98a, MPS98b, Mei91, Mof96a, Mof96b, Mor91a, Mor96a, Mor96b, MZ91, Mul99, Nai91, Net90, Neu99, OBW92, OH99b, Pas93, Ped93, Por97a, Por97b, PM99b, PT90, RS98a, RS98b, Rao93, Rei96a, Rei96b, RG96a, RG96b, RW98a, RW98b, Ros96a, Ros96b, Rub91, Rub94, RF99, Rud91, RS90b, San98b, San98a, Sch93a, Sch90a, Sch99a, Sch95c, Sch95d, Sco91, ST96a, ST96b, Sha91, SC99a, SC99b, Sim93a, SHL98a, SHL98b, Spe91, SF97a, SF97b, Ste99a, Ste99b, SM93b, Stu99, TS92, Tou90, TL95a, TL95b, Var98a, Var98b, WS91, Wac93, WW98a]. Comment [WW98b, Was99a, WPD98, Wel93, Wil99, Wol95a, Wol95b, Ylv93, ZL91, ZL99, Zha98a, Zha98c, vdL99, Efr94b, Ano99h]. comments [Goo91b, HW94, Las94b, SCM91b]. Committee [GBA^{+}99]. Common [Agr98a, Agr98b, Ano99h, BB98b, BB98c, DD90, GF96a, HPS98, MPS98a, MPS98b, RN97b, RV98, Sto99, SW98b, SW98c, SW98d, VHE91]. Comparability [GH95]. Comparative [CC96b, HH91, WM91]. Compare [Llo98]. Comparing [AKM94, Bac90, Bra90, CS92a, DDV99, Hol95, Jac93, MH99, PW96, SRR95, Sob93, SK90, TL97]. Comparison [AD90, Bur94, CG92, CJ99, CHW92, FT94, HMN94, HA94, HA95, JD91, JS97a, Kul95b, Kun90, Las94a, LD91, LTW96, Mac94, MKSB90, PM90, Sev91, SN90, SKF98, SO94, WJ93, WPK99a, WPK99b, Las94b]. Comparisons [BP99, GB98, LTW96, ME96, Pen97, SMF94]. Competing [FW96, FG99, HW95b, Kra90, Mei94]. Competitions [Est92]. Competitive [Giv99]. Complementary [Ano96a, Cox95]. Completely [Pai97]. Complex [ALM97, BP94b, GK93, Riz92, Sit92, SR96, SRSD94]. Compliance [Bak98b, Bak98a, BP97, CCC96, EF91a, EF91b, GM96, Mei91, OMF^{+}93, Rub91, SCF93, ZL91]. Component [PWK99, Sch91c, SHGL96, Tar99, Wes95]. Components [Ahn97, BR94a, CD91, EL90, GP92, HA94, HS94, LB96, McC94, Roe94, She93, Vel98a, Wee91, HA95]. Composite [CS96a, DR91, HL98, KL97, LKM98, SS99, Wel90b]. Composition [NSS97]. Compositional [BR94a]. Compound [CW94c, WR94]. Compromise [Eas91, Goo92]. Computable [WR94]. Computation [Dha93, Hab95b, KY96, MG97a, Str93a, VHE91, WZ97a]. Computational [HH91]. Computations [BS98, Efr90, Liu94, WK94, WT90b]. Computed [SC94, Wei97]. Computer [CMMY91, KR76, KR90a, Sut93, Ye98b]. Computer-Intensive [Sut93]. Computerized [BWC98a, BWC98b]. Computing [DKRW97, ES90a, Hir92, LK92b, MBR91, VW95, WK92]. Concatenated [AMCS98]. Concave [GMPS99]. Concept [Soo94]. Conception [HW90b]. Conclusion [KSS96]. Concurrent [KL90]. Condition [HW90a]. Conditional [Ano99i, BBGL97, CW99, FK99, FP95, For94, GJ99b, GFT^{+}93, GH94, HWY99, HK92, Joe95, Joh92, KA95, KT94b, KT94d, LL97c, MBR91, PC97, Sch96c, Sev95, Soo92, SC97c, Val90, Val93, WM98, Xio95, KT94c]. Conditionally [AC97b, CMW97, Riz92, Yam90]. Conditioning [Mil92]. Conditions [Ros95c, Ros95b]. Cones [BM96]. Confidence [Ano91f, Ano95e, BS96a, BHS98, Ber90b, BB94, BHL93, BS96b, BD93, Bur94, CG90, CJ95a, CJ95b, CJ96, CH95b, CS96b, CW90, CHP96, ES93, FS95a, FH89, FHJW96, Gri92, Har90a, HH94, HMY97, Hol99a, HB99, LH91, Llo90, Mar90c, MZ97, Mee90, MP99, Mur95, Owe95, Pan97, PC96, Pes93, Roe90, Sat95, SG95b, Spu99, SW97a, Tho90a, TF90, VHE91, Wee93, WCH98, ZD97, BCH95, FH90, SG95a, Wee94]. Confirmation [Fun93]. Conflict [Luc93a]. Congenital [SHBF94]. Congressional [Sch91b]. Conjecture [SMF94]. Conjugate [Ano96h, CV92, GPS95, JJ93]. Consequences [GK90]. Considerations [PSVS92]. Consistency [GS98a, Tar99]. Consistent [FG95, Jia98, LB93, Riz92, ZL95]. Constrain [BHS98]. Constrained [BMR99, CC96a, CW94c, GSL92, Gey91, Gho92, MM93, RMS90]. Constraint [RS93]. Constraints [Ano98a, AG96, HT97d, RG93, Wes97b, Wes98]. Constructed [LK95]. Constructing [FHJW96]. Construction [BHL93, Hay94, Jac93, JMS97, MST97, WW91a, Wel90a]. Consumer [GS93a, Zie90]. Contagious [Hil93]. Contain [BS96a]. Containing [Pan97]. Containment [GLB98]. Contaminated [Gle93]. Content [CM91b]. Context [DF98]. Contextual [Pre96b]. Contextually [WM91]. Contingency [AMP90, Alb97, BS97a, BS97c, CS92b, DD90, Fri94, Goo94b, KP97, KA95, LW91, Qui98a, RG93, YF90]. Contingent [CH91a]. Continuous [CR92, FL95, GJR98, JV93, KC96]. Continuum [BS94]. Contoured [CN94]. Contracts [Ano98a]. Contrast [Wes97b]. Contrasts [Mar92b, Rub90, Spu99, Wes98]. Contributions [BR94a, SRR95]. Control [Ano90q, Ano92e, BC90, BHB96, BKP90, Bre96a, CGL93, CWW95, CNF95, CS92a, Cox90, Dor90, FIB90, FS95b, GJ94, GHMW94, GS94, HT90a, HT90b, KA91, Kra90, Liu95, LT96b, LCW98, Net90, PT90, RCL96, Rol90, RP96, SMP95a, SMP95b, SN90, TL97]. Controlled [GM96, LB94b]. Controlling [Owe94]. Controls [IRC98a, IRC98b, Jac93]. Controversy [Kra99]. Convergence [BPR99, Cha91a, CC96b, DG92, GJR98, Joh96b, Joh98b, Min98, Ros95c, ZM95, Ros95b]. Convex [FHTW97, Gey91, MP99]. Coordinates [Weg90b]. Copulas [GR93]. Correct [WPD98]. Correcting [MW98a, MW98b, MW98c, PMW93, RW98a, RW98b, Ye98a]. Correction [Ano91e, Ano91f, Ano91g, Ano91h, Ano91i, Ano91j, Ano92e, Ano92f, Ano95e, Ano96f, Ano96g, Ano97i, Ano98h, Ano98i, Ano98j, Cre90, FH90, HA95, Jen95, KR90a, KT94c, KT94d, LB96, LB94c, LB94d, Mar90c, Raz90a, Ros95b, Rub90, SKC94, SS92, Sil97b, SG95a, Sta90, Sta94a, Sta94b, SCW97, Val90, Wee94, Whi92a, hHMT93]. Corrections [Ano96h, Ano96i, Ano97j, Ano97k, Ano99h, Ano99i, MS95b]. Correlated [Alt90, Ano99a, BR94b, DLW95, KPR91, Llo90, ML94, QH98a, QH98b, Wan98b, ZT97]. Correlation [AP96, BY95, BJB95, CTC96, DBB^{+}94, GS98a, OV95, Sch94d, Sch91c, WEST91]. Correlations [Ano98a, Ber91a, BG91a, Gri92, Owe94, Wes97b, Wes98]. Correspondence [Goo96, RG93]. Corrupted [OP96]. Coseismic [KM96a]. Cost [Bak98b, Bak98a, GJ94, GLB98]. Cost-Effective [GJ94]. Cost-Effectiveness [Bak98b, Bak98a]. Count [McC98b, McC98c, MR97a, MR97b, PB98, TW98a, TW98b, TW98c, Var98a, Var98b]. Countries [GMWZ99]. Counts [CL95, DB97, Gro90, PB98, TDW95]. County [CCG^{+}97a, CCG^{+}97c, CCG^{+}97b, Cre97a, Cre97b, GMS97a, GMS97b, SF97a, SF97b]. Coupled [Joh96b]. Coupling [Joh98b]. Coupling-Regeneration [Joh98b]. Covariables [RCL96]. Covariance [AS90b, CD91, CD95, CLT96, CH99, DK99, GS91a, Hou95, JB92, MR91, Rig97, Roc92, SG92, Sie93, Ver93, WHG98, XLR91, YB97, ZB92]. Covariate [Fon92, FS95b, HHS97, HAS97, JT97, LY93, LIZ99, PF91, SH90, WWZtO97]. Covariate-Matched [HHS97]. Covariates [BL93, BZ92, BD93, CGL93, CL99a, CN94, FSE98, GWG90, HDG97, HH96, IRC98a, IRC98b, Neu98, PC93, Roc95, WB96, Zuc98, vR98]. Coverage [CL92b, CT98, CH93a, GWG98, Mar90c, Pas93, RAGW93a, RAGW93b]. Cox [BD93, Bur94, DwSH94, Lin91, LY93, Peñ98, Sas93, ST99a, ST99b, SH90].

Creation [DLW95]. Credence [O’H90]. Crime [Sta91]. Criminology [RLN99]. Criteria [BS91a, Jen91, Jen95, MT95a, ZM95, Zha92]. Criterion [Bly93b, Bly93a, CW93a, GKS93, HZ90, KW95, MNS99, Ped93, Rao93, RHS93a, RHS93b]. Critical [JB92, ZB92]. Criticism [CPP96]. Cross [And91, BCdF91, Ben91b, CR91, Cox91, ET97, Goo91a, Goo96, Hab91, HY98, KAT91, Par93a, RFB97, SBS94, Sha93b, Tho91a, Wan91, Goo91b]. Cross-Classified [And91, BCdF91, Ben91b, CR91, Cox91, Goo91a, Goo96, Hab91, Goo91b]. Cross-Sectional [Wan91]. Cross-Validated [Tho91a]. Cross-Validation [ET97, HY98, KAT91, RFB97, SBS94, Sha93b]. Cross-Validatory [Par93a]. Crossed [BR98b, BR98c, BR98a, FZ98a, FZ98b, Kne98a, Kne98b, Lan98a, Lan98b, RS98a, RS98b, WW98a, WW98b]. Crossover [Gui93]. Crystal [SW90]. Cumulant [PSS93]. Cumulative [AKM94, BK99a, BK99b, BS98, Chr99a, Chr99b, Han99a, Han99b, HLC92, IT94, LKCH99a, LKCH99b, LKCH99c, SC99a, SC99b, Ste99a, Ste99b, SW97a]. Cumulatives [RV90c]. Current [BF92, GS93b, NS93, RT96b, SK93b, vR98]. Current-Status [GS93b]. Curtailment [LYY99]. Curve [AC95a, AGM94, FHMP96, GK95, MP95, OO97, RS90a, RS93, Wan98a]. Curves [BR92, BR98b, BR98c, BR98a, CM97a, CM99, CW90, DBB^{+}94, DLW95, FZ98a, FZ98b, FL98, HYP97, Kne98a, Kne98b, KA91, Kul95b, KW97, Lan98a, Lan98b, Lee96a, Llo98, Pru93, RS98a, RS98b, WW98a, WW98b]. Customer [BZ99]. Cut [dE97]. Cut-Points [dE97]. Cycle [ZWWB96]. Cyclic [CH95b]. Cyclical [Wes95].

D [Kas97b]. Daily [Gro90, NCDF96]. Danish [HHH92]. Darkly [Bar91]. Data [AML95, AB97b, AC93, Alb94, Alt90, AC95b, And91, ASL91, ASL96, AD90, Ano97k, AS90b, BCdF91, BPS96, Bak98b, Bak98a, Ben91b, BF92, Bin96a, Bin96b, BW92, BCG99, BO97, BO95, BZ99, Bri92, BRW99, Car94b, CL92a, CH95a, CH97a, CW99, CS93, Cha96b, CTC96, CCD97, CIS99, CDS99, CW94a, Che94b, CMM93, CR91, Coo91a, CCC96, Cox91, Cro95, DwSH94, DH91, DAO94, Daw95, DS97a, DS97b, DM97, DLW95, DG92, DA96c, dE97, Efr98, Efr94a, EP94, EP99, Elt96a, Elt96b, EV92, Eub97, FL98, FSE98, Fay96a, Fay96c, Fay96b, FH89, FHJW96, FHTW97, For92, For94, GFT^{+}93, GHRPS90, GSL92, GB97a, GB97b, GC91, GK96, GM99, GBA^{+}99, GH94, GJP96, GWG90, Goo91a, Goo96, GRB^{+}95, GK93, Gra92, Gra95, GL96, GS93b, GY93, GR92a, GS95]. Data [Gür96, GT91, Hab91, HS98b, HW95a, Har92a, HL98, HW90b, HYP97, HBH93, HR90, Hei91, HLC92, Hir92, Hod94, HP92, HOS95, HL96, HW95b, Ibr90, IRC98a, IRC98b, JD91, JMV94, Joh96a, JV93, Jud96a, Jud96b, JS97b, KL97, KL99a, KM91b, KSW92, KP92, KLW94, KC96, LT97, LI98a, Lav91b, LR96, Led94, LD91, LWY93, LLE93, LS95, LRR91, Li92, Li95, LQT97, LL97a, LL97b, LYY99, LB94c, LIZ99, Lit93c, LS93, LGWC93, Liu96, LS97, LW99, LWT94, MS99a, MZ94, Man90, Mar92b, McC94, McC98b, McC98c, Mel99, Min98, ML94, MR97a, Moy99, MSK96, MR97b, MST97, Mul99, NA99, NP95, NT95, Niu96, O’S93, OP96, OK96, Pai97, PM90, PB94, PB98, PF91, PC97, PSVS92, PML98, Pre96b, Pru93, PHL98, QN98, Rao96c, Rao96a, Rao96b, RRZ95, RR95]. Data [RW96, RT96b, RB90, Rub94, Rub96c, Rub96b, Sas97, SDW98, Sch99b, Sch96c, SGC95, Seg92, SLS97, SS96b, SCC98, SS99, SBD97a, SBD97b, SBD97c, SC94, SJ92, SKW99a, SKW99b, Shu97, SG96b, SR95, Sin93, SD97, Sit92, SW93, SEH95, SM92, SM90b, SL98a, SL98b, SKF98, SC97c, SK93b, TCS94, TW98a, TW98b, TW98c, TDW95, TMP93, Var96, Var98a, Var98b, Vel95, VFR97, WH90, Wan91, Wan92, WSJ92, WWZtO97, Weg90b, WT90a, WT90b, WSLP90, Wei97, WZ97a, WY96, WW91b, WHG98, WKL95, WCH98, XXP93, YKB97, Ye98a, ZDL96, Zas93, ZLRS98, vR98, Efr94b, FH90, Goo91b, KT94a, LB94d, LB94e]. Data-Based [BRW99, GM99, MS99a, Mul99, SKW99a, SKW99b]. Data-Dependent [NP95, WSLP90]. Data-Driven [Efr98, KL97, KL99a, Led94, PM90]. Dave [CB94]. David [Eld96]. Dealing [PF91]. Death [Hof92, SKF98]. Death-Penalty [SKF98]. Decaying [Ber91a, BG91a]. Deciding [HPS98]. Decision [Giv99, HH94, LB94a, SM90b]. Decision-Theoretic [LB94a]. Decisions [Kad90b, SA92]. Decisive [GKB98]. Decomposition [BC96, KO96]. Decompositions [Amm93]. Deconvolution [CT97, EL97, LC95b, MO98, OO97, WW93]. Decriminalization [Mod93]. Default [BM99]. Defects [Sho93]. Definite [JS94]. Definiteness [GL92]. Definition [SW95a]. Deformable [AGP91]. Degrees [OR92, PWM92]. Delaney [Bro92a, Bro92b]. Delay [DXP92]. Delayed [BBH96]. Delays [Har90b, Wan92]. Deletion [Per97]. Demand [HK93, HK92]. Deming [Man94]. Demographic [Bel93a, CH93a, Pas93, Pol91, RAGW93a, RAGW93b]. Demography [DKW97]. Dennis [Hab95a]. Densities [Fre94, GR96a, GS90a, GR96b, Joh92, LK95, RV90c, SBS94]. Density [Atk91, BF93a, BF90, CCG^{+}97a, CCG^{+}97c, CCG^{+}97b, Che94a, Cho93, CB90, Cre97a, Cre97b, Dua91, Efr97, Efr98, EL97, EW95, FJ90, Gou97, Gu93, GMS97a, GMS97b, HYP97, Hon99, Ize91, JMS96, JS97a, Min98, MZ91, OP96, PO94, Roe90, RW97, Rud91, Rup97, SS96a, Sco91, SF97a, SF97b, Sto93, Ter90, VW95, WMR91a, WJ93, WY96, jW97, YM99, ZKJ99, WMR91b]. Density-Based [BF93a]. Dependence [Goo94b, HZ98, LR97b, QN98, RS94a, SV93, SJ92]. Dependent [AME96, Ada98, ALHH92, CLWZ95, FT94, FSE98, GW99, GV96, HDG97, HS94, LT96b, NP95, Pep91, PC93, SC97a, SC97b, SSK90, WSLP90, vR98]. Depends [Ye93]. Deposition [Haa90, Haa95b, Oeh93b]. Depressive [TFAH94]. Depth [CRS99a, CRS99b, He99a, He99b, Koe99a, Koe99b, LS93, LS97, LS99a, LS99b, MS99b, MS99c, OH99a, OH99b, PM99a, PM99b, RH99, RVH99, RHV99]. Derivation [GH95]. Derivative [HMT92, Sto93, hHMT93]. Derivatives [jW97]. Design [BdL99, CM93, CD92, CC96a, CM91b, CMMY91, EK93, Fan92, HT97a, HT97b, HT97c, Min93, MP95, Pen97, Por97a, Por97b, RG95, STR97, Tom96]. Design-adaptive [Fan92]. Designed [FR90, IL94]. Designing [Ano91h]. Designs [Akr90, Akr91a, AA94l, AAB97, AB97b, AC95c, BS91a, BPS95, BDM97, CC96a, CW94c, DL92, DPWH99, HYP97, Jac93, JMS97, JR96, KW96b, Kus98, LB94a, MS92, PR93, SR96, TA90, Tho91b, Tom94, VK92, Vos93, WZ97b, Wie98]. Destination [Var96]. Detect [RMS90]. Detecting [DR98, Sto99]. Detection [Ano92e, Atk94, BSPK98, BC90, FNP95, IT94, KM96a, LPT91, LZ98a, LZ98b, Sho93]. Determination [Ano95e, Fin90, JMS97, SG95b, SG95a]. Determine [BHS98, Est92]. Determining [Cre90, Fer98b, Roe94, Sch94e, VB97]. Deterministic [Buc95a, Buc95b, CMMY91, RGZ95a, RGZ95b, RGZ95c, Sch95c, Sch95d, TL95a, TL95b, Wol95a, Wol95b]. Deterrent [Gro90]. Developed [PH90b]. Development [Giv99, KS96]. Developmental [BO95, GMR99, SGC95]. Developments [Ize91]. Deviance [CJ96]. Deviation [Jen91, Jen95, RC93]. Deviation-Type [Jen91, Jen95]. Deviations [SB90]. Device [Pet92]. Diagnosing [Joh98b, dJP98]. Diagnosis [Bac90, TMP93, XXP93]. Diagnostic [Ano99a, CPP96, Eas94, LTW96, Llo98, QH98a, QH98b, SL93b, Wei95b]. Diagnostics [AC97b, BAP95, BF97, BVT96, CP91, CJ95a, Cho92, CT90, CHW92, CC96b, FMSP90, FM92, Fun95, GS92, Gu92, HLC92, LR95b, LR92, MSH90, PY99, SC92a, Soo92, SS91b, Tho91a, Vel95]. Dichotomous [CDS99, For94]. Dickey [VW95]. Dictionaries [Hau92]. Differ [LW91]. Difference [Ano98i, HH90, HHS97, Llo90, MS98b, Pes93, MS98a]. Differenced [PH90a]. Differences [GW99, YK96]. Different [Alb99, BRL99a, BRL99b, Kad99, Sch99a, SKF98]. Differential [RG94, RV90c]. Differentiation [AM92]. Diffusions [GT98]. Diltiazem [OMF^{+}93]. Dilution [ZDL96]. Dimension [Bri91a, CL92a, CW91b, CL99b, Fer98b, HT91, Ken91, Li91a, Li92, WR94, Li91b]. Dimensional [Daw95, DH99, DDD94, GS98a, O’S90, dCDdSO94]. Dimensionality [Bre92, Sch94e]. Dimensionless [Cho93, Gu93]. Dipoles [RTF92]. Direct [ALJ99, HH96]. Directional [FH89, FHJW96, PML98, Pre96b, FH90]. Directions [Coo98a, Coo98b, Coo98c, Li92, Li98b, Li98c]. Dirichlet [Esc94, GB98, MG97b]. Discharge [KR96]. Discharges [GM93a]. Disclosure [BKP90, DDD94, dCDdSO94]. Discontinuity [BdL99]. Discounts [CP93]. Discovery [Bil97]. Discrete [BPS96, BRMZR97, BD97, CR92, FW96, FL95, Hil93, HH96]. Discretization [GJR98, Moj99]. Discretized [DR96, O’S91]. Discriminant [BC96, BS97b, Fun95, HTB94, HM97]. Discrimination [BFH99, FBF97, KST98, KM91a, Li96, O’N94, PR93, Rig97]. Discussion [CRS99a, Gel94a, Gel94b, Koe99a, LS99a, MS99b, OH99a, PM99a, DG93b, Efr96c, Fab91b, Li91b, RHS93b, WMR91b]. Disease [LZ98a, LZ98b, LTP94, Mel99, PS93, SK93a, SH90, SHBF94, WCXG97, WHG98, YL97]. Diseases [BBH96]. Disorder [TFAH94]. Dispersion [GH95, LB96]. Displays [And91, BCdF91, Ben91b, CR91, Cox91, Fri94, Goo91a, Hab91, PC93, Goo91b].

Distance [CCB96, EL90, Har92a, Yam90, Dha93]. Distances [BPL97]. Distinguishability [SEH95]. Distribution [BS93, BK99a, BK99b, BS98, CH97a, Chr99a, Chr99b, CT97, FS98a, Fin90, FP95, FK94, Gel94a, Gel94b, GRB^{+}95, Hal94, HWY99, Han99a, Han99b, HRW90, HL97, Kai94, KM96b, KS93, LKCH99a, LKCH99b, LKCH99c, LMR96, LK92b, LRR91, Lie94, MZ96, ME96, McW90, Mer94a, Mer94b, Mer94c, ML94, MSK96, OV95, Owe95, Ryu93, SC99a, SC99b, SAK90, SP91, Ste99a, Ste99b, Sut93, Wan95]. Distribution-Free [Kai94, McW90]. Distributional [LK91, Zha92]. Distributions [ALHH92, AHY93, Ano98j, BPS96, BM92, Buc95c, Che95a, CF95, Cra94, DXP92, DJ98, EGGS97, FOS95, GBJ96, Hay94, Hir92, JS94, KO95, KW96a, KR93, LA94, LM93, NCDF96, RS94a, Sat95, SV93, Ste90c, VM96, WBB93, dHY93]. Divisional [BH93b]. DNA [Bal95, BGH97, DRR92, KM91b, NA99]. Do [CB94, CH94]. Dogma [Moy99]. Domain [HW92b, HZ90, Pap96, RTF92, SC92a, Ste95, Wan95]. Donald [Eld96, Efr94b]. Donoho [MY95]. Don’t [RSV95]. Dose [CNF95, GK96, HYP97, HB99, Lee96a, OK96, PSVS92, RG94, Rub90, Rub91, ZL91]. Dose- [HYP97]. Dose-Related [GK96, OK96]. Dose-Response [Lee96a, PSVS92, Rub91, ZL91]. Double [Ano91g, Buo90, Dor94, NW97]. Double-Scan [NW97]. Doubly [AML95, EP99, JMV94]. Down [Fin93, RCH94]. Draft [Ang91]. Drinking [MSL93]. Driven [Efr98, KL97, KL99a, Led94, PM90]. Driving [MSL93]. Drop [Dig99, FZ99, Fre99a, LP99, Lit95, LR99, SRR99a, SRR99b, vdL99]. Drop-Out [Dig99, FZ99, Fre99a, LP99, Lit95, LR99, SRR99a, SRR99b, vdL99]. Dropout [PHL98]. Dropouts [HMC^{+}93]. Drug [GJ94, Mod93, YK96]. Drug-Use [YK96]. Drugs [CP93]. Dual [AMWK93, BM96, SR96, Zas93]. Dual-System [AMWK93, Zas93]. Duration [FW96, Ryu94]. Durbin [SAK90]. Durbin-Watson [SAK90]. During [MSMP92]. Dynamic [ALJ99, BBGL97, Bol95, CMW97, Eas94, Fah92, FT94, Gri94, LC98a, LC98b, Now92, OKS97b, PSS99, SK91, SS91a, SS92, Wes95]. Dynamical [Gri90]. Dynamics [Buc95a, Buc95b, RGZ95a, RGZ95b, RGZ95c, Sch95c, Sch95d, TL95a, TL95b, Wol95a, Wol95b].

Each [BD91a, PB98]. Early [Ano92e, BC90, LZ98a, LZ98b, NGC^{+}95, SB90]. Earnings [RBD93]. Earthquakes [PS96]. Eaton [DH93]. Ecological [FKOR98, Kin99]. Economic [JS97b]. Edge [CGGM98a, CGGM98b, CGGM98c, HW91, Joh98a, Joh98c, MT98a, MT98b, SHL98a, SHL98b]. Edge-Preserving [CGGM98b, Joh98c, MT98b, SHL98b]. Editor [AG94, Ano91o, Ano91p, Ano92k, Ano92l, Ano92m, Ano93i, Ano93j, Ano94i, Dav96, Sek90, Ano98k, Ano99o, Ano99p, LHA94, LHAL95]. Editorial [Ano95f, Ano97a, Ano98p]. Editors [Ano95k, Ano99j, CCG90, CCWG91, CLWG92, LHA96, LCK97, LWG93]. Educational [AK92, ZT97]. Edwards [Man94]. EEG [WPK99a, WPK99b]. Effect [AK92, BW94, Gro90, HC96, MSMP92, MW98a, MW98b, MW98c, Mod93, Neu99, OMF^{+}93, Raz90a, Raz90b, RCH94, RGH99b, RGH99a, RW98a, RW98b, Ros93a, RF99, Was99a, WPD98, YKB97, Yas95, ZL99]. Effective [Ano96f, GJ94, Rub90, RSW95]. Effectiveness [Bak98b, Bak98a, SG96b]. Effects [AL95, AGM98a, AGM98b, AI95, AIR96a, AIR96b, AIR96c, Ano96g, BP97, Bro92b, BD91b, Car94b, CL93, CLM93, Coo91a, CM91b, Cro95, DW99, FW96, GHCF94, HW90a, HW91, Har90a, Hec96a, Hec96b, HDG97, Hol95, JTC99, LB94b, LB94c, MM95, MH99, Mof96a, Mof96b, MPCB95, RG96a, RG96b, Ros96a, Ros96b, Sch91b, SLS97, ST99a, ST99b, SH90, VL96, Wac93, WL93, Wei92, WM91, Ye98a, ZDL96, ZK91, ZWWB96, LB94d, LB94e]. Efficacy [Ano99a, GBA^{+}99, QH98a, QH98b]. Efficiencies [TA90]. Efficiency [Ano97j, BL93, DB97, LLR95, Neu98, RR95, Ruk93, STR97, Ski91, Wel90b]. Efficient [CH93b, Efr98, Efr90, ES97, Kus98, PS94a, Rob94, RK96, Sär96, SP91, ZT97, vR98]. Effort [Ode91]. Eigenvalue [BC96]. Elasticities [BG91b]. Elderly [CM91b]. Electoral [GK90]. Electorates [Gel94a, Gel94b, Mer94a, Mer94b, Mer94c]. Electrical [RTF92]. Electricity [HK93, HK92]. Electron [QTC96]. Elemental [Haw93a]. Elementary [DG92]. ELISA [GHMW94]. Elliptical [KBB93]. Elliptically [CN94]. Elongation [Gle93]. Emergency [Mod93]. Emission [MO98, Wei97, WEST91]. Emotions [Boc99]. Empirical [AC95a, Ano91h, ALM97, BD91b, CG90, CLWZ95, DH99, Efr96a, Efr96b, Efr96c, FP95, Gel96a, Gel96b, Geo96a, Geo96b, HMN94, Hon99, Las94a, Mac94, MS96, Mor96a, Mor96b, NS93, Rag93, Rei96a, Rei96b, Rup97, ST96a, ST96b, Sob93, SHBF94, VS93, Yan99, Las94b]. Empirical-Bias [Rup97]. Employer [Kad90b]. Employment [Yam92]. Endogenous [BJB95]. Endpoint [Pep91]. Endpoints [LTM92]. Ends [BHS98]. Enhancing [Wal93]. Enlistment [Ang91]. Enriching [Wal93]. Entropy [GJP96]. Entry [AK92]. Enumeration [AMWK93, BDM^{+}93a, BDM^{+}93b, Bel93a, HS93b, HS93c, Hog93, Lit93b, Sch93a, Wac93, Ylv93]. Envelope [Sto99]. Environmental [Niu96]. Epidemic [GJDT98a, GJDT98b, TMP93, XXP93]. Epidemiologic [BSB98, Mel99]. Epidemiology [Bre96a]. Episodes [Mod93]. Equal [Rig97, Ros93b]. Equation [LIZ99, Pai97]. Equations [BG91b, CRW98, HS98b, MBL98, MN93, O’N94, RV90c]. Equivalence [CC96a, CW94c, GS94, Mun96, Wan97]. Equivalent [PWM92]. Equivariant [MP99]. Era [Ang91]. Eras [Alb99, BRL99a, BRL99b, Kad99, Sch99a]. Errata [MS98a]. Erratic [Hal94]. Error [AS90a, Ano91g, BD97, Bre92, Buo90, Buo96, BS96c, CL92a, CKLS96, CS94, CNS90, Efr97, EKWS91, FS95b, GMP97, HJ92, HDG97, HTP95, JTC99, LR95a, MW98a, MW98b, MW98c, MT95a, MP96, MS91b, MS91c, O’N94, PR90b, PH90b, RW98a, RW98b, SR95, SB95, SC95, TDW95, WS91, WLGC98, WJ92, WW91b]. Errors [Alt90, Ano97k, BR94b, BH98, CGL93, CH97b, FS95a, Hil93, HOS95, HZ90, LS95, Now92, PSVS92, RBD93, RCL96, Sch94d, SK91, SS90a, SS95b, Wan98b, Wie98]. Errors-in- [Now92]. Errors-in-Variables [CH97b, LS95, PSVS92]. Essay [Joh96a]. Establishing [DMLC97]. Estimands [Rub91, ZL91]. Estimate [MZ96, Sta90, Zas93]. Estimated [FS98a, GL92]. Estimates [BS97c, BH94, CH91a, CW94a, Cox90, Dor90, Efr92, EKWS91, FKYZ99, Hab95b, HT90a, HT90b, Hös94a, LT96a, MSMP92, MY91, MS91a, MS95b, MS91b, MS91c, Net90, PT90, RN97b, Ros93a, Sev96, SRC92, SR95, WS91, WC91, MS95a]. Estimating [AMWK93, Bac90, Bak98b, Bak98a, Bal95, BBH96, BP94b, BBC98, BF90, BF93b, BR98d, CRW98, CL92b, CG92, DXP92, DA96c, Esc94, Est92, FG91, Fre94, GK90, GKB98, GR92a, GT91, Haa90, HS98a, HS98b, HWY99, HPS98, HW90b, HRW90, HW95b, HOS94b, Joh92, JK93, KS93, Lel94, LR97c, LW90a, LW90b, LIZ99, MEGN92, Mil92, MPCB95, NNR95, NCDF96, O’N94, Pag90, Pai97, RZ98, Raz90a, Raz90b, RS90a, RCH94, Riv98, RS93, Sch99b, SMP95a, SMP95b, TMP93, Var96, WPD98, XXP93, Yan97, YL97, ZDL96]. Estimation [AC97a, AR90, Akr91b, Alb97, AMWK93, AF97, Amr95, ASL96, AGM98a, AGM98b, AI95, AGM94, AS90b, ALM97, Atk91, BRR92, BHB96, BVT96, BDM^{+}93a, BDM^{+}93b, Bel93a, BM95, BP94b, Bin96a, Bin96b, BG91b, BCG99, BJB95, BRW99, BS96c, CD91, CD95, CK92, CS90a, CHR94, CB94, CDW93, CL95, CT98, CL93, Che94a, CS96a, Che94b, CD92, Cho93, CLM93, CH93a, CB90, CS94, CCB96, DK99, DFGJ93, DLML99, DD90, Dor94, Dua91, Efr97, Efr98, Efr99, EL97, Elt96a, Elt96b, EW95, Fah92, FJ90, Fay96c, Fay96b, FHTW97, Fon92, For94, FR90, GM99, Gho92, GNK96, GNSC98, GMPS99, GH95, GW99, GV96, GC95, GM94, GS93a, GMCM96, Gou97, Gri94, GL96, Gu93, Gür96, HMT92, HJ92, He90, Hoo93, HMC^{+}93, HH96, HL96, HR97, HM91, HH91, Ize91]. Estimation [JMV94, JMS96, Jud96a, Jud96b, KP91, KT95, KPR91, KP90, KA91, KS97, KS94, LR95a, LK97, LT97, LS95, Lee96a, Li95, LCG91, Lit93b, LM93, MS99a, Mau95, McC94, MSH93, MV91, MN93, Mei94, Min98, MP96, MS94, Muk96, MZ91, Mul99, MRvdV97, NS93, NSS97, Neu98, Neu99, O’S93, OP96, OR96, OKS97b, Par93a, PS94a, Pas93, PO94, PC96, PSS99, Pol91, PR90b, Rag93, RS96a, Rao96c, Rao96a, Rao96b, RTF92, Ric97a, RH96, RRZ94, RGH99b, RGH99a, RAGW93a, RAGW93b, Roe90, RW97, Rol94, RK96, Rub96c, Rub96b, RF99, Rud91, RV98, Rup97, SS96a, SR94, Sam93a, SG92, Sco91, SS94, SS99, She93, SC94, SKW99a, SKW99b, Sit97, Ski91, SR96, SS96c, SRSD94, Sta94b, SB95, SW98a, SW91a, Sto93, SI90, TFP90, Ter94, Ter90, Val93]. Estimation [Vel95, VFR97, Wac93, WMR91a, WJ93, Wan91, WR95, WWZtO97, Wan98a, WC99, Was99a, WJ92, WY96, WM90, WR94, jW97, YM99, YR95, ZL99, Zie90, hHMT93, vR98, Sta94a, WMR91b]. Estimator [AML95, AHY93, CKLS96, Cha91a, CH93b, Dor94, ES97, GS91a, GS91b, GR92b, LK92a, LR97c, Sev91, Sev96, Sit97, SM92, SO94, Tab90, WZ97a, YF90, ZKJ99, dHY93, MY95, YF95]. Estimators [AM92, Ber91a, BPR99, BGH93, DS92, DS91, Eas91, FHMP96, FS95b, GS91b, HW91, Hal94, Har92a, Jia98, JDP94, JS97a, KBB93, KCHH93, KAT91, LR95a, Lin91, MKSB90, PR90b, Riz92, Ryu94, SW95a, Sär96, Sas93, SM90a, SO94, SS90b, Thé99, Val90, VS93, WR94, Zha99, CRH94, Dha93].

Ethnic [YK96]. Ethnicity [Joh90]. Euclidean [Vil90]. Evaluating [Ano92e, Ano99a, BC90, BSB98, MSL93, QH98a, QH98b, SG96b]. Evaluation [BdL99, CS92b, CS90b, DW99, HZS^{+}90, Joh90, LB94a, Luc93b, SAK90, SS95b, SHBF94, Ter94, WPK99a, WPK99b, WC91, Zas93]. Event [DB97, HL96, Sin93, YK96]. Events [DR96, GKB98, HMC^{+}93, JTC99, NW97, Pep91, SH92, SPM99, Sol95]. Evidence [BGH97, CH94, GCW96, LS91b]. Evolution [PS96]. Exact [AMP90, Agr98a, Agr98b, Ano91g, Ano99h, BOP92, BB98b, BB98c, Buo90, Gui93, Hir92, KA95, Koo97, MS92, Mau95, MPS98a, MPS98b, SMP95a, SMP95b, ST92, Spu99, SK90, SW98b, SW98c, SW98d, VHE91, WS96, Hjo94]. Examination [BHB96, CW91a]. Examinations [LZ98a, LZ98b]. Example [Sta91]. Examples [GFT^{+}93, ZT97]. Excess [FP95, MS91a, MS95a, MS95b]. Exchangeable [BO95, GK96, OK96, QN98]. Exemplified [Gey91, Roe90]. Exercises [Hod94]. Exhaustion [GM93a]. Exhibitions [CH94]. Existence [LK92a]. Expansion [LW99]. Expect [BD91a]. Expectations [DM97, Erk94, Joe95, WK92]. Expected [CP91, Est92]. Expenditure [SM90b, Zie90]. Experiment [FR90]. Experimental [EGMP93, EK93, Min93, PR93, Tom96]. Experiments [CM93, CV95, CMMY91, DL92, HW95a, HPS98, HH91, IL94, MP95, NA99, Roc92, SGC95, ST92, SN90, Tom94, Ye98b]. Expert [CS95a, GMD95, Ueb93]. Experts [PJT96]. Explained [DBB^{+}94]. Explanatory [BJB95, EF91a, EF91b, Jon96, Mei91, Rub91, ZL91]. Exploration [CM99, WHGA96]. Exploratory [Sol91, Sol95]. Exploring [Alb94, HZ98, Sam93a]. Explosive [Sol91]. Exponent [MEGN92]. Exponential [Ano96h, CP99, CV92, CT90, GPS95, Hay94, KT94b, KT94d, LCG91, LM93, PSS93, Ryu93, Sla94, dCP99b, KT94c]. Exponentiality [dCP99a]. Exponentially [RP96]. Exponentials [CB90, HRW90]. Exposure [BRR92, BL93, CCG^{+}97a, CCG^{+}97c, CCG^{+}97b, Cre97a, Cre97b, GMS97a, GMS97b, Neu99, RGH99b, RGH99a, RF99, SK93a, SF97a, SF97b, SCF93, Was99a, ZL99]. Extended [Fah92, HS98b, Yan97]. Extension [LW90a, LW90b, Ryu93, Zuc98]. Extensions [Thé99]. Extra [NM98]. Extraction [CCD97, GWB99, KA91]. Extrapolation [CS94, DH91, Lav91b, SC95]. Extreme [HW92b, OV95, Zel93].

F [Ano91h, DE96]. Fabrication [Ano98h, HOLDG98]. Face [Fie90, Hah90, KR90b, Sch90a, SJ90a, SJ90b, Tou90]. Face-to-Face [Fie90, Hah90, KR90b, Sch90a, SJ90a, SJ90b, Tou90]. Faces [PS94b]. Facilitating [RT92]. Facilities [KR96]. Factor [Akr90, BP96, BD92, DL92, DH99, GL92, MS96]. Factorial [AA94l, AAB97, AB97b, BDM97, Tom94, Vos93]. Factors [BSB98, BM99, BPR95, Bre91, DKRW97, GF96b, Har91, Hoc91, Hsi97, HTP95, KR95, LR97c, PWK99, Pet92, SCM91a, Spe91, VW95, Vid98, SCM91b]. Failure [AMCS98, CKLH92, GFT^{+}93, GS95, HR97, LWY93, LW92, PC93, SDW98, SEH95, SL98a, Yam92]. Failure-Time [Yam92, HR97]. Failures [The93a]. Fairness [SKF98]. False [BR95]. False-Match [BR95]. Families [Ano96h, CV92, GNK96, GPS95, KT94b, KT94d, Sla94, KT94c]. Family [AGM98a, AGM98b, CW94b, CT90, Gle93, LK95, MSMP92, PSS93, WHG98]. Far [HHM92]. Fast [Atk94, FR99, GKK91, LB93, Luc93b, PY99, SAK90, VHE91]. Fat [FS98b]. Fatigue [CH95a]. FDG [OO97]. Feasible [MS94, SM92]. Features [AF97, BR98d, DR98]. Fecundability [HW90b]. Federal [Cox90, Dor90, HT90a, HT90b, Net90, PT90]. Feedback [RH96]. Feedforward [Whi89, Whi92a, Fab94, Whi92b]. Feeding [GS93b]. Fertility [SA92, XP92, ZWWB96]. Few [HA94, HA95]. Fields [Cre94, DKS97, Gut94b, Han94, KDG94, KO96, OKS97a, HW94]. Fieller [SW93]. Filter [GM94, SW91a]. Filtered [DJ98]. Filtered-Variate [DJ98]. Filtering [Fah92, Koo97, Li96, NNR95, PS99]. Filters [PS99]. Finetti [BM92]. Finetti-type [BM92]. Fingerprints [DRR92]. Finite [ALM97, BBH94, CDW93, Cra94, CCB96, HS98a, Lou96a, MV91, NS93, SW95a, SG96a, WS96]. Finite-Sample [SW95a, SG96a]. Finitely [KO95]. Firm [HT97d]. First [DM95, HW90b, Sch96c]. First-Order [DM95, Sch96c]. Fish [HAS97, Pol91]. Fisher [Goo96, Leh93, Par96]. Fishing [Mil92]. Fit [ACH99, AP99, BOP92, BPL97, CW94b, Chr91, ES90b, Eub97, HP92, Hua97, KL96, KM99b, LaR91a, Lan96, Led94, Lin91, MNS99, OR92, Pap96, Peñ98, SW91b, Wel90b]. Fits [CHW92]. Fitting [Ano91h, CH97a, HYP97, JV93, MS96, MP95, OR98, Sch96c, WCR94]. Fixed [Bre92, Qui98a, Ste95]. Fixed-Domain [Ste95]. Flat [LMR96]. Flexible [GKK91, Gra92, HTB94]. Floe [BR92]. Flow [SR95]. Fluctuations [PM97]. Focus [Weg91]. Follow [GLR93, MZ94]. Follow-Up [MZ94]. Follow-ups [GLR93]. Following [FR90]. Food [BD97, PH90b]. Football [CH94, GS98b]. Force [SG96b, SR95]. Forecast [CNS90, SS90a]. Forecasting [Alh92, CMW97, DR91, HK93, KL90, LCMA92, LC92, McN92, MZTT98]. Forecasts [AS90a, LT94b, SMD93]. Foregone [KSS96]. Forensic [Bal95, DRR92]. Forensics [Gei92]. Forest [RC94a, RC94b]. Form [CS90b, JMV94]. Formal [KW96a]. Forms [JS94, Lie94]. Formulation [Soo92]. Four [GNK96, HZS^{+}90]. Four-Person [GNK96]. Fourier [Bea91, Bri91b, LK91, Mor91a, O’S91, Sto91a, Sto91b]. Fraction [CIS99, Yam92]. Fractionally [LL97c, PH90a]. Fractionated [HW95a]. Fractions [KR93, SS99]. Frailty [LK97, Sas97]. Frame [FG91, Ski91, SR96]. Free [Kai94, McW90, RMS90]. Freedman [Kin99]. Freedom [OR92, PWM92]. Freeze [SC96a]. Frequencies [GC95, SHBF94]. Frequency [Coo91a, HZ90, Pap96, RTF92, SC92a]. Frequentist [CS91, HH91, SR94]. Frequently [GS93a]. Freshman [MH94b]. Friends [Yam90]. Front [Ano90e, Ano90f, Ano90g, Ano90h, Ano91k, Ano91l, Ano91m, Ano91n, Ano92g, Ano92h, Ano92i, Ano92j, Ano93e, Ano93f, Ano93g, Ano93h, Ano94e, Ano94f, Ano94g, Ano94h, Ano95g, Ano95h, Ano95i, Ano95j, Ano96j, Ano96k, Ano96l, Ano96m, Ano97l, Ano97m, Ano97n, Ano97o, Ano98l, Ano98m, Ano98n, Ano98o, Ano99k, Ano99l, Ano99m, Ano99n]. Frontier [GMPS99, PS94a]. Full [MS96, SC96a]. Full-Information [MS96, SC96a]. Fully [AA94l, OR98, Qui98b, Wei97]. Function [ACH99, Bjø96, BBC98, BRW99, BD93, Cha91a, CT97, FHTW97, GM99, GS93a, HWY99, Hon99, HMC^{+}93, KL99b, LaR91a, MS99a, Mul99, Owe95, Pap96, Raz90a, Raz90b, SKW99a, SKW99b, Sta90, SW91b, WC99, YL97, ZKJ99]. Function-Indexed [KL99b]. Functional [CT93a, GRB^{+}95, HHS97, Sam93a]. Functional-Coefficient [CT93a]. Functionals [Che94b, LC95a]. Functioning [CM91b]. Functions [AKM94, BPS96, BP94b, BS97b, BK99a, BK99b, BS98, CW94b, CM98, Chr99a, Chr99b, CV92, CH99, CMMY91, DS91, FW96, FHW95, FS98a, Fre94, GMPS99, GH95, Han99a, Han99b, HMY97, Hou95, HL96, KS93, LKCH99a, LKCH99b, LKCH99c, Lel94, LTW96, LW90a, LW90b, LWT94, MS94, Muk96, NNR95, OB93, Rue95, SC99a, SC99b, Sla94, Ste99a, Ste99b, SW97a, SG97, WK94, Yan97, Yan99]. Future [DXP92, Gla91, Kin99, Pol91]. Fuzzy [Hab95a].

G [FKOR98, Hab95a, Sta94a]. Galaxies [Roe90]. Games [EGMP93]. Gamma [EP94]. Gamma-Ray [EP94]. Gaussian [BC96, Ber91a, CMW97, CMM93, DKS97, Gu90, OKS97a, SW91a]. Gaussians [MZ96]. Gear [Mil92]. Gender [YK96]. Gene [LR96, Liu94]. General [Ano98a, Ano98p, Ano98q, BL92, Buo96, DDD94, GR96a, Goo96, GW91, HJ92, HR94, LYY99, Luc93b, MY91, MLRR92, SSW99, SC94, Vel98a, Wes97b, dCDdSO94, Wes98]. Generalizable [Soo92]. Generalization [Ano91f, Bjø96, EM90, Mee90, MSK96, VW95]. Generalizations [WM91]. Generalized [Ano91i, BCJ96, BdL99, BH98, BC93, BS96c, CFGW97, CH97a, CCD98, CRH94, DG99, DSS93, ES97, Fah92, FHW95, FM92, GS99, GNSC98, HS98b, HMM98, HLC92, Hon99, HN92, Ibr90, IL91, JGC95, Jia98, LR95b, LB96, LM93, MO98, McC97, MP96, MG97b, Neu98, Pai97, RSY93, RS94a, SMF94, SW91b, Tho90a, WL93, WLGC98, Wee91, Wee93, Wee94, YF90, ZK91]. Generated [LJK93]. Generating [BPS96]. Generation [AP99, KR76, KR90a]. Genetic [HZS^{+}90]. Geocode [GBN96]. Geographic [SS96c]. Geographically [MST97]. Geometric [Cha96b, Riv98]. Geometrical [HW91]. Geometry [CS93, Daw95]. Georgia [RC94a, RC94b]. Gibbs [KT94c, KT94d, Cha93a, Chi95, GHRPS90, GSL92, GS99, GM93b, HC96, KT94b, Liu94, LNL95, RT92, ZK91, ZM95]. Gilks [Kas97b]. Give [Sat95]. Glass [Bar91]. Global [GJ99a, WKL95]. Globally [FKYZ99]. GM [SRC92]. Good [GS97]. Goodness [AP99, BOP92, BPL97, CW94b, ES90b, Eub97, HP92, Hua97, KL96, KM99b, LaR91a, Lan96, LC93, Lin91, OR92, Peñ98, Wel90b]. Goodness-of-Fit [AP99, BOP92, BPL97, HP92, Hua97, LaR91a, Lan96, Lin91, OR92, Peñ98, Wel90b]. Gradient [JJ93]. Grading [Joh96a, NGH^{+}95]. Graduation [MSMP92, MH94b]. Graphical [And91, BCdF91, Ben91b, CR91, Cox91, Cre90, Eas94, GC91, Goo91a, Hab91, Lau92b, MR94, PC93, Roe94, Goo91b].

Graphics [Coo96, CW97]. Gray [GC95]. Griddy [RT92]. Griddy-Gibbs [RT92]. Groenen [O’C99]. Gross [SR95]. Ground [Niu96]. Ground-Level [Niu96]. Group [Ano92e, BC90, HOS94b, JT97, LB94a, PSS99, Rom90, Ryu94, STR97]. Group- [STR97]. Group-Sequential [JT97]. Group-Testing [HOS94b]. Grouped [HLC92, VFR97]. Groups [GF96a, Gra95, PW96, SRR95]. Growth [AC95a, CH92]. Guatemala [GR92a]. Guided [ES90a]. Guidelines [HTP95]. Gumbel [Wan95].

H [Fab94, Hab95a]. H. [Ano91h]. Haenszel [YF95]. Haenzel [YF90]. Halcion [GBA^{+}99]. Half [RS96a]. Half-Sample [RS96a]. Handcock [HW94]. Handling [Kot94, RSV95]. Harris [CCG^{+}97a, CCG^{+}97c, CCG^{+}97b, Cre97a, Cre97b, GMS97a, GMS97b, SF97a, SF97b]. Having [CV92]. Hazard [AME96, AKM94, BAP95, FW96, GS93a, GMCM96, KST95, KS97, Peñ98, Ryu94, SW97a, Yan99]. Hazard-Based [BAP95]. Hazards [CL99a, FG99, GR92a, HN92, Peñ98, SDW98, Sch99b]. Head [Bro92b, Wei92]. Headache [DPWH99]. Health [CM91b, DG99, GWG98, MSML97]. Heaping [HR90]. Heart [SHBF94]. Hellinger [CCB96]. Helmet [Bro92b, Wei92]. Hematopoiesis [NGC^{+}95]. Hemisphere [Sol91]. Heptathlon [DAO94]. Hessian [Coo98a, Coo98b, Coo98c, Li92, Li98b, Li98c]. Heterogeneity [AMWK93, GK96, GS93a, HS93b, HS93c, KMT99, KR96, OK96, Sch93a, VL96, Ylv93]. Heterogeneous [DFGJ93, DPWH99, PSS99]. Heteroscedastic [Ano98i, AC95c, DM96, FS95a, Hoo93, MS98b, WCR94, Wie98, MS98a]. Heteroscedasticity [LL97c, Roc92]. Heterosexual [GJDT98a, GJDT98b]. Hidden [Whi89, Whi92a, Fab94, Whi92b]. Hidden-Layer [Whi92a, Fab94, Whi92b]. Hierarchical [AC97b, BDM^{+}93a, BDM^{+}93b, BMR99, BZ99, CS96a, CM97c, CNWW98a, CNWW98b, DG99, DK99, DLML99, HK92, HC96, Joh92, LCG^{+}92b, LCG92a, Lit93b, Lou92a, MR97a, MR97b, PJT96, PS94b, Sil97a, Sta91, TS92, Wac93, WCXG97, Wes97a, WM91]. Hierarchy [SKF98]. High [CMNS99, CH93b, DH99, GS98a, HM97, Hös94a, LT94b, MSMP92, MH94a, RG94, Str93a, WR94]. High- [RG94]. High-Breakdown [CMNS99, HM97]. High-Dimensional [DH99]. Higher [JS97a, Min98]. Higher-Order [JS97a, Min98]. Highly [LWY93, Sam90b]. Hispanic [Joh90]. Histogram [DJ98, OP96]. Histograms [SC96b]. Historic [EV92]. Historical [GT91, IRC98a, IRC98b, VB97]. Hitting [Alb93a, Alb94, AASM93, Alb93c, SM93b]. HIV [DLD91, DA96c, GJ94, GHMW94, LI98a, LCG^{+}92b, LCG92a, Lou92a, SJ92, TS92]. HIV-1 [GHMW94]. HIV-Positive [LI98a]. HMOs [GLB98]. Hodges [Ros93a]. Hodges-Lehmann [Ros93a]. Home [GM93a]. Homicide [Gro90]. Homogeneity [Ano91j, CF95, JGC95, JB92, MM93, Mit91, NC99, Qui98a, Shi91, YF90, ZB92]. Homophily [Yam90]. Horizon [SC96a]. Hormone [GWB99]. Horvitz [SO94]. Hospital [Mod93, SRR95]. Hospitalization [LI98a]. Hot [JK93]. Hourly [HK93, RBD93]. Hours [RBD93]. Household [AL94, SA92]. HRS [JS97b]. Human [Ano92f, RTF92]. Hybrid [LW97, WZ97a]. Hybridization [ST92]. Hypercube [Owe94]. Hypercubes [Tan93, Ye98b]. Hyperdimensional [Weg90b]. Hypnotic [GBA^{+}99]. Hypotheses [AA94l, AAB97, Ano91i, Ano97j, BPR95, Bre90, GCW96, KW95, LK91, LTM92, Leh93, MBR98, Rob94, Ruk93, Sil92]. Hypothesis [AME96, BM99, BS97c, GK93, HRW90, Hon99, KL97, Kra99, SSC94, SC97a, SC97b, Tom96, Tro95a, Wel90b]. Hypothesis-Testing [Wel90b].

Ice [BR92]. Identifiability [GS99, Now92, Pru93]. Identification [AIR96a, AIR96b, AIR96c, Bal95, BR92, Bar93b, BG99, BWC98a, BWC98b, DG93a, Fie93a, GH99, HS93a, Hec96a, Hec96b, Mof96a, Mof96b, RG96a, RG96b, RW96, Ros96a, Ros96b, Sim93a, TA94a, TA94b, Wel93, DG93b]. Identified [LTM92]. Identifying [HYP97, LT98, Rub90]. Identities [Hou95]. IGF [HDG97]. ignorable [RRS98]. Ignored [KMT99]. Ignoring [Wac93]. iid [SM92]. Illicit [CP93]. Illustrated [DH91, Lav91b]. Illustration [GHRPS90]. Image [AGP91, AZ99, BO97, CGGM98a, CGGM98b, CGGM98c, FR99, GC95, Joh98a, Joh98c, Kol96, MT98a, MT98b, O’S90, O’S93, QTC96, SHL98a, SHL98b]. Images [BR92, CHR94, FS94, Joh94, Pre96b]. Imaginary [Pet92]. Imaging [Rue95]. Imbedding [Lou96a]. Immigration [GMWZ99]. Immunoassay [ZD97]. Immunodeficiency [Ano92f]. Immunologic [Hei91]. Impact [BHB96, BSB98, Kad90b]. Imperfect [Ano99a, BP97, BW94, PHS94, QH98a, QH98b]. Implementation [MNS99, WT90a]. Implication [STR97]. Implications [BAP95, Cla96]. Importance [Che94a, GR96b, Gle92, LCW98, OB93, Per97, Zha96]. Importance-Weighted [Che94a]. Improper [GS99, HC96]. Improved [CS92a, DH93, FHJW96, KA95, ZKJ99, Zha99]. Improvement [NGH^{+}95]. Improvements [ET97, YB97]. Improving [BS97a, CJ95b, JS97b, Pea91, WC91]. Impulse [SG97]. Imputation [BDM^{+}93a, BDM^{+}93b, Bin98b, Bin98a, Bre98b, Bre98a, CRS^{+}91, Efr94a, GKL98c, GKL98d, GKL98a, GKL98b, GLR93, HR90, HL94, Jud98b, Jud98a, Kad98b, Kad98a, Lit93b, Rub94, Rub96a, San98b, San98a, SCC98, SS99, Wac93, Efr94b]. Imputations [KLW94, LC95b]. Imputed [Bin96a, Bin96b, Elt96a, Elt96b, Fay96a, Fay96c, Fay96b, Jud96a, Jud96b, LRR91, Rao96c, Rao96a, Rao96b, Rub96c, Rub96b, SDW98, SS96b]. Incidence [AKM94, BBH96, DXP92, Har90b, MST97]. Including [LCG91, MLRR92]. Income [DDV99, DM97, GNK96]. Incomplete [AC95b, Car94b, FLZ96, Ibr90, KW96b, LY93, Lit93c, Liu96]. Incorporating [CS95a, JT97]. Increase [RZ98]. Increasing [GS95]. Incubation [Bac90, DA96c, HMC^{+}93]. Independence [BBGL97, CB97, GR97a, GK96, GR97b, Goo94b, KL99a, OK96]. Independent [AC97b, GF96a, JK96, LT96b, MSZ93, ZM93]. Index [Ano97p, BVT96, CFGW97, DR91, DMLC97, GRB^{+}95, HH96, LS93, XL99].

Index-Mortality [DMLC97]. Indexed [KL99b]. Indicator [Jon96]. Indicators [EH91]. Indices [CG92]. Indirect [OP96]. Individual [GBN96, Wan99]. Industry [CRS^{+}91]. Inequalities [Fin90, Hou95]. Inequality [Ede90, ME96]. Infarction [OMF^{+}93]. Infection [Ano92f, Bac90, DLD91, DXP92, LCG^{+}92b, LCG92a, Lou92a, TS92]. Infectivity [SJ92]. Inference [AMP90, Agr98a, Agr98b, Ahn97, Ano99h, AG96, BM96, BB98b, BB98c, BC93, Buc95a, Buc95b, Buc95c, CS91, CS95a, Cha90, CM93, CS93, DF98, DM95, DRR92, DKR95, EW95, FIB90, FOS95, FOS97, Fra91, FKOR98, GHRPS90, GR93, GT95, GM96, Gui93, HR90, KBB93, KA95, Kin99, KM99b, KT94b, KT94d, LWW91, LR96, LWT94, MSML97, MT93, McC98b, McC98c, MPS98a, MPS98b, Mel99, NCC96, O’H90, PJT96, Pep91, PSS93, PWM92, Qui98b, RGZ95a, RGZ95b, RGZ95c, RT96a, SDW98, Sch95c, Sch95d, Sev94, Sev95, SN90, SW98b, SW98c, SW98d, TW98a, TW98b, TW98c, TL95a, TL95b, Var98a, Var98b, Vil90, WH90, WSLP90, WS96, Wes95, Wol95a, Wol95b, YP99, KT94c]. Inferences [CH97b, Gus96, LQT97, MT95b, RK90, SK93a, SRC92]. Infinite [Cha90]. Infinitesimal [WZ97b]. Influence [BF97, BL92, CP91, CH93b, DS91, GMWZ99, HR94, LA93, LKC97, MH90, MH94a, Ric97a, Tab90, Tho90a, Tho91a, Wei95b]. Information [Ano90r, Ano91z, Ano92v, Ano93r, Ano94p, Ano95r, Ano96u, Ano97u, Ano98r, Ano99q, Bel93a, CB94, Cla96, DPWH99, GJP96, JT97, MS96, Par96, SC96a, SK93a, Sev95, Soo94, SEH95, VK92, WPD98, WC92]. Informative [FS94, GBJ96, PHL98, VS93]. Ingwer [O’C99]. Inhomogeneities [KM91b]. Inhomogeneous [AZ99]. Initial [Koo97]. Injuries [Bro92b, Wei92]. Inland [KSJ94]. Inmate [BdL99]. Innovations [DM95]. Insights [GS93a]. Institute [GBA^{+}99]. Instrumental [AK92, AIR96a, AIR96b, AIR96c, BJB95, BS96c, Hec96a, Hec96b, Mof96a, Mof96b, RG96a, RG96b, Ros96a, Ros96b, SB95]. Instruments [BJB95, ZT97]. Insurance [GWG98]. Intake [NCDF96]. Intangible [Soo94]. Integrals [CS90b]. Integrated [Cha90, KM99a, LL97c, SL93a]. Integration [Ano98i, MS98b, MG97a, OB93, MS98a]. Intensities [Var96]. Intensity [AI95, CHP96]. Intensive [Sut93]. Intent [Roc95]. Intent-to-Treat [Roc95]. Intentions [Man90]. Interaction [Ano96g, MM95]. Interactional [Fie90, Hah90, KR90b, Sch90a, SJ90a, SJ90b, Tou90]. Interactions [CH91b, Fab91a, Nai91, Sha91, Fab91b]. Interblock [TA90]. Interchangeability [ES99]. Interdependence [VB97]. Interdirection [JR96]. Interest [GT98, Ye93]. Intermittently [FSE98]. Interpoint [BPL97]. Interpolation [Ano97k, GM94, HT97a, HT97b, HT97c, HOS95, Por97a, Por97b]. Interpretation [Coo94, Gar94, MSH93]. Interpreting [Bre91, Har91, Hoc91, SCM91b, SCM91a, Spe91]. Interval [HR97, LH91, MV91, Mun96, Pes93, SDW98]. Intervals [Ano91f, Ano95e, BS96a, BHS98, BS96b, Bur94, CG90, Che90, CJ95a, CJ95b, CJ96, CH95b, DH99, Gri92, Har90a, HB99, HD97, Llo90, Mar90c, Mee90, Mur95, Sat95, SG95b, SW93, TS90, TF90, Wee93, ZD97, SG95a, Wee94]. Intervention [MSL93]. Interview [MSML97]. Interviews [Fie90, Hah90, KR90b, Sch90a, SJ90a, SJ90b, Tou90]. Intrablock [BS91a]. Intrasubject [Wan97]. Intrinsic [BP96, MBR98]. Introduction [Roc91, Rol90, Weg91]. Invariance [Rom90]. Invariant [CP99, HMO97a, HMO97b, Rol94, SAK90, dCP99b]. Inverse [Bri91a, CW91b, Fer98b, HT91, Ken91, Li91a, Sch94e, Li91b]. Investigating [Daw95]. Investigation [MS96]. Investment [HT97d]. Involving [CL95, PB98, PSS93, Sie93, ST99a, ST99b]. IQ [MW98a, MW98b, MW98c, RW98a, RW98b, WPD98]. Iraqi [DKW97]. Irrelevant [LT97]. Isotonic [Gey91]. Issues [Bro92d, EFS92a, EFS92b, Gei92, GS93a, Gre92, Har92b, Lou92b, NGG97, OBW92]. Item [LCG91, MS96]. Iterated [YM99]. Iteration [Mar90c]. Iterative [Hoo93, O’S90]. Iteratively [DS91].

J [DE96, Eld96, Fab94, Kas97b, LB94d, MS98a, MS95a, SG95a, Sta94a, Wee94, Wes98, Whi92b]. Jackknife [SC95]. James [GS91b]. Japan [Yam92]. Jeffreys [IL91]. Joint [BS94, CL93, CLM93, JS94, LA94]. Judgment [BW94]. Jurors [SKF98].

Kalman [Fah92, GM94, Koo97, SW91a]. KAP [GJDT98a, GJDT98b]. Kenneth [Gla91, Hab95a]. Kernel [Alt90, Bly93b, BGH93, FHW95, Gou97, HW91, HW92a, JDP94, JS97a, MBR91, Sch95b, SM90a, Sta90, WJ93, jW97, WCH98, YM99]. Kernel- [HW91]. Kernel-Smoothed [MBR91]. Kernel-Type [JDP94]. King [FKOR98, FORK99]. Know [RSV95]. Knowledge [CS95a, Pea91]. Known [LW91]. Kolassa [KT94d]. Kriging [HMN94, Las94b, Las94a, Mac94]. Kurdish [DKW97].

L [Ano99h, Gla91]. Labeling [GRB^{+}95]. Labor [SG96b]. Labour [SR95]. Lack [Chr91, MNS99, SW91b]. Lack-of-Fit [Chr91, MNS99, SW91b]. Lacking [Ano91e]. Lag [CHR94, NP95]. Lags [TA94b]. Laplace [Cra94, Erk94, LR97c, Shu97]. Laplace-Metropolis [LR97c]. Laplacian [O’S90, O’S91]. Large [HW92a, Jen91, Jen95, LRR91, NM98, OR92, PY99, RB90]. Large-Sample [LRR91]. Largest [Wan95]. Last [Sch99b]. Last-Move [Sch99b]. Latency [DXP92]. Latent [BRMZR97, BZ99, CR92, CJ99, For92, For94, Joh90, LR97a, LCG91, RLN99, WPK99a, WPK99b, ZT97]. Latent-Class [Joh90]. Latin [Owe94, Tan93, Ye98b]. Law [Ano91h]. Laws [Roj96]. Layer [Whi89, Whi92a, Fab94, Whi92b]. Layout [PB98]. LDU [CD96]. LDU-Based [CD96]. Lead [MW98a, MW98b, MW98c, RW98a, RW98b, WPD98]. Leading [DR91]. League [BH93b, CH94, GS98b]. Learning [EGMP93, Eld96, Whi89, Whi92a, Fab94, Whi92b]. Least [AI95, Ano96f, CD91, CW94a, CHW92, DS91, DS97c, ES90a, Gar94, GF96b, Hoo93, HM91, RSY93, RSW95]. Leaves [Ros93b]. Left [GL96]. Left- [GL96]. Legislative [GK90]. Lehmann [Mun96, Ros93a]. Lemma [Li92]. Length [ME96, SS90a]. Less [Bar91, BS96a]. Lessons [JS97b]. Letter [Dav96]. Letters [AG94, Ano91o, Ano91p, Ano92k, Ano92l, Ano92m, Ano93i, Ano93j, Ano94i, Ano95k, Ano99o, Ano99p, Sek90]. Level [BMR99, KP91, KM96a, MP99, Niu96, PM97, Tom94]. Levels [CO90, Fin90, GC95, LRR91, SKF98, YKB97]. Leverage [CH90, EM92, Fun93, KM90, Rv90a, Rv90b, RvZ90, RS90b, SC92b]. Liberal [Moo98a, Moo98b]. Life [BM92, BCG99, Zuc98]. Lifetime [CH95a]. Lifetimes [HHH92]. Likelihood [AC97a, AF97, Ano99i, AS90b, Ber91b, Bjø96, Bre90, CM93, CJ96, Chi95, CM98, DF98, DKR95, Eas91, Efr92, FHW95, For94, Fra91, FWW99, Gey91, Hab95b, HL98, HMY97, Hun94a, Jen91, Jen95, Jun96, KBB93, KC93, KT95, KAT91, KS94, LK97, LT97, LMS99, Lav91c, LR96, LK92a, Li95, LQT97, LIZ99, Luc93b, MBL98, Mau95, McC94, McC97, Mel99, Mur95, MRvdV97, Owe95, PO94, RH96, Sas93, Sev91, SSW99, She93, Sta90, Sta94b, SW91a, TL97, VM96, Wan96, WM98, WZ97a, Yas95, CS90a, SS94, Sta94a]. Likelihood-Based [Hun94a]. Likelihoods [Efr96a, Efr96b, Efr96c, Gel96a, Gel96b, Geo96a, Geo96b, Mor96a, Mor96b, Rei96a, Rei96b, ST96a, ST96b, TR95]. Likely [BS96a]. limacon [BCH95]. Limit [WBB93]. Limitations [Akr91a]. Limiting [KSJ94, LS97, MBR98]. Limits [LPT91, SW97a, VHE91]. Lindstrom [LB94c]. Linear [AC95c, BF97, BCJ96, BH98, BC93, Buo96, BS96c, CFGW97, CH97b, CTC96, CB97, CCD98, DG99, DM93, DH99, DM96, Efr99, EMM95, ES90b, Fah92, FG94, FG95, FHW95, FS95a, FK99, For92, Fun95, GS99, GNSC98, GLG92, GH94, HS93a, HMM98, HJ92, HM97, HLC92, HC96, Hof92, Hoo93, Hös94a, Ibr90, IL91, JGC95, JMS97, Jia98, Jon96, KL96, KS96, KC93, KT95, KM99a, LR95b, LWY93, LS95, LW92, LB96, LB94c, Liu96, LJK93, Luc93b, MH90, MH94a, MKSB90, McC97, MT95b, MG97b, MLRR92, Neu98, Now92, Per97, PML98, RMH97, Ric97a, Riv98, RFB97, SC92a, SB97, Sha93b, SK91, SS91a, SS92, Sil92, SRC92, SW91b, SL93b, Tho90a, Vel98a, VL96, WL93, WLGC98, Wes95, WM91, WEST91, YJ98, ZK91, ZL95, ZM93, DH93, LB94d, LB94e]. Linearization [CJ95a, CJ95b]. Lines [APR92, ES90a, Spu99]. Link [CL92a, CDS99, CW94b, CM98, McC98b, McC98c, TW98a, TW98b, TW98c, Var96, Var98a, Var98b]. Linkage [BR95]. Linked [BD91b]. Linking [Are92, NFL92a, NFL92b]. Liouville [RS94a]. Liquidity [HT97d]. List [Ano90i, Ano90j, Ano90k, Ano90l, Ano91q, Ano91r, Ano91s, Ano91t, Ano92n, Ano92o, Ano92p, Ano92q, Ano93k, Ano93l, Ano93m, Ano93n, Ano94j, Ano94k, Ano94l, Ano94m, Ano95l, Ano95m, Ano95n, Ano95o, Ano96n, Ano96o, Ano96p, Ano96q, Ano97q, Ano97r, Ano97s, Ano97t]. Literacy [Wal93]. Little [Bre92]. Liver [SH90]. Local [AC97a, Ano96f, Ano98i, CRW98, CNF95, DBB^{+}94, FG94, FHW95, FHMP96, GR96b, Gus96, Haa95b, LNL95, MS98b, RSW95, Rup97, SG96a, YJ98, MS98a].

Localized [AM92]. Locally [BGH93, FKYZ99, SS96a, vR98]. Locating [CG97]. Location [Ber91a, BHL93, Eas91, O’H90, Par93a, PR90a, WR94]. Log [FK99, GH94, KL99b, Yam90, Yan97]. Log-Bilinear [Yam90]. Log-Linear [FK99, GH94]. Log-Rank [KL99b, Yan97]. Logarithmic [CLT96, KM99a]. Logical [Ano98a, Wes97b, Wes98]. Logistic [AL94, BL93, BDM^{+}93a, BDM^{+}93b, CWW95, CRS^{+}91, CS90b, FLZ96, For92, Gey91, KR96, KM91a, Lit93b, O’N94, ST99a, ST99b, Wac93]. Logistic-Normal [CS90b]. Logit [MH94b, Soo92]. Lognormal [Haa90]. Long [LR97b, Luc93b]. Long-Range [LR97b]. Longest [Lou96a]. Longitudinal [BF97, CCC96, DS97a, DS97b, DLW95, GB97a, GB97b, HZ98, KC96, LCG^{+}92b, LCG92a, LL97a, LL97b, Lou92a, PC97, PHL98, Sch96c, Seg92, SBD97a, SBD97b, SBD97c, SL98b, SC97c, TS92, TCS94, TDW95, WCH98, ZLRS98]. Longleaf [RC94a, RC94b]. Look [APR92, Shu97]. Lorelogram [HZ98]. Lorenz [KM96b]. Loss [BL93, Rue95]. Lost [DB97]. Lottery [Ang91, JK93]. Low [LT94b, RG94]. Low-Dose [RG94]. Lower [BF90, CS96b, Sat95]. Lupus [AME96]. Lyapunov [MEGN92].

M [Ano99h, DE96]. Machine [Eld96]. Main [Ano96g, MM95]. Major [BH93b]. Mallow [RS94b]. Malpractice [Coo91a, GHCF94]. Man [WT90a]. Management [Giv99]. Managerial [GS93a]. Mann [Ano91f, BW94, Mee90]. Manteiga [Stu99, Wil99]. Mantel [YF90, YF95]. Mantel-Haenzel [YF90]. Manton [Hab95a]. Many [BFH99, MBR91]. Mapping [WCXG97]. Maps [MST97]. Margin [Qui98a]. Marginal [Che94a, Chi95, GR96a, GS90a, HZ96, Joh92, LK97, LA94, LMS99, ML94, Neu99, PC93, RGH99b, RGH99a, RF99, SL98a, Was99a, ZL99]. Marginals [Gus96, LWW91, LK95]. Margins [LW91]. Marijuana [Mod93]. Marital [XP92]. Markers [LTW96, SL93b]. Market [CW91a, PS93]. Marketing [ALJ99]. Markov [Ros95b, BBGL97, BM95, CLWZ95, CC96b, DwSH94, FK94, GT95, GRS98, GJR98, Hig98a, Joh96b, Joh98b, Lou96a, MTY95, Ros95c, RP96, Sam90b, Kas97b]. Markov-Dependent [CLWZ95]. Markovian [LNL99]. Marriage [BB90]. Marrow [DwSH94]. MARS [LS91a]. Marshall [Ryu93]. Martingale [Bro97]. Masking [BG99]. Mass [DMLC97, MS91a, MS95b, MS95a]. Match [BR95]. Matched [Dab90, FS95b, HHS97]. Mathematical [BR92]. Mating [Shu97]. Matrices [CD91, CD95, DK99, GL92, JB92, MR91, Rig97, XLR91, ZB92]. Matrix [CLT96, CD96, GS91a, HW90a, LKC97, Sch91c, ZT97]. Matrix-Logarithmic [CLT96]. Matter [Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano91a, Ano91b, Ano91c, Ano91d, Ano91k, Ano91l, Ano91m, Ano91n, Ano92a, Ano92b, Ano92c, Ano92d, Ano92g, Ano92h, Ano92i, Ano92j, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95a, Ano95b, Ano95c, Ano95d, Ano95g, Ano95h, Ano95i, Ano95j, Ano96b, Ano96c, Ano96d, Ano96e, Ano96j, Ano96k, Ano96l, Ano96m, Ano97b, Ano97c, Ano97d, Ano97e, Ano97l, Ano97m, Ano97n, Ano97o, Ano98b, Ano98c, Ano98d, Ano98e, Ano98l, Ano98m, Ano98n, Ano98o, Ano99b, Ano99c, Ano99d, Ano99e, Ano99k, Ano99l, Ano99m, Ano99n, CH94]. Maturation [AHB97]. Max [Hab95a, KSW92]. Maximal [Ter90]. Maximin [MNS99]. Maximized [BB94]. Maximizing [VK92]. Maximum [AF97, AS90b, CJ96, Eas91, Efr92, For94, GJ99a, Gey91, GJP96, Hab95b, KBB93, KT95, KAT91, KS94, LT97, LMS99, LK92a, LIZ99, Mau95, McC94, McC97, MRvdV97, RH96, Sas93, Sev91, She93, SW91a, WZ97a]. Maxwell [Bro92a, Bro92b]. May [GS95]. Mean [Ano99i, AS90b, Che95a, Che94b, CHW92, HJ92, HL96, LR95a, MT93, NS93, Neu99, PSS99, PR90b, RGH99b, RGH99a, RF99, RV98, SMF94, Sev91, SW91b, SY95, Sut93, The93a, WM98, Was99a, YB97, ZL99, Zuc98]. Means [ALM97, BHS98, BM96, Cro97, Esc94, FNP95, Fon92, GEG99, HH90, HHS97, Kun90, Lau92b, MM93, PC97, RV98, SP91, SC97c]. Measure [EL90, Hua97, MKSB90]. Measured [CJ99, HTP95, TDW95]. Measurement [Ano91g, Buo90, Buo96, BS96c, CL92a, CKLS96, CS94, Efr97, FS95b, GMP97, Gle92, HDG97, JTC99, Joh90, MW98a, MW98b, MW98c, RW98a, RW98b, Sho93, SK91, SB95, SC95, WLGC98]. Measurement-Error [MW98a, MW98b, MW98c, RW98a, RW98b]. Measurements [Agr97, Con92, GW99, HZ96, JD91, Kus98, LD91, LD92, LY93, LT96b, OR96]. Measures [Akr91a, AA94l, AHB97, And91, ALHH92, Ano91e, BCdF91, BL92, Ben91b, CPP96, Car94b, CR91, Cox91, DBB^{+}94, Goo91b, Goo91a, Hab91, JR96, KW96b, LB94c, Lit95, Lju93a, Lju93b, MLRR92, OMF^{+}93, Roc92, Sev96, TA90, Tho91b, LB94d, LB94e]. measures-data [LB94d, LB94e]. Measuring [For94, HHH92, Ye98a]. Mechanism [Lit95]. Mechanisms [SKF98]. Median [BD93, CHW92, ES90a, GNK96, Jun96, RC93, SW98a, WH90, Yan99, YJW95]. Medians [BHNO92]. Medicaid [GLB98]. Medical [Coo91a, GHCF94, NGG97, Ueb93]. Medicare [GM93a, KR96]. Medicine [Gei92, GBA^{+}99]. Medium [LT94b]. Melanoma [BHB96, BSB98]. Membership [RLN99]. memoriam [Man94]. Memory [CKLH92]. Men [MPCB95]. Merging [Giv99]. Meta [DPWH99, LB94b, Sil97a, Gla91]. Meta- [Sil97a]. Meta-Analysis [DPWH99, LB94b, Gla91]. Meteorological [Cre94, Gut94b, Han94, KDG94, HW94]. Method [Akr90, BR95, BCG99, BO97, CCD97, CV95, Cra94, Eas94, Efr92, ET97, GKK91, Goo96, Gou97, GW91, HW91, HRW90, Hol95, KP91, LB93, Lou96a, OR98, Pap96, PF91, Pes93, Rag93, RH96, RB90, SC97a, SC97b, SS97, SMP95a, SMP95b, Tro95a, WS94, XP92, YL97, Zuc98, YF95]. Method-of-Moments [BCG99]. Methodological [WHGA96]. Methodology [SHBF94]. Methods [AB97b, AGM94, Atk94, BOP92, BBGL97, BBH94, Bre92, BPS95, CG92, CMM93, DM93, Efr96a, Efr96b, Efr96c, EP99, Fer98b, FHJW96, Gel96a, Gel96b, GC91, Geo96a, Geo96b, Goo96, Gra92, Haa90, HT97a, HT97b, HT97c, HWY99, HA94, Hig98a, Jon96, KO96, LKM98, LK91, LT96b, LC98a, LC98b, MSH90, Mor96a, Mor96b, NP95, NGG97, O’S91, Por97a, Por97b, Rei96a, Rei96b, Seg92, ST96a, ST96b, She97, SW90, SG96b, SHGL96, Sut93, Weg91, Yas95, Zie90, HA95]. Metropolis [LR97c]. Michie [Eld96]. Microdata [BKP90]. Microscopy [QTC96]. Migraine [DPWH99]. Min [KSW92]. Minimal [AP99, PHS94]. Minimax [Amr95, Sob91, Wie98]. Minimum [CCB96, Har92a, Rub90, Ryu94, Sie93, Dha93]. Mining [Ye98a]. Minorization [Ros95b, Ros95c]. Miron [Gla91]. Misclassification [AHB97]. Misclassified [GWG90]. Mismeasured [PF91]. Misregistered [CHR94]. Missing [AS90b, CL99a, Che94b, CLWZ95, Efr94b, Efr94a, Jon96, KLW94, LIZ99, Lit92, Pai97, RRZ95, RR95, Rub94, SLS97, SKF98, WSJ92, WWZtO97]. Missing-Data [SKF98]. Missingness [BCG99, CCC96]. Misspecification [CHW92]. Misspecified [GLG92]. Misspecifying [BL93]. Mixed [BH98, BC93, BD92, BD91b, GS91a, HDG97, HC96, Jia98, Lau92b, LB96, LB94c, MS92, McC97, MPCB95, MG97b, Ric97a, VL96, WLGC98, Wee91, ZDL96, ZLRS98, LB94d, LB94e]. Mixed-Effects [HDG97, MPCB95, VL96, ZDL96, LB94d, LB94e]. Mixing [CDP96, Gou97, HRW90, LK92b, MZ96]. Mixture [Ano91j, Cra94, CCB96, GLR93, HRW90, LMR96, LR92, Lit93c, MR97a, MR97b, O’S93, Roe94, RCL96, Wes97a, WC92]. Mixtures [BR94a, EW95, Joh92, LB93, MZ96, PJT96, Pri94, RW97]. Mixtures-of-Experts [PJT96]. MLE [LK92b]. Mobile [Sam90b]. Mode [Erk94, Fah92, Hsi97]. Model [Agr97, AP99, AMCS98, AC97b, ALJ99, AL95, Ano99a, BAP95, BP96, BMR99, Bol95, BO95, BZ99, BD92, Buc95a, Buc95b, Buo96, BD93, Bur94, CM97a, CKLH92, CPP96, CW91a, CL93, CIS99, CDS99, CGM98a, CGM98b, CGM98c, CLT96, CLM93, CDP96, CLWZ95, CHW92, Cro95, DwSH94, DR98, DSM98a, DSM98b, DM96, Eas94, ES97, ES90b, FG99, FBF97, Gel94a, Gel94b, GHCF94, GLG92, GS98b, GS93a, GR92a, HJ92, HLC92, HN92, Hös94a, HR97, HM91, Joh94, Kit98, KKT98a, KKT98b, LRM96, LK92a, Lin91, LCG91, MR94, Mer94a, Mer94b, Mer94c, MR97a, MBR98, MPCB95, MR97b, MRvdV97, Pap96, PS94a, Peñ98, Per97, PJ94, PH90a, PHS94, PR93, PHL98, QH98a, QH98b, RGZ95a, RGZ95b, RGZ95c, RMH97, Ric97a, RFB97, RT96b, RP96, Ryu94, SW95a, Sam90b, Sas93]. Model [Sas97, SDW98, Sch95c, Sch95d, SB97, SLS97, Sha93b, Sha96, SK91, Sil92, ST99a, ST99b, SH90, SA92, Ste94, SW98a, SGP98, SW91b, TCS94, TL95a, TL95b, VL96, WL93, WR95, WJ92, Wol95a, Wol95b, WM91, WEST91, WCH98, Yam90, Yam92, YP99, YR95, Yas95, Ye98a, Zha92, Zha98a, Zha98c, ZWWB96, Dha93, Hjo94]. Model-based [DR98]. Modeler [Lit93a]. Modeling [AME96, AHB97, Alh92, AL94, Ano92f, Ano98h, BB90, BMR99, BD97, CPS92, CH95a, CNF95, CM97c, CCC96, Cre94, DLD91, FOS95, FS98b, Fyg97, GMD95, GBJ96, GMR99, GS90b, GRB^{+}95, Gri90, GWB99, Gut94b, Han94, HMM98, HYP97, Hei91, HOLDG98, KDG94, LA94, LMS99, LMR96, LCMA92, LC92, LS91a, LR97b, Lit95, MH94b, McN92, Mel99, ML94, NGC^{+}95, NT95, OO97, PS93, PC97, PS94b, Pol91, RLN99, SA92, SC97c, Tsa98, TDW95, Ueb93, WB96, HW94]. Models [AC97a, AR90, Ahn97, AC97b, AS90a, And91, AI95, Ano96a, AC95c, AZ99, BCdF91, BF97, BH93b, BR94b, BCJ96, BL93, BDM^{+}93a, BDM^{+}93b, Ben91b, BBGL97, Boc99, BH98, BDR92, Bre90, BC93, BR98b, BR98c, BR98a, BD91b, BH97, BS96c, CD91, CMW97, CS90a, CKLS96, CFGW97, CH95a, CP99, CR92, CL95, CH97b, CT93a, CT93b, CWY97, Cho92, CB97, CM91a, CR91, Coo91a, CT90, CS94, CW97, CCD98, Cox91, Cox95, CS90b, Cro97, CNS90, CCB96, DG99, DK99, DM93, DR96, Dea92, DB97, Dig99, DH99, Fah92, FT94, FW96, FG95, FHW95, FZ98a, FZ98b, FZ99, FOS97, FL95, FLZ96, Fre99a, GHRPS90, GS99, GNSC98, GL92, GH94, GLR93, Goo91a, Gri94, Gu90, Gu92, Hab91, Hab95b, HS93a, HKS98, HMM98, HW92a, HZ96, HBH93, HK92, HR94, HDG97]. Models [HC96, Hoo93, HH96, HW95b, HS94, Hun94a, Ibr90, IL91, JGC95, Jia98, JTC99, Joh92, JV93, Jun96, KSJ94, KP97, KL96, KS96, KPR91, Kne98a, Kne98b, Kol99, Koo97, LP99, LK97, LR95b, Lan96, LCG^{+}92b, LCG92a, Lan98a, Lan98b, Lau92b, LMR96, LS95, LR97a, LB94b, LT98, LB96, LCG91, LR92, LB94c, LL97c, LW91, Lit93b, Lit93c, LR99, LNL95, Lju93a, Lju93b, LJK93, Lou92a, MR94, MO98, MH90, MH94a, MS92, Mau95, McC97, MS96, MG97b, MLRR92, NM98, Neu98, NP95, Niu96, Now92, OR98, OKS97b, OR92, PB94, PWK99, PJT96, Pet92, PWM92, PML98, Qui98b, RMH97, RS98a, RS98b, RG93, Riv98, RRZ95, RR95, RK96, SL93a, SRR99a, SRR99b, Sch96c, SGC95, SS94, SSW99, She93, SS91a, SS92, Sil97a, SL98a, Sta90, SS91b, Sta94b]. Models [Sta91, SKF98, SB95, Ste90c, SW91a, SHGL96, TR97, TS92, Ter94, Tho90a, Tsa98, Vil90, Wac93, Wak96, WB96, WLGC98, WW98a, Wan98b, WW98b, Wee91, WJ92, WCR94, Wes95, Wes97a, Whi89, Whi92a, WHG98, Wie98, XL99, Yam92, YJW95, ZDL96, ZK91, ZLRS98, ZL95, ZM93, dCP99b, vdL99, Fab94, Goo91b, LB94d, LB94e, Sta94a, Whi92b]. Modern [O’C99]. Modification [Zuc98]. Modified [Qui98b, Shu97]. Modular [TFP90]. Molecular [SRSD94]. Moment [CTC96, LRR91]. Moment-Based [LRR91]. Moments [AK92, BCG99, HRW90, HN92, HL97, LB93, SC94]. Monetary [PH90a]. Monitoring [Ano97k, GS90b, HOS95, Yas95]. Monotone [BS97b, GMPS99, HS98c, HS98d, Lee96a, MS94, VS93]. Monotonic [AHB97, SS97]. Monotonicity [Fin93]. Monte [Kas97b, Ros95b, BBGL97, BS96b, CPS92, CL95, CC96b, GT95, GRS98, GJR98, Hig98a, Joh96b, Joh98b, LC98a, LC98b, MP95, NG94, Ode91, OB93, Ros95c, SMP95a, SMP95b, WT90a]. Morphology [BR92, FR99]. Mortality [AS90a, Alh92, DMLC97, GWG90, HLC92, LCMA92, LC92, McN92, OMF^{+}93, SK91]. Mosaic [Fri94]. Motion [PC96, Riv98, Ste94]. Motorcycle [Bro92b, Wei92]. Move [Sch99b]. Moving [BDR92, Haa90, LL97c, RP96, SL93a, TR97, YR95]. Moving-Average [LL97c]. MR [Fab94, MS95a, SG95a, Wee94, Wes98, Whi92b]. Multi [Ano91h, Fri94]. Multi-Stage [Ano91h]. Multi-Way [Fri94]. Multiargument [MS94]. Multichannel [BO97, O’S93]. Multiclinic [BPS95]. Multicriterion [Giv99]. Multidimension [CS93].

Multidimensional [BPL97, DL92, Lav91a, O’C99]. Multilevel [MH94b]. Multimodal [OB93]. Multimodality [MS91a, MS95b, MS95a]. Multinomial [Ano95e, Boc99, CMW97, Eub97, GJP96, He90, HLC92, NM98, OR92, RMS90, SG95b, SG95a]. Multiparameter [AC97a]. Multiparty [Gel94a, Gel94b, Mer94a, Mer94b, Mer94c]. Multiple [Ano98a, Atk94, BRMZR97, Bar93b, BPR95, Bin98b, Bin98a, Boc99, BS91a, Bre98b, Bre98a, BS94, CT98, CRS^{+}91, CC96a, DFGJ93, DG93a, DT92, Fie93a, Fin90, Fin93, GKL98c, GKL98d, GKL98a, GKL98b, GLR93, GB98, GCW96, HW90a, HS93a, HR90, HL94, Hol99a, Jon96, Jud98b, Jud98a, Kad98b, Kad98a, Kun90, LTM92, LLR95, LR97a, LB96, LNL95, MZ97, ME96, MP96, PB98, Pen97, Pep91, Rub96a, SC97a, San98b, San98a, SC97b, SMF94, SLS97, Sim93a, Sin93, Ski91, Sto99, SL93b, Tom96, Tro95a, Wel93, Wes97b, Wes98, Yam90, Zha98b, DG93b]. Multiple- [HL94]. Multiple-Recapture [DFGJ93]. Multiple-System [CT98]. Multiplicity [HB99]. Multiply [LRR91]. Multiprocess [Bol95]. Multirater [Joh96a]. MultiSample [JSKK98]. Multiscale [Kol99]. Multistage [SCC98]. Multistate [Sch99b]. Multivariate [Agr97, AR90, AA94l, Ano91g, Ano96i, BL92, BG99, BDR92, BF93a, Buo90, BS98, CD91, Cha96b, CM97b, CH90, DM93, EM90, Fah92, FOS97, FLZ96, Fol96, GMR99, GWG98, GR96b, Gle92, GR92a, HMO97a, HMO97b, Hou95, Joe95, JSKK98, KST98, KM90, KL96, KP90, KM96b, LA94, Lan96, LMS99, LK92a, LS91a, LT98, LW92, LB93, Lit93c, LS93, Liu95, Liu96, MZ97, ME96, ML94, MLRR92, MP99, Now92, OV95, PR90a, PML98, Rig97, RR95, RW96, Rv90a, Rv90b, RvZ90, RP96, RS90b, SBS94, Sie93, Sob91, SL98a, Tan94, The93a, Tho91b, Tsa98, Vel95, WHG98, WR94, MY95]. Multivariate-t [ME96]. Multiyear [CD92]. Myocardial [OMF^{+}93].

Nadir [DMLC97]. Names [Are92, NFL92a, NFL92b]. National [CM91b, CH94, GS98b, HZS^{+}90, MSML97, Sta91, ZT97]. Natural [Ano96h, GPS95, KO95, Sla94]. Near [Cha90, Chr91, MNS99]. Near-Integrated [Cha90]. Nearest [BR98d]. Needed [LS91b, SM90b]. Negative [Boc99]. Neighbor [BR98d]. Neighborhoods [Pre96b]. Nephritis [AME96]. Nerve [Bri92]. ness [Hjo94]. Nested [BR98b, BR98c, BR98a, FZ98a, FZ98b, KW95, Kne98a, Kne98b, Lan98a, Lan98b, Mar92b, NG94, RS98a, RS98b, Sas97, WW98a, WW98b, WJ92, Yas95]. Net [PJ94]. Network [Ano96a, BH97, Cox95, McC98b, McC98c, NT92, TW98a, TW98b, TW98c, Var96, Var98a, Var98b, Whi89, Whi92a, Fab94, Whi92b]. Networks [HD97, SI90]. Neural [Eld96, HD97, PJ94]. Neuroanatomy [DLB91]. Neurological [Wes97a]. Never [GKB98]. Newton [LB94d, LB94c, LB94e]. Neyman [Fan96, Led94, Leh93]. Nino [Sol95]. No [BZ99, Raz90a, Raz90b, YKB97, Fab94, LB94d, MS98a, MS95a, SG95a, Sta94a, Wee94, Wes98, Whi92b]. No-Observed-Adverse-Effect [YKB97]. Noisy [FHTW97, GC95, Joh94, Pre96b]. Nominal [CJ99]. Nomination [KS93]. Non [CMM93, DF98, Gu90, HH91, RRS98, SM92]. Non-Bayes [Goo92]. Non-Bayesian [HH91]. Non-Gaussian [CMM93, Gu90]. Non-ignorable [RRS98]. Non-iid [SM92]. Non-parametric [DF98]. Non-response [RRS98]. Nonadditivity [GH90]. Nonavailability [PMW93]. Noncentral [WK94]. Nonconjugate [DK99]. Nondetects [LPT91]. None [Bak98b, Bak98a]. Nonequivariant [BS96a]. Nonexperimental [DW99]. Nonhomogeneity [Pri96]. Nonhomogeneous [KY96]. Nonignorable [BCG99, Con92, CCC96, Dig99, FZ99, Fre99a, GLR93, LP99, LR99, PB94, SRR99a, SRR99b, Wac93, vdL99]. Noninformative [CW93b, DG95, MV91]. Nonlinear [CPS92, CKLS96, CM93, CT93b, CJ95a, CJ95b, CW90, CT90, FWW99, GR96b, GP92, Hei91, HDG97, LS95, LS91a, MPCB95, Niu96, OKS97b, SA92, SC92b, SR92, Str93a, TA94a, TA94b, Vid98]. Nonlinearity [LR97b]. Nonlinearly [EL97]. Nonmonotone [SV93]. Nonnegative [Ano99i, Rol94, WM98]. Nonnegligible [SS99]. Nonnested [BM99]. Nonnormal [CPS92, FWW99, WBB93]. Nonnormalities [LM93]. Nonnull [CM91a]. Nonparametric [AA94l, AAB97, AB97b, AF97, AC95a, AD90, BRR92, BP99, BZ92, BW92, BRW99, BPS95, BDM97, CDW93, CCD97, Che94b, Cho92, CMM93, DN98a, DN98b, EP99, EL97, ES90b, ES93, Fan92, FG95, FHMP96, GR97a, GM99, GR97b, GMCM96, Gou97, GL96, Gu92, HH90, HW91, HT97a, HT97b, HT97c, HMC^{+}93, Ize91, Jay96, JMV94, KS94, LT97, LK92b, Li95, LCWH97, MS99a, MZ96, MEGN92, MN93, MS94, Mul99, Nyc91, NGH^{+}95, Owe95, PO94, PF91, Por97a, Por97b, Qui98a, Raz90a, Raz90b, Rup97, Sam93a, SG92, SC94, SKW99a, SKW99b, SK97, SL98b, TFP90, TA94a, TA94b, Wan91, WS94, WR95, WC99, WZ97a, WK98, ZDL96, Zha96, Zha99]. Nonparametrics [SCF93]. Nonrandom [Ros95a]. Nonrandomly [DS97a, DS97b, GB97a, GB97b, LL97a, LL97b, SBD97a, SBD97b, SBD97c]. Nonresponse [Con92, Dig99, FZ99, Fre99a, GLR93, Kot94, LP99, LR99, PB94, SRR99a, SRR99b, Sta91, vdL99]. Nonseparable [CH99]. Nonstationarity [Cre90]. Nonstationary [Ada98, AR90, GM94, Koo97, Qui98b, Rob94, SG92, WPK99a, WPK99b, YR95]. Normal [AL94, Ano99i, BH96, CT90, CS90b, Cro97, Esc94, GHRPS90, Gle93, Hou95, Joe95, KM91a, KR76, KR90a, Kun90, Lie94, LB93, Luc93a, OV95, OR92, Pes93, Rig97, Rol94, Sie93, SY95, WM98, dCP99a, Hjo94]. Normality [BG91a, BF93a, BH96]. Normals [Roe94, RW97]. Norms [BM96]. Northeast [Sas97]. Northern [Sol91]. Note [Dor94, Kot94]. Notion [Cha96b, CS96a]. Notions [LS97]. Nuisance [Ano91i, BG91a, BB94, CW93b, LK91, Sil96, Sil97b]. Null [Kra99]. Number [BBC98, BF93b, CL92b, GC95, HS98a, HW90a, HL97, PMW93, Roe94, Vel98a]. Numbers [JK93, KO95, LCG^{+}92b, LCG92a, Lou92a, TS92]. Nursing [GM93a, KR96]. Nutrient [KSJ94, MN93].

O [DH91, Lav91b]. O-Ring [DH91, Lav91b]. Objective [CC96a]. Objects [LJK93]. Observational [RK90, Ros93a, Ros95a, SH92]. Observations [CLWZ95, Fre94, KS97, Kus98, LW92, MBR91, OV95, Pet92, SB97, WSJ92]. Observed [FSE98, GT98, RRZ94, Roc95, YKB97]. Obtain [MR91, XLR91]. Obtained [RV90c]. Occam [MR94]. Occasion [NS93]. Occupation [CRS^{+}91]. Occurred [GKB98]. Occurrence [BRR92, HW95b, Sol91]. Occurrence/Exposure [BRR92]. Occurs [Erk94, Hsi97]. Odds [Agr98a, Agr98b, Ano99h, BB98b, BB98c, DD90, FS95b, HR97, MPS98a, MPS98b, MRvdV97, RT96b, Shi91, SW98b, SW98c, SW98d, VHE91, WKL95, YP99]. Offered [CH91a]. Official [AS90a]. Olkin [Ryu93]. Olympic [DAO94]. Omitted [MW98a, MW98b, MW98c, Neu98, RW98a, RW98b, WPD98]. Omitted-Variables [MW98a, MW98b, MW98c, RW98a, RW98b]. Omitting [BL93]. Oncology [GFT^{+}93]. One [Ano96i, CS92a, CF95, Fol96, HHS97, HY98, Hay90, HMO97a, HMO97b, Leh93, MH99, PR90a, Sil92, SS95a, SRC92, Tab90, Tan94, WH90]. One-Sample [HMO97a, HMO97b, PR90a, WH90]. One-Sided [Ano96i, CS92a, CF95, Fol96, HHS97, HY98, Hay90, MH99, Sil92, SS95a, Tan94]. One-Step [SRC92, Tab90]. Only [ASL96, HA94, Riz92, HA95]. Onset [BBH96]. Operating [Llo98]. Operations [Hog93]. Opinion [GMD95]. Optical [HYP97]. Optimal [Bly93b, BCH95, CM93, CW94c, Ede90, EK93, GLG92, GV96, Gri94, HTB94, HBH93, HTP95, JMS97, KM91a, KP91, KPR91, Kus98, MP95, Par93b, SC96a, Tam95, Tom94, Wan97]. Optimality [KW97]. Optimization [MLMT94]. Optimum [AC95c, Fin90, GS97, HHM92, HYP97, MS92]. Order [AP99, DM95, Ede90, GH99, GT91, JS97a, MM93, Min98, MT95b, NSS97, Par96, QN98, RG93, RH96, Sch96c, Shi91, Wan95, Zel93]. Order-Constrained [MM93]. Order-Restricted [MT95b, NSS97]. Ordered [AMP90, AC95b, BM96, Cro95, DS97a, DS97b, dE97, EGGS97, FT94, GB97a, GB97b, GM96, Hay90, JSKK98, LL97a, LL97b, LGWC93, RK96, SBD97a, SBD97b, SBD97c]. Ordering [AG96, DKR95, Muk96, Pan97, RS93, Wan96]. Orderings [DF98, PW96]. Ordinal [AHB97, BP99, BZ99, CCC96, Goo94b, HZ96, Joh96a, KSW92, KC96, ML94, WKL95]. Organizing [Kit98]. Orthogonal [AM92, CB90, Mar92b, Rol94, Tan93, WW91a, Ye98b]. Orthogonalization [SG97]. Orthogonalized [CDP96]. Oscillation [KO96]. Other [BD91a, Bre92, Bre90, CS92a, SAK90, WM91]. Our [Wal93]. Out-Patients [TFAH94]. Outcome [FL95, GM96, Neu99, RGH99b, RGH99a, RF99, VK92, Was99a, ZL99]. Outcomes [BRMZR97, CR92, DDV99, GMR99, LLR95, LR97a, RRZ95, RRS98, SRR95]. Outlier [BG99, Bra90, BS98, CL93, CLM93, GR92b, PY99, Pet92]. Outliers [Atk94, Bar93b, BWC98a, BWC98b, CH90, DG93a, Fie93a, Fun93, HS93a, KM90, LMR96, LRM96, MZ94, O’H90, RW96, Rv90a, Rv90b, RvZ90, RS90b, Sim93a, Wel93, DG93b]. Output [Chi95, DR91]. Overdispersed [Bre90, DB97, KT94a]. Overdispersion [Dea92, Efr92, GS92, LR95b]. Ozone [CCG^{+}97a, CCG^{+}97c, CCG^{+}97b, Cre97a, Cre97b, GMS97a, GMS97b, NT95, Niu96, SF97a, SF97b].

p [FWW99, SMP95a]. Pain [PHL98]. Paired [AD90, FT94, LLE93]. Paired-Comparison [AD90]. Pairs [Dab90]. Panel [GH94, HT97d, LI98a]. Panels [PSS99]. Paper [Ano98q]. Paradigms [Fay96a, Ima95]. Paradox [He90, KMT99, Mit91, Sam93b]. Paradoxes [Hau92]. Parallel [Weg90b]. Parameter [BB94, BMR99, CD95, Che90, EL97, Erk94, GSL92, GMCM96, KW97, Lav91a, Lin91, LW99, NT92, O’H90, Par93a, Sev91, SW98a, Tho91a, Ye93]. Parameterizations [Ano96h, GPS95, Sla94, WJ93]. Parameters [Ano91i, BG91a, BHL93, CL93, CLM93, CW93b, FS95b, GT91, HHM92, KT95, KAT91, LK91, Pag90, Pol91, RTF92, Sil96, Sil97b, Sob93, Str93a, WJ92, Yas95]. Parametric [ACH99, CW94b, CS94, DN98a, DN98b, HMM98, HR94, Li96, Mei94, NM98, YK96, DF98]. Pareto [BVT96, CH97a]. Part [Akr91a]. Partial [AM92, CCD98, DS97c, Gar94, GMD95, GF96b, Gri92, Sas93]. Partially [AR90, BM96, CFGW97, CJ99, GT98, QN98, Qui98b, Rol94, YR95]. Participation [BD97]. Particle [PS99]. Particular [GK93]. Partition [Cro97, GS97]. Partitioning [Lan96]. Partitions [POH97]. Partly [PC97, SC97c]. Partner [SJ92]. Party [Gel94a, Gel94b, Mer94a, Mer94b, Mer94c]. Pascal [BCH95]. Past [Pol91]. Path [ME96]. Paths [ASL96, Joh96b]. Patient [SRR95]. Patients [KR96, LI98a, TFAH94, TDW95]. Patrick [O’C99]. Pattern [GS98a, Lit93c]. Pattern-Mixture [Lit93c]. Patterns [Bac90, BB90, CS91, DLB91, KO96, XP92, YK96]. Payment [PH90b]. PCP [RG94]. Peaks [Pan97]. Pearson [EL90, Goo96, Leh93]. Pediatric [DXP92, DA96c]. Penalized [PO94]. Penalty [GH95, SKF98]. People [EGG95]. Performance [GHMW94, HMN94, KAT91, Las94a, Mac94, PH90b, Las94b]. Period [Bac90, DA96c, SS90a]. Periodic [GHL96, LZ98a, LZ98b, LR97b]. Periodograms [Ste95]. Permanent [Yam92]. Permutation [BPS96, ES99, FS98a, Ode91, RK90, Sch94d, SP91, Wel90a]. Permutations [Ste90c]. Permuted [LR96]. Person [GNK96]. Personal [Are92, NFL92a, NFL92b]. Perspective [BCJ96, Hay94, Lit93a, Lju93a, Lju93b, Wei95b]. PES [EKWS91, MS91b, MS91c, WS91]. PET [OO97]. Pharmacokinetic [GBJ96, Wak96, WB96]. Pharmacokinetics [HDG97]. Phase [Sit97, TS96]. Phenomena [LR97b, Sam93b]. Phi [EL90]. Phi-Squared [EL90]. Photon [Wei97]. Physiological [GBJ96]. Piecewise [MPCB95]. Pilot [CS96b]. Pine [RC94a, RC94b]. Pitman [Bly93a, CW93a, GKS93, MKSB90, Ped93, Rao93, RHS93b, RHS93a]. Pivotal [FHJW96]. Placebo [GM96]. Placebo-Controlled [GM96]. Placement [AGM98a, AGM98b]. Plackett [ML94]. Planning [AGM98a, AGM98b]. Plebiscite [EV92, RSV95]. Plot [Hol95, MS92]. Plots [BVT96, BH96, Coo94, CCD98, EM90, Gal94]. Plurality [Est92]. Point [AF97, BH93a, BG99, BR98d, CS95a, CHP96, DR98, DLB91, Hös94a, MH94a, MY91, RC94a, RC94b, Ros93a, SW95a, SR94, SK93b]. Points [CH90, dE97, Fun93, KM90, Rv90a, Rv90b, RvZ90, RS90b]. Poisson [Bol95, Bre90, CS95a, CM97c, Coo91a, CHP96, Dea92, DB97, Efr92, Kol99, KY96, NC99, VM96]. Policy [Ano91y, Kra90, Moy99, Nor90, Pea91, Roc91, Zel92]. Poly [BS93]. Poly-Weibull [BS93]. Polychotomous [AC93, KBS97]. Polymerase [NA99]. Polynomial [AC97a, AM92, FHW95, HW92a, Jay96, Rup97]. Polynomials [SG96a]. Polytomous [For92, Hir92]. Pooled [PSS99]. Pooling [GS95, LTP94]. Poor [WT90a]. Population [ALM97, Bac90, BF92, Buc95a, Buc95b, CCG^{+}97a, CCG^{+}97c, CCG^{+}97b, CH93a, Cre97a, Cre97b, DKW97, DFGJ93, EKWS91, FG91, GBJ96, GMS97a, GMS97b, HS98a, LT94b, MV91, MR97a, MR97b, MS91b, MS91c, NS93, NSS97, Pas93, RGZ95a, RGZ95b, RGZ95c, RZ98, RAGW93a, RAGW93b, Sch95c, Sch95d, SS90a, SS95b, SF97a, SF97b, TL95a, TL95b, VL96, WS91, Wak96, WB96, WS96, Wol95a, Wol95b, WC91, Zas93, PWM92]. Populations [BBH94, CDW93, HS94, KBB93, LW91, Pol91, Rig97, Wan96]. Positive [DM95, LI98a]. Positively [SC97a, SC97b]. Positron [MO98, WEST91]. Possible [BGH97]. Possibly [GS91b, SLS97]. Post [Bel93a, HS93b, HS93c, Hog93, Lit93a, Sch93a, Ylv93]. Post-Enumeration [Bel93a, HS93b, HS93c, Hog93, Sch93a, Ylv93]. Post-Strata [HS93b, HS93c, Sch93a, Ylv93]. Post-Stratification [Lit93a]. Posterior [Che94a, Erk94, Fah92, Joh92, LR97c, PSS93, Sev91, WT90b]. Postrandomization [Roc95]. Poststratification [Val93]. Potency [KCHH93]. Potential [GJ94]. Power [Ano91e, Ano98i, BHS98, Bly93b, EMM95, KW97, Kun90, LLR95, MS98b, MNS99, MLRR92, Ode91, RV90c, MS98a]. Powerful [Ano96i, Tan94]. Practical [Ano97i, RW97, SCW96, SCW97, YB97]. Practice [Kas97b]. Preaveraged [AM92]. Predict [Coo91a, Man90]. Predictands [BS94]. Predicting [BH93b, CKLH92, CWY97, GHCF94, HZS^{+}90, The93a]. Prediction [Ber90a, BP96, Ber91b, BBC98, BH98, Bre92, BK99a, BK99b, Chr99a, Chr99b, CDP96, CMMY91, DKS97, DH99, DLW95, GM94, Haa95b, Han99a, Han99b, HJ92, HA94, HD97, KL96, LKCH99a, LKCH99b, LKCH99c, MT93, MP96, OKS97a, OKS97b, PJ94, SC99a, SC99b, Ste99a, Ste99b, TS90, WL93, WW91b, HA95]. Predictions [CS95a, DDV99]. Predictive [CO90, CS91, HMN94, IL94, KO96, Las94a, Mac94, Las94b]. Predictor [Ver93]. Predictors [CS90a, SRR95]. Pregnancy [SPM99]. Preliminary [GLG92]. Premia [CP93]. Presence [AC95b, Ano91i, Car94b, Fon92, JMS97, LTP94, RRZ95, Sil96, Sil97b, Wan92]. Present [Pol91]. Preserved [Hei95]. Preserving [CGGM98a, CGGM98b, CGGM98c, Joh98a, Joh98c, MT98a, MT98b, SHL98a, SHL98b]. Presidential [Moo98a, Ket97a]. Pressure [HL94]. Prevalent [HW95b]. Preventative [BHB96]. Price [BG91b, CG92, DN98a, DN98b]. Prices [CG97]. Primary [BHB96]. Principal [BR92, Coo98a, Coo98b, LT94a, Li92, Li98c, Sch91c, Tar99]. Principle [Bjø96, Coo98c, Li98b, Ter90]. Prior [Ano98j, AZ99, BRW99, Cla96, DJ98, Esc94, GBJ96, GM99, Gus96, IL91, KW96a, LWW91, Lav91c, MS99a, Mul99, RH96, SKW99a, SKW99b, SHBF94, SY95]. priority [MS95a]. Priors [BCJ96, CW93b, Cla96, CV92, DG95, FS94, GS99, GB98, HC96, MR97a, MR97b, Ye93]. Privacy [Lju93a, Lju93b]. Probabilistic [GMD95, Gel94a, Gel94b, Mer94a, Mer94b, Mer94c, SW90]. Probabilities [AMWK93, Ano91f, AG96, CWY97, Ede90, GR96a, GFT^{+}93, GW91, He90, Joe95, Lau92b, Li95, Mee90, Pes93, Riz92, Ros93b, Sta91, WK94]. Probability [BF90, BS98, GKB98, HPS98, KO95, LPT91, Roj96, Sär96, SK93a, SMD93, SO94, WK94, Xio95]. Probability-of-Exposure [SK93a]. Probit [GHCF94, GWG98, LK92a, RK96, SA92]. Problem [Akr91b, Ano96i, Bro97, CH91b, DKR95, EL90, Fab91a, Fin93, JK96, Liu94, LWT94, Nai91, PR90a, PSVS92, QTC96, SC96a, Sha91, SM90b, Tan94, Vel98a, Wel90b, WHGA96, Fab91b, FKOR98]. Problems [Ano91g, Ano97i, BH93a, BJB95, Buo90, CM91a, CC96a, CS92a, DH91, EL97, GJ99b, GSL92, GK93, JMV94, KM91b, KLW94, Lav91b, LA93, LTW96, PY99, SCW96, SCW97, WPD98, Zha99]. Procedure [Akr91a, BCG99, Cre90, DT92, Giv99, HOS94b, LW90a, LW90b, MBR98, PY99, Sit92]. Procedures [BW92, BW94, BD92, DSS93, Fin93, GR93, GS97, HS93a, HH94, HH91, MH94a, MSZ93, MH99, SG96b, Sob91, Sob93, WH90, WC91]. Process [AP96, Ano98h, Coo91a, CHP96, Esc94, GB98, Haa95b, HOLDG98, Luc93b, MLMT94, MG97b, RC94a, RC94b, The93a]. Processes [AF97, Ber91a, BR98d, CS95a, DR98, DG92, Hof92, KM99a, KM99b, Kol99, KR93, KY96, LRM96, Liu95, NC99, PS93, SK93b]. Processing [CGGM98a, CGGM98b, CGGM98c, Joh98a, Joh98c, MT98a, MT98b, SHL98a, SHL98b]. Product [CTC96, Cro97, GS93a, Pag90]. Product-Moment [CTC96]. Products [Bal95, SY95]. Profession [Hun94b, Ima95]. Profile [CJ96]. Profiles [Bal95, BS97c]. Profiling [NGG97]. Prognosis [Gra92]. Program [AGM98a, AGM98b, Bak98b, Bak98a, HZS^{+}90]. Programs [Ano90q, DW99, FIB90, GMWZ99, HT90a, Rol90, Cox90, Dor90, HT90b, Net90, PT90]. Progress [Ste94, ZT97]. Progression [Ano92f, Bri92, DLD91, LCG^{+}92b, LCG92a, Lou92a, TS92, YK96]. Projected [PML98]. Projectile [SM92]. Projecting [DKW97, DXP92]. Projection [Amm93, Dua90, SS95b, YF95]. Projection-method [YF95]. Projections [EV92, TA94a]. Promotion [BMR99]. Promotional [BG91b]. Propagation [Lau92b]. Properties [AP99, BAP95, BW94, CW91a, CD96, GEG99, HH91, LIZ99, RS94a, SK90, Val90, Zha92]. Property [CG92]. Prophylaxis [RG94]. Proportional [CL99a, FG99, GR92a, HN92, HR97, MRvdV97, Pag90, Peñ98, RT96b, Ryu94, SDW98, YP99]. Proportional-Odds [HR97]. Proportions [Amr95, Ano95e, HOS94b, Llo90, SG95b, SK90, WS96, SG95a]. Prospective [CWW95, DA96c]. Prostate [MPCB95]. Protocol [Hei95]. Provided [PSVS92]. Providers [NGG97]. Provisional [SK91]. Proxy [GBN96]. Pseudo [AS90b, KBB93]. Pseudo-Maximum [AS90b, KBB93]. Pseudolikelihood [GMR99, SI90]. Psychology [Kra99]. Public [Ano91y, CRS^{+}91, Moy99, Nor90, Roc91, Zel92]. Public-use [CRS^{+}91]. Publications [Ano90m, Ano90n, Ano90o, Ano90p, Ano91u, Ano91v, Ano91w, Ano91x, Ano92r, Ano92s, Ano92t, Ano92u, Ano93o, Ano93p, Ano93q, Ano94n, Ano94o]. Pulses [GWB99]. Punishment [Gro90]. Purchase [AL94, ALJ99, GS93a]. Purposive [AGM98a, AGM98b]. Pursuit [Amm93, Dua90].

Quadratic [CV92, GK93, JS94, Lie94, Rol94, Sla94]. Qualitative [MT95a]. Quality [Ano90q, BF92, BCG99, CP93, Cox90, Dor90, FIB90, GJ94, GHMW94, HT90a, HT90b, JS97b, Kra90, LS93, Net90, NGH^{+}95, PT90, Riv98, Rol90]. Quality-of-Life [BCG99]. Quantal [CDS99]. Quantile [BVT96, DG92, EM90, JSKK98, KM99b, LaR91a, LTW96, OR96, SM90a, YJ98]. Quantile-Quantile [EM90]. Quantiles [Cha96b, HK92, MBR91, Ros95a, RV90c, Sat95, SM92]. Quantitative [HTP95]. Quantity [CP93]. Quasi [Bre90, CS90a, CM98, Efr99, FHW95, Goo94b, Jun96, Kul95b, Rag93, SS94, Sta94a, Sta94b]. Quasi-Dependence [Goo94b]. Quasi-Empirical [Rag93]. Quasi-Independence [Goo94b]. Quasi-Likelihood [Bre90, CM98, FHW95, Jun96, Sta94b, CS90a, SS94, Sta94a]. Quasi-Linear [Efr99]. Quasi-Residuals [Kul95b]. Questionnaire [RG95]. Quindimil [Stu99, Wil99].

R [Ano99h, Kas97b, GBA^{+}99]. Racial [YK96]. Racial/Ethnic [YK96]. Radial [Gal94, MG97a]. Radiation [PSVS92]. Radiotherapy [CNF95]. Rail [Sho93]. Rainfall [KP91]. Raking [DSS93, Ski91]. Random [AP99, Amr95, AL95, Bre91, BD91b, Che94b, Coo91a, Cro95, DKS97, dE97, GHCF94, Har91, Hoc91, JTC99, KR76, KR90a, KO96, LB94b, MS92, OKS97a, Pai97, RS96a, SC96a, SCM91a, SLS97, SSW99, Spe91, SW97a, VL96, WL93, Wan98b, Yas95, ZK91, ZWWB96, DH93, SCM91b]. Random-Effects [Cro95, GHCF94, VL96, ZWWB96]. Randomization [Raz90a, Raz90b, Rom90]. Randomized [Ano92e, BC90, Bak98b, Bak98a, Lju93a, Lju93b, RG94, SO94]. Randomly [CLWZ95, HP92, KS97]. Randomness [SB90]. Range [BHL93, Fin90, Hay90, LR97b, Nyc91]. Rank [Akr90, Akr91a, AAB97, Ano96g, BH94, BD92, CMNS99, CM97b, CM91a, CD96, Dab90, DM93, DM96, GH99, GL92, GW99, Gri92, HMO97a, HMO97b, Hös94a, Kai94, KL99a, KW96b, KL99b, LD92, LS93, Mar92b, MM95, MSH90, PR90a, Tab90, TA90, Tho91b, Yan97]. Rank-Based [DM93, DM96, GH99, Hös94a, MSH90]. Ranked [AD90, BW92, BW94, KS94]. Ranked-Set [BW92, BW94]. Ranking [Fon92]. Rankings [BW94]. Raphson [LB94d, LB94c, LB94e]. Rare [SH92]. Rasch [LCG91]. Rate [AMCS98, BRR92, BPR99, DA96c, GT98, GMCM96, HL96, HW95b, MN93, MZTT98, RZ98, O’N94]. Rates [AKM94, BR95, BBH96, CNWW98a, CNWW98b, DLML99, GS95, KR96, Min98, PH90b, RG94, Ros95c, WCXG97, YL97, Ros95b]. Rather [DB97, SRR95]. Rating [BP94a]. Ratings [Ueb93]. Ratio [AME96, Agr98b, Ano99h, Ano99i, BB98c, DD90, DKR95, FS95b, GR92b, HMY97, Jen91, Jen95, KC93, Li95, LQT97, Lie94, MPS98b, Mur95, Pen97, Ski91, SW98b, SW98d, TL97, VW95, VHE91, VM96, Wan96, WM98, WKL95, Xio95, Yas95]. Ratio-Based [Mur95]. Ration [Agr98a, BB98b, MPS98a, SW98c]. Rationality [SSC94]. Ratios [Shi91, WC92]. Raw [HYP97]. Ray [EP94]. Reach [LS91b]. Reaction [NA99]. Real [DR91]. Real- [DR91]. Really [CB94]. Reasonable [Bly93a, CW93a, GKS93, Ped93, Rao93, RHS93a, RHS93b]. Reasoning [KSS96]. Recall [RMS90]. Recapture [DFGJ93, KMT99, Pol91]. Received [Ano90m, Ano90n, Ano90o, Ano90p, Ano91u, Ano91v, Ano91w, Ano91x, Ano92r, Ano92s, Ano92t, Ano92u, Ano93o, Ano93p, Ano93q, Ano94n, Ano94o]. Receiver [Llo98]. Recensoring [LLE93]. Recognition [GS98a, PJT96]. Reconciliation [Goo96, SR94]. Reconnaissance [Hod94]. Reconstruction [AZ99, Kol96]. Reconstructions [Wei97]. Record [BR95, Sol95]. Records [Are92, NFL92a, NFL92b]. Recovering [GJP96]. Rectangle [Joe95]. Recurrence [CW99]. Recurrent [HL96, PC93, SPM99, WC99]. Recursive [Gri94, KS96]. Recycling [NG94]. Red [RCH94]. Redescending [BH94]. Redistricting [GK90]. Reduced [SS97]. Reduction [Bri91a, CL92a, CW91b, CL99b, HT91, Ken91, Li91a, Li92, Li91b]. Reevaluating [DW99]. Reference [Ano99a, Cla96, Goo94b, GK93, KW95, LRR91, QH98a, QH98b, SY95, Ye93]. Referendum [CH91a]. Refinements [PWM92]. Refining [Ber90b]. Reflections [Nor90]. Reform [DN98a, DN98b]. Regeneration [GRS98, Joh98b, MTY95]. Region [BO97, HW92a, PC96, SS96c]. Region-Based [BO97]. Regional [Oeh93b]. Regions [Ano91f, Ber90a, CHP96, FH89, FHJW96, MZ97, Mee90, Tho90a, WCH98, FH90]. Regression [AL94, AHY93, Ano96f, Ano97i, BRMZR97, BL92, BR94b, BL93, BVT96, BDM^{+}93a, BDM^{+}93b, BdL99, BZ92, Boc99, Bre92, Bre90, Bri91a, BGH93, BS94, BRW99, CL92a, CRS99a, CRS99b, CW99, CMNS99, CJ95a, CJ95b, CTC96, CL99a, CW94a, CM98, Cho92, CM97c, CMM93, CRS^{+}91, CH93b, CW90, CW91b, Coo94, CN94, CW97, CL99b, Cox90, DwSH94, Dea92, DB97, DBB^{+}94, Dor90, Dor94, Dua90, DS97c, ES90a, EM92, ES90b, ES93, Fan92, FG94, FG95, FHW95, FS95a, FOS97, Fer98b, FKYZ99, FL95, FS95b, FWW99, GMP97, GM99, Gey91, GS91a, GLG92, Gle92, GF96b, GK93, GL96, GS93b, Gu90, Gu92, GY93, GP92, HH90, HW91, HT97a, HT97b, HT97c, HT90a, HT90b, HT91, HHM92, HW92a, Haw93a, He99a, He99b, HZ96, HZ98, Hil93, HR97, Hun94a, HZ90, Jay96]. Regression [JTC99, JDP94, Jon96, Jun96, KR96, KM91a, Ken91, KC93, Koe99a, KM99b, Koe99b, KST95, KBS97, Kul95b, KW97, LC95a, LT96a, LWY93, LS91a, Li91a, Lin91, LW92, LY93, LCWH97, Lit92, Lit93b, Liu96, LS99a, LS99b, LJK93, MS99a, MW98a, MW98b, MW98c, MY91, MEGN92, MSH90, MS99b, MS99c, MNS99, MT95b, Mul99, Net90, NCC96, Nyc91, NGH^{+}95, OH99a, OH99b, PY99, PC93, Por97a, Por97b, PM99a, PM99b, PT90, RMH97, RSY93, Raz90a, Raz90b, RRZ94, RRZ95, RR95, RW98a, RW98b, Rol94, RT96b, RRS98, RH99, RVH99, RHV99, RSW95, Rup97, SW95a, Sam93a, SS97, Sch94d, Sch94e, Sch95b, Sev96, SSW99, She93, She97, SAK90, SKW99a, SKW99b, SRC92, Sit97, SH90, SK97, SL98a, Spu99, SC92b, Sta90, SS91b, SL98b, SB95, SCW96, SCW97, SR92]. Regression [Str93a, SGP98, Tab90, Tho90a, Vel98a, Wac93, WS94, WR95, WWZtO97, WJ92, WT90b, WCR94, Wie98, WK98, XL99, Yam92, Yan97, Yan99, YP99, YL97, YJW95, YJ98, Zha99, dHY93, Li91b]. Regressions [Coo96, DM90, KP90]. Regressor [CO90]. Regressors [RRZ94]. Regular [BH94]. Regularity [Vos93]. Regularized [BC96]. Regulation [LR96, Liu94]. Reinfarction [OMF^{+}93]. Reinterview [BF92, SG96b]. Rejection [LCW98]. Rejoinder [Alb93c, AIR96b, AIR96c, BDM^{+}93b, BRL99b, BR98b, BR98a, CCG^{+}97c, CCG^{+}97b, CGM98b, CGM98c, CGGM98b, CGGM98c, Coo98b, Coo98c, DG93a, EF91b, Efr94a, Efr96b, Efr96c, EFS92b, Fab91a, Fay96c, Fay96b, GKL98d, GKL98b, Goo91a, HT97b, HT97c, Han94, HT90b, HS93c, LKCH99b, LKCH99c, LCG92a, Las94a, LC92, Li91a, MW98b, MW98c, Mer94b, Mer94c, MS91c, NFL92b, RGZ95b, RGZ95c, Rao96c, Rao96b, RHS93a, RGH99a, RAGW93b, Rv90a, RvZ90, RVH99, RHV99, Rub96c, Rub96b, SCM91a, SRR99b, SBD97b, SBD97c, SKW99b, Sto91b, SW98c, SW98d, SJ90b, TW98b, TW98c, WMR91a, Efr94b, Fab91b, Goo91b, HW94, Las94b, Li91b, SCM91b, WMR91b]. Related [Ber91a, Fer98b, FMSP90, Fon92, GK96, KM99b, LCG91, MO98, OK96, Sam93b]. Relations [KSJ94]. Relationship [DMLC97, HH94, KW95, SS90a, TDW95]. Relationships [GMP97, GR96b, HH91, PSS93]. Relative [GC95, KCHH93, Wel90b]. Relatives [BGH97]. Relevant [HA94, HA95]. Reliability [AMCS98, CP99, CS96a, DC94, Gle92, KY96, Mei94, dCP99b]. Remarks [DG95]. Remedian [RB90]. Remedication [PHL98]. REML [CD95]. Removal [BR98d, Pol91]. Removing [HW91]. Renewal [DwSH94]. Repair [AP99, PHS94]. Repeat [CG92]. Repeated [Agr97, Akr91a, AA94l, AHB97, Ano91e, Car94b, Con92, GW99, JD91, JR96, JTC99, KW96b, Kus98, LD91, LD92, LB94c, Lit95, Luc93b, MLRR92, Neu99, OR96, RRZ95, RGH99b, RGH99a, Roc92, RRS98, RF99, SCC98, TA90, Tho91b, Was99a, ZDL96, ZL99, LB94d, LB94e]. Repeated-Measures [Ano91e, JR96, KW96b, Lit95]. Replacement [BBC98]. Replicate [Chr91, SC96b, Vos93]. Replicated [DLB91]. Replicates [MNS99]. Replication [SCC98]. Reply [Stu99, DG93b, Efr96c, RHS93b]. Report [Ano98k, Ano99j, CCG90, CCWG91, CLWG92, LHA96, LCK97, LWG93, LHA94, LHAL95]. Reported [GJDT98a, GJDT98b]. Reporting [DXP92, Har90b, Wan92]. Reports [RBD93]. Representations [BM92].

Reproductive [SGC95]. Reprojecting [GT98]. Resampling [DC94, Sit92, Tro95a]. Rescue [SMP95a, SMP95b]. Research [Kin99]. Residential [CG92]. Residual [CCD98, LR92, SG97]. Residuals [BAP95, CHW92, KS96, Kul95b, MSH93]. Resistant [Bra90, GR92b, Jac93]. Resolution [BR94a]. Respect [GS94]. Response [Agr97, AC93, BD97, Buo96, CO90, CM97a, CDS99, Coo96, CL99b, FORK99, Fyg97, GV96, GJP96, HYP97, Hil93, Hir92, HB99, KPR91, KC96, Lan96, Lee96a, LCG91, Lju93a, Lju93b, PSVS92, Rub91, SW93, SG97, ZL91, RRS98]. Response-Dependent [GV96]. Responses [AHB97, BP99, BZ99, FLZ96, HZ98, LA94, LMS99, RSV95, Zha98b]. Restoration [AGP91, O’S90]. Restricted [CD95, KT95, MT95b, NSS97, PF97, Zuc98]. Restrictions [GLG92, KT95, RH96, Shi91]. Results [CV95, GBA^{+}99, Hog93, LB94b, NM98, Whi89, Whi92a, Fab94, Whi92b]. Retrospective [GV96, IT94, SPM99]. Retrospectively [Wan92]. Returns [FP95]. Reverse [GS95]. Review [AA94a, AA94b, AA94c, AA94d, AA94e, AA94f, AA94g, AA94h, AA94i, AA94j, AA94k, AA95a, AA95b, AA95c, AA95d, AA95e, AA96, AB97a, AJM99, Ack91, Ack93, Ack98, Ahn99, Ahs90, Alb93b, Alb95, Ale98, Alt94, Ame99, Ame92, And95, Ang93, Ang90, Ano98q, Anr96, App97, Asm93, AS93, Ath99, Atr94, BB98a, BBB98, Bag90, Bag91, Bal94, Bal96, Bal97, Ban90, Ban93, Bar93a, Bar97, Bar92, Bar98, Bec91, Bec98, Bec99, Bel93b, Ben97, Ben91a, Ber94a, Ber94b, Ber93, Ber95, Ber92a, Ber92b, Bha94, Bha90, Bie96, Big98, Bil91, Bil96, Bla92, Blo98, Boi90, Bol91, Bol94, Boo94, Boo99, Bov98, Bra98, Bre96b, Bri98, Bri99, BGH94, Bro99, Bro90, Bro91, Bro92c, Bry97, Bud96, BF93b, Bun96, Bur92, BS91b, Buz97, Car90a, Car94a, Car90b, Cha96a, Cha92, Cha91b]. Review [Cha93b, Cha91c, Che91, Che92b, Che92a, Che95b, Che98, Chi91, Chi90, CS95b, Cla90b, Cla90a, Cla92, Cla98, Cla93, Cla99, Cle99, Cli99, Clo94, Col91, Col93, Col97, Con97, Coo91b, Coo93, Coo95, Cor99, Cor92, Cow91, CC96b, Cow98, Cox99, Cra95, CL94, Cry99, Cwi94, D’A93, D’A94a, D’A94b, D’A94c, D’A94d, D’A95a, D’A95b, D’A95c, D’A96a, D’A96b, DSS97a, DSS97b, DSS97c, DSS98, DSS99, Dab92, Das95, Dav92, Dav99, Dav91, Dav98, DeG95, Dev95, DiC95, Dia99, DE96, Dic96, Die93, Die97, Dog95, Dra94a, Dra94b, EJD97a, EJD97b, EJD98a, EJD98b, Eas92, Edd90, Ela91, Eld96, Eli98, Emb95, Emb96a, Emb96b, Emb97a, Emb97b, Emb98a, Emb98b, Emb98c, Emb98d, Eub93, FH95, Far97, Fau90, Fau94, Fer98a, Fie93b, Fis91, Fla90, Fla91]. Review [Fle90, Fle91, Fou92, Fou93, Fra99, FM93, FKOR98, Fre95, Fre98, Fre99b, Fri98, Fun90, Fun92, Gas91, Gas95, Gas96, Gav91, Gen93, Gen99, Gho90, Gho94, Gho96, Gho98, Ghy98, GH91, Gil94, Gin95, Gla91, Gly98, Gol95, Gol93, Goo92, Goo94a, Gra91, Gre91, Gro92, Gru93, Gru98, Gue96, Guo93, Gut90a, Gut90b, Gut91a, Gut91b, Gut92a, Gut93a, Gut99, Gut90c, Gut92b, Gut92c, Gut93b, Haa95a, Haa96, Hab95a, Hag97, HRF96, Hal97, Ham90a, Ham93, Ham90b, Har96, Hau97a, Hau97b, Hau90a, Hau90b, Hav95, Haw93b, Haw96, Haz95, Hea90, Hei90, Her98, Her93, Her97, Her91, Hig98b, Hil92, Hil91, Hil95, Hip95, Hod91, Hol98, Hol99b, Hol99c, Hos97, Hsi90, Igl91, Igl92, Iye97a, Iye97b, JEB98, JFM98, JGB99, Jam98, Jan95, Jas92a, Jas92b, Joh95]. Review [Jon90, Jor96, KB97a, KB97b, KK94a, KK94b, KK94c, KK95a, KK95b, KK95c, KK95d, KK95e, KK96a, KK96b, KK96c, KK97a, KK97b, KK97c, KK97d, KK97e, KK97f, KK98a, KK98b, KK98c, KK98d, KK98e, KK98f, KK98g, KK98h, KNB98, Kab98, Kad90a, Kad91, Kaf94, Kaf95, Kaf97a, Kaf97b, Kaf97c, Kaf97d, Kaf97e, Kaf97f, Kaf97g, Kaf97h, Kaf99a, Kaf99b, Kan90, Kar90, Kas97a, Kas90, Kas97b, Kat97a, Kat97b, Kaw99, Kep97, Kfa97, Khu94, Kia90, KS96, Kir98, Klu94, Koe97, Koo91, Kot93, Koy91, Kri94, Kro90, Kul95a, Kup90, LD96a, LD96b, LD97a, LD97b, LD98a, LD98b, LDH99, LH96, LH97a, LH97b, LH97c, LH97d, LH97e, LaV90, Lac90a, Lac90b, Lac90c, Lac90d, Lac91, Lac92, Lac93a, Lac93b, Lac93c, Lam97, LKM98, Lan93, Lan95a, Lan95b, Lan97, Lar91b, Lau92a, Lee93, Lee96b, Leh97]. Review [Leh98, Lel99, Len98, Len99a, Len99b, Len99c, Len99d, Leo92a, Leo92b, Lev94, Li98d, Li99, Lig98, Lig94, Lin98, Lin94, Lin99a, Lin99b, Lip95, Lit94, Lit96, Lit92, Liu97, Lju95, Loc90, Loh96, Lon91, Lor92, Lou99, Lou96b, LuV91, Luc98, Lun97a, Lun97b, Lun98b, Lun98a, Lun98c, Lun98d, Lun99, Lye97, Lyn98, MTW96a, MTW96b, MTW96c, MTW96d, MTW97a, MTW97b, MTW97c, MTW98a, MTW98b, MTW98c, MTW98d, MTW98e, MTW98f, MTW99a, MTW99b, MTW99c, MTW99d, MTW99e, MTW99f, MTW99g, MTW99h, MTW99i, MTW99j, MTW99k, MTW99l, MTW99m, MTW99n, MW95a, MW95b, MW95c, MW96a, MW96b, MW96c, MW96d, MW97a, MW97b, MW97c, MW97d, MW97e, MW98d, MW98e, Mad93, Mal95, Mar94, Mar90a, Mar92a, Mar98a, Mar95a, MA95, Mar90b, Mar95b, Mar98b, Mar98c, Mar92c, Mar98d, Mas92, Mas98, Mat99, Mat94, Mat93, McC98a, McD90, McD92, McD97, McG90]. Review [McK92, McK95, McN97, McN98a, McN98b, Mee95, MS90, Mel90, Meu95, Mey99, Mic97, Mic98b, Mic98a, Mik93, Mik94, Mik95a, Mik95b, Mik97, Mik98, Mit94, Mod95, Moo91, Moo99, Mor90, Mor91b, Mor93a, Mor93b, Mou95, Muk90, Muk92, Mul91, Mul93, Mul98, Mur97, Mur90, Mur91, NL94a, NL94b, NL94c, NL95, Nac91, Nak92, Nar93, Nei97, Nel92, Nel97, Nel98a, Nel98b, Neu92, New92, New90, Nol92, Not90, Not94, O’C99, O’R91, Oeh93a, Ogu99, Olk91, Olk92a, Olk92b, Olk93, Olk97, Olk99, Ols90, Omo95, Ove91, Ove94, Owe92, PAL96a, PAL96b, PD99, PE94a, PE94b, PE94c, PE94d, PE95a, PE95b, PE95c, PE95d, PE95e, PE95f, PE95g, PE95h, PE96a, PE96b, PE97a, PE97b, PE98a, PE98b, PE98c, PE98d, PE99, Pad93, Par91, Pau96, Ped94, Per96, dBPI97, Per90, Pfe99, Pic97, Pie93]. Review [Pin97, Pit99a, Pit99b, Ple91, Pod94, Pol90, Pot96, Pre94, Pre96a, Pre92, Puh98, Qua90, RAJ96, RBD96, RBL97a, RBL97b, RBL99a, RBL99b, RBL99c, RC99, RJ95, RJ96, RL96, RL98a, RL98b, RL98c, RLS99a, RLS99b, RS95a, RS95b, RS95c, RS95d, RS95e, RS95f, RS96b, RVL97a, RVL97b, RVL97c, RVL97d, RVL97e, RVL97f, RVL99a, RVL99b, RVL99c, Rad95, Ram93, Rat98, Rat94, Rau91, Rav97, Ree98, RN97a, Rei90, Ric90, Ric97b, Ric98, Rob90, Rob91, Roc93, Rod90, Roe92, Ros94, Ros99, Ros97a, Ros97b, Ruk98, Run99, Rut91, SRJ99a, SRJ99b, SRJ99c, SRJ99d, SRJ99e, SRJ99f, SRJ99g, SW94a, SW94b, SW95b, SW95c, SW95d, SW96, SW97b, SW97c, SW98e, SW98f, SW99, Sal93, Sam90a, Sam97, San91, San90, Sap91, Sar91a, Sar94, Sar91b, Sat90, Sca92, Sch96a, Sch91a, Sch94a, Sch90b, Sch96b, Sch95a, Sch94b]. Review [Sch90c, Sch94c, Sch98a, Sch98b, Sch99c, Sch97, Sch91d, Sch95e, Sco99, Sco90, Sea98, See95, See94, SM93a, SM99, Sel95, Sel96, Sen92, Sen93, Sen94, Ser93, Ser91, Set95, Sey97, Sey99, Sha92a, Sha99, Sha98, Sha92b, Sha93a, She95, Shl90, Sho92, Shu90, Shu93, Sim94, Sim93b, Sin97, Ski96, Sla93, Smi94, Smi96b, Smi96a, Smi98, Sob96, Sob98, Sol94, Spe93, Spe98, Sri93, Sta98a, Sta99, Sta98b, Sta93, Ste90a, Ste98, Ste90b, Ste92, Sti95, Sti92, Sto94, Sto97, Sto90, Str98a, Str99a, Str99b, Str99c, Str99d, Str95, SZ99, Str93b, Str94, Str98b, Stu92, Sud90, Sun99, Swa99, Swi90, Sys96, Tah92, Tak92, Tan90, Tas98, Tay90, Tei90, The93b, Thi94, Tho90b, Tiw99, Tro95b, Tsa90, Tsu94, Tur98, Tur90, Utt93, Utt94]. Review [Vau90, Vel98b, Vel99, Vin93, Vog92, Vog99, WP92, Wak91, Wal91, Wal95, Wal96, Wal98, War91, War97, Was97, Was98a, Was98b, Was90, Was91, Was93, Was94, Was99b, Wat90, Weg90a, Weh94, Wei90a, Wei91a, Wei98, Wei91b, Wei90b, Wei94, Wei95a, Wel99, Wel94, Wil91, Win96, Wol90, Wol93, Wol91, Won93, Won96, Woo90, Woo91, Woo93, Wu98, Wue96, Yan90, Yan98, You97, Zaj93, Zeh95, Zer96, Zer99, Zim92, dL93]. Reviews [Ano90i, Ano90j, Ano90k, Ano90l, Ano91q, Ano91r, Ano91s, Ano91t, Ano92n, Ano92o, Ano92p, Ano92q, Ano93k, Ano93l, Ano93m, Ano93n, Ano94j, Ano94k, Ano94l, Ano94m, Ano95l, Ano95m, Ano95n, Ano95o, Ano95p, Ano95q, Ano96n, Ano96o, Ano96p, Ano96q, Ano96r, Ano96s, Ano96t, Ano97f, Ano97g, Ano97h, Ano97q, Ano97r, Ano97s, Ano97t, Ano98f, Ano98g, Ano99f, Ano99g]. Revising [XP92]. Revisited [Coo98a, Coo98b, Coo98c, Li98b, Li98c]. Reweighted [DS91]. Reweighting [CN94]. Rheumatology [BS97b]. Rhythm [PF97]. Richardson [Kas97b]. Richness [HAS97]. Right [GL96, Gür96, LWT94, Qui98a]. Right-Censored [GL96, LWT94]. Right-Truncated [Gür96]. Rigid [PC96, Riv98]. Ring [DH91, Lav91b]. Risk [BRR92, BSB98, BF90, FG99, HM91, HTP95]. Risks [FW96, HW95b, Mei94, Pep91]. Robust [Akr91b, Amm93, Atk94, CDW93, FKYZ99, HR94, LR95a, Lav91a, LRM96, Liu96, LM93, MH90, MSH93, RS94b, RFB97, RB90, Sil92, SRSD94, Tam95, TR95, Vel95, VFR97, WS94, WM90, WZ97b, Wie98, WK98, WR94, MY95]. Robustification [EMM95]. Robustness [FOS97, GEG99, HSP90, KC93]. Rodent [HZS^{+}90]. Roles [SB97]. Room [Mod93]. Root [CJ96, MBL98, NP95, SL93a, TR97, jW97]. Roots [GHL96, YR95]. Rotation [CD92]. Roughness [Hua97]. Rows [HW90a]. Rubin [Efr94b]. Rule [Pen97, Ye93]. Rules [BG99, Bra90, KW96a, MG97a, MP99, Vid98, WSLP90]. Run [Lou96a]. Runs [FK94, Lou96a, McW90, RMS90].

S [Kas97b, Lit92]. Saddlepoint [BS98, GR96a, GJ99b, Lie94, WBB93]. Safety [GBA^{+}99]. Salamander [Shu97]. Sales [CG92]. Sample [Akr91b, Amr95, Ano91h, Ano95e, Ano97i, BP99, BW92, BPS95, CL92b, CT98, Chr91, CM91a, Cla96, DFGJ93, DKR95, EL90, HMO97a, HMO97b, Joh96b, Kai94, KM91a, KSW92, Kot94, LRR91, LTW96, LWT94, MSL93, PR90a, PWM92, RS96a, RN97b, Riz92, RK90, SW95a, SG96a, SG95b, Sob91, SN90, SCW96, SCW97, TF90, WH90, Zie90, SG95a]. Sample-Admissible [Sob91]. Sampled [LW91]. Sampler [Cha93a, KT94b, Liu94, RT92, ZM95, KT94c, KT94d]. Samplers [MTY95]. Samples [AK92, BBH96, BW92, BW94, BR98b, BR98c, BR98a, CH95b, CRS^{+}91, CS96b, FZ98a, FZ98b, GK95, Hau92, Kne98a, Kne98b, KS93, KS94, LT97, Lan98a, Lan98b, LTP94, Owe94, RS98a, RS98b, RSY93, Ros95a, VS93, WW98a, WW98b]. Sampling [Ano91g, BBC98, Buo90, BS98, CD92, DS92, DSS93, DD90, Dor94, GHRPS90, GS90a, GSL92, GS99, GM93b, GR96b, GV96, Hei95, HC96, KM91a, LKM98, LS91b, LNL95, LCW98, MV91, MS96, OB93, Per97, RS96a, Ros93b, Sär96, SS99, Sit97, SO94, Tam95, Thé99, Tho90c, TS96, Val90, Ver93, ZK91, ZT97, Zha96]. Sampling-Based [GS90a]. Satellite [BR92, FS94, NT95]. Satisfaction [BZ99]. Saturated [BO95]. Savage [VW95]. Savage-Dickey [VW95]. Scalar [Ahn97]. Scale [AL95, CF95, Gle93, Har92a, MY91, WM90]. Scaled [CW94a]. Scaling [O’C99]. Scan [HL97, NW97, POH97]. Scheduling [LZ98a, LZ98b]. Scheme [DC94, Joh98b]. Schemes [APR92]. School [AK92, MSMP92]. Schwarz [KW95]. Science [Zel92]. Score [NC99, SS95a]. Scores [BH96, GS98b, Ste94, Tab90]. Scoring [HTB94]. Scorings [KSW92]. Screening [Ano92e, BC90, Bak98b, Bak98a, GJ94, HTP95, LTP94, Par93b]. Sea [LR97b, PM97]. Search [CGM98a, CGM98b, CGM98c, DSM98a, DSM98b, KKT98a, KKT98b, MBL98, Zha98a, Zha98c]. Searching [GK95]. Seasonal [FMSP90, GHL96, PH90a, TR97]. Second [GT91]. Second-Order [GT91]. Secondary [BHB96, WSJ92]. Section [Sch93b]. Sectional [Wan91]. Secular [HL94]. Seemingly [BG91b]. Segmentation [BO97, Joh94]. Seismic [PB92]. Select [MNS99]. Selected [CB94, WK94]. Selecting [GS94, TA94b]. Selection [Ano98j, BP96, BZ92, Bre92, BRW99, CKLH92, EL97, GM93b, GM99, GLB98, GMCM96, HZ90, JMS96, KW96a, KW97, LRM96, LA93, MS99a, MR94, MBR98, Mul99, NP95, OP96, OR98, Riz92, RFB97, SH92, Sha93b, Sha96, SKW99a, SKW99b, Sil97a, Sob91, Wan95, Wei95b, Ye98a, Zha92, ZL95]. selectivity [Mil92]. Selector [Ano96f, RSW95]. Selectors [PM90, jW97]. Self [BHB96, GJDT98a, GJDT98b, Kit98, Tar99, WK94]. Self-Consistency [Tar99]. Self-Examination [BHB96]. Self-Organizing [Kit98]. Self-Reported [GJDT98a, GJDT98b]. Self-Validating [WK94]. SEM [MR91, XLR91]. Semiconductor [Ano98h, HOLDG98]. Semilinear [ES97]. Semiparametric [CWY97, Dig99, FZ99, For94, Fre99a, GMP97, HMM98, HH96, Hun94a, JTC99, KS97, LP99, LC95a, LQT97, LCG91, LR99, NCC96, NCDF96, PS94a, RRZ95, RR95, RCL96, RT96b, RRS98, STR97, SRR99a, SRR99b, SKC94, SS94, Sin93, SD97, Sta94b, WWZtO97, YP99, Zel93, ZLRS98, vdL99, Sta94a]. Sen [AML95]. Sensitivity [Ano99a, BPS96, CH91a, Gus96, Lav91c, QH98a, QH98b, RK90, Sev96]. Separate [Ano97j, BHB96, Ruk93]. Separating [GS97]. Sequence [KM91b, LNL95, LNL99, YK96]. Sequences [CLWZ95]. Sequencing [Hil93]. Sequential [EGMP93, GY93, HTP95, JD91, JT97, KLW94, LD91, LD92, LB94a, LC95b, LCW98, LC98a, LC98b, STR97, SMD93, TS96, Xio95]. Sequentially [FR90, HPS98]. Sera [LTP94]. Serial [AP96, Ber91a, Hei91, ZDL96]. Series [Ada98, CMW97, Cha90, CL95, CW91a, CL93, CLM93, CB90, Cre90, DR96, EH91, GNK96, GM94, GS90b, GWB99, Hon99, KST98, KL90, Koo97, LK91, LMR96, LS91a, Li96, LT98, LL97c, Luc93b, MT93, MP96, MSL93, Niu96, SAK90, SK91, Sto99, The93a, TA94a, TA94b, VHE91, WPK99a, WPK99b, dJP98]. Service [GWG98]. Set [BB94, BW92, BW94, Daw95, FHTW97, Haw93a, Jac93, KS94]. Sets [Ber90b, DD90, HH94, Hol99a, Roe90, RB90, BCH95, Hab95a]. Several [Agr98a, Agr98b, Ano99h, BHL93, BB98b, BB98c, CV95, MPS98a, MPS98b, MH99, SW98b, SW98c, SW98d, TL97, Wan96]. Severini [Sta94a, Sta94b]. Severity [Bro92b, Wei92, YKB97]. Shape [WR94]. Shaping [Ket97a, Ket97b]. Shares [Zas93]. Shift [FNP95, LWT94]. Shifts [MT93, The93a]. Shocks [dJP98]. Short [KL90]. Shrinkage [BG91b, CKM97, DJ95, KCHH93, Kol96, Vid98]. Shuttle [DH91, Lav91b]. Sided [Ano96i, CS92a, CF95, Fol96, HHS97, HY98, Hay90, MH99, Sil92, SS95a, Tan94]. Sides [WR95, KT94a]. Sieve [HR97, SSW99]. Sign [BHS98, BHNO92, CS93]. Signal [BSPK98, CCD97, GWB99, KA91]. Signals [Sto99]. Signed [CJ96, Dab90, HMO97a, HMO97b, PR90a]. Signed-Rank [Dab90, HMO97a, HMO97b, PR90a]. Significance [Ano91i, Fan96, FL98, Fin90, Fra91, LRR91]. Significant [TA94b]. Simes [SC97a, SC97b]. SIMEX [CKLS96, WLGC98]. Similar [LIZ99]. Similarities [HHH92]. Simple [Eas94, Fol96, FWW99, Hay90, JR96, KC93, LCG91, MM93, Sär96, Wel90b]. Simplex [RS94a]. Simpson [Mit91, Sam93b]. Simulating [Zel93]. Simulation [CS94, DKRW97, DD90, LR97c, Ode91, PS99, SC95]. Simulation-Extrapolation [CS94, SC95]. Simultaneous [Ano95e, BS96a, Ber90b, BM96, FS95a, FNP95, MY91, Sat95, SG95a, SG95b, SA92, SN90]. Simultaneously [LA94]. Single [CFGW97, Goo96, HH96, Sev91, Vos93, Wei97, Whi89, Whi92a, XL99, Fab94, Whi92b]. Single-Index [CFGW97, HH96, XL99]. Single-Parameter [Sev91]. Single-Photon [Wei97]. Singly [dCP99a]. Singular [Amm93]. Situ [ST92]. Size [Amr95, Ano95e, Cla96, GMP97, GLG92, Mil92, PWM92, RZ98, SG95b, SN90, SG95a]. Size-selectivity [Mil92]. SiZer [CM99]. Skating [BP94a]. Skewed [Che95a, CDS99]. Skewness [FS98b]. Skilled [KR96]. Skin [BHB96]. Sliced [Bri91a, CW91b, Fer98b, HT91, Ken91, Li91b, Li91a, Sch94e]. Sliding [FMSP90]. Sliding-Spans [FMSP90]. Slovenian [RSV95]. Slowly [Ber91a, BG91a]. Small [Ano97i, Chr91, GNSC98, HOS94b, HH91, LR95a, MSML97, PR90b, Rag93, SCW96, SCW97, TF90, WC91]. Small-Area [HH91, PR90b]. Small-Sample [Chr91, SCW96, SCW97, TF90]. Smith [Ano91h, DE96]. Smooth [DLW95, GF96b, KL97, LaR91a, Led94, MZ96, Peñ98, Raz90a, Raz90b, Ter94]. Smoothed [EL97, Llo98, MBR91]. Smoothers [CGGM98a, CGGM98b, CGGM98c, Joh98a, Joh98c, MT98a, MT98b, SC92a, SHL98a, SHL98b]. Smoothing [Alt90, AM92, BF93a, BR98b, BR98c, BR98a, Cho93, DJ98, EL97, FW96, FHMP96, FZ98a, FZ98b, GKK91, GMCM96, Gu90, Gu93, HHM92, HS98c, HS98d, Kne98a, Kne98b, KAT91, Koo97, KW97, Lan98a, Lan98b, MT95a, MST97, NNR95, Nyc91, O’S90, O’S91, Par93a, RS98a, RS98b, Sto93, Ter90, Tho91a, WJ93, WW98a, Wan98b, WW98b, WCH98]. Smoothness [DJ95]. Soccer [RCH94]. Social [Ano91y, GMWZ99, Pea91, Roc91, SI90, Yam90]. Society [Wal93]. Socioeconomic [GBN96]. Software [AMCS98, KY96]. Soldering [MLMT94]. Solution [FKOR98]. Solutions [SG96a]. Some [Akr90, AKM94, Ano91g, Ano98i, Buo90, CS92b, DG95, DR96, Fab94, Fin90, Gal94, Gei92, Goo96, HMN94, Hou95, HH91, KM91b, Las94a, LH91, Mac94, MSZ93, MS98b, PC93, Riv98, RRZ94, SK90, Val90, Whi89, Whi92a, Whi92b, WBB93, Las94b, MS98a]. Sort [Ver93]. Sound [BSPK98]. Source [BR94a, GMWZ99, Var96]. Source-Destination [Var96]. Southern [RC94a, RC94b]. Space [ASL91, CPS92, Erk94, GL92, GS98b, Kit98, RC94a, RC94b, SW91a, DH91, Lav91b]. Space-Time [RC94a, RC94b]. Spaced [JV93]. Spaces [CD95, Lav91a]. Spacings [Kai94]. Spans [FMSP90]. Sparse [HT97a, HT97b, HT97c, Por97a, Por97b]. Spatial [AF97, Ano97k, Ano98h, BK99a, BK99b, BR98d, Chr99a, Chr99b, CH92, Cre94, DR98, DLB91, Gel94a, Gel94b, Gut94b, Han94, Han99a, Han99b, HL98, HAS97, HOS95, HOLDG98, JV93, KDG94, KP92, KO96, LKCH99a, LKCH99b, LKCH99c, Mer94a, Mer94b, Mer94c, MST97, POH97, SG92, SC94, SC99a, SC99b, Ste95, Ste99a, Ste99b, WEST91, YL97, HW94]. Spatial-Temporal [Cre94, Gut94b, Han94, KDG94, KO96, HW94]. Spatially [BR94b]. Spatio [CH99, Haa95b, WCXG97]. Spatio-Temporal [CH99, Haa95b, WCXG97]. Special [Goo94b, Sch93b]. Species [BBC98, BF93b, HAS97]. Specific [AKM94, BRR92, GFT^{+}93, PH90b, Sch91c, WM91]. Specification [DM90, GMD95, HN92, Ryu94, Ter94]. Specificity [Ano99a, QH98a, QH98b]. Specified [Cha91a, LWW91, RS90a, Wie98]. Specimens [Hei95]. Spectra [SRSD94]. Spectral [Ada98, Efr98, Hon99, PO94, SS90b, Sto99]. Speech [PJT96, SB90]. Sperm [NA99]. Spherical [FOS95, MG97a]. Spherical-Radial [MG97a]. Sphericity [Ano91e]. Spiegelhalter [Eld96, Kas97b]. Spike [Bri92]. Spline [ASL96, BR98b, BR98c, BR98a, Cho93, FZ98a, FZ98b, Gu90, Gu93, HS98c, HS98d, HK92, Kne98a, Kne98b, KAT91, Lan98a, Lan98b, RS98a, RS98b, Tho91a, WW98a, Wan98b, WW98b]. Splines [DS97c, Gra92, HMN94, HK93, Las94a, LS91a, LW97, Mac94, SH90, Las94b]. Split [LC93, MS92, RG95]. Split-Plot [MS92]. Sports [Alb99, BRL99a, BRL99b, Gut94a, Kad99, Sch99a, Ste94]. Spread [Mel99]. Squared [EL90, HJ92, HP92, LR95a, PR90b, Ryu94]. Squares [AI95, Ano96f, CD91, CW94a, CHW92, DS91, DS97c, ES90a, Gar94, GF96b, Hoo93, HM91, IT94, RSY93, RSW95]. Stability [SRC92]. Stable [Buc95c]. Stage [AI95, Ano91h, HOS94b, RSY93]. Stagewise [Hol99a]. Stahel [MY95]. Stahel-Donoho [MY95]. Stamp [BD97, PH90b]. Standard [BH98]. Standardized [LKC97]. Standards [PH90b, SCF93]. Staniswalis [Sta94a]. State [ASL91, CPS92, GL92, GS98b, Kit98, PH90b, SSK90, SW91a]. State-Dependent [SSK90]. State-Space [ASL91, CPS92, GL92, GS98b, SW91a]. State-Specific [PH90b]. States [DLML99, CH93a, LT94b, Pas93, RAGW93a, RAGW93b]. Stationarity [Ano98i, MS98b, MS98a]. Stationary [BM95, CH99, Mau95, PR94]. Statist [Fab94, LB94d, MS98a, MS95a, SG95a, Sta94a, Wee94, Wes98, Whi92b]. Statistic [AP99, Ano91f, BW94, Ede90, HL97, McW90, Mee90, NW97, POH97, SC94]. Statistical [Alb93a, AASM93, Alb93c, Bea91, Bri91b, Bro92d, CH92, Cre94, DDD94, Eld96, EFS92a, EFS92b, Fra91, Gei92, GR93, GHMW94, Gre92, Gut94b, Han94, Har92b, Hau92, Kad90b, KSJ94, KBB93, KM91b, KDG94, Lou92b, MSZ93, Mor91a, NGG97, OBW92, OKS97b, RV90c, SJ92, SM93b, Sto91a, Sto91b, Ueb93, Wal93, WSLP90, dCDdSO94, DH93, HW94, Hab95a]. Statistics [AAB97, Alb94, Ano91y, Ano97j, Bre96a, CM91a, Gai96, GK93, Gut94a, Hun94b, Ima95, Ket97a, Ket97b, KL99b, Kra90, Lan96, Lav91c, LRR91, Moo98a, Moo98b, Nor90, Par96, Pea91, Pol91, Roc91, Ruk93, RV98, SC97a, SC97b, SAK90, SK90, TA90, Wan95, Wel90b, WBB93, Zel92, Zel93]. Status [BDM^{+}93a, BDM^{+}93b, GS93b, Lit93b, RT96b, SK93b, Wac93, vR98]. Stein [GS91b, Li92, LW90a, LW90b]. Step [DT92, Fin93, HDG97, SRC92, Tab90]. Step-Down [Fin93]. Step-Up [DT92]. Stepwise [HH94, HB99, Tro95a]. Stochastic [AG96, BH97, DF98, FT94, KR93, LT94b, LYY99, Muk96, NGC^{+}95, PS94a, POH97, RS93, She93, TCS94, Wan96, ZLRS98]. Stochastically [EGGS97]. Stock [CG97]. Stopper [RT92]. Stopping [SC96a, Ye93]. Straf [Gla91]. Strata [HS93b, HS93c, Sch93a, Ylv93]. Strategies [EGG95, HBH93, LNL95, Tam95]. Strategy [BSB98]. Stratification [Jon96, Lit93a]. Stratified [BPS95, KM91a, LWY93, RS96a, SCC98, Val90, WS96]. Stratum [KMT99]. Strawderman [Ano99h]. Streaks [Alb93a, AASM93, Alb93c, SM93b]. Strength [Pri96]. Stretches [LMR96]. Stroke [KR96]. Strong [GR96b]. Structural [AGP91]. Structure [GK95, MSMP92, Sam93a, SG92, SV93, SW90, WJ92, WPK99a, WPK99b, WHG98, YB97]. Structured [Seg92, SK91]. Structures [AS90b, CM99, FK99, GHL96, LNL99, Roc92, Vil90]. Student [MH94b]. Student-Athletes [MH94b]. Studentized [Hay90]. Studied [HZS^{+}90]. Studies [AGM98a, Ano99a, Bak98b, BP97, Bin98a, BO95, BWC98a, Bre98a, CS91, CGL93, CWW95, CNWW98a, DN98a, DW99, FS95b, GKL98a, GKL98b, GJDT98a, HB99, IRC98a, JK96, Jud98a, Kad98a, KMT99, LZ98a, LB94b, Lit95, OO97, Pep91, QH98a, QH98b, RCL96, RK90, Ros93a, Ros95a, SH92, San98a, STR97, WHGA96, ZWWB96]. Study [Akr91a, BHB96, Bre96a, Coo91a, DM97, DD90, DC94, DMLC97, EGMP93, GWG90, Hei91, Hod94, RLN99, Sas97, SJ92, SHBF94, TFAH94, Zas93]. Studying [AHB97, Joh96b]. Stute [Wil99]. Subdistribution [FG99]. Subgroups [CJ99, For94]. Subject [Con92, GW99, Gui93, PHL98, RG93, RS93]. Subjective [CKLH92, DM97]. Subpopulations [Mit91]. Subsampling [BPR99, BK99a, BK99b, Chr99a, Chr99b, Han99a, Han99b, LKCH99a, LKCH99b, LKCH99c, SC99a, SC99b, Ste99a, Ste99b]. Subseries [She97]. Subset [Pan97, SH92]. Subspace [FBF97]. Substitutions [KL90]. Subsurvival [TFP90]. Subtidal [PM97]. Success [HPS98, Ket97a, Ket97b]. Sufficient [MZ94]. Sulfate [Haa95b, Oeh93b]. Summarize [Llo98]. Summarizing [BGH97]. Sums [IT94]. Super [PWM92]. Super- [PWM92]. Superclusters [Roe90]. Superior [BS97a]. Superleverage [SC92b]. Supersmooth [Efr97]. Supplementing [Roc95]. Support [FHTW97, Gel94a, Gel94b, Mer94a, Mer94b, Mer94c]. Suppression [Ano96a, Cox95]. Surface [AM92, LR97b, Sho93]. Surfaces [LT94a]. Surprising [Sie93]. Surrogate [CS90a]. Surveillance [APR92, TMP93, XXP93]. Survey [Bel93a, BF92, Bin96a, Bin96b, BZ99, CCD97, DS92, DSS93, Elt96a, Elt96b, Fay96a, Fay96c, Fay96b, Fie90, GK93, Hah90, HS93b, HS93c, Hog93, JMS96, Jud96a, Jud96b, KR90b, MSL93, PWM92, RG95, Rao96c, Rao96a, Rao96b, RBD93, RSV95, Rub96c, Rub96b, Sch93a, Sch90a, SS96b, SCC98, SS99, SG96b, Sit92, Sta91, SJ90a, SJ90b, Thé99, Tou90, Ylv93, Zie90, MSML97, SR95, Sta91]. Surveys [Ano91h, ALM97, BP94b, Bin98b, Bin98a, Bre98b, Bre98a, CD92, CM91b, GKL98c, GKL98d, GKL98a, GKL98b, GJDT98a, GJDT98b, Hil93, Jud98b, Jud98a, Kad98b, Kad98a, Kot94, PMW93, RN97b, Riz92, San98b, San98a, Ski91, SR96]. Survival [ALHH92, AL95, AG96, BRR92, Bak98b, Bak98a, BD93, BH97, CP99, Cha91a, CIS99, CWY97, DG92, EP94, FSE98, GWG90, Gra92, Gra95, GR92a, HMY97, HMC^{+}93, HW95b, LI98a, LC93, LYY99, LGWC93, MZ94, MSK96, Muk96, Mur95, Pru93, RC94a, RC94b, RS90a, RS93, Sas97, SD97, TDW95, TMP93, WH90, Wan91, WC99, XXP93, Yan99, YJW95, Zuc98, dCP99b]. Surviving [CIS99, Yam92]. Survivor [PSVS92, SW97a]. Suspects [BGH97]. Sweep [ES90a]. Switching [SS91a, SS92]. Symmetric [Yam90]. Symmetry [CNS90, Ede90, FG95, McW90]. Synthesis [Goo96, KO96]. Syringe [NGH^{+}95]. System [AMWK93, BdL99, CT98, MEGN92, RV90c, Zas93]. Systematic [SO94]. Systems [Ber91b, BD91b, FT94, GT98, Gri90, Lel94, LC98a, Llo98, LC98b, Mei94].

T [Sta94a, ME96, LCG^{+}92b, LCG92a, Lou92a, TS92]. T-Cell [LCG^{+}92b, LCG92a, Lou92a, TS92]. Table [Alb97]. Tables [AMP90, Agr98a, Agr98b, Ano99h, BS97a, BS97c, BB98b, BB98c, CS92b, DD90, DDD94, Fri94, Goo94b, KP97, KA95, LW91, MPS98a, MPS98b, Qui98a, RG93, SW98b, SW98c, SW98d, VB97, VHE91, YF90, dCDdSO94, YF95]. Tactical [Hod94]. Tail [BVT96, Ede90, GR96a, GW91, Roj96]. Tails [FS98b]. Tanner [KT94d]. Tanzania [AGM98a, AGM98b]. Target [LW91]. Tax [DN98a, DN98b]. Taylor [Eld96]. Team [CH94]. Technique [Bri92, Joh92, Roe94, Zel93]. Techniques [ES90b, Jay96, KW96b]. Technological [KL90]. Telegraphic [Ano95p, Ano95q, Ano96r, Ano96s, Ano96t]. Telemetry [ASL91]. Temperatures [LR97b]. Template [PS94b]. Templates [AGP91]. Temporal [Cre94, CH99, Gut94b, Haa95b, Han94, KDG94, KO96, WCXG97, HW94]. Temporally [DLW95]. Ten [RCH94]. Teratogenesis [LR97a]. Terms [Cho92, Gu92]. Tessellation [AF97]. Test [Ano97j, Ano99a, Ano99i, BPL97, DT92, FG95, Fan96, FL98, Fin93, Fol96, GJ99a, GR97a, GHMW94, GR97b, GLG92, GK93, HH90, HW92b, Hay90, HP92, HN92, Jac93, JSKK98, KW95, KC93, Led94, LH91, McW90, MH99, PR90a, PB92, PF97, QH98a, QH98b, RSY93, Ruk93, SC97a, SC97b, Sch91c, Shi91, SAK90, SS95a, Sil96, Sil97b, SK90, SW91b, TL97, VM96, Wan96, WM98, YF90, dCP99a]. Testing [AME96, ACH99, AP99, Alb97, AC95b, AP96, Ano91i, Ano98a, BG91a, BM99, BPR95, BDR92, BS98, Che95a, CG97, CW94b, CB97, CF95, Dea92, ES90b, Eub97, GJ94, GJ99b, GHL96, GS94, GL92, GK93, HMM98, HW92a, Hay90, HRW90, Hon99, Hua97, HOS94b, Jay96, JB92, JK93, Kra99, KW97, LTM92, Leh93, LGWC93, MZ94, MH94a, MM93, MBR98, Mun96, PHS94, Raz90a, Raz90b, SSC94, SL93a, SC97a, SC97b, SMD93, Tom96, Tro95a, Tsa98, Wan99, Wee91, Wel90b, Wes97b, WHG98, YR95, ZB92, Wes98]. Tests [AC97b, AKM94, Ano91j, Ano96g, Ano96i, Ano97j, Ano98i, Ano99a, BOP92, BS97a, BY95, BS96b, BF93a, BWC98a, BWC98b, Bre90, BHNO92, BH96, CS93, CM97b, Chr91, CS92b, CS92a, CD96, Dab90, DM90, EMM95, ES99, FS98a, Fin90, GK96, Gra95, HHS97, Hau92, HSP90, HR94, HMO97a, HMO97b, Hol99a, IRC98a, IRC98b, JGC95, JR96, Kai94, KL97, KL99a, KM99a, Kun90, LaR91a, LK91, LD92, LLR95, LS93, Lou96a, MM95, MH90, MS92, MSZ93, MS98b, MNS99, MS91a, MS95b, MP99, NP95, NC99, Ode91, OK96, OR92, Peñ98, QH98a, QH98b, Rob94, Rom90, RMS90, Ruk93, Ryu94, Sch94d, Sil92, ST92, Sut93, TR97, Tan94, Tho91b, Wan97, Wel90a, Xio95, ZM93, MS98a, MS95a]. Texas [CCG^{+}97a, CCG^{+}97c, CCG^{+}97b, Cre97a, Cre97b, GMS97a, GMS97b, SF97a, SF97b]. Theil [AML95]. Theil-Sen [AML95]. Their [BS97b, FG94, GC95, HMN94, HHM92, Las94a, Mac94, Ye98b, Las94b]. Theoretic [LB94a]. Theoretical [YB97]. Theories [Leh93]. Theory [BS98, CP99, DDV99, dE97, FK94, Leh93, LYY99, O’C99, dCP99b, DE96]. Therapy [RG94]. There [LTM92, WSJ92]. Thompson [SO94]. Those [Alb94]. Three [DFGJ93, KA95, Kun90, Spu99, Tom94]. Three-Level [Tom94]. Three-Sample [DFGJ93]. Three-Way [KA95]. Threshold [KP91, PB92, SGC95, Tsa98]. Thresholding [Fan96]. Tight [Ver93]. Time [AME96, Ada98, ALHH92, Boc99, BD93, CM97a, CMW97, Cha90, CL95, CW91a, CW99, CL93, CLM93, Cre90, DR91, EH91, FT94, FW96, FSE98, GNK96, GS90b, Gri90, GWB99, HK93, HDG97, Hon99, KST98, KL90, Koo97, KC96, LMR96, LWY93, LS91a, Li96, LT98, LW92, LL97c, MT93, MP96, MSL93, Neu99, Niu96, PC93, PS96, RC94a, RC94b, RGH99b, RGH99a, Roc92, RF99, SPM99, SJ92, SAK90, Sin93, SS96c, SL98a, SW98a, Sto99, The93a, TA94a, TA94b, Was99a, WPK99a, WPK99b, Yam92, ZL99, dJP98, vR98, HR97]. Time-Dependent [AME96, Ada98, ALHH92, FT94, FSE98, HDG97, vR98]. Time-Varying [FW96, Gri90, HK93, Neu99, RGH99b, RGH99a, RF99, Was99a, ZL99]. Times [DB97, HW90b, MPCB95, PC93, SDW98, WH90]. Timing [ALJ99]. Tobit [CCC96]. Tolerance [Ano91f, Mee90]. tolerate [Hjo94]. Tolley [Hab95a]. Tomographic [Kol96]. Tomography [MO98, Var96, Wei97, WEST91]. Topological [ES90a]. Total [EKWS91, FG91, Mil92, MS91b, MS91c, WS91]. Totals [Qui98a, Zie90]. Toxicity [BO95, HZS^{+}90, HB99]. Toxicological [HLC92]. Toxicology [GMR99, HZS^{+}90]. Traces [WPK99a, WPK99b]. Tracking [ASL96]. Tracts [SS95b]. Trade [BMR99]. Traffic [McC98b, McC98c, TW98a, TW98b, TW98c, Var96, Var98a, Var98b]. Train [Bri92]. Training [DW99, Hod94]. Traits [CJ99, ZT97]. Transfer [Pap96]. Transform [Akr90, Akr91a, KT94a, WR95]. Transform-Both-Sides [WR95, KT94a]. Transformation [CWY97, KW96b, NCDF96, ST99a, ST99b, WR95, YM99]. Transformation-kernel [YM99]. Transformations [Atk91, CJ95b, Dua91, KM99a, LCWH97, MZ91, Rud91, Sco91, Vel95, WMR91b, WMR91a]. Transformed [DKS97, OKS97a]. Transition [LMR96, MPCB95, Sch99b, Ter94]. Transmission [DA96c, Wes97a]. Transplant [DwSH94]. Trapping [WW91b]. Treat [Roc95]. Treatment [AI95, Ano92e, BC90, BP97, BS97a, Car94b, GK96, GW99, Hol95, KA91, LTM92, MH99, OK96, RG94, Ros93a, ST99a, ST99b, TFAH94, Ueb93, WSLP90]. Treatments [CS92a, DPWH99, GS97, Jac93, TL97]. Treaty [PB92]. Tree [Ros93b, Seg92]. Tree-Structured [Seg92]. Trees [LC93, Zha98b]. Trend [CS92b, GK96, IRC98a, IRC98b, Jac93, OK96, She93, SS96c]. Trend-Resistant [Jac93]. Trends [HL94, JMS97, Oeh93b]. Trial [Bak98b, Bak98a, CCC96, GM96, RG94, TFAH94]. Trials [Bro92d, BPS95, DS97a, DS97b, EF91a, EF91b, EFS92a, EFS92b, GB97a, GB97b, Gre92, Gui93, Har92b, LD92, LB94a, LL97a, LL97b, Lou92b, Mei91, OBW92, Roc95, Rub91, SBD97a, SBD97b, SBD97c, ZL91]. Triangular [Goo94b]. Trimmed [GEG99, SP91]. Triogram [HKS98]. Trispectrum [WW93]. Troubles [Fie90, Hah90, KR90b, Sch90a, SJ90a, SJ90b, Tou90]. True [EH91, GC95]. Truncated [CTC96, EP99, GSL92, GMP97, GL96, Gür96, HL96, Li95, LQT97, dCP99a]. Truncation [Fan96, NP95]. Trussell [XP92]. Trying [CH94]. TSMARS [LR97b]. Tukey [SMF94]. Tumor [CH92, Hei95]. Twins [HHH92]. Two [Akr90, Akr91b, Alb97, AK92, AI95, BP99, BHS98, BW92, BDR92, BD92, BPS95, CM91a, CF95, DL92, DDD94, DKR95, GR97a, GR97b, GH90, HPS98, HDG97, HOS94b, JMV94, Kai94, KSW92, Leh93, LTW96, Llo90, LWT94, MH99, NW97, O’S90, PB98, Pen97, Pes93, RSY93, RN97b, RK90, Sit97, SK90, Tom94, TS96, WH90, dCDdSO94, NM98, Rig97]. Two- [CF95, Tom94]. Two-Dimensional [DDD94, O’S90, dCDdSO94]. Two-Factor [Akr90, BD92, DL92]. Two-Phase [Sit97, TS96]. Two-Sample [Akr91b, BP99, BW92, BPS95, CM91a, DKR95, Kai94, KSW92, LTW96, LWT94, RK90, WH90]. Two-Sided [MH99]. Two-Stage [AI95, HOS94b, RSY93]. Two-Step [HDG97]. Two-Way [Alb97, GH90, PB98, Pen97]. Type [BDM97, GS91b, HW91, Jen91, Jen95, JDP94, Sil92, BM92]. Types [NW97]. Typing [Mel99, NA99].

U.S. [Alh92, DLML99, GMWZ99, Joh90, LCMA92, LC92, McN92, MZTT98]. Umbrella [PW96, Pan97]. Unbalanced [AAB97, AD90, MS92]. Unbiased [GS91a, GS91b, SW98a, Wan97]. Unbiasedness [SSC94]. Uncertainty [CH93a, MR94, RLN99]. Unconditional [For94]. Uncontrolled [LB94b]. Uncorrelated [Kus98]. Uncovering [EGG95]. Underclass [Pea91]. Undercount [BDM^{+}93a, BDM^{+}93b, CT98, Lit93b, MS93, Sch93b, Sch91b, Wac93]. Underground [KM96a]. Understanding [Gle93, Kra90]. Underwater [BSPK98]. Unemployment [DLML99, MZTT98]. Unequal [Sär96]. Unequally [JV93]. Unified [DF98, LT98, Lju93a, Lju93b, Tho91b]. Uniform [CF95, KO95, Muk96, RS93]. Uniformity [Ano98h, HOLDG98, JK93]. Uniformly [Ano96i, Tan94]. Uniqueness [LK92a]. Unit [GHL96, NP95, SL93a, TR97, YR95]. United [CH93a, LT94b, Pas93, RAGW93a, RAGW93b]. Univariate [dE97, WK94]. Unknown [BPR99, CL92a, CM98, DJ95, MPCB95, Pan97]. Unmasking [CH90, Fun93, KM90, Rv90a, Rv90b, RvZ90, RS90b]. Unobserved [KR96, KT94a]. Unrelated [BG91b]. Unresolved [BDM^{+}93a, BDM^{+}93b, Lit93b, Wac93]. Unseen [BBC98]. Unskilled [Sam90b]. Unsupervised [FR99]. untitled [AA94a, AA94b, AA94c, AA94d, AA94e, AA94f, AA94g, AA94h, AA94i, AA94j, AA94k, AA95a, AA95b, AA95c, AA95d, AA95e, AA96, AB97a, AJM99, Ack91, Ack93, Ack98, Ahn99, Ahs90, Alb93b, Alb95, Ale98, Alt94, Ame99, Ame92, And95, Ang93, Ang90, Anr96, App97, Asm93, AS93, Ath99, Atr94, BB98a, BBB98, Bag90, Bag91, Bal94, Bal96, Bal97, Ban90, Ban93, Bar93a, Bar97, Bar92, Bar98, Bec91, Bec98, Bec99, Bel93b, Ben97, Ben91a, Ber94a, Ber94b, Ber93, Ber95, Ber92a, Ber92b, Bha94, Bha90, Bie96, Big98, Bil91, Bil96, Bla92, Blo98, Boi90, Bol91, Bol94, Boo94, Boo99, Bov98, Bra98, Bre96b, Bri98, Bri99, BGH94, Bro99, Bro90, Bro91, Bro92c, Bry97, Bud96, Bun96, Bur92, BS91b, Buz97, Car90a, Car94a, Car90b, Cha96a, Cha92, Cha91b, Cha93b, Cha91c]. untitled [Che91, Che92b, Che92a, Che95b, Che98, Chi91, Chi90, CS95b, Cla90b, Cla90a, Cla92, Cla98, Cla93, Cla99, Cle99, Cli99, Clo94, Col91, Col93, Col97, Con97, Coo91b, Coo93, Coo95, Cor99, Cor92, Cow91, Cow98, Cox99, Cra95, CL94, Cry99, Cwi94, D’A93, D’A94a, D’A94b, D’A94c, D’A94d, D’A95a, D’A95b, D’A95c, D’A96a, D’A96b, DSS97a, DSS97b, DSS97c, DSS98, DSS99, Dab92, Das95, Dav92, Dav99, Dav91, Dav98, DeG95, Dev95, DiC95, Dia99, Dic96, Die93, Die97, Dog95, Dra94a, Dra94b, EJD97a, EJD97b, EJD98a, EJD98b, Eas92, Edd90, Ela91, Eli98, Emb95, Emb96a, Emb96b, Emb97a, Emb97b, Emb98a, Emb98b, Emb98c, Emb98d, Eub93, FH95, Far97, Fau90, Fau94, Fer98a, Fie93b, Fis91, Fla90, Fla91, Fle90, Fle91, Fou92, Fou93, Fra99]. untitled [FM93, Fre95, Fre98, Fre99b, Fri98, Fun90, Fun92, Gas91, Gas95, Gas96, Gav91, Gen93, Gen99, Gho90, Gho94, Gho96, Gho98, Ghy98, GH91, Gil94, Gin95, Gly98, Gol95, Gol93, Goo94a, Gra91, Gre91, Gro92, Gru93, Gru98, Gue96, Guo93, Gut90a, Gut90b, Gut91a, Gut91b, Gut92a, Gut93a, Gut99, Gut90c, Gut92b, Gut92c, Gut93b, Haa95a, Haa96, Hag97, HRF96, Hal97, Ham90a, Ham93, Ham90b, Har96, Hau97a, Hau97b, Hau90a, Hau90b, Hav95, Haw93b, Haw96, Haz95, Hea90, Hei90, Her98, Her93, Her97, Her91, Hig98b, Hil92, Hil91, Hil95, Hip95, Hod91, Hol98, Hol99b, Hol99c, Hos97, Hsi90, Igl91, Igl92, Iye97a, Iye97b, JEB98, JFM98, JGB99, Jam98, Jan95, Jas92a, Jas92b, Joh95, Jon90, Jor96, KB97a, KB97b, KK94a, KK94b, KK94c, KK95a, KK95b, KK95c]. untitled [KK95d, KK95e, KK96a, KK96b, KK96c, KK97a, KK97b, KK97c, KK97d, KK97e, KK97f, KK98a, KK98b, KK98c, KK98d, KK98e, KK98f, KK98g, KK98h, KNB98, Kab98, Kad90a, Kad91, Kaf94, Kaf95, Kaf97a, Kaf97b, Kaf97c, Kaf97d, Kaf97e, Kaf97f, Kaf97g, Kaf97h, Kaf99a, Kaf99b, Kan90, Kar90, Kas97a, Kas90, Kat97a, Kat97b, Kaw99, Kep97, Kfa97, Khu94, Kia90, Kir98, Klu94, Koe97, Koo91, Kot93, Koy91, Kri94, Kro90, Kul95a, Kup90, LD96a, LD96b, LD97a, LD97b, LD98a, LD98b, LDH99, LH96, LH97a, LH97b, LH97c, LH97d, LH97e, LaV90, Lac90a, Lac90b, Lac90c, Lac90d, Lac91, Lac92, Lac93a, Lac93b, Lac93c, Lam97, Lan93, Lan95a, Lan95b, Lan97, Lar91b, Lau92a, Lee93, Lee96b, Leh97, Leh98, Lel99, Len98, Len99a, Len99b, Len99c, Len99d, Leo92a, Leo92b, Lev94, Li98d, Li99, Lig98]. untitled [Lig94, Lin98, Lin94, Lin99a, Lin99b, Lip95, Lit94, Lit96, Liu97, Lju95, Loc90, Loh96, Lon91, Lor92, Lou99, Lou96b, LuV91, Luc98, Lun97a, Lun97b, Lun98b, Lun98a, Lun98c, Lun98d, Lun99, Lye97, Lyn98, MTW96a, MTW96b, MTW96c, MTW96d, MTW97a, MTW97b, MTW97c, MTW98a, MTW98b, MTW98c, MTW98d, MTW98e, MTW98f, MTW99a, MTW99b, MTW99c, MTW99d, MTW99e, MTW99f, MTW99g, MTW99h, MTW99i, MTW99j, MTW99k, MTW99l, MTW99m, MTW99n, MW95a, MW95b, MW95c, MW96a, MW96b, MW96c, MW96d, MW97a, MW97b, MW97c, MW97d, MW97e, MW98d, MW98e, Mad93, Mal95, Mar94, Mar90a, Mar92a, Mar98a, Mar95a, MA95, Mar90b, Mar95b, Mar98b, Mar98c, Mar92c, Mar98d, Mas92, Mas98, Mat99, Mat94, Mat93, McC98a, McD90, McD92, McD97, McG90, McK92, McK95, McN97, McN98a, McN98b, Mee95, MS90, Mel90, Meu95, Mey99, Mic97, Mic98b, Mic98a]. untitled [Mik93, Mik94, Mik95a, Mik95b, Mik97, Mik98, Mit94, Mod95, Moo91, Moo99, Mor90, Mor91b, Mor93a, Mor93b, Mou95, Muk90, Muk92, Mul91, Mul93, Mul98, Mur97, Mur90, Mur91, NL94a, NL94b, NL94c, NL95, Nac91, Nak92, Nar93, Nei97, Nel92, Nel97, Nel98a, Nel98b, Neu92, New92, New90, Nol92, Not90, Not94, O’R91, Oeh93a, Ogu99, Olk91, Olk92a, Olk92b, Olk93, Olk97, Olk99, Ols90, Omo95, Ove91, Ove94, Owe92, PAL96a, PAL96b, PD99, PE94a, PE94b, PE94c, PE94d, PE95a, PE95b, PE95c, PE95d, PE95e, PE95f, PE95g, PE95h, PE96a, PE96b, PE97a, PE97b, PE98a, PE98b, PE98c, PE98d, PE99, Pad93, Par91, Pau96, Ped94, Per96, dBPI97, Per90, Pfe99, Pic97, Pie93, Pin97, Pit99a, Pit99b, Ple91, Pod94, Pol90, Pot96, Pre94, Pre96a, Pre92, Puh98, Qua90, RAJ96]. untitled [RBD96, RBL97a, RBL97b, RBL99a, RBL99b, RBL99c, RC99, RJ95, RJ96, RL96, RL98a, RL98b, RL98c, RLS99a, RLS99b, RS95a, RS95b, RS95c, RS95d, RS95e, RS95f, RS96b, RVL97a, RVL97b, RVL97c, RVL97d, RVL97e, RVL97f, RVL99a, RVL99b, RVL99c, Rad95, Ram93, Rat98, Rat94, Rau91, Rav97, Ree98, RN97a, Rei90, Ric90, Ric97b, Ric98, Rob90, Rob91, Roc93, Rod90, Roe92, Ros94, Ros99, Ros97a, Ros97b, Ruk98, Run99, Rut91, SRJ99a, SRJ99b, SRJ99c, SRJ99d, SRJ99e, SRJ99f, SRJ99g, SW94a, SW94b, SW95b, SW95c, SW95d, SW96, SW97b, SW97c, SW98e, SW98f, SW99, Sal93, Sam90a, Sam97, San91, San90, Sap91, Sar91a, Sar94, Sar91b, Sat90, Sca92, Sch96a, Sch91a, Sch94a, Sch90b, Sch96b, Sch95a, Sch94b, Sch90c, Sch94c, Sch98a, Sch98b, Sch99c, Sch97, Sch91d, Sch95e, Sco99, Sco90, Sea98, See95, See94]. untitled [SM93a, SM99, Sel95, Sel96, Sen92, Sen93, Sen94, Ser93, Ser91, Set95, Sey97, Sey99, Sha92a, Sha99, Sha98, Sha92b, Sha93a, She95, Shl90, Sho92, Shu90, Shu93, Sim94, Sim93b, Sin97, Ski96, Sla93, Smi94, Smi96b, Smi96a, Smi98, Sob96, Sob98, Sol94, Spe93, Spe98, Sri93, Sta98a, Sta99, Sta98b, Sta93, Ste90a, Ste98, Ste90b, Ste92, Sti95, Sti92, Sto94, Sto97, Sto90, Str98a, Str99a, Str99b, Str99c, Str99d, Str95, SZ99, Str93b, Str94, Str98b, Stu92, Sud90, Sun99, Swa99, Swi90, Sys96, Tah92, Tak92, Tan90, Tas98, Tay90, Tei90, The93b, Thi94, Tho90b, Tiw99, Tro95b, Tsa90, Tsu94, Tur98, Tur90, Utt93, Utt94, Vau90, Vel98b, Vel99, Vin93, Vog92, Vog99, WP92, Wak91, Wal91, Wal95, Wal96, Wal98, War91, War97]. untitled [Was97, Was98a, Was98b, Was90, Was91, Was93, Was94, Was99b, Wat90, Weg90a, Weh94, Wei90a, Wei91a, Wei98, Wei91b, Wei90b, Wei94, Wei95a, Wel99, Wel94, Wil91, Win96, Wol90, Wol93, Wol91, Won93, Won96, Woo90, Woo91, Woo93, Wu98, Wue96, Yan90, Yan98, You97, Zaj93, Zeh95, Zer96, Zer99, Zim92, dL93]. Updating [NT92]. Upper [BHL93, Sat95]. ups [GLR93]. Uptake [MN93]. Urban [Pea91]. Use [Ano91h, Are92, BP94b, Bro92b, CM91a, DK99, GFT^{+}93, GK93, Hod94, IT94, LW90a, LW90b, Man90, Mar92b, MZ97, MSH93, NFL92a, NFL92b, OMF^{+}93, Wei92, YK96, CRS^{+}91]. Used [Cha91a, RS90a]. Useful [Zuc98]. Using [AF97, ASL96, AL95, AIR96a, AIR96b, AIR96c, Ano92e, Ano98a, Ano98h, BSPK98, BC90, Bal95, BR92, Bel93a, BBH96, BdL99, BRW99, BK99a, BK99b, CDW93, CD92, Chr99a, Chr99b, CRS^{+}91, CB90, CS96b, Coo91a, CHW92, Dig99, DM97, EW95, FZ99, FG91, FS98a, For94, Fre99a, GHRPS90, GSL92, GBJ96, GM99, GBN96, GB98, Gra92, GR92a, Hab95a, Han99a, Han99b, HK93, Hec96a, Hec96b, HMC^{+}93, HL97, HOLDG98, IL91, IRC98a, IRC98b, Jay96, Joh92, Joh96b, KO95, Kul95b, KS93, LKCH99a, LKCH99b, LKCH99c, LP99, LCG^{+}92b, LCG92a, LK91, LMR96, LS95, LR97a, LS91a, LR97b, LRR91, LR99, Llo98, Lou92a, MS99a, MR94, MZ96, MKSB90, McC98b, McC98c, MR91, Mof96a, Mof96b, ML94, MPCB95, Mul99, PO94, PM97, Pet92, PWM92, Pri96, Riz92]. Using [RG96a, RG96b, RW97, Ros96a, Ros96b, Sam93a, SK93a, SRR99a, SRR99b, SC99a, SC99b, SKW99a, SKW99b, Ste99a, Ste99b, Sto99, TS92, TW98a, TW98b, TW98c, Var98a, Var98b, VW95, WPD98, Weg90b, Wes97b, WKL95, WR94, XLR91, Yan99, vdL99, Wes98]. Usual [NCDF96]. Utilitarian [Lju93a, Lju93b]. Utilities [SSK90]. Utility [CP91]. Utilization [DG99, GWG98].

Validated [Tho91a]. Validating [WK94, WC92]. Validation [ET97, HY98, KAT91, LS95, RFB97, SBS94, Sha93b]. Validatory [Par93a]. Validity [GBN96, Har90a, SMD93]. Valuation [CH91a]. Value [Amm93, BF90, CG92, HW92b]. Values [Ano91g, Ano91i, BB94, Buo90, CH91a, FWW99, GC95, JB92, LS97, SMP95a, SMP95b, Wee91, ZB92]. Variability [Gui93, MO98, Per97, Wan97]. Variable [AI95, BRMZR97, BZ92, BJB95, BZ99, BRW99, BS96c, CR92, EF91a, EF91b, GM93b, GM99, Hig98a, LA93, LR97a, MS99a, Mei91, Mul99, Rub91, SKW99a, SKW99b, SB95, SO94, Wei95b, ZL91, ZL95]. Variables [AK92, AIR96a, AIR96b, AIR96c, BJB95, BFH99, CH97b, EGGS97, GH95, Hec96a, Hec96b, Hou95, Jon96, KR76, KR90a, LS95, Lie94, MSML97, MW98a, MW98b, MW98c, Mof96a, Mof96b, Now92, PSVS92, RN97b, RG96a, RG96b, RW98a, RW98b, Ros96a, Ros96b, Sie93, WPD98, DH93]. Variance [Ano96g, Bin96a, Bin96b, CD91, CH91b, Cha90, CG97, CM98, CM97b, CM91a, CV92, DLB91, DBB^{+}94, Dor94, Elt96a, Elt96b, Fab91a, Fay96c, Fay96b, GH90, GP92, HS94, IT94, Jud96a, Jud96b, KPR91, MM95, MT93, McC94, MR91, MP96, Nai91, PWK99, RS96a, Rao96c, Rao96a, Rao96b, Rub96c, Rub96b, Sär96, SG96a, Sha91, SS99, Sit97, Sla94, SO94, SW98a, SHGL96, Val93, Wee91, XLR91, Fab91b]. Variance-Covariance [MR91, XLR91]. Variances [Ano91h, FNP95, Lau92b]. Variate [DJ98]. Variates [GF96a, SMP95a, SMP95b]. Variation [Gra95, NM98, Pag90]. Variations [DG99]. Various [GMWZ99, SKF98]. Vary [SRR95]. Varying [FW96, Gri90, HK93, Neu99, RGH99b, RGH99a, RF99, SW98a, Was99a, WCH98, ZL99]. Varying-Coefficient [WCH98]. Vector [Ahn97, Ano99i, Mau95, Qui98b, SG97, WM98, YR95]. Vectors [BDR92, GR97a, GR97b, SMF94]. Verification [LT97, PB92]. Versatility [KL99b]. Version [Led94, RS94b]. Versions [JDP94]. Versus [HMM98, RG94, VM96]. Vertical [DA96c]. Very [Atk94]. Via [BPS96, CDP96, ES90b, GM93b, HR90, KT94b, KT94d, NNR95, CL92b, CCD97, DR98, DJ95, Gou97, Hon99, KT94c, LR97c, LC95b, Moj99, MP95, PS94b, PS99, Wan98a]. Viability [ZWWB96]. Vietnam [Ang91]. Views [Kra90]. Viral [ZDL96]. Virus [Ano92f]. Visual [MT95a]. Visualization [CL92a, Li92]. Vitro [HZS^{+}90]. Voids [Roe90]. Volcanicity [GT91]. Volcanism [Sol91]. Volume [Ano90r, Ano91z, Ano92v, Ano93r, Ano94p, Ano95r, Ano96u, Ano97p, Ano97u, Ano98r, Ano99q]. Voluntary [Ang91]. Voronoi [AF97]. Vote [GKB98]. Votes [Est92]. Voyage [Bil97].

W [Gla91, Kas97b]. W. [Man94]. Wachter [Gla91]. Wages [RBD93]. Waiting [HW90b]. Wald [LH91, Sil92]. Wald-Type [Sil92]. Walsh [Bea91, Bri91b, Mor91a, Sto91a, Sto91b]. Warning [APR92]. Water [KM96a]. Waters [KSJ94]. Watson [SAK90]. Wave [MLMT94]. Wave-Soldering [MLMT94]. Wavelet [AGM94, CKM97, DJ95, Efr99, Fan96, Kol96, Vid98]. Wavelets [BSPK98, PM97, Wan98a]. Way [Alb97, Fri94, GH90, KA95, PB98, Pen97]. Weak [BJB95, Cha91a, DG92, RT96a]. Weibull [BS93, MSK96]. Weighted [Che94a, CC96a, Hoo93, HM91, KL99b, LIZ99, MBL98, RV98, RP96, Sas93, WWZtO97, Yan97, Yan99]. Weighting [PMW93, Zie90]. Weights [Bly93b, Per97, PWM92, Wie98]. Welfare [Ano90q, Cox90, Dor90, FIB90, HT90a, HT90b, Kra90, Net90, PT90, Rol90]. Wells [Ano99h]. Wet [Haa95b, Oeh93b]. Whale [RZ98]. Whales [Buc95a, Buc95b, RGZ95a, RGZ95b, RGZ95c, Sch95c, Sch95d, TL95a, TL95b, Wol95a, Wol95b]. Whaling [Giv99]. Which [Est92, SRR95]. White [Fab94]. Whitney [BW94, Ano91f, Mee90]. Whittle [PO94]. Wilcoxon [BW94, Tab90]. Wildlife [ASL91, Pol91]. Windfield [CB94]. Window [Haa90, WHGA96, MR94]. Winner [Est92]. Winners [BH93b]. Within [Gui93, HS93b, HS93c, Sch93a, WJ92, Ylv93]. Within-Subject [Gui93]. Without [Ano92e, BC90, EV92, Rom90, HB99]. Woodbury [Hab95a]. Worked [RBD93]. Workers [Sam90b]. Working [CKLH92]. World [Nor90].

Years [Rub96a]. York [DXP92]. Yule [Goo96, KMT99].

Zero [BS96a]. Zonoid [KM96b].

References

AA:1994:BRBa

[AA94a] AA. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1150, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290968.

AA:1994:BRBb

[AA94b] AA. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1150, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290969.

AA:1994:BRBc

[AA94c] AA. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1150, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290970.

AA:1994:BRBd

[AA94d] AA. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1150, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290971.

AA:1994:BRBe

[AA94e] AA. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1150, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290972.

AA:1994:BRBf

[AA94f] AA. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1150, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290973.

AA:1994:BRBg

[AA94g] AA. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1150, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290974.

AA:1994:BRBh

[AA94h] AA. Book review: [untitled]. Journal of the American Statistical Association, 89(428):1569, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291032.

AA:1994:BRBi

[AA94i] AA. Book review: [untitled]. Journal of the American Statistical Association, 89(428):1569, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291033.

AA:1994:BRBj

[AA94j] AA. Book review: [untitled]. Journal of the American Statistical Association, 89(428):1569, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291034.

AA:1994:BRBk

[AA94k] AA. Book review: [untitled]. Journal of the American Statistical Association, 89(428):1570, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291038.

Akritas:1994:FNH

[AA94l] Michael G. Akritas and Steven F. Arnold. Fully nonparametric hypotheses for factorial designs I: Multivariate repeated measures designs. Journal of the American Statistical Association, 89(425): 336–343, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291230.

AA:1995:BRBa

[AA95a] AA. Book review: [untitled]. Journal of the American Statistical Association, 90(429):398, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291187.

AA:1995:BRBb

[AA95b] AA. Book review: [untitled]. Journal of the American Statistical Association, 90(429):398, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291188.

AA:1995:BRBc

[AA95c] AA. Book review: [untitled]. Journal of the American Statistical Association, 90(429):398, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291189.

AA:1995:BRBd

[AA95d] AA. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1135–1136, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291368.

AA:1995:BRBe

[AA95e] AA. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1136, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291370.

AA:1996:BRB

[AA96] AA. Book review: [untitled]. Journal of the American Statistical Association, 91(433):441, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291439.

Akritas:1997:NHR

[AAB97] Michael G. Akritas, Steven F. Arnold, and Edgar Brunner. Nonparametric hypotheses and rank statistics for unbalanced factorial designs. Journal of the American Statistical Association, 92(437):258–265, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/march/akritas.htm;
 http://www.jstor.org/stable/2291470.

Albright:1993:SAHa

[AASM93] S. Christian Albright, Jim Albert, Hal S. Stern, and Carl N. Morris. A statistical analysis of hitting streaks in baseball. Journal of the American Statistical Association, 88(424):1175–1183, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291254.

AB:1997:BRB

[AB97a] AB. Book review: [untitled]. Journal of the American Statistical Association, 92(438):799, June 1997. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965750.

Akritas:1997:NMF

[AB97b] Michael G. Akritas and Edgar Brunner. Nonparametric methods for factorial designs with censored data. Journal of the American Statistical Association, 92(438):568–576, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965705.

Albert:1993:BAB

[AC93] James H. Albert and Siddhartha Chib. Bayesian analysis of binary and polychotomous response data. Journal of the American Statistical Association, 88(422):669–679, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290350.

Altman:1995:NEB

[AC95a] N. S. Altman and G. Casella. Nonparametric empirical Bayes growth curve analysis. Journal of the American Statistical Association, 90(430):508–515, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2291061.

Alvo:1995:TOA

[AC95b] Mayer Alvo and Paul Cabilio. Testing ordered alternatives in the presence of incomplete data. Journal of the American Statistical Association, 90(431):1015–1024, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291338.

Atkinson:1995:ODH

[AC95c] A. C. Atkinson and R. D. Cook. D-optimum designs for heteroscedastic linear models. Journal of the American Statistical Association, 90(429):204–212, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291144.

Aerts:1997:LPE

[AC97a] Marc Aerts and Gerda Claeskens. Local polynomial estimation in multiparameter likelihood models. Journal of the American Statistical Association, 92(440):1536–1545, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_97/
december/AERTS.HTM;   http://www.jstor.org/stable/2965424.

Albert:1997:BTM

[AC97b] Jim Albert and Siddhartha Chib. Bayesian tests and model diagnostics in conditionally independent hierarchical models. Journal of the American Statistical Association, 92(439):916–925, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_97/sept/albert.htm;   http://www.jstor.
org/stable/2965555.

Aerts:1999:TFP

[ACH99] Marc Aerts, Gerda Claeskens, and Jeffrey D. Hart. Testing the fit of a parametric function. Journal of the American Statistical Association, 94(447):869–879, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670002.

Ackerman:1991:BRB

[Ack91] Terry A. Ackerman. Book review: [untitled]. Journal of the American Statistical Association, 86(414):549, June 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290615.

Ackerman:1993:BRB

[Ack93] Terry A. Ackerman. Book review: [untitled]. Journal of the American Statistical Association, 88(422):707, June 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290371.

Ackerson:1998:BRB

[Ack98] Lynn Ackerson. Book review: [untitled]. Journal of the American Statistical Association, 93(441):408–409, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669653.

Andrews:1990:NAU

[AD90] D. M. Andrews and H. A. David. Nonparametric analysis of unbalanced paired-comparison or ranked data. Journal of the American Statistical Association, 85(412):1140–1146, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289613.

Adak:1998:TDS

[Ada98] Sudeshna Adak. Time-dependent spectral analysis of nonstationary time series. Journal of the American Statistical Association, 93(444):1488–1501, December 1998. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
dec/adak.html;   http://www.jstor.org/stable/2670062.

Allard:1997:NML

[AF97] Denis Allard and Chris Fraley. Nonparametric maximum likelihood estimation of features in spatial point processes using Voronoi tessellation. Journal of the American Statistical Association, 92(440):1485–1493, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/december/ALLARD.HTM;   http://www.jstor.
org/stable/2965419.

Aitkin:1994:LE

[AG94] Murray Aitkin and Seymour Geisser. Letters to the editor. Journal of the American Statistical Association, 89(425):364, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291250.

Arjas:1996:BIS

[AG96] Elja Arjas and Dario Gasbarra. Bayesian inference of survival probabilities, under stochastic ordering constraints. Journal of the American Statistical Association, 91(435):1101–1109, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291729.

Antoniadis:1994:WMC

[AGM94] A. Antoniadis, G. Grégoire, and I. W. McKeague. Wavelet methods for curve estimation. Journal of the American Statistical Association, 89(428):1340–1353, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290996.

Angeles:1998:ACS

[AGM98a] Gustavo Angeles, David K. Guilkey, and Thomas A. Mroz. Applications and case studies — purposive program placement and the estimation of family planning program effects in Tanzania. Journal of the American Statistical Association, 93(443): 884–899, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/sept/angeles.html.

Angeles:1998:PPP

[AGM98b] Gustavo Angeles, David K. Guilkey, and Thomas A. Mroz. Purposive program placement and the estimation of family planning program effects in Tanzania. Journal of the American Statistical Association, 93(443):884–899, September 1998. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2669827.

Amit:1991:SIR

[AGP91] Yali Amit, Ulf Grenander, and Mauro Piccioni. Structural image restoration through deformable templates. Journal of the American Statistical Association, 86(414):376–387, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290581.

Agresti:1997:MRM

[Agr97] Alan Agresti. A model for repeated measurements of a multivariate binary response. Journal of the American Statistical Association, 92(437):315–321, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/march/
agresti.htm;   http://www.jstor.org/stable/2291476.

Agresti:1998:AEI

[Agr98a] Alan Agresti. Approximately exact inference for the common odds ration in several 2 × 2 tables: Comment. Journal of the American Statistical Association, 93(444):1307–1310, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670045.

Agresti:1998:CAE

[Agr98b] Alan Agresti. Comment: Approximately exact inference for the common odds ratio in several 2×2 tables. Journal of the American Statistical Association, 93(444):1307–??, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Albert:1997:MRM

[AHB97] Paul S. Albert, Sally A. Hunsberger, and Frank M. Biro. Modeling repeated measures with monotonic ordinal responses and misclassification, with applications to studying maturation. Journal of the American Statistical Association, 92(440):1304–1311, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_97/december/ALBERTPS.HTM;   http://www.
jstor.org/stable/2965400.

Ahn:1997:IVA

[Ahn97] Sung K. Ahn. Inference of vector autoregressive models with cointegration and scalar components. Journal of the American Statistical Association, 92(437):350–356, March 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/march/
ahn.htm;   http://www.jstor.org/stable/2291480.

Ahn:1999:BRB

[Ahn99] Sung K. Ahn. Book review: [untitled]. Journal of the American Statistical Association, 94(445):346, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669718.

Ahsanullah:1990:BRB

[Ahs90] Mohammad Ahsanullah. Book review: [untitled]. Journal of the American Statistical Association, 85(411):908–909, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290048.

Angelis:1993:ABA

[AHY93] D. De Angelis, Peter Hall, and G. A. Young. Analytical and bootstrap approximations to estimator distributions in L^{1} regression. Journal of the American Statistical Association, 88(424): 1310–1316, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291271.

Angrist:1995:TSL

[AI95] Joshua D. Angrist and Guido W. Imbens. Two-stage least squares estimation of average causal effects in models with variable treatment intensity. Journal of the American Statistical Association, 90(430):431–442, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291054.

Angrist:1996:ICEa

[AIR96a] Joshua D. Angrist, Guido W. Imbens, and Donald B. Rubin. Identification of causal effects using instrumental variables. Journal of the American Statistical Association, 91(434):444–455, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291629.

Angrist:1996:ICEb

[AIR96b] Joshua D. Angrist, Guido W. Imbens, and Donald B. Rubin. Identification of causal effects using instrumental variables: Rejoinder. Journal of the American Statistical Association, 91(434):468–472, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291634.

Angrist:1996:RIC

[AIR96c] Joshua D. Angrist, Guido W. Imbens, and Donald B. Rubin. Rejoinder: Identification of causal effects using instrumental variables. Journal of the American Statistical Association, 91(434): 444–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

AJM:1999:BRB

[AJM99] AJM. Book review: [untitled]. Journal of the American Statistical Association, 94(448):1389, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669964.

Angrist:1992:EAS

[AK92] Joshua D. Angrist and Alan B. Krueger. The effect of age at school entry on educational attainment: An application of instrumental variables with moments from two samples. Journal of the American Statistical Association, 87(418):328–336, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290263.

Aly:1994:STC

[AKM94] Emad-Eldin A. A. Aly, Subhash C. Kochar, and Ian W. McKeague. Some tests for comparing cumulative incidence functions and cause-specific hazard rates. Journal of the American Statistical Association, 89(427):994–999, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290925.

Akritas:1990:RTM

[Akr90] Michael G. Akritas. The rank transform method in some twofactor designs. Journal of the American Statistical Association, 85(409):73–78, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289527.

Akritas:1991:LRT

[Akr91a] Michael G. Akritas. Limitations of the rank transform procedure: a study of repeated measures designs, Part I. Journal of the American Statistical Association, 86(414):457–460, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290594.

Akritas:1991:RET

[Akr91b] Michael G. Akritas. Robust M estimation in the two-sample problem. Journal of the American Statistical Association, 86(413): 201–204, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289731.

Allenby:1994:MHP

[AL94] Greg M. Allenby and Peter J. Lenk. Modeling household purchase behavior with logistic normal regression. Journal of the American Statistical Association, 89(428):1218–1231, December 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290986.

Anderson:1995:SAU

[AL95] Jon E. Anderson and Thomas A. Louis. Survival analysis using a scale change random effects model. Journal of the American Statistical Association, 90(430):669–679, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291080.

Albert:1993:SAH

[Alb93a] Jim Albert. A statistical analysis of hitting streaks in baseball: Comment. Journal of the American Statistical Association, 88 (424):1184–1188, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291255.

Albert:1993:BRB

[Alb93b] Paul S. Albert. Book review: [untitled]. Journal of the American Statistical Association, 88(421):378–379, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290740.

Albright:1993:SAHb

[Alb93c] S. Christian Albright. A statistical analysis of hitting streaks in baseball: Rejoinder. Journal of the American Statistical Association, 88(424):1194–1196, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291257.

Albert:1994:EBH

[Alb94] Jim Albert. Exploring baseball hitting data: What about those breakdown statistics? Journal of the American Statistical Association, 89(427):1066–1074, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290936.

Albert:1995:BRB

[Alb95] Paul S. Albert. Book review: [untitled]. Journal of the American Statistical Association, 90(430):801–802, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291099.

Albert:1997:BTE

[Alb97] James H. Albert. Bayesian testing and estimation of association in a two-way contingency table. Journal of the American Statistical Association, 92(438):685–693, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965716.

Albert:1999:BDE

[Alb99] Jim Albert. Bridging different eras in sports: Comment. Journal of the American Statistical Association, 94(447):677–680, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669974.

Alexopoulos:1998:BRB

[Ale98] Christos Alexopoulos. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1244–1245, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669884.

Alho:1992:MFU

[Alh92] Juha M. Alho. Modeling and forecasting U.S. mortality: Comment. Journal of the American Statistical Association, 87(419): 673–674, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290203.

Anderson:1992:TDA

[ALHH92] Jon E. Anderson, Thomas A. Louis, Niels V. Holm, and Bent Harvald. Time-dependent association measures for bivariate survival distributions. Journal of the American Statistical Association, 87 (419):641–650, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290199.

Allenby:1999:DMP

[ALJ99] Greg M. Allenby, Robert P. Leone, and Lichung Jen. A dynamic model of purchase timing with application to direct marketing. Journal of the American Statistical Association, 94 (446):365–374, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/allenby.pdf;   http://www.jstor.org/
stable/2670153.

Arora:1997:EBE

[ALM97] Vipin Arora, P. Lahiri, and Kanchan Mukherjee. Empirical Bayes estimation of finite population means from complex surveys. Journal of the American Statistical Association, 92(440): 1555–1562, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/december/ARORA.HTM;
 http://www.jstor.org/stable/2965426.

Altman:1990:KSD

[Alt90] N. S. Altman. Kernel smoothing of data with correlated errors. Journal of the American Statistical Association, 85(411):749–759, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290011.

Altham:1994:BRB

[Alt94] P. M. E. Altham. Book review: [untitled]. Journal of the American Statistical Association, 89(426):718–719, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290886.

Azari:1992:PLO

[AM92] A. S. Azari and Hans-Georg Müller. Preaveraged localized orthogonal polynomial estimators for surface smoothing and partial differentiation. Journal of the American Statistical Association, 87 (420):1005–1017, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290638.

Al-Mutairi:1998:ACF

[AMCS98] Dhaifalla Al-Mutairi, Yiping Chen, and Nozer D. Singpurwalla. An adaptive concatenated failure rate model for software reliability. Journal of the American Statistical Association, 93(443): 1150–1163, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/sept/al-mutairi.html;
http://www.jstor.org/stable/2669858.

Ames:1992:BRB

[Ame92] Michael H. Ames. Book review: [untitled]. Journal of the American Statistical Association, 87(418):589, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290306.

Abrahamowicz:1996:TDH

[AME96] Michal Abrahamowicz, Todd MacKenzie, and John M. Esdaile. Time-dependent hazard ratio: Modeling and hypothesis testing with application in lupus nephritis. Journal of the American Statistical Association, 91(436):1432–1439, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291569.

Amemiya:1999:BRB

[Ame99] Yasuo Amemiya. Book review: [untitled]. Journal of the American Statistical Association, 94(445):337–338, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669709.

Akritas:1995:TSE

[AML95] Michael G. Akritas, Susan A. Murphy, and Michael P. LaValley. The Theil-Sen estimator with doubly censored data and applications to astronomy. Journal of the American Statistical Association, 90(429):170–177, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291140.

Ammann:1993:RSV

[Amm93] Larry P. Ammann. Robust singular value decompositions: a new approach to projection pursuit. Journal of the American Statistical Association, 88(422):505–514, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290330.

Agresti:1990:EIC

[AMP90] Alan Agresti, Cyrus R. Mehta, and Nitin R. Patel. Exact inference for contingency tables with ordered categories. Journal of the American Statistical Association, 85(410):453–458, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289783.

Amrhein:1995:MEP

[Amr95] Peter Amrhein. Minimax estimation of proportions under random sample size. Journal of the American Statistical Association, 90 (431):1107–1111, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291349.

Alho:1993:EHP

[AMWK93] Juha M. Alho, Mary H. Mulry, Kent Wurdeman, and Jay Kim. Estimating heterogeneity in the probabilities of enumeration for dual-system estimation. Journal of the American Statistical Association, 88(423):1130–1136, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290810.

Andersen:1991:MMG

[And91] E. B. Andersen. Measures, models, and graphical displays in the analysis of cross-classified data: Comment. Journal of the American Statistical Association, 86(416):1111–1112, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290529.

Andrich:1995:BRB

[And95] David Andrich. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1131, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291361.

Angus:1990:BRB

[Ang90] John E. Angus. Book review: [untitled]. Journal of the American Statistical Association, 85(412):1174–1175, December 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289629.

Angrist:1991:DLV

[Ang91] Joshua D. Angrist. The draft lottery and voluntary enlistment in the Vietnam era. Journal of the American Statistical Association, 86(415):584–595, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290386.

Angers:1993:BRB

[Ang93] Jean-François Angers. Book review: [untitled]. Journal of the American Statistical Association, 88(421):387, March 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290751.

Anonymous:1990:BMa

[Ano90a] Anonymous. Back matter. Journal of the American Statistical Association, 85(409):??, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289585.

Anonymous:1990:BMb

[Ano90b] Anonymous. Back matter. Journal of the American Statistical Association, 85(410):??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289834.

Anonymous:1990:BMc

[Ano90c] Anonymous. Back matter. Journal of the American Statistical Association, 85(411):??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290055.

Anonymous:1990:BMd

[Ano90d] Anonymous. Back matter. Journal of the American Statistical Association, 85(412):1183–1186, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289642.

Anonymous:1990:FMa

[Ano90e] Anonymous. Front matter. Journal of the American Statistical Association, 85(409):??, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289518.

Anonymous:1990:FMb

[Ano90f] Anonymous. Front matter. Journal of the American Statistical Association, 85(410):??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289760.

Anonymous:1990:FMc

[Ano90g] Anonymous. Front matter. Journal of the American Statistical Association, 85(411):??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289991.

Anonymous:1990:FMd

[Ano90h] Anonymous. Front matter. Journal of the American Statistical Association, 85(412):??, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289587.

Anonymous:1990:LBRa

[Ano90i] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 85(409):254, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289557.

Anonymous:1990:LBRb

[Ano90j] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 85(410):590, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289803.

Anonymous:1990:LBRc

[Ano90k] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 85(411):900, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290034.

Anonymous:1990:LBRd

[Ano90l] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 85(412):1165, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289617.

Anonymous:1990:PRa

[Ano90m] Anonymous. Publications received. Journal of the American Statistical Association, 85(409):270–271, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289581.

Anonymous:1990:PRb

[Ano90n] Anonymous. Publications received. Journal of the American Statistical Association, 85(410):606–607, June 1990. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289830.

Anonymous:1990:PRc

[Ano90o] Anonymous. Publications received. Journal of the American Statistical Association, 85(411):913–914, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290054.

Anonymous:1990:PRd

[Ano90p] Anonymous. Publications received. Journal of the American Statistical Association, 85(412):1180–1181, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289639.

Anonymous:1990:QCW

[Ano90q] Anonymous. Quality control of welfare programs. Journal of the American Statistical Association, 85(411):848–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1990:VI

[Ano90r] Anonymous. Volume information. Journal of the American Statistical Association, 85(412):1187–1192, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289586.

Anonymous:1991:BMa

[Ano91a] Anonymous. Back matter. Journal of the American Statistical Association, 86(413):??, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289759.

Anonymous:1991:BMb

[Ano91b] Anonymous. Back matter. Journal of the American Statistical Association, 86(414):??, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290624.

Anonymous:1991:BMc

[Ano91c] Anonymous. Back matter. Journal of the American Statistical Association, 86(415):??, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290443.

Anonymous:1991:BMd

[Ano91d] Anonymous. Back matter. Journal of the American Statistical Association, 86(416):1155–1158, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290554.

Anonymous:1991:CAP

[Ano91e] Anonymous. Correction: Approximate power for repeatedmeasures ANOVA lacking sphericity. Journal of the American Statistical Association, 86(413):255–256, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289757.

Anonymous:1991:CCI

[Ano91f] Anonymous. Correction: Confidence intervals for probabilities and tolerance regions based on a generalization of the Mann– Whitney statistic. Journal of the American Statistical Association, 86(413):255, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289756.

Anonymous:1991:CDS

[Ano91g] Anonymous. Correction: Double sampling for exact values in some multivariate measurement error problems. Journal of the American Statistical Association, 86(415):837, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290440.

Anonymous:1991:CFH

[Ano91h] Anonymous. Correction: Fitting H. F. Smith’s empirical law to cluster variances for use in designing multi-stage sample surveys. Journal of the American Statistical Association, 86(415): 837, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290441.

Anonymous:1991:CGV

[Ano91i] Anonymous. Correction: Generalized p-values in significance testing of hypotheses in the presence of nuisance parameters. Journal of the American Statistical Association, 86(413):256, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289758.

Anonymous:1991:CHT

[Ano91j] Anonymous. Correction: Homogeneity tests against centralmixture alternatives. Journal of the American Statistical Association, 86(415):837, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290442.

Anonymous:1991:FMa

[Ano91k] Anonymous. Front matter. Journal of the American Statistical Association, 86(413):??, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289705.

Anonymous:1991:FMb

[Ano91l] Anonymous. Front matter. Journal of the American Statistical Association, 86(414):??, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2290555.

Anonymous:1991:FMc

[Ano91m] Anonymous. Front matter. Journal of the American Statistical Association, 86(415):??, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290382.

Anonymous:1991:FMd

[Ano91n] Anonymous. Front matter. Journal of the American Statistical Association, 86(416):??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290494.

Anonymous:1991:LEa

[Ano91o] Anonymous. Letters to the editor. Journal of the American Statistical Association, 86(413):255, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289755.

Anonymous:1991:LEb

[Ano91p] Anonymous. Letters to the editor. Journal of the American Statistical Association, 86(416):1153–1154, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290553.

Anonymous:1991:LBRa

[Ano91q] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 86(413):239, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289734.

Anonymous:1991:LBRb

[Ano91r] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 86(414):542, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290605.

Anonymous:1991:LBRc

[Ano91s] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 86(415):822, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290421.

Anonymous:1991:LBRd

[Ano91t] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 86(416):1139, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290536.

Anonymous:1991:PRa

[Ano91u] Anonymous. Publications received. Journal of the American Statistical Association, 86(413):253–254, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289754.

Anonymous:1991:PRb

[Ano91v] Anonymous. Publications received. Journal of the American Statistical Association, 86(414):554–555, June 1991. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2290623.

Anonymous:1991:PRc

[Ano91w] Anonymous. Publications received. Journal of the American Statistical Association, 86(415):836, September 1991. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2290439.

Anonymous:1991:PRd

[Ano91x] Anonymous. Publications received. Journal of the American Statistical Association, 86(416):1151–1152, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290552.

Anonymous:1991:SSP

[Ano91y] Anonymous. Social statistics and public policy for the 1990s. Journal of the American Statistical Association, 86(414):486–??, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1991:VI

[Ano91z] Anonymous. Volume information. Journal of the American Statistical Association, 86(416):1159–1163, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290493.

Anonymous:1992:BMa

[Ano92a] Anonymous. Back matter. Journal of the American Statistical Association, 87(417):??, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290492.

Anonymous:1992:BMb

[Ano92b] Anonymous. Back matter. Journal of the American Statistical Association, 87(418):??, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290314.

Anonymous:1992:BMc

[Ano92c] Anonymous. Back matter. Journal of the American Statistical Association, 87(419):914, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290253.

Anonymous:1992:BMd

[Ano92d] Anonymous. Back matter. Journal of the American Statistical Association, 87(420):1253–1255, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290684.

Anonymous:1992:CES

[Ano92e] Anonymous. Correction: Evaluating screening for the early detection and treatment of cancer without using a randomized control group. Journal of the American Statistical Association, 87(417): 258, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290490.

Anonymous:1992:CMP

[Ano92f] Anonymous. Correction: Modeling the progression of human immunodeficiency virus infection. Journal of the American Statistical Association, 87(417):258, March 1992. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290491.

Anonymous:1992:FMa

[Ano92g] Anonymous. Front matter. Journal of the American Statistical Association, 87(417):??, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2290444.

Anonymous:1992:FMb

[Ano92h] Anonymous. Front matter. Journal of the American Statistical Association, 87(418):??, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2290254.

Anonymous:1992:FMc

[Ano92i] Anonymous. Front matter. Journal of the American Statistical Association, 87(419):??, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290191.

Anonymous:1992:FMd

[Ano92j] Anonymous. Front matter. Journal of the American Statistical Association, 87(420):??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290626.

Anonymous:1992:LEa

[Ano92k] Anonymous. Letters to the editor. Journal of the American Statistical Association, 87(417):258, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290489.

Anonymous:1992:LEb

[Ano92l] Anonymous. Letters to the editor. Journal of the American Statistical Association, 87(418):594–596, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290313.

Anonymous:1992:LEc

[Ano92m] Anonymous. Letters to the editor. Journal of the American Statistical Association, 87(419):912–913, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290251.

Anonymous:1992:LBRa

[Ano92n] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 87(417):248, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290476.

Anonymous:1992:LBRb

[Ano92o] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 87(418):584, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290298.

Anonymous:1992:LBRc

[Ano92p] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 87(419):896, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290230.

Anonymous:1992:LBRd

[Ano92q] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 87(420):1238, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290665.

Anonymous:1992:PRa

[Ano92r] Anonymous. Publications received. Journal of the American Statistical Association, 87(417):256–257, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290488.

Anonymous:1992:PRb

[Ano92s] Anonymous. Publications received. Journal of the American Statistical Association, 87(418):592–593, June 1992. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2290312.

Anonymous:1992:PRc

[Ano92t] Anonymous. Publications received. Journal of the American Statistical Association, 87(419):910–911, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290250.

Anonymous:1992:PRd

[Ano92u] Anonymous. Publications received. Journal of the American Statistical Association, 87(420):1250–1251, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290682.

Anonymous:1992:VI

[Ano92v] Anonymous. Volume information. Journal of the American Statistical Association, 87(420):1256–1260, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290625.

Anonymous:1993:BMa

[Ano93a] Anonymous. Back matter. Journal of the American Statistical Association, 88(421):??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290754.

Anonymous:1993:BMb

[Ano93b] Anonymous. Back matter. Journal of the American Statistical Association, 88(422):??, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290381.

Anonymous:1993:BMc

[Ano93c] Anonymous. Back matter. Journal of the American Statistical Association, 88(423):??, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290828.

Anonymous:1993:BMd

[Ano93d] Anonymous. Back matter. Journal of the American Statistical Association, 88(424):1480–1483, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291312.

Anonymous:1993:FMa

[Ano93e] Anonymous. Front matter. Journal of the American Statistical Association, 88(421):??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2290685.

Anonymous:1993:FMb

[Ano93f] Anonymous. Front matter. Journal of the American Statistical Association, 88(422):??, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2290315.

Anonymous:1993:FMc

[Ano93g] Anonymous. Front matter. Journal of the American Statistical Association, 88(423):??, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290755.

Anonymous:1993:FMd

[Ano93h] Anonymous. Front matter. Journal of the American Statistical Association, 88(424):??, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291253.

Anonymous:1993:LEa

[Ano93i] Anonymous. Letters to the editor. Journal of the American Statistical Association, 88(422):716–717, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290379.

Anonymous:1993:LEb

[Ano93j] Anonymous. Letters to the editor. Journal of the American Statistical Association, 88(424):1478–1479, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291311.

Anonymous:1993:LBRa

[Ano93k] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 88(421):374, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290734.

Anonymous:1993:LBRb

[Ano93l] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 88(422):694, June 1993. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290353.

Anonymous:1993:LBRc

[Ano93m] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 88(423):1167, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290816.

Anonymous:1993:LBRd

[Ano93n] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 88(424):1466, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291294.

Anonymous:1993:PRa

[Ano93o] Anonymous. Publications received. Journal of the American Statistical Association, 88(422):713–715, June 1993. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2290378.

Anonymous:1993:PRb

[Ano93p] Anonymous. Publications received. Journal of the American Statistical Association, 88(423):1173–1174, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290827.

Anonymous:1993:PRc

[Ano93q] Anonymous. Publications received. Journal of the American Statistical Association, 88(424):1476–1477, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291310.

Anonymous:1993:VI

[Ano93r] Anonymous. Volume information. Journal of the American Statistical Association, 88(424):1484–1488, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291252.

Anonymous:1994:BMa

[Ano94a] Anonymous. Back matter. Journal of the American Statistical Association, 89(425):??, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291251.

Anonymous:1994:BMb

[Ano94b] Anonymous. Back matter. Journal of the American Statistical Association, 89(426):??, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290896.

Anonymous:1994:BMc

[Ano94c] Anonymous. Back matter. Journal of the American Statistical Association, 89(427):??, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290975.

Anonymous:1994:BMd

[Ano94d] Anonymous. Back matter. Journal of the American Statistical Association, 89(428):1573–1582, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291045.

Anonymous:1994:FMa

[Ano94e] Anonymous. Front matter. Journal of the American Statistical Association, 89(425):??, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2291194.

Anonymous:1994:FMb

[Ano94f] Anonymous. Front matter. Journal of the American Statistical Association, 89(426):??, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2290829.

Anonymous:1994:FMc

[Ano94g] Anonymous. Front matter. Journal of the American Statistical Association, 89(427):??, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290897.

Anonymous:1994:FMd

[Ano94h] Anonymous. Front matter. Journal of the American Statistical Association, 89(428):??, December 1994. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290977.

Anonymous:1994:LE

[Ano94i] Anonymous. Letters to the editor. Journal of the American Statistical Association, 89(426):724–725, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290894.

Anonymous:1994:LBRa

[Ano94j] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 89(425):353, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291232.

Anonymous:1994:LBRb

[Ano94k] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 89(426):710, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290876.

Anonymous:1994:LBRc

[Ano94l] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 89(427):1135, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290944.

Anonymous:1994:LBRd

[Ano94m] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 89(428):1558, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291019.

Anonymous:1994:PRa

[Ano94n] Anonymous. Publications received. Journal of the American Statistical Association, 89(425):363–364, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291249.

Anonymous:1994:PRb

[Ano94o] Anonymous. Publications received. Journal of the American Statistical Association, 89(426):723, June 1994. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290893.

Anonymous:1994:VI

[Ano94p] Anonymous. Volume information. Journal of the American Statistical Association, 89(428):1577–1581, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290976.

Anonymous:1995:BMa

[Ano95a] Anonymous. Back matter. Journal of the American Statistical Association, 90(429):??, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291193.

Anonymous:1995:BMb

[Ano95b] Anonymous. Back matter. Journal of the American Statistical Association, 90(430):??, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291121.

Anonymous:1995:BMc

[Ano95c] Anonymous. Back matter. Journal of the American Statistical Association, 90(431):??, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291376.

Anonymous:1995:BMd

[Ano95d] Anonymous. Back matter. Journal of the American Statistical Association, 90(432):1495–1499, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291561.

Anonymous:1995:CSC

[Ano95e] Anonymous. Correction: Simultaneous confidence intervals and sample size determination for multinomial proportions. Journal of the American Statistical Association, 90(432):1494, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291560.

Anonymous:1995:EC

[Ano95f] Anonymous. Editorial collaborators. Journal of the American Statistical Association, 90(432):1495–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1995:FMa

[Ano95g] Anonymous. Front matter. Journal of the American Statistical Association, 90(429):??, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2291122.

Anonymous:1995:FMb

[Ano95h] Anonymous. Front matter. Journal of the American Statistical Association, 90(430):??, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2291046.

Anonymous:1995:FMc

[Ano95i] Anonymous. Front matter. Journal of the American Statistical Association, 90(431):??, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291313.

Anonymous:1995:FMd

[Ano95j] Anonymous. Front matter. Journal of the American Statistical Association, 90(432):??, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291505.

Anonymous:1995:LE

[Ano95k] Anonymous. Letters to the editors. Journal of the American Statistical Association, 90(430):811–812, June 1995. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2291119.

Anonymous:1995:LBRa

[Ano95l] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 90(429):383, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291165.

Anonymous:1995:LBRb

[Ano95m] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 90(430):796, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291092.

Anonymous:1995:LBRc

[Ano95n] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 90(431):1122, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291351.

Anonymous:1995:LBRd

[Ano95o] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 90(432):1482, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291541.

Anonymous:1995:TRa

[Ano95p] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 90(430):809–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1995:TRb

[Ano95q] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 90(432):1492–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1995:VI

[Ano95r] Anonymous. Volume information. Journal of the American Statistical Association, 90(432):1500–1504, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291504.

Anonymous:1996:ANM

[Ano96a] Anonymous. Addendum: Network models for complementary cell suppression. Journal of the American Statistical Association, 91 (436):1757, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291623.

Anonymous:1996:BMa

[Ano96b] Anonymous. Back matter. Journal of the American Statistical Association, 91(433):??, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291442.

Anonymous:1996:BMb

[Ano96c] Anonymous. Back matter. Journal of the American Statistical Association, 91(434):??, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291708.

Anonymous:1996:BMc

[Ano96d] Anonymous. Back matter. Journal of the American Statistical Association, 91(435):??, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291767.

Anonymous:1996:BMd

[Ano96e] Anonymous. Back matter. Journal of the American Statistical Association, 91(436):1758–1762, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291626.

Anonymous:1996:CEB

[Ano96f] Anonymous. Correction: An effective bandwidth selector for local least squares regression. Journal of the American Statistical Association, 91(435):1380, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291766.

Anonymous:1996:CRT

[Ano96g] Anonymous. Correction: Rank tests for main and interaction effects in analysis of variance. Journal of the American Statistical Association, 91(434):916, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291707.

Anonymous:1996:CCP

[Ano96h] Anonymous. Corrections: Conjugate parameterizations for natural exponential families. Journal of the American Statistical Association, 91(436):1757, December 1996. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291625.

Anonymous:1996:CUM

[Ano96i] Anonymous. Corrections: Uniformly more powerful tests in a onesided multivariate problem. Journal of the American Statistical Association, 91(436):1757, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291624.

Anonymous:1996:FMa

[Ano96j] Anonymous. Front matter. Journal of the American Statistical Association, 91(433):??, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2291377.

Anonymous:1996:FMb

[Ano96k] Anonymous. Front matter. Journal of the American Statistical Association, 91(434):??, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2291627.

Anonymous:1996:FMc

[Ano96l] Anonymous. Front matter. Journal of the American Statistical Association, 91(435):??, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291709.

Anonymous:1996:FMd

[Ano96m] Anonymous. Front matter. Journal of the American Statistical Association, 91(436):??, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291563.

Anonymous:1996:LBRa

[Ano96n] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 91(433):429, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291422.

Anonymous:1996:LBRb

[Ano96o] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 91(434):905, June 1996. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291684.

Anonymous:1996:LBRc

[Ano96p] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 91(435):1371, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291753.

Anonymous:1996:LBRd

[Ano96q] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 91(436):1749, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291605.

Anonymous:1996:TRa

[Ano96r] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 91(433):440–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1996:TRb

[Ano96s] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 91(434):915–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1996:TRc

[Ano96t] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 91(435):1380–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1996:VI

[Ano96u] Anonymous. Volume information. Journal of the American Statistical Association, 91(436):1763–1768, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291562.

Anonymous:1997:EC

[Ano97a] Anonymous. 1997 editorial collaborators. Journal of the American Statistical Association, 92(440):1658–??, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
97editors.html.

Anonymous:1997:BMa

[Ano97b] Anonymous. Back matter. Journal of the American Statistical Association, 92(437):??, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291503.

Anonymous:1997:BMb

[Ano97c] Anonymous. Back matter. Journal of the American Statistical Association, 92(438):??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2965760.

Anonymous:1997:BMc

[Ano97d] Anonymous. Back matter. Journal of the American Statistical Association, 92(439):??, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965621.

Anonymous:1997:BMd

[Ano97e] Anonymous. Back matter. Journal of the American Statistical Association, 92(440):1658–1661, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965465.

Anonymous:1997:BRa

[Ano97f] Anonymous. Book reviews. Journal of the American Statistical Association, 92(437):383–??, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1997:BRb

[Ano97g] Anonymous. Book reviews. Journal of the American Statistical Association, 92(438):787–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1997:BRc

[Ano97h] Anonymous. Book reviews. Journal of the American Statistical Association, 92(439):1213–??, September 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1997:CCP

[Ano97i] Anonymous. Correction: Correction to practical small sample asymptotics for regression problems. Journal of the American Statistical Association, 92(440):1657, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965464.

Anonymous:1997:CBE

[Ano97j] Anonymous. Corrections: Bahadur efficiency of tests of separate hypotheses and adaptive test statistics. Journal of the American Statistical Association, 92(437):390, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291501.

Anonymous:1997:CSI

[Ano97k] Anonymous. Corrections: Spatial interpolation errors for monitoring data. Journal of the American Statistical Association, 92(437):390, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291502.

Anonymous:1997:FMa

[Ano97l] Anonymous. Front matter. Journal of the American Statistical Association, 92(437):??, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2291443.

Anonymous:1997:FMb

[Ano97m] Anonymous. Front matter. Journal of the American Statistical Association, 92(438):??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2965682.

Anonymous:1997:FMc

[Ano97n] Anonymous. Front matter. Journal of the American Statistical Association, 92(439):??, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965544.

Anonymous:1997:FMd

[Ano97o] Anonymous. Front matter. Journal of the American Statistical Association, 92(440):??, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965389.

Anonymous:1997:IV

[Ano97p] Anonymous. Index to volume 92. Journal of the American Statistical Association, 92(440):1662–??, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1997:LBRa

[Ano97q] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 92(437):383, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291484.

Anonymous:1997:LBRb

[Ano97r] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 92(438):788, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965728.

Anonymous:1997:LBRc

[Ano97s] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 92(439):1213, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965587.

Anonymous:1997:LBRd

[Ano97t] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 92(440):1639, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965435.

Anonymous:1997:VI

[Ano97u] Anonymous. Volume information. Journal of the American Statistical Association, 92(440):1662–1667, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965388.

Anonymous:1998:AMT

[Ano98a] Anonymous. Addendum: “multiple testing of general contracts using logical constraints and correlations”. Journal of the American Statistical Association, 93(441):413, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669670.

Anonymous:1998:BMa

[Ano98b] Anonymous. Back matter. Journal of the American Statistical Association, 93(441):??, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669671.

Anonymous:1998:BMb

[Ano98c] Anonymous. Back matter. Journal of the American Statistical Association, 93(442):??, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2670150.

Anonymous:1998:BMc

[Ano98d] Anonymous. Back matter. Journal of the American Statistical Association, 93(443):??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669910.

Anonymous:1998:BMd

[Ano98e] Anonymous. Back matter. Journal of the American Statistical Association, 93(444):??, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670087.

Anonymous:1998:BRa

[Ano98f] Anonymous. Book reviews. Journal of the American Statistical Association, 93(441):397–??, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1998:BRb

[Ano98g] Anonymous. Book reviews. Journal of the American Statistical Association, 93(442):832–??, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1998:CAU

[Ano98h] Anonymous. Correction: Achieving uniformity in a semiconductor fabrication process using spatial modeling. Journal of the American Statistical Association, 93(443):1252, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669908.

Anonymous:1998:CPS

[Ano98i] Anonymous. Correction: The power of some tests for difference stationarity under local heteroscedastic integration. Journal of the American Statistical Association, 93(443):1252, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669909.

Anonymous:1998:CSP

[Ano98j] Anonymous. Correction: “the selection of prior distributions”. Journal of the American Statistical Association, 93(441):412, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2669669.

Anonymous:1998:ER

[Ano98k] Anonymous. Editor’s report for 1997. Journal of the American Statistical Association, 93(443):845, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669818.

Anonymous:1998:FMa

[Ano98l] Anonymous. Front matter. Journal of the American Statistical Association, 93(441):397, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669596.

Anonymous:1998:FMb

[Ano98m] Anonymous. Front matter. Journal of the American Statistical Association, 93(442):832, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2670088.

Anonymous:1998:FMc

[Ano98n] Anonymous. Front matter. Journal of the American Statistical Association, 93(443):1231, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669817.

Anonymous:1998:FMd

[Ano98o] Anonymous. Front matter. Journal of the American Statistical Association, 93(444):1512, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670039.

Anonymous:1998:GEC

[Ano98p] Anonymous. General — editorial collaborators. Journal of the American Statistical Association, 93(444):1531–??, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1998:GRP

[Ano98q] Anonymous. General — review paper. Journal of the American Statistical Association, 93(443):1216–??, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1998:VI

[Ano98r] Anonymous. Volume information. Journal of the American Statistical Association, 93(444):1531–1541, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670038.

Anonymous:1999:AME

[Ano99a] Anonymous. Acknowledgement: a model for evaluating sensitivity and specificity for correlated diagnostic tests in efficacy studies in an imperfect reference test. Journal of the American Statistical Association, 94(446):660, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670214.

Anonymous:1999:BMa

[Ano99b] Anonymous. Back matter. Journal of the American Statistical Association, 94(445):??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669736.

Anonymous:1999:BMb

[Ano99c] Anonymous. Back matter. Journal of the American Statistical Association, 94(446):??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2670217.

Anonymous:1999:BMc

[Ano99d] Anonymous. Back matter. Journal of the American Statistical Association, 94(447):??, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670037.

Anonymous:1999:BMd

[Ano99e] Anonymous. Back matter. Journal of the American Statistical Association, 94(448):??, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669971.

Anonymous:1999:BRa

[Ano99f] Anonymous. Book reviews. Journal of the American Statistical Association, 94(445):333–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1999:BRb

[Ano99g] Anonymous. Book reviews. Journal of the American Statistical Association, 94(446):645–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1999:CCA

[Ano99h] Anonymous. Corrections: Comment on “Approximately Exact Inference for the Common Odds Ratio in Several 2×2 Tables” by R. L. Strawderman and M. L. Wells. Journal of the American Statistical Association, 94(446):660, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670216.

Anonymous:1999:CCL

[Ano99i] Anonymous. Corrections: Conditional likelihood ratio test for a nonnegative normal mean vector. Journal of the American Statistical Association, 94(446):660, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670215.

Anonymous:1999:ER

[Ano99j] Anonymous. Editors’ report for 1998. Journal of the American Statistical Association, 94(446):661–??, June 1999. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/index.html#98tm.

Anonymous:1999:FMa

[Ano99k] Anonymous. Front matter. Journal of the American Statistical Association, 94(445):333, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669672.

Anonymous:1999:FMb

[Ano99l] Anonymous. Front matter. Journal of the American Statistical Association, 94(446):645, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2670151.

Anonymous:1999:FMc

[Ano99m] Anonymous. Front matter. Journal of the American Statistical Association, 94(447):979, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669972.

Anonymous:1999:FMd

[Ano99n] Anonymous. Front matter. Journal of the American Statistical Association, 94(448):1382, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669912.

Anonymous:1999:LEa

[Ano99o] Anonymous. Letters to the Editor. Journal of the American Statistical Association, 94(445):351–??, March 1999. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1999:LEb

[Ano99p] Anonymous. Letters to the Editor. Journal of the American Statistical Association, 94(446):659–??, June 1999. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1999:VI

[Ano99q] Anonymous. Volume information. Journal of the American Statistical Association, 94(448):1391–1401, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669911.

Anroch:1996:BRB

[Anr96] Jaromir Anroch. Book review: [untitled]. Journal of the American Statistical Association, 91(433):438, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291434.

Andrews:1996:TSC

[AP96] Donald W. K. Andrews and Werner Ploberger. Testing for serial correlation against an ARMA(1, 1,) process. Journal of the American Statistical Association, 91(435):1331–1342, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291751.

Agustin:1999:OSP

[AP99] Ma. Zenia N. Agustin and Edsel A. Peña. Order statistic properties, random generation, and goodness-of-fit testing for a minimal repair model. Journal of the American Statistical Association, 94 (445):266–272, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2669701.

Appel:1997:BRB

[App97] M. Appel. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1216–1217, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965592.

Assaf:1992:NLW

[APR92] David Assaf, Moshe Pollak, and Ya’acov Ritov. A new look at warning and action lines of surveillance schemes. Journal of the American Statistical Association, 87(419):889–895, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290229.

Ahn:1990:EPN

[AR90] Sung K. Ahn and Gregory C. Reinsel. Estimation for partially nonstationary multivariate autoregressive models. Journal of the American Statistical Association, 85(411):813–823, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290020.

Arellano:1992:UNL

[Are92] Max. G. Arellano. The use of names for linking personal records: Comment. Journal of the American Statistical Association, 87 (420):1204–1206, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290661.

Alho:1990:EMO

[AS90a] Juha M. Alho and Bruce D. Spencer. Error models for official mortality forecasts. Journal of the American Statistical Association, 85(411):609–616, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289992.

Arminger:1990:PML

[AS90b] Gerhard Sobel Arminger and Michael E. Sobel. Pseudo-maximum likelihood estimation of mean and covariance structures with missing data. Journal of the American Statistical Association, 85(409): 195–203, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289545.

Assuncao:1993:BRB

[AS93] Renato Assuncao and Paul D. Sampson. Book review: [untitled]. Journal of the American Statistical Association, 88(421): 383, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290746.

Anderson-Sprecher:1991:SSA

[ASL91] Richard Anderson-Sprecher and Johannes Ledolter. State-space analysis of wildlife telemetry data. Journal of the American Statistical Association, 86(415):596–602, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290387.

Anderson-Sprecher:1996:SEP

[ASL96] Richard Anderson-Sprecher and Russell V. Lenth. Spline estimation of paths using bearings-only tracking data. Journal of the American Statistical Association, 91(433):276–283, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291405.

Asmussen:1993:BRB

[Asm93] Soren Asmussen. Book review: [untitled]. Journal of the American Statistical Association, 88(424):1474, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291306.

Athreya:1999:BRB

[Ath99] K. B. Athreya. Book review: [untitled]. Journal of the American Statistical Association, 94(448):1386–1387, December 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669956.

Atkinson:1991:TDE

[Atk91] A. C. Atkinson. Transformations in density estimation: Comment. Journal of the American Statistical Association, 86(414): 354–355, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290571.

Atkinson:1994:FVR

[Atk94] A. C. Atkinson. Fast very robust methods for the detection of multiple outliers. Journal of the American Statistical Association, 89 (428):1329–1339, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290995.

Atrostic:1994:BRB

[Atr94] B. K. Atrostic. Book review: [untitled]. Journal of the American Statistical Association, 89(426):721–722, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290892.

Aykroyd:1999:IPM

[AZ99] R. G. Aykroyd and S. Zimeras. Inhomogeneous prior models for image reconstruction. Journal of the American Statistical Association, 94(447):934–946, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670008.

Bacchetti:1990:EIP

[Bac90] Peter Bacchetti. Estimating the incubation period of AIDS by comparing population infection and diagnosis patterns. Journal of the American Statistical Association, 85(412):1002–1008, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289596.

Bagchi:1990:BRB

[Bag90] Parthasarathy Bagchi. Book review: [untitled]. Journal of the American Statistical Association, 85(412):1168–1169, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289622.

Bagghi:1991:BRB

[Bag91] Parthasarathy Bagghi. Book review: [untitled]. Journal of the American Statistical Association, 86(413):244–245, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289739.

Baker:1998:ASD

[Bak98a] Stuart G. Baker. Analysis of survival data from a randomized trial with all-or-none compliance: Estimating the cost-effectiveness of a cancer screening program. Journal of the American Statistical Association, 93(443):929–934, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669831.

Baker:1998:ACS

[Bak98b] Stuart G. Baker. Applications and case studies — analysis of survival data from a randomized trial with all-or-none compliance: Estimating the cost-effectiveness of a cancer screening program. Journal of the American Statistical Association, 93(443): 929–934, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/sept/baker.html.

Ballerini:1994:BRB

[Bal94] Rocco Ballerini. Book review: [untitled]. Journal of the American Statistical Association, 89(425):357, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291238.

Balding:1995:EPF

[Bal95] David J. Balding. Estimating products in forensic identification using DNA profiles. Journal of the American Statistical Association, 90(431):839–844, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291317.

Ballerini:1996:BRB

[Bal96] Rocco Ballerini. Book review: [untitled]. Journal of the American Statistical Association, 91(434):913–914, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291695.

Balls:1997:BRB

[Bal97] Kim G. Balls. Book review: [untitled]. Journal of the American Statistical Association, 92(437):386–387, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291489.

Banks:1990:BRB

[Ban90] David L. Banks. Book review: [untitled]. Journal of the American Statistical Association, 85(409):264, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289571.

Banks:1993:BRB

[Ban93] David L. Banks. Book review: [untitled]. Journal of the American Statistical Association, 88(422):708–710, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290373.

Baltazar-Aban:1995:PHB

[BAP95] Inmaculada Baltazar-Aban and Edsel A. Peña. Properties of hazard-based residuals and implications in model diagnostics. Journal of the American Statistical Association, 90(429):185– 197, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291142.

Barabba:1991:TGL

[Bar91] Vincent P. Barabba. Through a glass less darkly. Journal of the American Statistical Association, 86(413):1–8, March 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289706.

Bartko:1992:BRB

[Bar92] John J. Bartko. Book review: [untitled]. Journal of the American Statistical Association, 87(419):907–908, September 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290246.

Barbour:1993:BRB

[Bar93a] A. D. Barbour. Book review: [untitled]. Journal of the American Statistical Association, 88(424):1474–1475, December 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291307.

Barnett:1993:IMO

[Bar93b] Vic Barnett. The identification of multiple outliers: Comment. Journal of the American Statistical Association, 88(423):795–796, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290766.

Barrett:1997:BRB

[Bar97] J. Douglas Barrett. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1217–1218, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2965593.

Bartleson:1998:BRB

[Bar98] Becki Bucher Bartleson. Book review: [untitled]. Journal of the American Statistical Association, 93(441):408, March 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669652.

Bloom:1990:MAM

[BB90] David E. Bloom and Neil G. Bennett. Modeling American marriage patterns. Journal of the American Statistical Association, 85 (412):1009–1017, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289597.

Berger:1994:VMC

[BB94] Roger L. Berger and Dennis D. Boos. P values maximized over a confidence set for the nuisance parameter. Journal of the American Statistical Association, 89(427):1012–1016, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290928.

BB:1998:BRB

[BB98a] BB. Book review: [untitled]. Journal of the American Statistical Association, 93(441):412, March 1998. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669664.

Booth:1998:AEI

[BB98b] James G. Booth and Ronald W. Butler. Approximately exact inference for the common odds ration in several 2 × 2 tables: Comment. Journal of the American Statistical Association, 93(444): 1310–1313, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2670046.

Booth:1998:CAE

[BB98c] James G. Booth and Ronald W. Butler. Comment: Approximately exact inference for the common odds ratio in several 2 × 2 tables. Journal of the American Statistical Association, 93(444): 1310–??, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

BBB:1998:BRB

[BBB98] BBB. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1530, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670085.

Boneh:1998:EPF

[BBC98] Shahar Boneh, Arnon Boneh, and R. J. Caron. Estimating the prediction function and the number of unseen species in sampling with replacement. Journal of the American Statistical Association, 93(441):372–379, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

amstat.org/publications/jasa/abstracts_98/BONEH.HTM;
http://www.jstor.org/stable/2669633.

Berzuini:1997:DCI

[BBGL97] Carlo Berzuini, Nicola G. Best, Walter R. Gilks, and Cristiana Larizza. Dynamic conditional independence models and Markov chain Monte Carlo methods. Journal of the American Statistical Association, 92(440):1403–1412, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_97/
december/BERZUINI.HTM;   http://www.jstor.org/stable/
2965410.

Booth:1994:BMF

[BBH94] James G. Booth, Ronald W. Butler, and Peter Hall. Bootstrap methods for finite populations. Journal of the American Statistical Association, 89(428):1282–1289, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290991.

Berger:1996:EIR

[BBH96] Agnes Berger, Carol A. Bodian, and Warren M. Hirsch. On estimating incidence rates of diseases with delayed onset using biased samples. Journal of the American Statistical Association, 91(434): 831–840, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291678.

Baker:1990:ESE

[BC90] Stuart G. Baker and Kenneth C. Chu. Evaluating screening for the early detection and treatment of cancer without using a randomized control group. Journal of the American Statistical Association, 85(410):321–327, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289767.

Breslow:1993:AIG

[BC93] N. E. Breslow and D. G. Clayton. Approximate inference in generalized linear mixed models. Journal of the American Statistical Association, 88(421):9–25, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290687.

Bensmail:1996:RGD

[BC96] Halima Bensmail and Gilles Celeux. Regularized Gaussian discriminant analysis through eigenvalue decomposition. Journal of the American Statistical Association, 91(436):1743–1748, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291604.

Baccini:1991:MMG

[BCdF91] Alain Baccini, Henri Caussinus, and Antoine de Falguerolles. Measures, models, and graphical displays in the analysis of crossclassified data: Comment. Journal of the American Statistical Association, 86(416):1115–1117, December 1991. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2290531.

Bonetti:1999:MME

[BCG99] Marco Bonetti, Bernard F. Cole, and Richard D. Gelber. A method-of-moments estimation procedure for categorical qualityof-life data with nonignorable missingness. Journal of the American Statistical Association, 94(448):1025–1034, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669916.

Brown:1995:OCS

[BCH95] Lawrence D. Brown, George Casella, and J. T. Gene Hwang. Optimal confidence sets, bioequivalence, and the limaçon of Pascal. Journal of the American Statistical Association, 90(431):880–889, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291322.

Bedrick:1996:NPP

[BCJ96] Edward J. Bedrick, Ronald Christensen, and Wesley Johnson. A new perspective on priors for generalized linear models. Journal of the American Statistical Association, 91(436):1450–1460, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291571.

Bayarri:1991:WBE

[BD91a] M. J. Bayarri and M. H. DeGroot. What Bayesians expect of each other. Journal of the American Statistical Association, 86(416): 924–932, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290508.

Bryant:1991:EBA

[BD91b] John Bryant and Roger Day. Empirical Bayes analysis for systems of mixed models with linked autocorrelated random effects. Journal of the American Statistical Association, 86(416):1007–1012, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290518.

Brunner:1992:RPT

[BD92] Edgar Brunner and Holger Dette. Rank procedures for the twofactor mixed model. Journal of the American Statistical Association, 87(419):884–888, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290228.

Burr:1993:CBM

[BD93] Deborah Burr and Hani Doss. Confidence bands for the median survival time as a function of the covariates in the Cox model. Journal of the American Statistical Association, 88(424):1330– 1340, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291274.

Bollinger:1997:MDC

[BD97] Christopher R. Bollinger and Martin H. David. Modeling discrete choice with response error: Food stamp participation. Journal of the American Statistical Association, 92(439):827– 835, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/sept/bollinger.htm;
 http://www.jstor.org/stable/2965547.

Berk:1999:ECI

[BdL99] Richard A. Berk and Jan de Leeuw. An evaluation of California’s inmate classification system using a generalized regression discontinuity design. Journal of the American Statistical Association, 94(448):1045–1052, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669918.

Belin:1993:HLRa

[BDM^{+}93a] Thomas R. Belin, Gregg J. Diffendal, Steve Mack, Donald B. Rubin, Joseph L. Schafer, and Alan M. Zaslavsky. Hierarchical logistic regression models for imputation of unresolved enumeration status in undercount estimation. Journal of the American Statistical Association, 88(423):1149–1159, September 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290812.

Belin:1993:HLRb

[BDM^{+}93b] Thomas R. Belin, Gregg J. Diffendal, Steve Mack, Donald B. Rubin, Joseph L. Schafer, and Alan M. Zaslavsky. Hierarchical logistic regression models for imputation of unresolved enumeration status in undercount estimation: Rejoinder. Journal of the American Statistical Association, 88(423):1163–1166, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290815.

Brunner:1997:BTA

[BDM97] Edgar Brunner, Holger Dette, and Axel Munk. Box-type approximations in nonparametric factorial designs. Journal of the American Statistical Association, 92(440):1494–1502, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2965420.

Boudjellaba:1992:TCB

[BDR92] Hafida Boudjellaba, Jean-Marie Dufour, and Roch Roy. Testing causality between two vectors in multivariate autoregressive moving average models. Journal of the American Statistical Association, 87(420):1082–1090, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290645.

Beauchamp:1991:WFA

[Bea91] K. G. Beauchamp. Walsh–Fourier analysis and its statistical applications: Comment. Journal of the American Statistical Association, 86(414):480, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290596.

Becker:1991:BRB

[Bec91] Mark P. Becker. Book review: [untitled]. Journal of the American Statistical Association, 86(415):828–829, September 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290427.

Becker:1998:BRB

[Bec98] Mark P. Becker. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1526, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670073.

Becker:1999:BRB

[Bec99] Betsy Jane Becker. Book review: [untitled]. Journal of the American Statistical Association, 94(445):343–344, March 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669715.

Bell:1993:UID

[Bel93a] William R. Bell. Using information from demographic analysis in post-enumeration survey estimation. Journal of the American Statistical Association, 88(423):1106–1118, September 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290805.

Bellhouse:1993:BRB

[Bel93b] David Bellhouse. Book review: [untitled]. Journal of the American Statistical Association, 88(424):1469–1470, December 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291299.

Bentley:1991:BRB

[Ben91a] Donald L. Bentley. Book review: [untitled]. Journal of the American Statistical Association, 86(415):835, September 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290437.

Benzecri:1991:MMG

[Ben91b] J. P. Benzecri. Measures, models, and graphical displays in the analysis of cross-classified data: Comment. Journal of the American Statistical Association, 86(416):1112–1115, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290530.

Benjamini:1997:BRB

[Ben97] Yoav Benjamini. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1218–1219, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965594.

Beran:1990:CPR

[Ber90a] Rudolf Beran. Calibrating prediction regions. Journal of the American Statistical Association, 85(411):715–723, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290007.

Beran:1990:RBS

[Ber90b] Rudolf Beran. Refining bootstrap simultaneous confidence sets. Journal of the American Statistical Association, 85(410):417–426, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289778.

Beran:1991:ELG

[Ber91a] Jan Beran. M estimators of location for Gaussian and related processes with slowly decaying serial correlations. Journal of the American Statistical Association, 86(415):704–708, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290401.

Berliner:1991:LBP

[Ber91b] L. Mark Berliner. Likelihood and Bayesian prediction of chaotic systems. Journal of the American Statistical Association, 86(416): 938–952, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290510.

Berk:1992:BRBa

[Ber92a] Kenneth N. Berk. Book review: [untitled]. Journal of the American Statistical Association, 87(419):900–901, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290236.

Berk:1992:BRBb

[Ber92b] Kenneth N. Berk. Book review: [untitled]. Journal of the American Statistical Association, 87(419):901, September 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290237.

Beran:1993:BRB

[Ber93] Rudolf Beran. Book review: [untitled]. Journal of the American Statistical Association, 88(421):375–376, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290736.

Bera:1994:BRBa

[Ber94a] Anil K. Bera. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1144–1145, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290954.

Bera:1994:BRBb

[Ber94b] Anil K. Bera. Book review: [untitled]. Journal of the American Statistical Association, 89(428):1567–1569, December 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291031.

Beran:1995:BRB

[Ber95] Rudolf Beran. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1128–1129, September 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291358.

Brown:1990:LBR

[BF90] Lawrence D. Brown and Roger H. Farrell. A lower bound for the risk in estimating the value of a probability density. Journal of the American Statistical Association, 85(412):1147–1153, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289614.

Biemer:1992:QRD

[BF92] Paul P. Biemer and Gösta Forsman. On the quality of reinterview data with application to the current population survey. Journal of the American Statistical Association, 87(420):915–923, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290627.

Bowman:1993:ASD

[BF93a] A. W. Bowman and P. J. Foster. Adaptive smoothing and densitybased tests of multivariate normality. Journal of the American Statistical Association, 88(422):529–537, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290333.

Bunge:1993:ENS

[BF93b] J. Bunge and M. Fitzpatrick. Estimating the number of species: a review. Journal of the American Statistical Association, 88(421):

364–373, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290733.

Banerjee:1997:IDL

[BF97] Mousumi Banerjee and Edward W. Frees. Influence diagnostics for linear longitudinal models. Journal of the American Statistical Association, 92(439):999–1005, September 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/sept/
banerjee.htm;   http://www.jstor.org/stable/2965564.

Brown:1999:DMV

[BFH99] P. J. Brown, T. Fearn, and M. S. Haque. Discrimination with many variables. Journal of the American Statistical Association, 94(448):1320–1329, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669945.

Beran:1991:SDC

[BG91a] Jan Beran and Sucharita Ghosh. Slowly decaying correlations, testing normality, nuisance parameters. Journal of the American Statistical Association, 86(415):785–791, September 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290413.

Blattberg:1991:SEP

[BG91b] Robert C. Blattberg and Edward I. George. Shrinkage estimation of price and promotional elasticities: Seemingly unrelated equations. Journal of the American Statistical Association, 86(414): 304–315, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290562.

Becker:1999:MBP

[BG99] Claudia Becker and Ursula Gather. The masking breakdown point of multivariate outlier identification rules. Journal of the American Statistical Association, 94(447):947–955, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670009.

Brockmann:1993:LAB

[BGH93] Michael Brockmann, Theo Gasser, and Eva Herrmann. Locally adaptive bandwidth choice for kernel regression estimators. Journal of the American Statistical Association, 88(424):1302–1309, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291270.

Brock:1994:BRB

[BGH94] Dwight B. Brock, Jack M. Guralnik, and Richard J. Havlik. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1148–1149, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290960.

Belin:1997:SDE

[BGH97] Thomas R. Belin, David W. Gjertson, and Ming Yi Hu. Summarizing DNA evidence when relatives are possible suspects. Journal of the American Statistical Association, 92(438):706–716, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2965719.

Barry:1993:BAC

[BH93a] Daniel Barry and J. A. Hartigan. A Bayesian analysis for change point problems. Journal of the American Statistical Association, 88(421):309–319, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290726.

Barry:1993:CMP

[BH93b] Daniel Barry and J. A. Hartigan. Choice models for predicting divisional winners in major league baseball. Journal of the American Statistical Association, 88(423):766–774, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290761.

Brown:1994:RRR

[BH94] B. M. Brown and T. P. Hettmansperger. Regular redescending rank estimates. Journal of the American Statistical Association, 89(426):538–542, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290856.

Brown:1996:NSN

[BH96] B. M. Brown and T. P. Hettmansperger. Normal scores, normal plots, and tests for normality. Journal of the American Statistical Association, 91(436):1668–1675, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291594.

Butler:1997:SNM

[BH97] Ronald W. Butler and Aparna V. Huzurbazar. Stochastic network models for survival analysis. Journal of the American Statistical Association, 92(437):246–257, March 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/march/
butler.htm;   http://www.jstor.org/stable/2291469.

Booth:1998:SEP

[BH98] James G. Booth and James P. Hobert. Standard errors of prediction in generalized linear mixed models. Journal of the American Statistical Association, 93(441):262–272, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/BOOTH.HTM;   http://www.jstor.org/stable/
2669622.

Bhat:1990:BRB

[Bha90] U. Narayan Bhat. Book review: [untitled]. Journal of the American Statistical Association, 85(412):1176–1177, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289632.

Bhapkar:1994:BRB

[Bha94] Vasant P. Bhapkar. Book review: [untitled]. Journal of the American Statistical Association, 89(426):720–721, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290891.

Begg:1996:SEP

[BHB96] Colin B. Begg, Ying Huang, and Marianne Berwick. Separate estimation of primary and secondary cancer preventative impact: Analysis of a case-control study of skin self-examination and melanoma. Journal of the American Statistical Association, 91 (436):1381–1387, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291564.

Bofinger:1993:CUC

[BHL93] Eve Bofinger, Anthony J. Hayter, and Wei Liu. The construction of upper confidence bounds on the range of several location parameters. Journal of the American Statistical Association, 88 (423):906–911, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290780.

Brown:1992:CBS

[BHNO92] B. M. Brown, T. P. Hettmansperger, J. Nyblom, and H. Oja. On certain bivariate sign tests and medians. Journal of the American Statistical Association, 87(417):127–135, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290460.

Benjamini:1998:CIM

[BHS98] Y. Benjamini, Y. Hochberg, and P. B. Stark. Confidence intervals with more power to determine the sign: Two ends constrain the means. Journal of the American Statistical Association, 93 (441):309–317, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/HOCHBERG.HTM;   http://
www.jstor.org/stable/2669627.

Bierens:1996:BRB

[Bie96] Herman J. Bierens. Book review: [untitled]. Journal of the American Statistical Association, 91(435):1373, September 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291755.

Biggerstaff:1998:BRB

[Big98] Brad J. Biggerstaff. Book review: [untitled]. Journal of the American Statistical Association, 93(441):400–401, March 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669640.

Billard:1991:BRB

[Bil91] Lynne Billard. Book review: [untitled]. Journal of the American Statistical Association, 86(416):1150–1151, December 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290550.

Billard:1996:BRB

[Bil96] Lynne Billard. Book review: [untitled]. Journal of the American Statistical Association, 91(433):430, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291423.

Billard:1997:VD

[Bil97] Lynne Billard. A voyage of discovery. Journal of the American Statistical Association, 92(437):1–12, March 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/Billard/
billard.html;   http://www.jstor.org/stable/2291444.

Binder:1996:CVE

[Bin96a] David A. Binder. Comment: On variance estimation with imputed survey data. Journal of the American Statistical Association, 91(434):499–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Binder:1996:VEI

[Bin96b] David A. Binder. On variance estimation with imputed survey data: Comment. Journal of the American Statistical Association, 91(434):510–512, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291639.

Binder:1998:ACS

[Bin98a] David A. Binder. Applications and case studies — comment: Not asked and not answered: Multiple imputation for multiple surveys. Journal of the American Statistical Association, 93(443): 858–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Binder:1998:AAM

[Bin98b] David A. Binder. Not asked and not answered: Multiple imputation for multiple surveys: Comment. Journal of the American Statistical Association, 93(443):858–859, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669820.

Bound:1995:PIV

[BJB95] John Bound, David A. Jaeger, and Regina M. Baker. Problems with instrumental variables estimation when the correlation between the instruments and the endogenous explanatory variable is weak. Journal of the American Statistical Association, 90(430): 443–450, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291055.

Bjornstad:1996:GLF

[Bjø96] Jan F. Bjørnstad. On the generalization of the likelihood function and the likelihood principle. Journal of the American Statistical Association, 91(434):791–806, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291674.

Buhlmann:1999:CPS

[BK99a] Peter Bühlmann and Hans R. Künsch. Comment: Prediction of spatial cumulative distribution functions using subsampling. Journal of the American Statistical Association, 94(445):97–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Buhlmann:1999:PSC

[BK99b] Peter Buhlmann and Hans R. Kunsch. Prediction of spatial cumulative distribution functions using subsampling: Comment. Journal of the American Statistical Association, 94(445):97–99, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669681.

Bethlehem:1990:DCM

[BKP90] Jelke G. Bethlehem, Wouter J. Keller, and Jeroen Pannekoek. Disclosure control of microdata. Journal of the American Statistical Association, 85(409):38–45, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289523.

Barrett:1992:GCI

[BL92] Bruce E. Barrett and Robert F. Ling. General classes of influence measures for multivariate regression. Journal of the American Statistical Association, 87(417):184–191, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290468.

Begg:1993:LEC

[BL93] Melissa Dowd Begg and Stephen Lagakos. Loss in efficiency caused by omitting covariates and misspecifying exposure in logistic regression models. Journal of the American Statistical Association, 88(421):166–170, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2290710.

Blattenberger:1992:BRB

[Bla92] Gail Blattenberger. Book review: [untitled]. Journal of the American Statistical Association, 87(417):252–253, March 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290482.

Bloomfield:1998:BRB

[Blo98] Peter Bloomfield. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1527, December 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670076.

Blyth:1993:PCR

[Bly93a] Colin R. Blyth. Is Pitman closeness a reasonable criterion?: Comment. Journal of the American Statistical Association, 88(421): 72–74, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290697.

Blyth:1993:OKW

[Bly93b] Stephen Blyth. Optimal kernel weights under a power criterion. Journal of the American Statistical Association, 88(424):1284– 1286, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291268.

Barlow:1992:FTR

[BM92] Richard E. Barlow and Max B. Mendel. De Finetti-type representations for life distributions. Journal of the American Statistical Association, 87(420):1116–1122, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290649.

Billard:1995:ESM

[BM95] L. Billard and M. R. Meshkani. Estimation of a stationary Markov chain. Journal of the American Statistical Association, 90(429): 307–315, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291155.

Berk:1996:DCD

[BM96] Robert Berk and Ruth Marcus. Dual cones, dual norms, and simultaneous inference for partially ordered means. Journal of the American Statistical Association, 91(433):318–328, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291410.

Berger:1999:DBF

[BM99] J. O. Berger and J. Mortera. Default Bayes factors for nonnested hypothesis testing. Journal of the American Statistical Association, 94(446):542–554, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2670175.

Boatwright:1999:ALM

[BMR99] Peter Boatwright, Robert McCulloch, and Peter Rossi. Accountlevel modeling for trade promotion: An application of a constrained parameter hierarchical model. Journal of the American Statistical Association, 94(448):1063–1073, December 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669920.

Bowman:1995:SMA

[BO95] Dale Bowman and E. Olusegun George. A saturated model for analyzing exchangeable binary data: Applications to clinical and developmental toxicity studies. Journal of the American Statistical Association, 90(431):871–879, September 1995. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2291321.

Bose:1997:RBS

[BO97] Smarajit Bose and Finbarr O’Sullivan. A region-based segmentation method for multichannel image data. Journal of the American Statistical Association, 92(437):92–106, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_97/
march/bose.htm;   http://www.jstor.org/stable/2291453.

Bockenholt:1999:AME

[Boc99] Ulf Bockenholt. Analyzing multiple emotions over time by autoregressive negative multinomial regression models. Journal of the American Statistical Association, 94(447):757–765, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669988.

Boik:1990:BRB

[Boi90] Robert J. Boik. Book review: [untitled]. Journal of the American Statistical Association, 85(409):259–260, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289562.

Bolstein:1991:BRB

[Bol91] Richard Bolstein. Book review: [untitled]. Journal of the American Statistical Association, 86(414):551–552, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290618.

Bolthauser:1994:BRB

[Bol94] Erwin Bolthauser. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1145–1146, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290955.

Bolstad:1995:MDP

[Bol95] William M. Bolstad. The multiprocess dynamic Poisson model. Journal of the American Statistical Association, 90(429):227– 232, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291147.

Booth:1994:BRB

[Boo94] James G. Booth. Book review: [untitled]. Journal of the American Statistical Association, 89(425):354–355, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291234.

Booth:1999:BRB

[Boo99] James G. Booth. Book review: [untitled]. Journal of the American Statistical Association, 94(445):343, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669714.

Baglivo:1992:MEG

[BOP92] Jenny Baglivo, Donald Olivier, and Marcello Pagano. Methods for exact goodness-of-fit tests. Journal of the American Statistical Association, 87(418):464–469, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290278.

Bove:1998:BRB

[Bov98] Frank J. Bove. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1236, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669870.

Bassett:1994:RS

[BP94a] Gilbert W. Bassett, Jr. and Joseph Persky. Rating skating. Journal of the American Statistical Association, 89(427):1075–1079, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290937.

Binder:1994:UEF

[BP94b] David A. Binder and Zdenek Patak. Use of estimating functions for estimation from complex surveys. Journal of the American Statistical Association, 89(427):1035–1043, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290931.

Berger:1996:IBF

[BP96] James O. Berger and Luis R. Pericchi. The intrinsic Bayes factor for model selection and prediction. Journal of the American Statistical Association, 91(433):109–122, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291387.

Balke:1997:BTE

[BP97] Alexander Balke and Judea Pearl. Bounds on treatment effects from studies with imperfect compliance. Journal of the American Statistical Association, 92(439):1171–1176, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_97/sept/balke.htm;   http://www.jstor.
org/stable/2965583.

Bajorski:1999:NTS

[BP99] Peter Bajorski and John Petkau. Nonparametric two-sample comparisons of changes on ordinal responses. Journal of the American Statistical Association, 94(447):970–978, September 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670011.

Bartoszynski:1997:MGF

[BPL97] Robert Bartoszynski, Dennis K. Pearl, and John Lawrence. A multidimensional goodness-of-fit test based on interpoint distances. Journal of the American Statistical Association, 92(438): 577–586, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2965706.

Bertolino:1995:MBF

[BPR95] Francesco Bertolino, Ludovico Piccinato, and Walter Racugno. Multiple Bayes factors for testing hypotheses. Journal of the American Statistical Association, 90(429):213–219, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291145.

Bertail:1999:SEU

[BPR99] Patrice Bertail, Dimitris N. Politis, and Joseph P. Romano. On subsampling estimators with unknown rate of convergence. Journal of the American Statistical Association, 94(446):569–579, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670177.

Brunner:1995:NMS

[BPS95] Edgar Brunner, Madan L. Puri, and Shan Sun. Nonparametric methods for stratified two-sample designs with application to multiclinic trials. Journal of the American Statistical Association, 90 (431):1004–1014, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291337.

Baglivo:1996:PDG

[BPS96] Jenny Baglivo, Marcello Pagano, and Cathie Spino. Permutation distributions via generating functions with applications to sensitivity analysis of discrete data. Journal of the American Statistical Association, 91(435):1037–1046, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291723.

Banfield:1992:IFI

[BR92] Jeffrey D. Banfield and Adrian E. Raftery. Ice floe identification in satellite images using mathematical morphology and clustering about principal curves. Journal of the American Statistical Association, 87(417):7–16, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290446.

Bandeen-Roche:1994:RAM

[BR94a] Karen Bandeen-Roche. Resolution of additive mixtures into source components and contributions: a compositional approach. Journal of the American Statistical Association, 89(428):1450– 1458, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291006.

Basu:1994:RMS

[BR94b] Sabyasachi Basu and Gregory C. Reinsel. Regression models with spatially correlated errors. Journal of the American Statistical Association, 89(425):88–99, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2291204.

Belin:1995:MCF

[BR95] Thomas R. Belin and Donald B. Rubin. A method for calibrating false-match rates in record linkage. Journal of the American Statistical Association, 90(430):694–707, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291082.

Brumback:1998:SSMb

[BR98a] Babette Brumback and John Rice. Smoothing spline models for the analysis of nested and crossed samples of curves: Rejoinder. Journal of the American Statistical Association, 93(443):991–994, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2669843.

Brumback:1998:RSS

[BR98b] Babette A. Brumback and John A. Rice. Rejoinder: Smoothing spline models for the analysis of nested and crossed samples of curves. Journal of the American Statistical Association, 93(443): 991–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brumback:1998:SSMa

[BR98c] Babette A. Brumback and John A. Rice. Smoothing spline models for the analysis of nested and crossed samples of curves. Journal of the American Statistical Association, 93(443):961– 976, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/sept/brumback.html;
 http://www.jstor.org/stable/2669837. With comments and

a rejoinder by the authors.

Byers:1998:NNC

[BR98d] Simon Byers and Adrian E. Raftery. Nearest neighbor clutter removal for estimating features in spatial point processes. Journal of the American Statistical Association, 93(442):577–584, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_98/june/BYERS.HTM;   http://www.jstor.
org/stable/2670109.

Brant:1990:CCR

[Bra90] Rollin Brant. Comparing classical and resistant outlier rules. Journal of the American Statistical Association, 85(412):1083– 1090, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289605.

Branford:1998:BRB

[Bra98] Alan J. Branford. Book review: [untitled]. Journal of the American Statistical Association, 93(441):404, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669646.

Breslow:1990:THO

[Bre90] Norman Breslow. Tests of hypotheses in overdispersed Poisson regression and other quasi-likelihood models. Journal of the American Statistical Association, 85(410):565–571, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289799.

Bremer:1991:IBR

[Bre91] Ronald H. Bremer. Interpreting blocks and random factors: Comment. Journal of the American Statistical Association, 86(415): 815–817, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290419.

Breiman:1992:LBO

[Bre92] Leo Breiman. The little bootstrap and other methods for dimensionality selection in regression: X-fixed prediction error. Journal of the American Statistical Association, 87(419):738–754, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290212.

Breslow:1996:SEC

[Bre96a] N. E. Breslow. Statistics in epidemiology: The case-control study. Journal of the American Statistical Association, 91(433): 14–28, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291379.

Breslow:1996:BRB

[Bre96b] Norman Breslow. Book review: [untitled]. Journal of the American Statistical Association, 91(434):908–909, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291687.

Brehm:1998:ACS

[Bre98a] John Brehm. Applications and case studies — comment: Not asked and not answered: Multiple imputation for multiple surveys. Journal of the American Statistical Association, 93(443): 859–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brehm:1998:AAM

[Bre98b] John Brehm. Not asked and not answered: Multiple imputation for multiple surveys: Comment. Journal of the American Statistical Association, 93(443):859–860, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669821.

Brillinger:1991:SIR

[Bri91a] David R. Brillinger. Sliced inverse regression for dimension reduction: Comment. Journal of the American Statistical Association, 86(414):333, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290565.

Brillinger:1991:WFA

[Bri91b] David R. Brillinger. Walsh–Fourier analysis and its statistical applications: Comment. Journal of the American Statistical Association, 86(414):481–482, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290597.

Brillinger:1992:NCS

[Bri92] David R. Brillinger. Nerve cell spike train data analysis: a progression of technique. Journal of the American Statistical Association, 87(418):260–271, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290256.

Briggs:1998:BRB

[Bri98] William L. Briggs. Book review: [untitled]. Journal of the American Statistical Association, 93(441):410, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669656.

Briggs:1999:BRB

[Bri99] William M. Briggs. Book review: [untitled]. Journal of the American Statistical Association, 94(445):345–346, March 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669717.

Berry:1999:BDEa

[BRL99a] Scott M. Berry, C. Shane Reese, and Patrick D. Larkey. Bridging different eras in sports. Journal of the American Statistical Association, 94(447):661–676, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669973.

Berry:1999:BDEb

[BRL99b] Scott M. Berry, C. Shane Reese, and Patrick D. Larkey. Bridging different eras in sports: Rejoinder. Journal of the American Statistical Association, 94(447):684–686, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669977.

Bandeen-Roche:1997:LVR

[BRMZR97] Karen Bandeen-Roche, Diana L. Miglioretti, Scott L. Zeger, and Paul J. Rathouz. Latent variable regression for multiple discrete outcomes. Journal of the American Statistical Association, 92(440):1375–1386, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/december/BANDEEN.HTM;   http://www.jstor.
org/stable/2965407.

Bromberg:1990:BRB

[Bro90] Judith Bromberg. Book review: [untitled]. Journal of the American Statistical Association, 85(411):912–913, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290053.

Brown:1991:BRB

[Bro91] Byron William Brown, Jr. Book review: [untitled]. Journal of the American Statistical Association, 86(415):832, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290431.

Brody:1992:CMD

[Bro92a] Janet Brody. Comment on Maxwell and Delaney. Journal of the American Statistical Association, 87(417):57–??, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Bro92b, Wei92].

Brody:1992:EHU

[Bro92b] Janet Brody. The effects of helmet use on the severity of head injuries in motorcycle accidents: Comment on Maxwell and Delaney. Journal of the American Statistical Association, 87(417):

57, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2290451. See [Wei92, Bro92a].

Brown:1992:BRB

[Bro92c] Morton B. Brown. Book review: [untitled]. Journal of the American Statistical Association, 87(420):1239–1240, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290667.

Brown:1992:SIA

[Bro92d] Byron W. Brown, Jr. Statistical issues arising in AIDS clinical trials: Comment. Journal of the American Statistical Association, 87(418):569–571, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290292.

Brostrom:1997:MAC

[Bro97] Goran Brostrom. A martingale approach to the changepoint problem. Journal of the American Statistical Association, 92(439): 1177–1183, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/sept/brostrom.htm;
 http://www.jstor.org/stable/2965584.

Broffitt:1999:BRB

[Bro99] James D. Broffitt. Book review: [untitled]. Journal of the American Statistical Association, 94(445):341–343, March 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669713.

Babu:1992:NES

[BRR92] Gutti Jogesh Babu, C. Radhakrishna Rao, and M. Bhaskara Rao. Nonparametric estimation of specific occurrence/exposure rate in risk and survival analysis. Journal of the American Statistical Association, 87(417):84–89, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290455.

Brumback:1999:VSF

[BRW99] Babette A. Brumback, David Ruppert, and M. P. Wand. Variable selection and function estimation in additive nonparametric regression using a data-based prior: Comment. Journal of the American Statistical Association, 94(447):794–797, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669991.

Bryson:1997:BRB

[Bry97] Maurice C. Bryson. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1222, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965600.

Bradley:1991:IAD

[BS91a] Ralph A. Bradley and Frances P. Stewart. Intrablock analysis of designs with multiple blocking criteria. Journal of the American Statistical Association, 86(415):792–797, September 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290414.

Butler:1991:BRB

[BS91b] Michael D. Butler and Jolayne Service. Book review: [untitled]. Journal of the American Statistical Association, 86(414):543–544, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290607.

Berger:1993:BAP

[BS93] James O. Berger and Dongchu Sun. Bayesian analysis for the Poly-Weibull distribution. Journal of the American Statistical Association, 88(424):1412–1418, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291285.

Brooks:1994:JCR

[BS94] Rodney Brooks and Mervyn Stone. Joint continuum regression for multiple predictands. Journal of the American Statistical Association, 89(428):1374–1377, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290999.

Benjamini:1996:NSC

[BS96a] Yoav Benjamini and Philip B. Stark. Nonequivariant simultaneous confidence intervals less likely to contain zero. Journal of the American Statistical Association, 91(433):329–337, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291411.

Bolviken:1996:CIM

[BS96b] Erik Bølviken and Eva Skovlund. Confidence intervals from Monte Carlo tests. Journal of the American Statistical Association, 91 (435):1071–1078, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291726.

Buzas:1996:IVE

[BS96c] Jeffrey S. Buzas and Leonard A. Stefanski. Instrumental variable estimation in generalized linear measurement error models. Journal of the American Statistical Association, 91(435):999–1006, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291719.

Berger:1997:ITS

[BS97a] Vance Berger and Harold Sackrowitz. Improving tests for superior treatment in contingency tables. Journal of the American Statistical Association, 92(438):700–705, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965718.

Bloch:1997:MDF

[BS97b] Daniel A. Bloch and Bernard W. Silverman. Monotone discriminant functions and their applications in rheumatology. Journal of the American Statistical Association, 92(437):144–153, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_97/march/bloch.htm;   http://www.jstor.
org/stable/2291458.

Blyth:1997:HEA

[BS97c] Colin R. Blyth and Robert G. Staudte. Hypothesis estimates and acceptability profiles for 2 × 2 contingency tables. Journal of the American Statistical Association, 92(438):694–699, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2965717.

Butler:1998:SAM

[BS98] Ronald W. Butler and Richard K. Sutton. Saddlepoint approximation for multivariate cumulative distribution functions and probability computations in sampling theory and outlier testing.

Journal of the American Statistical Association, 93(442):596–604, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_98/june/BUTLER.HTM;   http://www.jstor.
org/stable/2670111.

Begg:1998:NSE

[BSB98] Colin B. Begg, Jaya M. Satagopan, and Marianne Berwick. A new strategy for evaluating the impact of epidemiologic risk factors for cancer with application to melanoma. Journal of the American Statistical Association, 93(442):415–426, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
june/BEGG.HTM;   http://www.jstor.org/stable/2670089.

Bailey:1998:SDU

[BSPK98] Trevor C. Bailey, Theofanis Sapatinas, Kenneth J. Powell, and Wojtek J. Krzanowski. Signal detection in underwater sound using wavelets. Journal of the American Statistical Association, 93(441):73–83, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/BAILEY.HTM;   http:
//www.jstor.org/stable/2669604.

Buckland:1995:CID

[Buc95a] Stephen T. Buckland. Comment: Inference from a deterministic population dynamics model for bowhead whales. Journal of the American Statistical Association, 90(430):417–??, ???? 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Buckland:1995:IDP

[Buc95b] Stephen T. Buckland. Inference from a deterministic population dynamics model for bowhead whales: Comment. Journal of the American Statistical Association, 90(430):417–420, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291049.

Buckle:1995:BIS

[Buc95c] D. J. Buckle. Bayesian inference for stable distributions. Journal of the American Statistical Association, 90(430):605–613, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291072.

Budescu:1996:BRB

[Bud96] David V. Budescu. Book review: [untitled]. Journal of the American Statistical Association, 91(436):1753–1754, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291613.

Bunge:1996:BRB

[Bun96] John Bunge. Book review: [untitled]. Journal of the American Statistical Association, 91(435):1372–1373, September 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291754.

Buonaccorsi:1990:DSE

[Buo90] John P. Buonaccorsi. Double sampling for exact values in some multivariate measurement error problems. Journal of the American Statistical Association, 85(412):1075–1082, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289604.

Buonaccorsi:1996:MER

[Buo96] John P. Buonaccorsi. Measurement error in the response in the general linear model. Journal of the American Statistical Association, 91(434):633–642, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291659.

Burchinal:1992:BRB

[Bur92] Margaret Burchinal. Book review: [untitled]. Journal of the American Statistical Association, 87(418):588–589, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290305.

Burr:1994:CCB

[Bur94] Deborah Burr. A comparison of certain bootstrap confidence intervals in the Cox model. Journal of the American Statistical Association, 89(428):1290–1302, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290992.

Buzas:1997:BRB

[Buz97] Jeffrey Buzas. Book review: [untitled]. Journal of the American Statistical Association, 92(437):385–386, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291487.

Beirlant:1996:TIE

[BVT96] Jan Beirlant, Petra Vynckier, and Jozef L. Teugels. Tail index estimation, Pareto quantile plots, and regression diagnostics. Journal of the American Statistical Association, 91(436):1659–1667, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291593.

Bohn:1992:NTS

[BW92] Lora L. Bohn and Douglas A. Wolfe. Nonparametric two-sample procedures for ranked-set samples data. Journal of the American Statistical Association, 87(418):552–561, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290290.

Bohn:1994:EIJ

[BW94] Lora L. Bohn and Douglas A. Wolfe. The effect of imperfect judgment rankings on properties of procedures based on the ranked-set samples analog of the Mann–Whitney–Wilcoxon statistic. Journal of the American Statistical Association, 89(425):168–176, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291213.

Bradlow:1998:ACS

[BWC98a] Eric T. Bradlow, Robert E. Weiss, and Meehyung Cho. Applications and case studies — Bayesian identification of outliers in computerized adaptive tests. Journal of the American Statistical Association, 93(443):910–919, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/sept/bradlow.html.

Bradlow:1998:BIO

[BWC98b] Eric T. Bradlow, Robert E. Weiss, and Meehyung Cho. Bayesian identification of outliers in computerized adaptive tests. Journal of the American Statistical Association, 93(443):910–919, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669829.

Bewley:1995:TCB

[BY95] Ronald Bewley and Minxian Yang. Tests for cointegration based on canonical correlation analysis. Journal of the American Statistical Association, 90(431):990–996, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291335.

Bickel:1992:VSN

[BZ92] Peter Bickel and Ping Zhang. Variable selection in nonparametric regression with categorical covariates. Journal of the American Statistical Association, 87(417):90–97, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290456.

Bradlow:1999:HLV

[BZ99] Eric T. Bradlow and Alan M. Zaslavsky. A hierarchical latent variable model for ordinal data from a customer satisfaction survey with “no answer” responses. Journal of the American Statistical Association, 94(445):43–52, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669676.

Carlin:1990:BRB

[Car90a] Bradley P. Carlin. Book review: [untitled]. Journal of the American Statistical Association, 85(412):1173–1174, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289628.

Carlson:1990:BRB

[Car90b] Roger Carlson. Book review: [untitled]. Journal of the American Statistical Association, 85(410):606, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289828.

Carlin:1994:BRB

[Car94a] Bradley P. Carlin. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1142–1143, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290952.

Carriere:1994:IRM

[Car94b] Keumhee Chough Carriere. Incomplete repeated measures data analysis in the presence of treatment effects. Journal of the American Statistical Association, 89(426):680–686, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290871.

Clutton-Brock:1990:DEU

[CB90] M. Clutton-Brock. Density estimation using exponentials of orthogonal series. Journal of the American Statistical Association, 85(411):760–764, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290012.

Casella:1994:ESB

[CB94] George Casella and Roger L. Berger. Estimation with selected binomial information or do you really believe that Dave Windfield is batting .471 ? Journal of the American Statistical Association, 89(427):1080–1090, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290938.

Christensen:1997:TIA

[CB97] Ronald Christensen and Edward J. Bedrick. Testing the independence assumption in linear models. Journal of the American Statistical Association, 92(439):1006–1016, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/sept/christensen.htm;   http://www.jstor.
org/stable/2965565.

Clyde:1996:ECW

[CC96a] Merlise Clyde and Kathryn Chaloner. The equivalence of constrained and weighted designs in multiple objective design problems. Journal of the American Statistical Association, 91(435): 1236–1244, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291742.

Cowles:1996:MCM

[CC96b] Mary Kathryn Cowles and Bradley P. Carlin. Markov chain Monte Carlo convergence diagnostics: a comparative review. Journal of the American Statistical Association, 91(434):883–904, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291683.

Cutler:1996:MHD

[CCB96] Adele Cutler and Olga I. Cordero-Braña. Minimum Hellinger distance estimation for finite mixture models. Journal of the American Statistical Association, 91(436):1716–1723, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291601.

Cowles:1996:BTM

[CCC96] Mary Kathryn Cowles, Bradley P. Carlin, and John E. Connett. Bayesian tobit modeling of longitudinal ordinal clinical trial compliance data with nonignorable missingness. Journal of the American Statistical Association, 91(433):86–98, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291385.

Chen:1997:NMB

[CCD97] Zhao-Guo Chen, Pierre A. Cholette, and Estela Bee Dagum. A nonparametric method for benchmarking survey data via signal extraction. Journal of the American Statistical Association, 92(440):1563–1571, December 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_97/
december/CHEN_ZG.HTM;   http://www.jstor.org/stable/2965427.

Cook:1998:PRP

[CCD98] R. Dennis Cook and Rodney Croos-Dabrera. Partial residual plots in generalized linear models. Journal of the American Statistical Association, 93(442):730–739, June 1998. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
june/COOK.HTM;   http://www.jstor.org/stable/2670123.

Clogg:1990:ER

[CCG90] Clifford C. Clogg, Raymond J. Carroll, and Donald Guthrie. Editors’ report for 1989. Journal of the American Statistical Association, 85(410):273, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289761.

Carroll:1997:OEPa

[CCG^{+}97a] R. J. Carroll, R. Chen, E. I. George, T. H. Li, H. J. Newton, H. Schmiediche, and N. Wang. Ozone exposure and population density in Harris County, Texas. Journal of the American Statistical Association, 92(438):392–404, June 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965684. See [GMS97b,

GMS97a, SF97a, SF97b, Cre97b, Cre97a, CCG^{+}97c, CCG^{+}97b].

Carroll:1997:ROE

[CCG^{+}97b] R. J. Carroll, R. Chen, E. I. George, T. H. Li, H. J. Newton, H. Schmiediche, and N. Wang. Rejoinder: Ozone exposure and population density in Harris County, Texas. Journal of the American Statistical Association, 92(438):414–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CCG^{+}97a].

Carroll:1997:OEPb

[CCG^{+}97c] R. J. Carroll, R. Chen, E. I. George, T. H. Li, H. J. Newton, and N. Whang. Ozone exposure and population density in Harris County, Texas: Rejoinder. Journal of the American Statistical Association, 92(438):414–415, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965688. See [CCG^{+}97a, Cre97a].

Clogg:1991:ER

[CCWG91] Clifford C. Clogg, Raymond J. Carroll, Edward J. Wegman, and Donald Guthrie. Editors’ report for 1990. Journal of the American Statistical Association, 86(414):257, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290556.

Calvin:1991:LSE

[CD91] James A. Calvin and Richard L. Dykstra. Least squares estimation of covariance matrices in balanced multivariate variance components models. Journal of the American Statistical Association, 86(414):388–395, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290582.

Chhikara:1992:EUM

[CD92] Raj S. Chhikara and Lih-Yuan Deng. Estimation using multiyear rotation design sampling in agricultural surveys. Journal of the American Statistical Association, 87(420):924–932, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290628.

Calvin:1995:REC

[CD95] James A. Calvin and Richard L. Dykstra. REML estimation of covariance matrices with restricted parameter spaces. Journal of the American Statistical Association, 90(429):321–329, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291157.

Cragg:1996:APL

[CD96] John G. Cragg and Stephen G. Donald. On the asymptotic properties of LDU-based tests of the rank of a matrix. Journal of the American Statistical Association, 91(435):1301–1309, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291748.

Clyde:1996:POM

[CDP96] Merlise Clyde, Heather Desimone, and Giovanni Parmigiani. Prediction via orthogonalized model mixing. Journal of the American Statistical Association, 91(435):1197–1208, September 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291738.

Chen:1999:NSL

[CDS99] Ming-Hui Chen, Dipak K. Dey, and Qi-Man Shao. A new skewed link model for dichotomous quantal response data. Journal of the American Statistical Association, 94(448):1172–1186, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669933.

Chambers:1993:BRE

[CDW93] Raymond L. Chambers, Alan H. Dorfman, and Thomas E. Wehrly. Bias robust estimation in finite populations using nonparametric calibration. Journal of the American Statistical Association, 88(421):268–277, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2290722.

Cohen:1995:THU

[CF95] Arthur Cohen and Mendel Fygenson. Testing homogeneity of uniform scale distributions against two- sided and one-sided alternatives. Journal of the American Statistical Association, 90(431): 1062–1067, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291343.

Carroll:1997:GPL

[CFGW97] R. J. Carroll, Jianqing Fan, Irene Gijbels, and M. P. Wand. Generalized partially linear single-index models. Journal of the American Statistical Association, 92(438):477–489, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965697.

Carlin:1990:AEB

[CG90] Bradley P. Carlin and Alan E. Gelfand. Approaches for empirical Bayes confidence intervals. Journal of the American Statistical Association, 85(409):105–114, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289531.

Clapp:1992:EPI

[CG92] John M. Clapp and Carmelo Giaccotto. Estimating price indices for residential property: a comparison of repeat sales and assessed value methods. Journal of the American Statistical Association, 87(418):300–306, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290260.

Chen:1997:TLV

[CG97] Jie Chen and A. K. Gupta. Testing and locating variance changepoints with application to stock prices. Journal of the American Statistical Association, 92(438):739–747, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965722.

Chu:1998:EPSa

[CGGM98a] C. K. Chu, I. K. Glad, F. Godtliebsen, and J. S. Marron. Edge preserving smoothers for image processing. Journal of the American Statistical Association, 93(442):526–541, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
june/CHU.HTM;   http://www.jstor.org/stable/2670100. With

comments and a rejoinder by the authors.

Chu:1998:EPSb

[CGGM98b] C. K. Chu, I. K. Glad, F. Godtliebsen, and J. S. Marron. Edgepreserving smoothers for image processing: Rejoinder. Journal of the American Statistical Association, 93(442):553–556, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670104.

Chu:1998:REP

[CGGM98c] C. K. Chu, I. K. Glad, F. Godtliebsen, and J. S. Marron. Rejoinder: Edge preserving smoothers for image processing. Journal of the American Statistical Association, 93(442):553–??, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:1993:CCS

[CGL93] R. J. Carroll, M. H. Gail, and J. H. Lubin. Case-control studies with errors in covariates. Journal of the American Statistical Association, 88(421):185–199, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290713.

Chipman:1998:BCMa

[CGM98a] Hugh A. Chipman, Edward I. George, and Robert E. McCulloch. Bayesian CART model search. Journal of the American Statistical Association, 93(443):935–947, September 1998. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_98/sept/
chipman.html;   http://www.jstor.org/stable/2669832.

Chipman:1998:BCMb

[CGM98b] Hugh A. Chipman, Edward I. George, and Robert E. McCulloch. Bayesian CART model search: Rejoinder. Journal of the American Statistical Association, 93(443):957–960, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669836.

Chipman:1998:RBC

[CGM98c] Hugh A. Chipman, Edward I. George, and Robert E. McCulloch. Rejoinder: Bayesian CART model search. Journal of the American Statistical Association, 93(443):957–??, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cook:1990:UMO

[CH90] R. Dennis Cook and Douglas M. Hawkins. Unmasking multivariate outliers and leverage points: Comment. Journal of the American Statistical Association, 85(411):640–644, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289996.

Cameron:1991:RCV

[CH91a] Trudy Ann Cameron and Daniel D. Huppert. Referendum contingent valuation estimates: Sensitivity to the assignment of offered values. Journal of the American Statistical Association, 86(416): 910–918, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290504.

Casella:1991:PIA

[CH91b] George Casella and Jiunn Tzon Hwang. On the problem of interactions in the analysis of variance: Comment. Journal of the American Statistical Association, 86(414):372–373, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290579.

Cressie:1992:SSA

[CH92] Noel Cressie and Frederick L. Hulting. A spatial statistical analysis of tumor growth. Journal of the American Statistical Association, 87(418):272–283, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290257.

Clogg:1993:EPC

[CH93a] Clifford C. Clogg and Christing L. Himes. Estimation of population coverage in the 1990 United States census based on demographic analysis: Comment: Uncertainty in demographic analysis. Journal of the American Statistical Association, 88(423): 1072–1074, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290800.

Coakley:1993:BIH

[CH93b] Clint W. Coakley and Thomas P. Hettmansperger. A bounded influence, high breakdown, efficient regression estimator. Journal of the American Statistical Association, 88(423):872–880, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290776.

Craig:1994:TTD

[CH94] Lee A. Craig and Alastair R. Hall. Trying out for the team: Do exhibitions matter? evidence from the National Football League. Journal of the American Statistical Association, 89(427): 1091–1099, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290939.

Castillo:1995:MLD

[CH95a] Enrique Castillo and Ali S. Hadi. Modeling lifetime data with application to fatigue models. Journal of the American Statistical Association, 90(431):1041–1054, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291341.

Clayton:1995:CIA

[CH95b] Murray K. Clayton and Brian D. Hudelson. Confidence intervals for autocorrelations based on cyclic samples. Journal of the American Statistical Association, 90(430):753–757, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291088.

Castillo:1997:FGP

[CH97a] Enrique Castillo and Ali S. Hadi. Fitting the generalized Pareto distribution to data. Journal of the American Statistical Association, 92(440):1609–1620, December 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_97/
december/CASTILLO.HTM;   http://www.jstor.org/stable/
2965432.

Chang:1997:ILE

[CH97b] Yi-Ping Chang and Wen-Tao Huang. Inferences for the linear errors-in-variables with changepoint models. Journal of the American Statistical Association, 92(437):171–178, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/march/chang.htm;   http://www.jstor.org/
stable/2291461.

Cressie:1999:CNS

[CH99] Noel Cressie and Hsin-Cheng Huang. Classes of nonseparable, spatio-temporal stationary covariance functions. Journal of the American Statistical Association, 94(448):1330–1340, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669946.

Chan:1990:INI

[Cha90] Ngai Hang Chan. Inference for near-integrated time series with infinite variance. Journal of the American Statistical Association, 85 (412):1069–1074, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289603.

Chang:1991:WCE

[Cha91a] Myron N. Chang. On weak convergence of an estimator of the survival function when new is better than used of a specified age. Journal of the American Statistical Association, 86(413): 173–178, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289728.

Chang:1991:BRB

[Cha91b] Potter C. Chang. Book review: [untitled]. Journal of the American Statistical Association, 86(416):1149, December 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290548.

Chase:1991:BRB

[Cha91c] Gary A. Chase. Book review: [untitled]. Journal of the American Statistical Association, 86(413):248, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289745.

Chaloner:1992:BRB

[Cha92] Kathryn Chaloner. Book review: [untitled]. Journal of the American Statistical Association, 87(419):899, September 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290234.

Chan:1993:ABG

[Cha93a] K. S. Chan. Asymptotic behavior of the Gibbs sampler. Journal of the American Statistical Association, 88(421):320–326, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290727.

Chang:1993:BRB

[Cha93b] Potter C. Chang. Book review: [untitled]. Journal of the American Statistical Association, 88(422):706–707, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290370.

Chakraborti:1996:BRB

[Cha96a] Subha Chakraborti. Book review: [untitled]. Journal of the American Statistical Association, 91(436):1750, December 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291606.

Chaudhuri:1996:GNQ

[Cha96b] Probal Chaudhuri. On a geometric notion of quantiles for multivariate data. Journal of the American Statistical Association, 91(434):862–872, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291681.

Chen:1990:AAI

[Che90] Han Feng Chen. The accuracy of approximate intervals for a binomial parameter. Journal of the American Statistical Association, 85(410):514–518, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289792.

Chen:1991:BRB

[Che91] Chung Chen. Book review: [untitled]. Journal of the American Statistical Association, 86(414):552–553, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290620.

Chen:1992:BRBb

[Che92a] Chung Chen. Book review: [untitled]. Journal of the American Statistical Association, 87(417):256, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290487.

Chen:1992:BRBa

[Che92b] Rong Chen. Book review: [untitled]. Journal of the American Statistical Association, 87(417):254–255, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290485.

Chen:1994:IWM

[Che94a] Ming-Hui Chen. Importance-weighted marginal Bayesian posterior density estimation. Journal of the American Statistical Association, 89(427):818–824, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290907.

Cheng:1994:NEM

[Che94b] Philip E. Cheng. Nonparametric estimation of mean functionals with data missing at random. Journal of the American Statistical Association, 89(425):81–87, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291203.

Chen:1995:TMS

[Che95a] Ling Chen. Testing the mean of skewed distributions. Journal of the American Statistical Association, 90(430):767–772, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291090.

Chen:1995:BRB

[Che95b] Pinyuen Chen. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1130–1131, September 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291360.

Chen:1998:BRB

[Che98] Pinyuen Chen. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1239–1240, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669874.

Chiswick:1990:BRB

[Chi90] Barry R. Chiswick. Book review: [untitled]. Journal of the American Statistical Association, 85(411):903–904, September 1990.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290038.

Chinchilli:1991:BRB

[Chi91] Vernon M. Chinchilli. Book review: [untitled]. Journal of the American Statistical Association, 86(413):249–250, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289747.

Chib:1995:MLG

[Chi95] Siddhartha Chib. Marginal likelihood from the Gibbs output. Journal of the American Statistical Association, 90(432):1313– 1321, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291521.

Chong:1992:DNR

[Cho92] Gu Chong. Diagnostics for nonparametric regression models with additive terms. Journal of the American Statistical Association, 87(420):1051–??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chong:1993:SSD

[Cho93] Gu Chong. Smoothing spline density estimation: a dimensionless automatic algorithm. Journal of the American Statistical Association, 88(422):495–??, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cowling:1996:BCR

[CHP96] Ann Cowling, Peter Hall, and Michael J. Phillips. Bootstrap confidence regions for the intensity of a Poisson point process. Journal of the American Statistical Association, 91(436):1516–1524, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291577.

Christensen:1991:SSC

[Chr91] Ronald Christensen. Small-sample characterizations of near replicate lack-of-fit tests. Journal of the American Statistical Association, 86(415):752–756, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290408.

Carroll:1994:ELM

[CHR94] R. J. Carroll, Peter Hall, and D. Ruppert. Estimation of lag in misregistered, averaged images. Journal of the American Statistical Association, 89(425):219–229, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291218.

Christensen:1999:CPS

[Chr99a] Ronald Christensen. Comment: Prediction of spatial cumulative distribution functions using subsampling. Journal of the American Statistical Association, 94(445):99–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Christensen:1999:PSC

[Chr99b] Ronald Christensen. Prediction of spatial cumulative distribution functions using subsampling: Comment. Journal of the American Statistical Association, 94(445):99–100, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669682.

Cook:1992:CMM

[CHW92] R. D. Cook, D. M. Hawkins, and S. Weisberg. Comparison of model misspecification diagnostics using residuals from least mean of squares and least median of squares fits. Journal of the American Statistical Association, 87(418):419–424, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290272.

Chen:1999:NBM

[CIS99] Ming-Hui Chen, Joseph G. Ibrahim, and Debajyoti Sinha. A new Bayesian model for survival data with a surviving fraction. Journal of the American Statistical Association, 94(447):909–919, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2670006.

Chen:1995:DLC

[CJ95a] Jian Shen Chen and Robert I. Jennrich. Diagnostics for linearization confidence intervals in nonlinear regression. Journal of the American Statistical Association, 90(431):1068–1074, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291344.

Chen:1995:TIL

[CJ95b] Jian-Shen Chen and Robert I. Jennrich. Transformations for improving linearization confidence intervals in nonlinear regression. Journal of the American Statistical Association, 90(432): 1271–1276, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291517.

Chen:1996:SRD

[CJ96] Jian-Shen Chen and Robert I. Jennrich. The signed root deviance profile and confidence intervals in maximum likelihood analysis. Journal of the American Statistical Association, 91(435):993–998, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291718.

Cohen:1999:CPM

[CJ99] Jon D. Cohen and Tao Jiang. Comparison of partially measured latent traits across nominal subgroups. Journal of the American Statistical Association, 94(448):1035–1044, December 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669917.

Carlstein:1992:BE

[CK92] E. Carlstein and C. Krishnamoorthy. Boundary estimation. Journal of the American Statistical Association, 87(418):430–438, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290274.

Carlin:1992:PWM

[CKLH92] Bradley P. Carlin, Robert E. Kass, F. Javier Lerch, and Brian R. Huguenard. Predicting working memory failure: a subjective Bayesian approach to model selection. Journal of the American Statistical Association, 87(418):319–327, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290262.

Carroll:1996:ASE

[CKLS96] Raymond J. Carroll, Helmut Küchenhoff, F. Lombard, and Leonard A. Stefanski. Asymptotics for the SIMEX estimator in nonlinear measurement error models. Journal of the American Statistical Association, 91(433):242–250, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291401.

Chipman:1997:ABW

[CKM97] Hugh A. Chipman, Eric D. Kolaczyk, and Robert E. McCulloch. Adaptive Bayesian wavelet shrinkage. Journal of the American Statistical Association, 92(440):1413–1421, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/december/CHIPMAN.HTM;   http://www.jstor.
org/stable/2965411.

Carroll:1992:MER

[CL92a] Raymond J. Carroll and Ker-Chau Li. Measurement error regression with unknown link: Dimension reduction and data visualization. Journal of the American Statistical Association, 87(420): 1040–1050, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290641.

Chao:1992:ENC

[CL92b] Anne Chao and Shen-Ming Lee. Estimating the number of classes via sample coverage. Journal of the American Statistical Association, 87(417):210–217, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2290471.

Chen:1993:JEM

[CL93] Chung Chen and Lon-Mu Liu. Joint estimation of model parameters and outlier effects in time series. Journal of the American Statistical Association, 88(421):284–297, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290724.

Crowley:1994:BRB

[CL94] John Crowley and Ping-Yu Liu. Book review: [untitled]. Journal of the American Statistical Association, 89(425):356, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291236.

Chan:1995:MCE

[CL95] K. S. Chan and Johannes Ledolter. Monte Carlo EM estimation for time series models involving counts. Journal of the American Statistical Association, 90(429):242–252, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291149.

Chen:1999:PHR

[CL99a] Hua Yun Chen and Roderick J. A. Little. Proportional hazards regression with missing covariates. Journal of the American Statistical Association, 94(447):896–908, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670005.

Cook:1999:DRB

[CL99b] R. Dennis Cook and Hakbae Lee. Dimension reduction in binary response regression. Journal of the American Statistical Association, 94(448):1187–1200, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669934.

Clark:1990:BRBb

[Cla90a] Cynthia Z. F. Clark. Book review: [untitled]. Journal of the American Statistical Association, 85(411):905, September 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290040.

Clark:1990:BRBa

[Cla90b] Virginia Clark. Book review: [untitled]. Journal of the American Statistical Association, 85(410):600–601, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289821.

Clason:1992:BRB

[Cla92] Dennis L. Clason. Book review: [untitled]. Journal of the American Statistical Association, 87(419):907, September 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290245.

Clayton:1993:BRB

[Cla93] Murray K. Clayton. Book review: [untitled]. Journal of the American Statistical Association, 88(422):703, June 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290365.

Clarke:1996:IRP

[Cla96] Bertrand Clarke. Implications of reference priors for prior information and for sample size. Journal of the American Statistical Association, 91(433):173–184, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291393.

Clason:1998:BRB

[Cla98] Dennis L. Clason. Book review: [untitled]. Journal of the American Statistical Association, 93(441):407, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669650.

Clayton:1999:BRB

[Cla99] Murray K. Clayton. Book review: [untitled]. Journal of the American Statistical Association, 94(448):1383, December 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669950.

Cleary:1999:BRB

[Cle99] Richard Cleary. Book review: [untitled]. Journal of the American Statistical Association, 94(446):655–656, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670198.

Cliff:1999:BRB

[Cli99] Andrew D. Cliff. Book review: [untitled]. Journal of the American Statistical Association, 94(446):654, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670195.

Chung:1993:JEM

[CLM93] Chen Chung and Liu Lon-Mu. Joint estimation of model parameters and outlier effects in time series. Journal of the American Statistical Association, 88(421):284–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Clogg:1994:BRB

[Clo94] Clifford C. Clogg. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1141, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290950.

Chiu:1996:MLC

[CLT96] Tom Y. M. Chiu, Tom Leonard, and Kam-Wah Tsui. The matrixlogarithmic covariance model. Journal of the American Statistical Association, 91(433):198–210, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291396.

Clogg:1992:ER

[CLWG92] Clifford C. Clogg, Roderick J. A. Little, Edward J. Wegman, and Donald Guthrie. Editors’ report for 1991. Journal of the American Statistical Association, 87(418):259, June 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290255.

Cole:1995:EBM

[CLWZ95] Bernard F. Cole, Mei-Ling T. Lee, G. Alex Whitmore, and Alan M. Zaslavsky. An empirical Bayes model for Markovdependent binary sequences with randomly missing observations. Journal of the American Statistical Association, 90(432):1364– 1372, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291527.

Chung:1991:UNM

[CM91a] Lyinn Chung and John I. Marden. Use of nonnull models for rank statistics in bivariate, two-sample, and analysis of variance problems. Journal of the American Statistical Association, 86 (413):188–200, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289730.

Corder:1991:NSH

[CM91b] Larry S. Corder and Kenneth G. Manton. National surveys and the health and functioning of the elderly: The effects of design and content. Journal of the American Statistical Association, 86 (414):513–525, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290603.

Chaudhuri:1993:NEO

[CM93] Probal Chaudhuri and Per A. Mykland. Nonlinear experiments: Optimal design and inference based on likelihood. Journal of the American Statistical Association, 88(422):538–546, June 1993.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290334.

Capra:1997:ATM

[CM97a] William B. Capra and Hans-Georg Müller. An accelerated-time model for response curves. Journal of the American Statistical Association, 92(437):72–83, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

amstat.org/publications/jasa/abstracts_97/march/capra.
htm;   http://www.jstor.org/stable/2291451.

Choi:1997:AMR

[CM97b] K. Choi and J. Marden. An approach to multivariate rank tests in multivariate analysis of variance. Journal of the American Statistical Association, 92(440):1581–1590, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/december/CHOI_K.HTM;   http://www.jstor.
org/stable/2965429.

Christiansen:1997:HPR

[CM97c] Cindy L. Christiansen and Carl N. Morris. Hierarchical Poisson regression modeling. Journal of the American Statistical Association, 92(438):618–632, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2965709.

Chiou:1998:QLR

[CM98] Jeng-Min Chiou and Hans-Georg Müller. Quasi-likelihood regression with unknown link and variance functions. Journal of the American Statistical Association, 93(444):1376–1387, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_98/dec/chiou.html;   http://www.jstor.
org/stable/2670053.

Chaudhuri:1999:SES

[CM99] Probal Chaudhuri and J. S. Marron. SiZer for exploration of structures in curves. Journal of the American Statistical Association, 94(447):807–823, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669996.

Cleveland:1993:AMN

[CMM93] William S. Cleveland, Colin L. Mallows, and Jean E. McRae. ATS methods: Nonparametric regression for non-Gaussian data. Journal of the American Statistical Association, 88(423):821–835, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290771.

Currin:1991:BPD

[CMMY91] Carla Currin, Toby Mitchell, Max Morris, and Don Ylvisaker. Bayesian prediction of deterministic functions, with applications to the design and analysis of computer experiments. Journal of the American Statistical Association, 86(416):953–963, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290511.

Chang:1999:HBR

[CMNS99] William H. Chang, Joseph W. McKean, Joshua D. Naranjo, and Simon J. Sheather. High-breakdown rank regression. Journal of the American Statistical Association, 94(445):205–219, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669695.

Cargnoni:1997:BFM

[CMW97] Claudia Cargnoni, Peter Muller, and Mike West. Bayesian forecasting of multinomial time series through conditionally Gaussian dynamic models. Journal of the American Statistical Association, 92(438):640–647, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2965711.

Cook:1994:RAE

[CN94] R. Dennis Cook and Christopher J. Nachtsheim. Reweighting to achieve elliptically contoured covariates in regression. Journal of the American Statistical Association, 89(426):592–599, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290862.

Chappell:1995:MDL

[CNF95] Rick Chappell, David M. Nondahl, and John F. Fowler. Modeling dose and local control in radiotherapy. Journal of the American Statistical Association, 90(431):829–838, September 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291316.

Cryer:1990:FES

[CNS90] Jonathan D. Cryer, John C. Nankervis, and N. E. Savin. Forecast error symmetry in ARIMA models. Journal of the American Statistical Association, 85(411):724–728, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290008.

Cohen:1998:ACS

[CNWW98a] Jacqueline Cohen, Daniel Nagin, Garrick Wallstrom, and Larry Wasserman. Applications and case studies — hierarchical Bayesian analysis of arrest rates. Journal of the American Statistical Association, 93(444):1260–1270, December 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/dec/cohen.html.

Cohen:1998:HBA

[CNWW98b] Jacqueline Cohen, Daniel Nagin, Garrick Wallstrom, and Larry Wasserman. Hierarchical Bayesian analysis of arrest rates. Journal of the American Statistical Association, 93(444):1260–1270, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2670041.

Cain:1990:RLB

[CO90] Michael Cain and Roger J. Owen. Regressor levels for Bayesian predictive response. Journal of the American Statistical Association, 85(409):228–231, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289549.

Collings:1991:BRB

[Col91] Bruce Jay Collings. Book review: [untitled]. Journal of the American Statistical Association, 86(416):1143, December 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290542.

Collings:1993:BRB

[Col93] Bruce Jay Collings. Book review: [untitled]. Journal of the American Statistical Association, 88(422):699, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290359.

Collings:1997:BRB

[Col97] Bruce Jay Collings. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1648–1649, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2965442.

Conaway:1992:ARC

[Con92] Mark R. Conaway. The analysis of repeated categorical measurements subject to nonignorable nonresponse. Journal of the American Statistical Association, 87(419):817–824, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290220.

Conn:1997:BRB

[Con97] Judith M. Conn. Book review: [untitled]. Journal of the American Statistical Association, 92(437):388, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291491.

Cooil:1991:UMM

[Coo91a] Bruce Cooil. Using medical malpractice data to predict the frequency of claims: a study of Poisson process models with random effects. Journal of the American Statistical Association, 86(414): 285–295, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290560.

Cook:1991:BRB

[Coo91b] Annabel K. Cook. Book review: [untitled]. Journal of the American Statistical Association, 86(413):241, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289736.

Cook:1993:BRB

[Coo93] R. Dennis Cook. Book review: [untitled]. Journal of the American Statistical Association, 88(422):710, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290374.

Cook:1994:IRP

[Coo94] R. Dennis Cook. On the interpretation of regression plots. Journal of the American Statistical Association, 89(425):177–189, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291214.

Cook:1995:BRB

[Coo95] Dianne Cook. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1126–1127, September 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291355.

Cook:1996:GRB

[Coo96] R. Dennis Cook. Graphics for regressions with a binary response. Journal of the American Statistical Association, 91(435):983–992, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291717.

Cook:1998:PHDa

[Coo98a] R. Dennis Cook. Principal Hessian directions revisited. Journal of the American Statistical Association, 93(441):84–94, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/COOK.HTM;   http://www.jstor.org/stable/

2669605. With comments by Ker-Chau Li and a rejoinder by the author.

Cook:1998:PHDb

[Coo98b] R. Dennis Cook. Principal Hessian directions revisited: Rejoinder. Journal of the American Statistical Association, 93(441): 98–100, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2669607.

Cook:1998:RPH

[Coo98c] R. Dennis Cook. Rejoinder: Principle Hessian directions revisited. Journal of the American Statistical Association, 93(441):98–??, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cornell:1992:BRB

[Cor92] Richard G. Cornell. Book review: [untitled]. Journal of the American Statistical Association, 87(418):590, June 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290308.

Corcoran:1999:BRB

[Cor99] Jem N. Corcoran. Book review: [untitled]. Journal of the American Statistical Association, 94(447):986, September 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670022.

Cowan:1991:BRB

[Cow91] Charles D. Cowan. Book review: [untitled]. Journal of the American Statistical Association, 86(413):251–252, March 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289750.

Cowles:1998:BRB

[Cow98] M. K. Cowles. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1523–1524, December 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670069.

Cox:1990:REF

[Cox90] Thomas Cox. Regression estimates in Federal Welfare Quality Control Programs: Comment. Journal of the American Statistical Association, 85(411):868–869, September 1990. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2290029.

Cox:1991:MMG

[Cox91] D. R. Cox. Measures, models, and graphical displays in the analysis of cross-classified data: Comment. Journal of the American Statistical Association, 86(416):1121, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290533.

Cox:1995:NMC

[Cox95] Lawrence H. Cox. Network models for complementary cell suppression. Journal of the American Statistical Association, 90(432): 1453–1462, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291538.

Cox:1999:BRB

[Cox99] Dennis D. Cox. Book review: [untitled]. Journal of the American Statistical Association, 94(448):1388, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669959.

Carlin:1991:EUA

[CP91] Bradley P. Carlin and Nicholas G. Polson. An expected utility approach to influence diagnostics. Journal of the American Statistical Association, 86(416):1013–1021, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290519.

Caulkins:1993:QDQ

[CP93] Jonathan P. Caulkins and Rema Padman. Quantity discounts and quality premia for illicit drugs. Journal of the American Statistical Association, 88(423):748–757, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290759.

Castillo:1999:IEM

[CP99] Joan Del Castillo and Pedro Puig. Invariant exponential models applied to reliability theory and survival analysis. Journal of the American Statistical Association, 94(446):522–528, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670172.

Carota:1996:DMM

[CPP96] Cinzia Carota, Giovanni Parmigiani, and Nicholas G. Polson. Diagnostic measures for model criticism. Journal of the American Statistical Association, 91(434):753–762, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291670.

Carlin:1992:MCA

[CPS92] Bradley P. Carlin, Nicholas G. Polson, and David S. Stoffer. A Monte Carlo approach to nonnormal and nonlinear state-space modeling. Journal of the American Statistical Association, 87 (418):493–500, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290282.

Clogg:1991:MMG

[CR91] Clifford C. Clogg and C. R. Rao. Measures, models, and graphical displays in the analysis of cross-classified data: Comment. Journal of the American Statistical Association, 86(416):1118–1120, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290532.

Catalano:1992:BLV

[CR92] Paul J. Catalano and Louise M. Ryan. Bivariate latent variable models for clustered discrete and continuous outcomes. Journal of the American Statistical Association, 87(419):651–658, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290200.

Crawford:1994:ALM

[Cra94] Sybil L. Crawford. An application of the Laplace method to finite mixture distributions. Journal of the American Statistical Association, 89(425):259–267, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291222.

Cranston:1995:BRB

[Cra95] Michael Cranston. Book review: [untitled]. Journal of the American Statistical Association, 90(430):808, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291107.

Cressie:1990:CGP

[Cre90] Noel Cressie. Correction: a graphical procedure for determining nonstationarity in time series. Journal of the American Statistical Association, 85(409):272, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289584.

Cressie:1994:ASS

[Cre94] Noel Cressie. An approach to statistical spatial-temporal modeling of meteorological fields: Comment. Journal of the American Statistical Association, 89(426):379–382, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290833.

Cressie:1997:COE

[Cre97a] Noel Cressie. Comment: Ozone exposure and population density in Harris County, Texas. Journal of the American Statistical Association, 92(438):411–??, June 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CCG^{+}97a, CCG^{+}97c].

Cressie:1997:OEP

[Cre97b] Noel Cressie. Ozone exposure and population density in Harris County, Texas: Comment. Journal of the American Statistical Association, 92(438):411–413, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965687. See [CCG^{+}97a].

Croux:1994:GE

[CRH94] Christophe Croux, Peter J. Rousseeuw, and Ola Hössjer. Generalized S-estimators. Journal of the American Statistical Association, 89(428):1271–1281, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290990.

Crouchley:1995:REM

[Cro95] Robert Crouchley. A random-effects model for ordered categorical data. Journal of the American Statistical Association, 90(430): 489–498, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291059.

Crowley:1997:PPM

[Cro97] Evelyn M. Crowley. Product partition models for normal means. Journal of the American Statistical Association, 92 (437):192–198, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/march/crowley.htm;
 http://www.jstor.org/stable/2291463.

Clogg:1991:MII

[CRS^{+}91] Clifford C. Clogg, Donald B. Rubin, Nathaniel Schenker, Bradley Schultz, and Lynn Weidman. Multiple imputation of industry and occupation codes in census public-use samples using Bayesian logistic regression. Journal of the American Statistical Association, 86(413):68–78, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289716.

Carroll:1999:DRD

[CRS99a] Raymond J. Carroll, David Ruppert, and Leonard A. Stefanski. Discussion: Regression depth. Journal of the American Statistical Association, 94(446):410–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:1999:RDC

[CRS99b] Raymond J. Carroll, David Ruppert, and Leonard A. Stefanski. Regression depth: Comment. Journal of the American Statistical Association, 94(446):410–411, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670159.

Carroll:1998:LEE

[CRW98] Raymond J. Carroll, David Ruppert, and Alan H. Welsh. Local estimating equations. Journal of the American Statistical Association, 93(441):214–227, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

amstat.org/publications/jasa/abstracts_98/CARROLL.HTM;
 http://www.jstor.org/stable/2669618.

Cryer:1999:BRB

[Cry99] Jonathan D. Cryer. Book review: [untitled]. Journal of the American Statistical Association, 94(448):1384, December 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669952.

Carroll:1990:AQL

[CS90a] Raymond J. Carroll and Leonard A. Stefanski. Approximate quasi-likelihood estimation in models with surrogate predictors. Journal of the American Statistical Association, 85(411):652–663, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290000.

Crouch:1990:EIF

[CS90b] Edmund A. C. Crouch and Donna Spiegelman. The evaluation of integrals of the form ∫_{−∞}^{+∞} f(t) exp(−t^{2}) dt: Application to logistic-normal models. Journal of the American Statistical Association, 85(410):464–469, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289785.

Calvin:1991:BFP

[CS91] James A. Calvin and J. Sedransk. Bayesian and frequentist predictive inference for the patterns of care studies. Journal of the American Statistical Association, 86(413):36–48, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289713.

Cohen:1992:ITC

[CS92a] Arthur Cohen and H. B. Sackrowitz. Improved tests for comparing treatments against a control and other one-sided problems. Journal of the American Statistical Association, 87(420): 1137–1144, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290652.

Cohen:1992:EST

[CS92b] Arthur Cohen and Harold B. Sackrowitz. An evaluation of some tests of trend in contingency tables. Journal of the American Statistical Association, 87(418):470–475, June 1992. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2290279.

Chaudhuri:1993:STM

[CS93] Probal Chaudhuri and Debapriya Sengupta. Sign tests in multidimension: Inference based on the geometry of the data cloud. Journal of the American Statistical Association, 88(424):1363– 1370, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291278.

Cook:1994:SEE

[CS94] J. R. Cook and L. A. Stefanski. Simulation-extrapolation estimation in parametric measurement error models. Journal of the American Statistical Association, 89(428):1314–1328, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290994.

Campodonico:1995:IPP

[CS95a] Sylvia Campodonico and Nozer D. Singpurwalla. Inference and predictions from Poisson point processes incorporating expert knowledge. Journal of the American Statistical Association, 90 (429):220–226, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291146.

Chuang-Stein:1995:BRB

[CS95b] Christy Chuang-Stein. Book review: [untitled]. Journal of the American Statistical Association, 90(429):391, March 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291174.

Chen:1996:NCR

[CS96a] Jingxian Chen and Nozer D. Singpurwalla. The notion of “composite reliability” and its hierarchical Bayes estimation. Journal of the American Statistical Association, 91(436):1474–1484, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291573.

Cohen:1996:LCB

[CS96b] Arthur Cohen and H. B. Sackrowitz. Lower confidence bounds using pilot samples with an application to auditing. Journal of the American Statistical Association, 91(433):338–342, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291412.

Cook:1990:DAA

[CT90] R. Dennis Cook and Chih-Ling Tsai. Diagnostics for assessing the accuracy of normal approximations in exponential family nonlinear models. Journal of the American Statistical Association, 85 (411):770–777, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290014.

Chen:1993:FCA

[CT93a] Rong Chen and Ruey S. Tsay. Functional-coefficient autoregressive models. Journal of the American Statistical Association, 88 (421):298–308, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290725.

Chen:1993:NAA

[CT93b] Rong Chen and Ruey S. Tsay. Nonlinear additive ARX models. Journal of the American Statistical Association, 88(423):955–967, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290787.

Cordy:1997:DDF

[CT97] Clifford B. Cordy and David R. Thomas. Deconvolution of a distribution function. Journal of the American Statistical Association, 92(440):1459–1465, December 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_97/
december/CORDY.HTM;   http://www.jstor.org/stable/2965416.

Chao:1998:SCA

[CT98] Anne Chao and P. K. Tsay. A sample coverage approach to multiple-system estimation with application to census undercount. Journal of the American Statistical Association, 93(441): 283–293, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/CHAO.HTM;   http://www.
jstor.org/stable/2669624.

Chen:1996:PMC

[CTC96] Chen-Hsin Chen, Wei-Yann Tsai, and Wei-Hsiung Chao. The product-moment correlation coefficient and linear regression for truncated data. Journal of the American Statistical Association, 91(435):1181–1186, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2291736.

Consonni:1992:CPE

[CV92] Guido Consonni and Piero Veronese. Conjugate priors for exponential families having quadratic variance functions. Journal of the American Statistical Association, 87(420):1123–1127, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290650.

Consonni:1995:BMC

[CV95] Guido Consonni and Piero Veronese. A Bayesian method for combining results from several binomial experiments. Journal of the American Statistical Association, 90(431):935–944, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291328.

Cook:1990:CCN

[CW90] R. Dennis Cook and Sanford Weisberg. Confidence curves in nonlinear regression. Journal of the American Statistical Association, 85(410):544–551, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289796.

Chang:1991:ETS

[CW91a] Wei-Chien Chang and Donald E. Weiss. An examination of the time series properties of beta in the market model. Journal of the American Statistical Association, 86(416):883–890, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290501.

Cook:1991:SIR

[CW91b] R. Dennis Cook and Sanford Weisberg. Sliced inverse regression for dimension reduction: Comment. Journal of the American Statistical Association, 86(414):328–332, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290564.

Casella:1993:PCR

[CW93a] George Casella and Martin T. Wells. Is Pitman closeness a reasonable criterion?: Comment. Journal of the American Statistical Association, 88(421):70–71, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290696.

Clarke:1993:NPN

[CW93b] Bertrand Clarke and Larry Wasserman. Noninformative priors and nuisance parameters. Journal of the American Statistical Association, 88(424):1427–1432, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291287.

Cheng:1994:ALS

[CW94a] K. F. Cheng and J. W. Wu. Adjusted least squares estimates for the scaled regression coefficients with censored data. Journal of the American Statistical Association, 89(428):1483–1491, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291010.

Cheng:1994:TGF

[CW94b] K. F. Cheng and J. W. Wu. Testing goodness of fit for a parametric family of link functions. Journal of the American Statistical Association, 89(426):657–664, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290868.

Cook:1994:ECC

[CW94c] R. Dennis Cook and Weng Kee Wong. On the equivalence of constrained and compound optimal designs. Journal of the American Statistical Association, 89(426):687–692, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290872.

Cook:1997:GAA

[CW97] R. Dennis Cook and Sanford Weisberg. Graphics for assessing the adequacy of regression models. Journal of the American Statistical Association, 92(438):490–499, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965698.

Chang:1999:CRA

[CW99] Shu-Hui Chang and Mei-Cheng Wang. Conditional regression analysis for recurrence time data. Journal of the American Statistical Association, 94(448):1221–1230, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669936.

Cwiek:1994:BRB

[Cwi94] Charles Cwiek. Book review: [untitled]. Journal of the American Statistical Association, 89(428):1563–1564, December 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291024.

Carroll:1995:PAL

[CWW95] R. J. Carroll, Suojin Wang, and C. Y. Wang. Prospective analysis of logistic case-control studies. Journal of the American Statistical Association, 90(429):157–169, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291139.

Cheng:1997:PSP

[CWY97] S. C. Cheng, L. J. Wei, and Z. Ying. Predicting survival probabilities with semiparametric transformation models. Journal of the American Statistical Association, 92(437):227–235, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/march/cheng.htm;   http://www.jstor.org/
stable/2291467.

DAgostino:1993:BRB

[D’A93] Ralph B. D’Agostino. Book review: [untitled]. Journal of the American Statistical Association, 88(424):1473–1474, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291305.

DAgostino:1994:BRBa

[D’A94a] Ralph B. D’Agostino. Book review: [untitled]. Journal of the American Statistical Association, 89(425):359, March 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291241.

DAgostino:1994:BRBb

[D’A94b] Ralph B. D’Agostino. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1142, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290951.

DAgostino:1994:BRBc

[D’A94c] Ralph B. D’Agostino. Book review: [untitled]. Journal of the American Statistical Association, 89(428):1564–1565, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291026.

DAgostino:1994:BRBd

[D’A94d] Ralph B. D’Agostino. Book review: [untitled]. Journal of the American Statistical Association, 89(428):1565, December 1994.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291027.

DAgostino:1995:BRBa

[D’A95a] Ralph B. D’Agostino. Book review: [untitled]. Journal of the American Statistical Association, 90(429):390–391, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291173.

DAgostino:1995:BRBb

[D’A95b] Ralph B. D’Agostino. Book review: [untitled]. Journal of the American Statistical Association, 90(430):799–800, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291095.

DAgostino:1995:BRBc

[D’A95c] Ralph B. D’Agostino. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1125–1126, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291354.

DAgostino:1996:BRBa

[D’A96a] Ralph B. D’Agostino. Book review: [untitled]. Journal of the American Statistical Association, 91(434):910, June 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291689.

DAgostino:1996:BRBb

[D’A96b] Ralph B. D’Agostino. Book review: [untitled]. Journal of the American Statistical Association, 91(434):910–911, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291690.

Dunn:1996:EHV

[DA96c] D. T. Dunn and A. E. Ades. Estimating the HIV vertical transmission rate and the pediatric AIDS incubation period from prospective data. Journal of the American Statistical Association, 91 (435):935–943, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291712.

Dabrowska:1990:SRT

[Dab90] Dorota M. Dabrowska. Signed-rank tests for censored matched pairs. Journal of the American Statistical Association, 85(410): 478–485, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289787.

Dabrowska:1992:BRB

[Dab92] Dorota M. Dabrowska. Book review: [untitled]. Journal of the American Statistical Association, 87(420):1244, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290673.

Dawkins:1994:AOH

[DAO94] Brian P. Dawkins, Peter M. Andreae, and Paul M. O’Connor. Analysis of Olympic heptathlon data. Journal of the American Statistical Association, 89(427):1100–1106, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290940.

Dassios:1995:BRB

[Das95] Angelos Dassios. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1134–1135, September 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291366.

Davis:1991:BRB

[Dav91] Charles S. Davis. Book review: [untitled]. Journal of the American Statistical Association, 86(414):553–554, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290622.

David:1992:BRB

[Dav92] Martin David. Book review: [untitled]. Journal of the American Statistical Association, 87(420):1248–1249, December 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290679.

David:1996:LE

[Dav96] Herbert A. David. Letter to the editor. Journal of the American Statistical Association, 91(436):1756, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291622.

Davis:1998:BRB

[Dav98] Richard A. Davis. Book review: [untitled]. Journal of the American Statistical Association, 93(442):834–835, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670133.

David:1999:BRB

[Dav99] Herbert A. David. Book review: [untitled]. Journal of the American Statistical Association, 94(446):646–648, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670185.

Dawkins:1995:IGD

[Daw95] Brian P. Dawkins. Investigating the geometry of a p-dimensional data set. Journal of the American Statistical Association, 90(429): 350–359, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291160.

Dean:1997:ELA

[DB97] C. B. Dean and R. Balshaw. Efficiency lost by analyzing counts rather than event times in Poisson and overdispersed Poisson regression models. Journal of the American Statistical Association, 92(440):1387–1398, December 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_97/
december/DEAN.HTM;   http://www.jstor.org/stable/2965408.

Doksum:1994:CCL

[DBB^{+}94] Kjell Doksum, Stephen Blyth, Eric Bradlow, Xiao li Meng, and Hongyu Zhao. Correlation curves as local measures of variance explained by regression. Journal of the American Statistical Association, 89(426):571–582, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2290860.

Pereira:1997:BRB

[dBPI97] Carlos A. de B. Pereira and Telba Irony. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1651– 1652, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2965448.

Doss:1994:CRS

[DC94] Hani Doss and Yuang Chin Chiang. Choosing the resampling scheme when bootstrapping: a case study in reliability. Journal of the American Statistical Association, 89(425):298–308, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291226.

deCarvalho:1994:SDT

[dCDdSO94] Filipa Duarte de Carvalho, Nico P. Dellaert, and Margarida de Sanches Osorio. Statistical disclosure in two-dimensional tables: General tables. Journal of the American Statistical Association, 89(428):1547–1557, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291018.

delCastillo:1999:BTE

[dCP99a] Joan del Castillo and Pedro Puig. The best test of exponentiality against singly truncated normal alternatives. Journal of the American Statistical Association, 94(446):529–532, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670173.

delCastillo:1999:IEM

[dCP99b] Joan del Castillo and Pedro Puig. Invariant exponential models applied to reliability theory and survival analysis. Journal of the American Statistical Association, 94(446):522–??, June 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Donald:1990:SSA

[DD90] Alan Donald and Allan Donner. A simulation study of the analysis of sets of 2 × 2 contingency tables under cluster sampling: Estimation of a common odds ratio. Journal of the American Statistical Association, 85(410):537–543, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289795.

DuarteDeCarvalho:1994:SDT

[DDD94] Filipa Duarte De Carvalho, Nico P. Dellaert, and Margarida De Sanches Osorio. Statistical disclosure in two-dimensional tables: General tables. Journal of the American Statistical Association, 89(428):1547–1557, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Das:1999:CPO

[DDV99] Marcel Das, Jeff Dominitz, and Arthur Van Soest. Comparing predictions and outcomes: Theory and application to income changes. Journal of the American Statistical Association, 94(445): 75–85, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2669679.

Dickey:1996:BRBa

[DE96] James M. Dickey and Morris L. Eaton. Book review: Bayesian Theory, by J. M. Bernardo; Adrian F. M. Smith. Journal of the American Statistical Association, 91(434):906–907, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291685.

Edwardes:1997:URC

[dE97] Michael D. deB. Edwardes. Univariate random cut-points theory for the analysis of ordered categorical data. Journal of the American Statistical Association, 92(439):1114– 1123, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/sept/edwardes.htm;
 http://www.jstor.org/stable/2965576.

Dean:1992:TOP

[Dea92] C. B. Dean. Testing for overdispersion in Poisson and binomial regression models. Journal of the American Statistical Association, 87(418):451–457, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290276.

DeGruttola:1995:BRB

[DeG95] Victor DeGruttola. Book review: [untitled]. Journal of the American Statistical Association, 90(430):797–798, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291093.

Devlin:1995:BRB

[Dev95] Bernie Devlin. Book review: [untitled]. Journal of the American Statistical Association, 90(432):1483–1488, December 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291542.

Dardanoni:1998:UAL

[DF98] Valentino Dardanoni and Antonio Forcina. A unified approach to likelihood inference on stochastic orderings in a non-parametric context. Journal of the American Statistical Association, 93(443): 1112–1123, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/sept/dardanoni.html;
http://www.jstor.org/stable/2669854.

Darroch:1993:TSM

[DFGJ93] John N. Darroch, Stephen E. Fienberg, Gary F. V. Glonek, and Brian W. Junker. A three-sample multiple-recapture approach to census population estimation with heterogeneous catchability. Journal of the American Statistical Association, 88(423): 1137–1148, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290811.

Doss:1992:EAW

[DG92] Hani Doss and Richard D. Gill. An elementary approach to weak convergence for quantile processes, with applications to censored survival data. Journal of the American Statistical Association, 87 (419):869–877, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290226.

Davies:1993:IMOb

[DG93a] Laurie Davies and Ursula Gather. The identification of multiple outliers: Rejoinder. Journal of the American Statistical Association, 88(423):797–801, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290768.

Davies:1993:IMOa

[DG93b] Laurie Davies and Ursula Gather. The identification of multiple outliers. With discussion and a reply by the authors. Journal of the American Statistical Association, 88(423):782–792, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290763.

Datta:1995:SRN

[DG95] Gauri Sankar Datta and Malay Ghosh. Some remarks on noninformative priors. Journal of the American Statistical Association, 90 (432):1357–1363, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291526.

Daniels:1999:HGL

[DG99] Michael J. Daniels and Constantine Gatsonis. Hierarchical generalized linear models in the analysis of variations in health care utilization. Journal of the American Statistical Association, 94 (445):29–42, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2669675.

Dalal:1991:PEI

[DH91] Siddhartha R. Dalal and Bruce Hoadley. Problems in extrapolation illustrated with Space Shuttle O-ring data: Comment. Journal of the American Statistical Association, 86(416):921–922, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290506.

Dufour:1993:IEB

[DH93] Jean-Marie Dufour and Marc Hallin. Improved Eaton bounds for linear combinations of bounded random variables, with statistical applications. Journal of the American Statistical Association, 88 (423):1026–1033, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290795.

Ding:1999:PIF

[DH99] A. Adam Ding and J. T. Gene Hwang. Prediction intervals, factor analysis models, and high-dimensional empirical linear prediction. Journal of the American Statistical Association, 94(446):446–455, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/ding.pdf;   http://www.jstor.org/stable/2670165.

Dhar:1993:CCM

[Dha93] Sunil K. Dhar. Computation of certain minimum distance estimators in AR[k] model. Journal of the American Statistical Association, 88(421):278–283, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290723.

deAngelis:1993:ABA

[dHY93] D. de Angelis, Peter Hall, and G. A. Young. Analytical and bootstrap approximations to estimator distributions in L^{1} regression. Journal of the American Statistical Association, 88(424): 1310–1316, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Diaconis:1999:BRB

[Dia99] Persi Diaconis. Book review: [untitled]. Journal of the American Statistical Association, 94(447):984, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670018.

DiCiccio:1995:BRB

[DiC95] Thomas J. DiCiccio. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1129–1130, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291359.

Dickey:1996:BRBb

[Dic96] James M. Dickey. Book review: [untitled]. Journal of the American Statistical Association, 91(434):907–908, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291686.

Diebold:1993:BRB

[Die93] Francis X. Diebold. Book review: [untitled]. Journal of the American Statistical Association, 88(423):1171, September 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290822.

Dietz:1997:BRB

[Die97] E. Jacquelin Dietz. Book review: [untitled]. Journal of the American Statistical Association, 92(438):797, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965740.

Diggle:1999:AND

[Dig99] Peter J. Diggle. Adjusting for nonignorable drop-out using semiparametric nonresponse models: Comment. Journal of the American Statistical Association, 94(448):1128–1129, December 1999.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669927.

Donoho:1995:AUS

[DJ95] David L. Donoho and Iain M. Johnstone. Adapting to unknown smoothness via wavelet shrinkage. Journal of the American Statistical Association, 90(432):1200–1224, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291512.

Dickey:1998:FVP

[DJ98] James M. Dickey and Thomas J. Jiang. Filtered-variate prior distributions for histogram smoothing. Journal of the American Statistical Association, 93(442):651–662, June 1998. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_98/june/
DICKEY.HTM;   http://www.jstor.org/stable/2670116.

deJong:1998:DST

[dJP98] Piet de Jong and Jeremy Pinzer. Diagnosing shocks in time series. Journal of the American Statistical Association, 93 (442):796–806, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/june/DE_JONG.HTM;
http://www.jstor.org/stable/2670129.

Daniels:1999:NBE

[DK99] Michael J. Daniels and Robert E. Kass. Nonconjugate Bayesian estimation of covariance matrices and its use in hierarchical models. Journal of the American Statistical Association, 94(448): 1254–1263, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2669939.

Dykstra:1995:ILR

[DKR95] Richard Dykstra, Subhash Kochar, and Tim Robertson. Inference for likelihood ratio ordering in the two-sample problem. Journal of the American Statistical Association, 90(431):1034–1040, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291340.

DiCiccio:1997:CBF

[DKRW97] Thomas J. DiCiccio, Robert E. Kass, Adrian Raftery, and Larry Wasserman. Computing Bayes factors by combining simulation and asymptotic approximations. Journal of the American Statistical Association, 92(439):903–915, September 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/sept/
diciccio.htm;   http://www.jstor.org/stable/2965554.

DeOliveira:1997:BPT

[DKS97] Victor De Oliveira, Benjamin Kedem, and David A. Short. Bayesian prediction of transformed Gaussian random fields. Journal of the American Statistical Association, 92(440):1422– 1433, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/DEOLIVEI.HTM.

Daponte:1997:BDP

[DKW97] Beth Osborne Daponte, Joseph B. Kadane, and Lara J. Wolfson. Bayesian demography: Projecting the Iraqi Kurdish population, 1977–1990. Journal of the American Statistical Association, 92(440):1256–1267, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/december/DAPONTE.HTM;   http://www.jstor.
org/stable/2965396.

Dean:1992:MDT

[DL92] A. M. Dean and S. M. Lewis. Multidimensional designs for twofactor experiments. Journal of the American Statistical Association, 87(420):1158–1165, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290655.

deLeeuw:1993:BRB

[dL93] Jan de Leeuw. Book review: [untitled]. Journal of the American Statistical Association, 88(422):696–697, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290356.

Diggle:1991:AVR

[DLB91] Peter J. Diggle, Nicholas Lange, and Francine M. Benes. Analysis of variance for replicated spatial point patterns in clinical neuroanatomy. Journal of the American Statistical Association, 86(415):618–625, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290390.

DeGruttola:1991:MPH

[DLD91] Victor DeGruttola, Nicholas Lange, and Urania Dafni. Modeling the progression of HIV infection. Journal of the American Statistical Association, 86(415):569–577, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290384.

Datta:1999:HBE

[DLML99] G. S. Datta, P. Lahiri, T. Maiti, and K. L. Lu. Hierarchical Bayes estimation of unemployment rates for the states of the U.S. Journal of the American Statistical Association, 94(448):1074– 1082, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2669921.

Donnelly:1995:PCS

[DLW95] Christl A. Donnelly, Nan M. Laird, and James H. Ware. Prediction and creation of smooth curves for temporally correlated longitudinal data. Journal of the American Statistical Association, 90(431):984–989, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291334.

Davidson:1990:STB

[DM90] Russell Davidson and James G. MacKinnon. Specification tests based on artificial regressions. Journal of the American Statistical Association, 85(409):220–227, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289548.

Davis:1993:RBM

[DM93] James B. Davis and Joseph W. McKean. Rank-based methods for multivariate linear models. Journal of the American Statistical Association, 88(421):245–251, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290719.

Datta:1995:BIF

[DM95] Somnath Datta and William P. McCormick. Bootstrap inference for a first-order autoregression with positive innovations. Journal of the American Statistical Association, 90(432):1289–1300, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291519.

Dixon:1996:RBA

[DM96] Sherry L. Dixon and Joseph W. McKean. Rank-based analysis of the heteroscedastic linear model. Journal of the American Statistical Association, 91(434):699–712, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291665.

Dominitz:1997:UED

[DM97] Jeff Dominitz and Charles F. Manski. Using expectations data to study subjective income expectations. Journal of the American Statistical Association, 92(439):855–867, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/sept/dominitz.htm;   http://www.jstor.org/
stable/2965550.

Durazo-Arvizu:1997:ENB

[DMLC97] Ramon Durazo-Arvizu, Daniel McGee, Zhaohai Li, and Richard Cooper. Establishing the nadir of the body mass indexmortality relationship: a case study. Journal of the American Statistical Association, 92(440):1312–1319, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/december/DURAZO.HTM;   http://www.jstor.
org/stable/2965401.

Deaton:1998:ACS

[DN98a] Angus Deaton and Serena Ng. Applications and case studies — parametric and nonparametric approaches to price and tax reform. Journal of the American Statistical Association, 93(443): 900–909, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/sept/deaton.html.

Deaton:1998:PNA

[DN98b] Angus Deaton and Serena Ng. Parametric and nonparametric approaches to price and tax reform. Journal of the American Statistical Association, 93(443):900–909, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669828.

Doganaksoy:1995:BRB

[Dog95] Necip Doganaksoy. Book review: [untitled]. Journal of the American Statistical Association, 90(430):804, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291102.

Dorfman:1990:REF

[Dor90] Alan H. Dorfman. Regression estimates in Federal Welfare Quality Control Programs: Comment. Journal of the American Statistical Association, 85(411):869–870, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290030.

Dorfman:1994:NVE

[Dor94] Alan H. Dorfman. A note on variance estimation for the regression estimator in double sampling. Journal of the American Statistical Association, 89(425):137–140, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291209.

Dominici:1999:MAM

[DPWH99] Francesca Dominici, Giovanni Parmigiani, Robert L. Wolpert, and Vic Hasselblad. Meta-analysis of migraine headache treatments: Combining information from heterogeneous designs. Journal of the American Statistical Association, 94(445):16–28, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669674.

Diebold:1991:FOC

[DR91] Francis X. Diebold and Glenn D. Rudebusch. Forecasting output with the composite leading index: a real- time analysis. Journal of the American Statistical Association, 86(415):603–610, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290388.

Davison:1996:SMD

[DR96] A. C. Davison and N. I. Ramesh. Some models for discretized series of events. Journal of the American Statistical Association, 91(434):601–609, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291656.

Dasgupta:1998:DFS

[DR98] Abhijit Dasgupta and Adrian E. Raftery. Detecting features in spatial point processes with clutter via model-based clustering. Journal of the American Statistical Association, 93(441): 294–302, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/DASGUPTA.HTM;   http://
www.jstor.org/stable/2669625.

Draper:1994:BRBa

[Dra94a] Norman R. Draper. Book review: [untitled]. Journal of the American Statistical Association, 89(426):713, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290880.

Draper:1994:BRBb

[Dra94b] Norman R. Draper. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1140, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290948.

Devlin:1992:FID

[DRR92] B. Devlin, Neil Risch, and Kathryn Roeder. Forensic inference from DNA fingerprints. Journal of the American Statistical Association, 87(418):337–350, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2290264.

Dollinger:1991:IFI

[DS91] Michael B. Dollinger and Robert G. Staudte. Influence functions of iteratively reweighted least squares estimators. Journal of the American Statistical Association, 86(415):709–716, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290402.

Deville:1992:CES

[DS92] Jean-Claude Deville and Carl-Erik Särndal. Calibration estimators in survey sampling. Journal of the American Statistical Association, 87(418):376–382, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2290268.

Diggle:1997:ANC

[DS97a] Peter J. Diggle and David M. Smith. Analysis of nonrandomly censored ordered categorical longitudinal data from analgesic trials: Comment. Journal of the American Statistical Association, 92 (440):1251–1252, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2965394.

Diggle:1997:CAN

[DS97b] Peter J. Diggle and David M. Smith. Comment: Analysis of nonrandomly censored ordered categorical longitudinal data from analgesic trials. Journal of the American Statistical Association, 92(440):1251–??, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Durand:1997:ASP

[DS97c] Jean-François Durand and Robert Sabatier. Additive splines for partial least squares regression. Journal of the American Statistical Association, 92(440):1546–1554, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/december/DURAND.HTM;   http://www.jstor.
org/stable/2965425.

Denison:1998:BCM

[DSM98a] D. G. T. Denison, A. F. M. Smith, and B. K. Mallick. Bayesian CART model search: Comment. Journal of the American Statistical Association, 93(443):954–957, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669835.

Denison:1998:CBC

[DSM98b] D. G. T. Denison, A. F. M. Smith, and B. K. Mallick. Comment: Bayesian CART model search. Journal of the American Statistical Association, 93(443):954–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Deville:1993:GRP

[DSS93] Jean-Claude Deville, Carl-Erik Särndal, and Olivier Sautory. Generalized raking procedures in survey sampling. Journal of the American Statistical Association, 88(423):1013–1020, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290793.

DSS:1997:BRBa

[DSS97a] DSS. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1226, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965609.

DSS:1997:BRBb

[DSS97b] DSS. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1656, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965460.

DSS:1997:BRBc

[DSS97c] DSS. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1656, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965463.

DSS:1998:BRB

[DSS98] DSS. Book review: [untitled]. Journal of the American Statistical Association, 93(442):844, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670149.

DSS:1999:BRB

[DSS99] DSS. Book review: [untitled]. Journal of the American Statistical Association, 94(445):351, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669730.

Dunnett:1992:SMT

[DT92] Charles W. Dunnett and Ajit C. Tamhane. A step-up multiple test procedure. Journal of the American Statistical Association, 87(417):162–170, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290465.

Duan:1990:APP

[Dua90] Naihua Duan. The adjoint projection pursuit regression. Journal of the American Statistical Association, 85(412):1029–1038, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289599.

Duan:1991:TDE

[Dua91] Naihua Duan. Transformations in density estimation: Comment. Journal of the American Statistical Association, 86(414):355–356, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290572.

Dehejia:1999:CEN

[DW99] Rajeev H. Dehejia and Sadek Wahba. Causal effects in nonexperimental studies: Reevaluating the evaluation of training programs. Journal of the American Statistical Association, 94(448): 1053–1062, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2669919.

Dabrowska:1994:CRM

[DwSH94] Dorota M. Dabrowska, Guo wen Sun, and Mary M. Horowitz. Cox regression in a Markov renewal model: An application to the analysis of bone marrow transplant data. Journal of the American Statistical Association, 89(427):867–877, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290911.

Degruttola:1992:PAN

[DXP92] Victor Degruttola, Ming Tu Xin, and Marcello Pagano. Pediatric AIDS in New York City: Estimating the distributions of infection, latency, and reporting delay and projecting future incidence. Journal of the American Statistical Association, 87(419): 633–640, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290198.

Easton:1991:CML

[Eas91] George S. Easton. Compromise maximum likelihood estimators for location. Journal of the American Statistical Association, 86 (416):1051–1064, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290524.

Easterling:1992:BRB

[Eas92] Robert G. Easterling. Book review: [untitled]. Journal of the American Statistical Association, 87(419):905–906, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290244.

Easton:1994:SDG

[Eas94] George S. Easton. A simple dynamic graphical diagnostic method for almost any model. Journal of the American Statistical Association, 89(425):201–207, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2291216.

Eddy:1990:BRB

[Edd90] William F. Eddy. Book review: [untitled]. Journal of the American Statistical Association, 85(410):599, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289818.

Edelman:1990:IOO

[Ede90] David Edelman. An inequality of optimal order for the tail probabilities of the T statistic under symmetry. Journal of the American Statistical Association, 85(409):120–122, March 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289533.

Efron:1991:CEVa

[EF91a] B. Efron and D. Feldman. Compliance as an explanatory variable in clinical trials. Journal of the American Statistical Association, 86(413):9–17, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289707.

Efron:1991:CEVb

[EF91b] B. Efron and D. Feldman. Compliance as an explanatory variable in clinical trials: Rejoinder. Journal of the American Statistical Association, 86(413):25–26, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289711.

Efron:1990:MEB

[Efr90] Bradley Efron. More efficient bootstrap computations. Journal of the American Statistical Association, 85(409):79–89, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289528.

Efron:1992:POE

[Efr92] Bradley Efron. Poisson overdispersion estimates based on the method of asymmetric maximum likelihood. Journal of the American Statistical Association, 87(417):98–107, March 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290457.

Efron:1994:MDIb

[Efr94a] Bradley Efron. Missing data, imputation, and the bootstrap: Rejoinder. Journal of the American Statistical Association, 89(426): 478–479, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290848.

Efron:1994:MDIa

[Efr94b] Bradley Efron. Missing data, imputation, and the bootstrap. With a comment by Donald B. Rubin and a rejoinder by the author. Journal of the American Statistical Association, 89(426): 463–475, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290846.

Efron:1996:EBMa

[Efr96a] Bradley Efron. Empirical Bayes methods for combining likelihoods. Journal of the American Statistical Association, 91(434): 538–550, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291646.

Efron:1996:EBMb

[Efr96b] Bradley Efron. Empirical Bayes methods for combining likelihoods: Rejoinder. Journal of the American Statistical Association, 91(434):562–565, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291652.

Efron:1996:REB

[Efr96c] Bradley Efron. Rejoinder: Empirical Bayes methods for combining likelihoods. With discussion and a reply by the author. Journal of the American Statistical Association, 91(434):538–565, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efromovich:1997:DEC

[Efr97] Sam Efromovich. Density estimation for the case of supersmooth measurement error. Journal of the American Statistical Association, 92(438):526–535, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2965701.

Efromovich:1998:DDE

[Efr98] Sam Efromovich. Data-driven efficient estimation of the spectral density. Journal of the American Statistical Association, 93(442):762–769, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/june/EFROMOVI.HTM;
 http://www.jstor.org/stable/2670126.

Efromovich:1999:QLW

[Efr99] Sam Efromovich. Quasi-linear wavelet estimation. Journal of the American Statistical Association, 94(445):189–204, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669694.

Ellenberg:1992:SIAa

[EFS92a] Susan S. Ellenberg, Dianne M. Finkelstein, and David A. Schoenfeld. Statistical issues arising in AIDS clinical trials. Journal of the American Statistical Association, 87(418):562–569, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290291.

Ellenberg:1992:SIAb

[EFS92b] Susan S. Ellenberg, Dianne M. Finkelstein, and David A. Schoenfeld. Statistical issues arising in AIDS clinical trials: Rejoinder. Journal of the American Statistical Association, 87(418):581–583, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290297.

El-Gamal:1995:PBU

[EGG95] Mahmoud A. El-Gamal and David M. Grether. Are people Bayesian? uncovering behavioral strategies. Journal of the American Statistical Association, 90(432):1137–1145, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291506.

Evans:1997:BAS

[EGGS97] Michael Evans, Zvi Gilula, Irwin Guttman, and Tim Swartz. Bayesian analysis of stochastically ordered distributions of categorical variables. Journal of the American Statistical Association, 92(437):208–214, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

amstat.org/publications/jasa/abstracts_97/march/evans.
htm;   http://www.jstor.org/stable/2291465.

El-Gamal:1993:BSE

[EGMP93] Mahmoud A. El-Gamal, Richard D. McKelvey, and Thomas R. Palfrey. A Bayesian sequential experimental study of learning in games. Journal of the American Statistical Association, 88(422): 428–435, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290321.

Edwards:1991:TTS

[EH91] Cheryl L. Edwards and E. Philip Howrey. A “true” time series and its indicators: An alternative approach. Journal of the American Statistical Association, 86(416):878–882, December 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290500.

EJD:1997:BRBa

[EJD97a] EJD. Book review: [untitled]. Journal of the American Statistical Association, 92(438):801, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965758.

EJD:1997:BRBb

[EJD97b] EJD. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1227, September 1997. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965613.

EJD:1998:BRBa

[EJD98a] EJD. Book review: [untitled]. Journal of the American Statistical Association, 93(441):412, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669668.

EJD:1998:BRBb

[EJD98b] EJD. Book review: [untitled]. Journal of the American Statistical Association, 93(442):844, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670147.

Etzioni:1993:OED

[EK93] Ruth Etzioni and Joseph B. Kadane. Optimal experimental design for Another’s analysis. Journal of the American Statistical Association, 88(424):1404–1411, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291284.

Ericksen:1991:TEP

[EKWS91] Eugene P. Ericksen, Joseph B. Kadane, Kenneth W. Wachter, and Terence P. Speed. Total error in PES estimates of population: Comment. Journal of the American Statistical Association, 86 (416):855–857, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290496.

Eubank:1990:CPP

[EL90] R. L. Eubank and V. N. LaRiccia. Components of Pearson’s phi-squared distance measure for the k-sample problem. Journal of the American Statistical Association, 85(410):441–445, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289781.

Eggermont:1997:NSE

[EL97] P. P. B. Eggermont and V. N. Lariccia. Nonlinearly smoothed EM density estimation with automated smoothing parameter selection for nonparametric deconvolution problems. Journal of the American Statistical Association, 92(440):1451–1458, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_97/december/EGERMONT.HTM;   http://www.
jstor.org/stable/2965415.

Elashoff:1991:BRB

[Ela91] Janet D. Elashoff. Book review: [untitled]. Journal of the American Statistical Association, 86(415):832–833, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290432.

Elder:1996:BRB

[Eld96] John F. Elder, IV. Book review: Machine Learning, Neural, and Statistical Classification, by Donald Michie; David J. Spiegelhalter; Charles C. Taylor. Journal of the American Statistical Association, 91(433):436–438, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291432.

Eliason:1998:BRB

[Eli98] Scott R. Eliason. Book review: [untitled]. Journal of the American Statistical Association, 93(441):403, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669644.

Eltinge:1996:CVE

[Elt96a] John L. Eltinge. Comment: On variance estimation with imputed survey data. Journal of the American Statistical Association, 91 (434):499–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Eltinge:1996:VEI

[Elt96b] John L. Eltinge. On variance estimation with imputed survey data: Comment. Journal of the American Statistical Association, 91(434):513–515, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291640.

Easton:1990:MGQ

[EM90] George S. Easton and Robert E. McCulloch. A multivariate generalization of quantile-quantile plots. Journal of the American Statistical Association, 85(410):376–386, June 1990. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289773.

Ellis:1992:LBR

[EM92] Steven P. Ellis and Stephan Morgenthaler. Leverage and breakdown in L_{1} regression. Journal of the American Statistical Association, 87(417):143–148, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2290462.

Embrechts:1995:BRB

[Emb95] Paul Embrechts. Book review: [untitled]. Journal of the American Statistical Association, 90(430):806, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291105.

Embrechts:1996:BRBa

[Emb96a] Paul Embrechts. Book review: [untitled]. Journal of the American Statistical Association, 91(434):915, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291697.

Embrechts:1996:BRBb

[Emb96b] Paul Embrechts. Book review: [untitled]. Journal of the American Statistical Association, 91(436):1753, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291612.

Embrechts:1997:BRBa

[Emb97a] Paul Embrechts. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1649, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965443.

Embrechts:1997:BRBb

[Emb97b] Paul Embrechts. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1649–1650, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965444.

Embrechts:1998:BRBa

[Emb98a] Paul Embrechts. Book review: [untitled]. Journal of the American Statistical Association, 93(441):399, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669638.

Embrechts:1998:BRBb

[Emb98b] Paul Embrechts. Book review: [untitled]. Journal of the American Statistical Association, 93(442):840, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670141.

Embrechts:1998:BRBc

[Emb98c] Paul Embrechts. Book review: [untitled]. Journal of the American Statistical Association, 93(442):843, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670145.

Embrechts:1998:BRBd

[Emb98d] Paul Embrechts. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1243–1244, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669882.

Ehm:1995:PRA

[EMM95] W. Ehm, E. Mammen, and D. W. Müller. Power robustification of approximately linear tests. Journal of the American Statistical Association, 90(431):1025–1033, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291339.

Efron:1994:SAG

[EP94] Bradley Efron and Vahe Petrosian. Survival analysis of the gamma-ray burst data. Journal of the American Statistical Association, 89(426):452–462, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2290845.

Efron:1999:NMD

[EP99] Bradley Efron and Vahe Petrosian. Nonparametric methods for doubly truncated data. Journal of the American Statistical Association, 94(447):824–834, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669997.

Erkanli:1994:LAP

[Erk94] Alaattin Erkanli. Laplace approximations for posterior expectations when the mode occurs at the boundary of the parameter space. Journal of the American Statistical Association, 89(425): 250–258, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291221.

Edelsbrunner:1990:CLM

[ES90a] Herbert Edelsbrunner and Diane L. Souvaine. Computing least median of squares regression lines and guided topological sweep. Journal of the American Statistical Association, 85(409):115– 119, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289532.

Eubank:1990:TGF

[ES90b] R. L. Eubank and C. H. Spiegelman. Testing the goodness of fit of a linear model via nonparametric regression techniques. Journal of the American Statistical Association, 85(410):387–392, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289774.

Eubank:1993:CBN

[ES93] R. L. Eubank and P. L. Speckman. Confidence bands in nonparametric regression. Journal of the American Statistical Association, 88(424):1287–1301, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291269.

Emond:1997:EEG

[ES97] Mary Jane Emond and Steven G. Self. An efficient estimator for the generalized semilinear model. Journal of the American Statistical Association, 92(439):1033–1040, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_97/
sept/emond.htm;   http://www.jstor.org/stable/2965568.

Ernst:1999:CPT

[ES99] Michael D. Ernst and William R. Schucany. A class of permutation tests of bivariate interchangeability. Journal of the American Statistical Association, 94(445):273–284, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669702.

Escobar:1994:ENM

[Esc94] Michael D. Escobar. Estimating normal means with a Dirichlet process prior. Journal of the American Statistical Association, 89 (425):268–277, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291223.

Esty:1992:VCW

[Est92] Warren W. Esty. Votes or competitions which determine a winner by estimating expected plurality. Journal of the American Statistical Association, 87(418):373–375, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290267.

Efron:1997:ICV

[ET97] Bradley Efron and Robert Tibshirani. Improvements on crossvalidation: The .632+ bootstrap method. Journal of the American Statistical Association, 92(438):548–560, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965703.

Eubank:1993:BRB

[Eub93] Randall L. Eubank. Book review: [untitled]. Journal of the American Statistical Association, 88(422):697–698, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290357.

Eubank:1997:TGF

[Eub97] R. L. Eubank. Testing goodness of fit with multinomial data. Journal of the American Statistical Association, 92(439):1084– 1093, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/sept/eubank.htm;
http://www.jstor.org/stable/2965573.

Engel:1992:CPP

[EV92] Eduardo Engel and Achilles Venetoulias. The Chilean plebiscite: Projections without historic data. Journal of the American Statistical Association, 87(420):933–941, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290629.

Escobar:1995:BDE

[EW95] Michael D. Escobar and Mike West. Bayesian density estimation and inference using mixtures. Journal of the American Statistical Association, 90(430):577–588, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291069.

Fabian:1991:PIAb

[Fab91a] Vaclav Fabian. On the problem of interactions in the analysis of variance: Rejoinder. Journal of the American Statistical Association, 86(414):374–375, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290580.

Fabian:1991:PIAa

[Fab91b] Václav Fabian. On the problem of interactions in the analysis of variance. With discussion and a rejoinder by the author. Journal of the American Statistical Association, 86(414):362–367, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290576. See

[Sha91].

Fabian:1994:CSA

[Fab94] Václav Fabian. Comment on: “Some asymptotic results for learning in single hidden-layer feedforward network models” [J. Amer. Statist. Assoc. 84 (1989), no. 408, 1003–1013, MR 92e:62119] by H. White. Journal of the American Statistical Association, 89 (428):1571, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2291041. See [Whi89, Whi92b].

Fahrmeir:1992:PME

[Fah92] Ludwig Fahrmeir. Posterior mode estimation by extended Kalman filtering for multivariate dynamic generalized linear models. Journal of the American Statistical Association, 87(418):501–509, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290283.

Fan:1992:DAN

[Fan92] Jianqing Fan. Design-adaptive nonparametric regression. Journal of the American Statistical Association, 87(420):998–1004, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290637.

Fan:1996:TSB

[Fan96] Jianqing Fan. Test of significance based on wavelet thresholding and Neyman’s truncation. Journal of the American Statistical Association, 91(434):674–688, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291663.

Faraway:1997:BRB

[Far97] Julian J. Faraway. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1214–1215, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2965589.

Faulkenberry:1990:BRB

[Fau90] G. David Faulkenberry. Book review: [untitled]. Journal of the American Statistical Association, 85(412):1178, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289634.

Faulkenberry:1994:BRB

[Fau94] G. David Faulkenberry. Book review: [untitled]. Journal of the American Statistical Association, 89(428):1563, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291023.

Fay:1996:APA

[Fay96a] Robert E. Fay. Alternative paradigms for the analysis of imputed survey data. Journal of the American Statistical Association, 91 (434):490–498, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291636.

Fay:1996:VEI

[Fay96b] Robert E. Fay. On variance estimation with imputed survey data: Rejoinder. Journal of the American Statistical Association, 91 (434):517–519, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291642.

Fay:1996:RVE

[Fay96c] Robert E. Fay. Rejoinder: On variance estimation with imputed survey data. Journal of the American Statistical Association, 91 (434):499–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Flury:1997:DSM

[FBF97] Leah Flury, Benzion Boukai, and Bernard D. Flury. The discrimination subspace model. Journal of the American Statistical Association, 92(438):758–766, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965724.

Fernholz:1998:BRB

[Fer98a] Luisa T. Fernholz. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1527, December 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670075.

Ferre:1998:DDS

[Fer98b] Louis Ferré. Determining the dimension in sliced inverse regression and related methods. Journal of the American Statistical Association, 93(441):132–140, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

amstat.org/publications/jasa/abstracts_98/FERRE.HTM;
http://www.jstor.org/stable/2669610.

Faulkenberry:1991:EPT

[FG91] G. David Faulkenberry and Abderrazak Garoui. Estimating a population total using an area frame. Journal of the American Statistical Association, 86(414):445–449, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290592.

Fan:1994:CRL

[FG94] Jianqing Fan and Irène Gijbels. Censored regression: Local linear approximations and their applications. Journal of the American Statistical Association, 89(426):560–570, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290859.

Fan:1995:CNT

[FG95] Yanqin Fan and Ramazan Gencay. A consistent nonparametric test of symmetry in linear regression models. Journal of the American Statistical Association, 90(430):551–557, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291066.

Fine:1999:PHM

[FG99] Jason P. Fine and Robert J. Gray. A proportional hazards model for the subdistribution of a competing risk. Journal of the American Statistical Association, 94(446):496–509, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670170.

Fisher:1989:BCR

[FH89] Nicholas I. Fisher and Peter Hall. Bootstrap confidence regions for directional data. Journal of the American Statistical Association, 84(408):996–1002, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2290075. See correction [FH90].

Fisher:1990:CBC

[FH90] Nicholas I. Fisher and Peter Hall. Correction: “Bootstrap confidence regions for directional data”. Journal of the American Statistical Association, 85(410):608, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289832. See [FH89].

FH:1995:BRB

[FH95] FH. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1136, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291371.

Fisher:1996:IPM

[FHJW96] Nicholas I. Fisher, Peter Hall, Bing-Yi Jing, and Andrew T. A. Wood. Improved pivotal methods for constructing confidence regions with directional data. Journal of the American Statistical Association, 91(435):1062–1070, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291725.

Fan:1996:LSN

[FHMP96] Jianqing Fan, Peter Hall, Michael A. Martin, and Prakash Patil. On local smoothing of nonparametric curve estimators. Journal of the American Statistical Association, 91(433):258–266, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291403.

Fisher:1997:ECS

[FHTW97] Nicholas I. Fisher, Peter Hall, Berwin A. Turlach, and G. S. Watson. On the estimation of a convex set from noisy data on its support function. Journal of the American Statistical Association, 92(437):84–91, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/march/fisher.htm;
http://www.jstor.org/stable/2291452.

Fan:1995:LPK

[FHW95] Jianqing Fan, Nancy E. Heckman, and M. P. Wand. Local polynomial kernel regression for generalized linear models and quasilikelihood functions. Journal of the American Statistical Association, 90(429):141–150, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2291137.

Fairley:1990:IWQ

[FIB90] William B. Fairley, Alan J. Izenman, and Partha Bagchi. Inference for welfare quality control programs. Journal of the American Statistical Association, 85(411):874–890, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290032.

Fienberg:1990:ITF

[Fie90] Stephen E. Fienberg. Interactional troubles in face-to-face survey interviews: Comment. Journal of the American Statistical Association, 85(409):241–244, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289551.

Fieller:1993:IMO

[Fie93a] Nick Fieller. The identification of multiple outliers: Comment. Journal of the American Statistical Association, 88(423):794–795, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290765.

Fienberg:1993:BRB

[Fie93b] Stephen E. Fienberg. Book review: [untitled]. Journal of the American Statistical Association, 88(421):375, March 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290735.

Finner:1990:SNI

[Fin90] Helmut Finner. Some new inequalities for the range distribution, with application to the determination of optimum significance levels of multiple range tests. Journal of the American Statistical Association, 85(409):191–194, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289544.

Finner:1993:MPS

[Fin93] H. Finner. On a monotonicity problem in step-down multiple test procedures. Journal of the American Statistical Association, 88 (423):920–923, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290782.

Fishburn:1991:BRB

[Fis91] Peter C. Fishburn. Book review: [untitled]. Journal of the American Statistical Association, 86(415):823–824, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290423.

Faraway:1990:BCB

[FJ90] Julian J. Faraway and Myoungshic Jhun. Bootstrap choice of bandwidth for density estimation. Journal of the American Statistical Association, 85(412):1119–1122, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289609.

Fu:1994:DTR

[FK94] J. C. Fu and M. V. Koutras. Distribution theory of runs: a Markov chain approach. Journal of the American Statistical Association, 89(427):1050–1058, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290933.

Fienberg:1999:CCL

[FK99] Stephen E. Fienberg and Sung-Ho Kim. Combining conditional log-linear structures. Journal of the American Statistical Association, 94(445):229–239, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669697.

Freedman:1998:BRB

[FKOR98] D. A. Freedman, S. P. Klein, M. Ostland, and M. R. Roberts. Book review: A Solution to the Ecological Inference Problem, by G. King. Journal of the American Statistical Association, 93(444): 1518–1522, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2670067.

Ferretti:1999:CLG

[FKYZ99] Nelida Ferretti, Diana Kelmansky, Victor J. Yohai, and Ruben H. Zamar. A class of locally and globally robust regression estimates. Journal of the American Statistical Association, 94(445): 174–188, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2669693.

Fitzmaurice:1995:RMB

[FL95] Garrett M. Fitzmaurice and Nan M. Laird. Regression models for a bivariate discrete and continuous outcome with clustering. Journal of the American Statistical Association, 90(431):845–852, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291318.

Fan:1998:TSW

[FL98] Jianqing Fan and Sheng-Kuei Lin. Test of significance when data are curves. Journal of the American Statistical Association, 93(443):1007–1021, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_98/sept/
fan.html;   http://www.jstor.org/stable/2669845.

Flack:1990:BRB

[Fla90] Virginia F. Flack. Book review: [untitled]. Journal of the American Statistical Association, 85(410):599–600, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289819.

Flack:1991:BRB

[Fla91] Virginia F. Flack. Book review: [untitled]. Journal of the American Statistical Association, 86(413):252–253, March 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289753.

Fleiss:1990:BRB

[Fle90] Joseph L. Fleiss. Book review: [untitled]. Journal of the American Statistical Association, 85(411):907, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290045.

Fleiss:1991:BRB

[Fle91] Joseph L. Fleiss. Book review: [untitled]. Journal of the American Statistical Association, 86(416):1140, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290537.

Fitzmaurice:1996:MLM

[FLZ96] Garrett M. Fitzmaurice, Nan M. Laird, and Gwendolyn E. P. Zahner. Multivariate logistic models for incomplete binary responses. Journal of the American Statistical Association, 91(433): 99–108, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291386.

Fox:1992:GCD

[FM92] John Fox and Georges Monette. Generalized collinearity diagnostics. Journal of the American Statistical Association, 87(417): 178–183, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290467.

Freedman:1993:BRB

[FM93] Vicki A. Freedman and John E. Marcotte. Book review: [untitled]. Journal of the American Statistical Association, 88(423): 1169–1170, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290820.

Findley:1990:SSD

[FMSP90] David F. Findley, Brian C. Monsell, Holly B. Shulman, and Marian G. Pugh. Sliding-spans diagnostics for seasonal and related adjustments. Journal of the American Statistical Association, 85 (410):345–355, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289770.

Flury:1995:SDS

[FNP95] Bernard D. Flury, Daan G. Nel, and Inet Pienaar. Simultaneous detection of shift in means and variances. Journal of the American Statistical Association, 90(432):1474–1481, December 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291540.

Follmann:1996:SMT

[Fol96] Dean Follmann. A simple multivariate test for one-sided alternatives. Journal of the American Statistical Association, 91(434): 854–861, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291680.

Fong:1992:RER

[Fon92] Duncan K. H. Fong. Ranking and estimation of related means in the presence of a covariate — A Bayesian approach. Journal of the American Statistical Association, 87(420):1128–1136, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290651.

Formann:1992:LLL

[For92] Anton K. Formann. Linear logistic latent class analysis for polytomous data. Journal of the American Statistical Association, 87 (418):476–486, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290280.

Formann:1994:MCL

[For94] Anton K. Formann. Measuring change in latent subgroups using dichotomous data: Unconditional, conditional, and semiparametric maximum likelihood estimation. Journal of the American Statistical Association, 89(427):1027–1034, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290930.

Freedman:1999:RKC

[FORK99] D. A. Freedman, M. Ostland, M. R. Roberts, and S. P. Klein. Response to King’s comment. Journal of the American Statistical Association, 94(445):355–357, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669735.

Fernandez:1995:MIS

[FOS95] Carmen Fernández, Jacek Osiewalski, and Mark F. J. Steel. Modeling and inference with v-spherical distributions. Journal of the American Statistical Association, 90(432):1331–1340, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291523.

Fernandez:1997:CBI

[FOS97] Carmen Fernandez, Jacek Osiewalski, and Mark F. J. Steel. Classical and Bayesian inference robustness in multivariate regression models. Journal of the American Statistical Association, 92(440): 1434–1444, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/FERNANDE.HTM;
 http://www.jstor.org/stable/2965413.

Foutz:1992:BRB

[Fou92] Robert V. Foutz. Book review: [untitled]. Journal of the American Statistical Association, 87(418):585, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290299.

Foutz:1993:BRB

[Fou93] Robert V. Foutz. Book review: [untitled]. Journal of the American Statistical Association, 88(424):1475, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291308.

Foresi:1995:CDE

[FP95] Silverio Foresi and Franco Peracchi. The conditional distribution of excess returns: An empirical analysis. Journal of the American Statistical Association, 90(430):451–466, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291056.

Frees:1990:EFS

[FR90] Edward W. Frees and David Ruppert. Estimation following a sequentially designed experiment. Journal of the American Statistical Association, 85(412):1123–1129, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289610.

Forbes:1999:BMF

[FR99] Florence Forbes and Adrian E. Raftery. Bayesian morphology: Fast unsupervised Bayesian image analysis. Journal of the American Statistical Association, 94(446):555–568, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670176.

Fraser:1991:SIL

[Fra91] D. A. S. Fraser. Statistical inference: Likelihood to significance. Journal of the American Statistical Association, 86(414):258–265, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290557.

Franke:1999:BRB

[Fra99] Jurgen Franke. Book review: [untitled]. Journal of the American Statistical Association, 94(446):652–653, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670192.

Frees:1994:EDF

[Fre94] Edward W. Frees. Estimating densities of functions of observations. Journal of the American Statistical Association, 89(426): 517–525, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290854.

Freeny:1995:BRB

[Fre95] Anne Freeny. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1127–1128, September 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291357.

Frey:1998:BRB

[Fre98] Rudiger Frey. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1247, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669888.

Freedman:1999:AND

[Fre99a] D. A. Freedman. Adjusting for nonignorable drop-out using semiparametric nonresponse models: Comment. Journal of the American Statistical Association, 94(448):1121–1122, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669924.

Frey:1999:BRB

[Fre99b] Rudiger Frey. Book review: [untitled]. Journal of the American Statistical Association, 94(447):986, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670021.

Friendly:1994:MDM

[Fri94] Michael Friendly. Mosaic displays for multi-way contingency tables. Journal of the American Statistical Association, 89(425): 190–200, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291215.

Frick:1998:BRB

[Fri98] Robert W. Frick. Book review: [untitled]. Journal of the American Statistical Association, 93(441):406–407, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669649.

Frigessi:1994:IPB

[FS94] Arnoldo Frigessi and Julian Stander. Informative priors for the Bayesian classification of satellite images. Journal of the American Statistical Association, 89(426):703–709, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290875.

Faraway:1995:SCB

[FS95a] Julian J. Faraway and Jiayang Sun. Simultaneous confidence bands for linear regression with heteroscedastic errors. Journal of the American Statistical Association, 90(431):1094–1098, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291347.

Forbes:1995:EOR

[FS95b] Andrew B. Forbes and Thomas J. Santner. Estimators of odds ratio regression parameters in matched case- control studies with covariate measurement error. Journal of the American Statistical Association, 90(431):1075–1084, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291345.

Fay:1998:PTU

[FS98a] Michael P. Fay and Joanna H. Shih. Permutation tests using estimated distribution functions. Journal of the American Statistical Association, 93(441):387–396, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.amstat.org/publications/jasa/abstracts_98/FAY.HTM;
 http://www.jstor.org/stable/2669635.

Fernandez:1998:BMF

[FS98b] Carmen Fernández and Mark F. J. Steel. On Bayesian modeling of fat tails and skewness. Journal of the American Statistical Association, 93(441):359–371, March 1998. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
FERNANDC.HTM;   http://www.jstor.org/stable/2669632.

Faucett:1998:ACS

[FSE98] Cheryl L. Faucett, Nathaniel Schenker, and Robert M. Elashoff. Analysis of censored survival data with intermittently observed time-dependent binary covariates. Journal of the American Statistical Association, 93(442):427–437, June 1998. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_98/june/
FAUCETT.HTM;   http://www.jstor.org/stable/2670090.

Fahrmeir:1994:DSM

[FT94] Ludwig Fahrmeir and Gerhard Tutz. Dynamic stochastic models for time-dependent ordered paired comparison systems. Journal of the American Statistical Association, 89(428):1438–1449, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291005.

Fung:1990:BRB

[Fun90] Karen Y. Fung. Book review: [untitled]. Journal of the American Statistical Association, 85(409):261–262, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289567.

Funk:1992:BRB

[Fun92] Sandra G. Funk. Book review: [untitled]. Journal of the American Statistical Association, 87(419):908–909, September 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290248.

Fung:1993:UOL

[Fun93] Wing-Kam Fung. Unmasking outliers and leverage points: a confirmation. Journal of the American Statistical Association, 88 (422):515–519, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290331.

Fung:1995:DLD

[Fun95] Wing K. Fung. Diagnostics in linear discriminant analysis. Journal of the American Statistical Association, 90(431):952–956, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291330.

Fahrmeir:1996:SHF

[FW96] Ludwig Fahrmeir and Stefan Wagenpfeil. Smoothing hazard functions and time-varying effects in discrete duration and competing risks models. Journal of the American Statistical Association, 91 (436):1584–1594, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291584.

Fraser:1999:RAN

[FWW99] D. A. S. Fraser, Augustine Wong, and Jianrong Wu. Regression analysis, nonlinear or nonnormal: Simple and accurate p values from likelihood analysis. Journal of the American Statistical Association, 94(448):1286–1295, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669942.

Fygenson:1997:NAM

[Fyg97] Mendel Fygenson. A new approach in modeling a categorical response. I. Binary case. Journal of the American Statistical Association, 92(437):322–332, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/march/
fygenson.htm;   http://www.jstor.org/stable/2291477.

Fan:1998:CSS

[FZ98a] Jianqing Fan and Jin-Ting Zhang. Comment: Smoothing spline models for the analysis of nested and crossed samples of curves. Journal of the American Statistical Association, 93(443):980–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:1998:SSM

[FZ98b] Jianqing Fan and Jin-Ting Zhang. Smoothing spline models for the analysis of nested and crossed samples of curves: Comment. Journal of the American Statistical Association, 93(443):980–983, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2669839.

Fan:1999:AND

[FZ99] Jianqing Fan and Chunming Zhang. Adjusting for nonignorable drop-out using semiparametric nonresponse models: Comment. Journal of the American Statistical Association, 94(448):1122– 1125, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2669925.

Gail:1996:SA

[Gai96] Mitchell H. Gail. Statistics in action. Journal of the American Statistical Association, 91(433):1–13, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291378.

Galbraith:1994:SAR

[Gal94] Rex F. Galbraith. Some applications of radial plots. Journal of the American Statistical Association, 89(428):1232–1242, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290987.

Garthwaite:1994:IPL

[Gar94] Paul H. Garthwaite. An interpretation of partial least squares. Journal of the American Statistical Association, 89(425):122– 127, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291207.

Gastwirth:1991:BRB

[Gas91] Joseph L. Gastwirth. Book review: [untitled]. Journal of the American Statistical Association, 86(414):544–546, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290608.

Gastwirth:1995:BRB

[Gas95] Joseph L. Gastwirth. Book review: [untitled]. Journal of the American Statistical Association, 90(429):388–390, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291171.

Gastwirth:1996:BRB

[Gas96] Joseph L. Gastwirth. Book review: [untitled]. Journal of the American Statistical Association, 91(435):1379–1380, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291762.

Gaver:1991:BRB

[Gav91] Donald P. Gaver. Book review: [untitled]. Journal of the American Statistical Association, 86(413):248–249, March 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289746.

Gelman:1997:ANC

[GB97a] Andrew Gelman and Frederic Y. Bois. Analysis of nonrandomly censored ordered categorical longitudinal data from analgesic trials: Comment. Journal of the American Statistical Association, 92 (440):1248–1250, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2965393.

Gelman:1997:CAN

[GB97b] Andrew Gelman and Frederic Y. Bois. Comment: Analysis of nonrandomly censored ordered categorical longitudinal data from analgesic trials. Journal of the American Statistical Association, 92(440):1248–??, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gopalan:1998:BMC

[GB98] Ramanan Gopalan and Donald A. Berry. Bayesian multiple comparisons using Dirichlet process priors. Journal of the American Statistical Association, 93(443):1130–1139, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/sept/gopalan.html;   http://www.jstor.org/
stable/2669856.

Gibbons:1999:ASE

[GBA^{+}99] Robert D. Gibbons, Byron W. M. Brown, Daniel L. Azarnoff, William E. Bunney, Robert Cancro, John C. Gillin, Sandra Hullett, Keith F. Killam, John H. Krystal, David J. Kupfer, Paul D. Stolley, Andrew M. Pope, and Geoffrey S. French. Assessment of the safety and efficacy data for the hypnotic halcion (R): Results of an analysis by an Institute of Medicine Committee. Journal of the American Statistical Association, 94(448):993–1002, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669913.

Gelman:1996:PPA

[GBJ96] Andrew Gelman, Frederic Bois, and Jiming Jiang. Physiological pharmacokinetic analysis using population modeling and informative prior distributions. Journal of the American Statistical Association, 91(436):1400–1412, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291566.

Geronimus:1996:VUC

[GBN96] Arline T. Geronimus, John Bound, and Lisa J. Neidert. On the validity of using census geocode characteristics to proxy individual socioeconomic characteristics. Journal of the American Statistical Association, 91(434):529–537, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291645.

Gentleman:1991:GMC

[GC91] R. Gentleman and J. Crowley. Graphical methods for censored data. Journal of the American Statistical Association, 86(415):

678–683, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290397.

Godtliebsen:1995:ENT

[GC95] Fred Godtliebsen and Chih Kang Chu. Estimation of the number of true Gray levels, their values, and relative frequencies in a noisy image. Journal of the American Statistical Association, 90(431): 890–899, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291323.

Goutis:1996:AEM

[GCW96] Constantinos Goutis, George Casella, and Martin T. Wells. Assessing evidence in multiple hypotheses. Journal of the American Statistical Association, 91(435):1268–1277, September 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291745.

Garcia-Escudero:1999:RPK

[GEG99] Luis Angel Garcia-Escudero and Alfonso Gordaliza. Robustness properties of k means and trimmed k means. Journal of the American Statistical Association, 94(447):956–969, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670010.

Geisser:1992:SSI

[Gei92] Seymour Geisser. Some statistical issues in medicine and forensics. Journal of the American Statistical Association, 87(419):607–614, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290193.

Gelman:1994:DPM

[Gel94a] Andrew Gelman. Discussion: a probabilistic model for the spatial distribution of party support in multiparty electorates. Journal of the American Statistical Association, 89(428):1198–??, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gelman:1994:PMS

[Gel94b] Andrew Gelman. A probabilistic model for the spatial distribution of party support in multiparty electorates: Discussion. Journal of the American Statistical Association, 89(428):1198, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290982.

Gelfand:1996:CEB

[Gel96a] Alan E. Gelfand. Comment: Empirical Bayes methods for combining likelihoods. Journal of the American Statistical Association, 91(434):538–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gelfand:1996:EBM

[Gel96b] Alan E. Gelfand. Empirical Bayes methods for combining likelihoods: Comment. Journal of the American Statistical Association, 91(434):551–552, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291647.

Genius:1993:BRB

[Gen93] Margarida Genius. Book review: [untitled]. Journal of the American Statistical Association, 88(424):1468, December 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291296.

Gentz:1999:BRB

[Gen99] Barbara Gentz. Book review: [untitled]. Journal of the American Statistical Association, 94(447):986–988, September 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670023.

George:1996:CEB

[Geo96a] Edward I. George. Comment: Empirical Bayes methods for combining likelihoods. Journal of the American Statistical Association, 91(434):538–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

George:1996:EBM

[Geo96b] Edward I. George. Empirical Bayes methods for combining likelihoods: Comment. Journal of the American Statistical Association, 91(434):553–555, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291648.

Geyer:1991:CML

[Gey91] Charles J. Geyer. Constrained maximum likelihood exemplified by isotonic convex logistic regression. Journal of the American Statistical Association, 86(415):717–724, September 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290403.

Goria:1996:CCV

[GF96a] Mohammed N. Goria and Bernard D. Flury. Common canonical variates in k independent groups. Journal of the American Statistical Association, 91(436):1735–1742, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291603.

Goutis:1996:PLS

[GF96b] Constantinos Goutis and Tom Fearn. Partial least squares regression on smooth factors. Journal of the American Statistical Association, 91(434):627–632, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291658.

Gaynor:1993:UCS

[GFT^{+}93] Jeffrey J. Gaynor, Eric J. Feuer, Claire C. Tan, Danny H. Wu, Claudia R. Little, David J. Straus, Bayard D. Clarkson, and Murray F. Brennan. On the use of cause-specific failure and conditional failure probabilities: Examples from clinical oncology data. Journal of the American Statistical Association, 88(422):400–409, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290318.

Goodman:1990:ANT

[GH90] Leo A. Goodman and Shelby J. Haberman. The analysis of nonadditivity in two-way analysis of variance. Journal of the American Statistical Association, 85(409):139–145, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289536.

Gibbons:1991:BRB

[GH91] Robert R. Gibbons and Donald R. Hedeker. Book review: [untitled]. Journal of the American Statistical Association, 86(416): 1146–1149, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290546.

Gilula:1994:CLL

[GH94] Zvi Gilula and Shelby J. Haberman. Conditional log-linear models for analyzing categorical panel data. Journal of the American Statistical Association, 89(426):645–656, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290867.

Gilula:1995:DCV

[GH95] Zvi Gilula and Shelby J. Haberman. Dispersion of categorical variables and penalty functions: Derivation, estimation, and comparability. Journal of the American Statistical Association, 90(432): 1447–1452, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291537.

Garel:1999:RBA

[GH99] Bernard Garel and Marc Hallin. Rank-based autoregressive order identification. Journal of the American Statistical Association, 94 (448):1357–1371, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669948.

Gibbons:1994:REP

[GHCF94] Robert D. Gibbons, Donald Hedeker, Sara C. Charles, and Paul Frisch. A random-effects probit model for predicting medical malpractice claims. Journal of the American Statistical Association, 89(427):760–767, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290901.

Ghysels:1996:PST

[GHL96] Eric Ghysels, Alastair Hall, and Hahn Shik Lee. On periodic structures and testing for seasonal unit roots. Journal of the American Statistical Association, 91(436):1551–1559, December 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291580.

Gentleman:1994:SQC

[GHMW94] R. C. Gentleman, M. S. Hamada, D. E. Matthews, and A. R. Wilson. Statistical quality control of HIV-1 ELISA test performance. Journal of the American Statistical Association, 89(428): 1200–1208, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290984.

Ghosh:1990:BRB

[Gho90] Subir Ghosh. Book review: [untitled]. Journal of the American Statistical Association, 85(409):260, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289563.

Ghosh:1992:CBE

[Gho92] Malay Ghosh. Constrained Bayes estimation with applications. Journal of the American Statistical Association, 87(418):533–540, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290287.

Ghosh:1994:BRB

[Gho94] Malay Ghosh. Book review: [untitled]. Journal of the American Statistical Association, 89(426):712–713, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290879.

Ghosh:1996:BRB

[Gho96] Malay Ghosh. Book review: [untitled]. Journal of the American Statistical Association, 91(433):431–432, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291425.

Ghosh:1998:BRB

[Gho98] Sujit K. Ghosh. Book review: [untitled]. Journal of the American Statistical Association, 93(441):402, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669642.

Gelfand:1990:IBI

[GHRPS90] Alan E. Gelfand, Susan E. Hills, Amy Racine-Poon, and Adrian F. M. Smith. Illustration of Bayesian inference in normal data models using Gibbs sampling. Journal of the American Statistical Association, 85(412):972–985, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289594.

Ghysels:1998:BRB

[Ghy98] Eric Ghysels. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1522–1523, December 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670068.

Gilula:1994:BRB

[Gil94] Zvi Gilula. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1140–1141, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290949.

Ginebra:1995:BRB

[Gin95] Josep Ginebra. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1134, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291365.

Givens:1999:MDM

[Giv99] Geof H. Givens. Multicriterion decision merging: Competitive development of an Aboriginal whaling management procedure. Journal of the American Statistical Association, 94(448):1003– 1014, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2669914.

Gastwirth:1994:SCE

[GJ94] J. L. Gastwirth and Wesley O. Johnson. Screening with costeffective quality control: Potential applications to HIV and drug testing. Journal of the American Statistical Association, 89(427): 972–981, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290923.

Gan:1999:TGM

[GJ99a] Li Gan and Jiming Jiang. A test for global maximum. Journal of the American Statistical Association, 94(447):847–854, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669999.

Gatto:1999:CSA

[GJ99b] Riccardo Gatto and S. Rao Jammalamadaka. A conditional saddlepoint approximation for testing problems. Journal of the American Statistical Association, 94(446):533–541, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670174.

Gersovitz:1998:ACS

[GJDT98a] Mark Gersovitz, Hanan G. Jacoby, F. Seri Dedy, and A. Gozé Tapé. Applications and case studies — the balance of self-reported heterosexual activity in KAP surveys and the AIDS epidemic in Africa. Journal of the American Statistical Association, 93(443): 875–883, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/sept/gersovitz.html.

Gersovitz:1998:BSR

[GJDT98b] Mark Gersovitz, Hanan G. Jacoby, F. Seri Dedy, and A. Goze Tape. The balance of self-reported heterosexual activity in KAP surveys and the AIDS epidemic in Africa. Journal of the American Statistical Association, 93(443):875–883, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669826.

Golan:1996:MEA

[GJP96] Amos Golan, George Judge, and Jeffrey M. Perloff. A maximum entropy approach to recovering information from multinomial response data. Journal of the American Statistical Association, 91 (434):841–853, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291679.

Guihenneuc-Jouyaux:1998:DCM

[GJR98] Chantal Guihenneuc-Jouyaux and Christian P. Robert. Discretization of continuous Markov chains and Markov chain Monte Carlo convergence assessment. Journal of the American Statistical Association, 93(443):1055–1067, September 1998. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_98/sept/
guihenneuc-joyaux.html;   http://www.jstor.org/stable/
2669849.

Gelman:1990:EEC

[GK90] Andrew Gelman and Gary King. Estimating the electoral consequences of legislative redistricting. Journal of the American Statistical Association, 85(410):274–282, June 1990. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289762.

Graubard:1993:HTC

[GK93] Barry I. Graubard and Edward L. Korn. Hypothesis testing with complex survey data: The use of classical quadratic test statistics with particular reference to regression problems. Journal of the American Statistical Association, 88(422):629–641, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290345.

Gasser:1995:SSC

[GK95] Theo Gasser and Alois Kneip. Searching for structure in curve samples. Journal of the American Statistical Association, 90(432): 1179–1188, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291510.

George:1996:TIT

[GK96] E. Olusegun George and Ralph L. Kodell. Tests of independence, treatment heterogeneity, and dose-related trend with exchangeable binary data. Journal of the American Statistical Association, 91(436):1602–1610, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291586.

Gelman:1998:EPE

[GKB98] Andrew Gelman, Gary King, and W. John Boscardin. Estimating the probability of events that have never occurred: When is your vote decisive? Journal of the American Statistical Association, 93(441):1–9, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/GELMAN.HTM;   http:
//www.jstor.org/stable/2669597.

Gasser:1991:FFM

[GKK91] Theo Gasser, Alois Kneip, and Walter Köhler. A flexible and fast method for automatic smoothing. Journal of the American Statistical Association, 86(415):643–652, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290393.

Gelman:1998:ACSa

[GKL98a] Andrew Gelman, Gary King, and Chuanhai Liu. Applications and case studies — not asked and not answered: Multiple imputation for multiple surveys. Journal of the American Statistical Association, 93(443):846–857, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/sept/gelman.html.

Gelman:1998:ACSb

[GKL98b] Andrew Gelman, Gary King, and Chuanhai Liu. Applications and case studies — rejoinder: Not asked and not answered: Multiple imputation for multiple surveys. Journal of the American Statistical Association, 93(443):869–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gelman:1998:AAMa

[GKL98c] Andrew Gelman, Gary King, and Chuanhai Liu. Not asked and not answered: Multiple imputation for multiple surveys. Journal of the American Statistical Association, 93(443):846–857, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669819.

Gelman:1998:AAMb

[GKL98d] Andrew Gelman, Gary King, and Chuanhai Liu. Not asked and not answered: Multiple imputation for multiple surveys: Rejoinder. Journal of the American Statistical Association, 93(443): 869–874, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2669825.

Ghosh:1993:PCR

[GKS93] Malay Ghosh, J. P. Keating, and Pranab Kumar Sen. Is Pitman closeness a reasonable criterion?: Comment. Journal of the American Statistical Association, 88(421):63–66, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290693.

Gill:1992:TRD

[GL92] Len Gill and Arthur Lewbel. Testing the rank and definiteness of estimated matrices with applications to factor, state-space, and ARMA models. Journal of the American Statistical Association, 87(419):766–776, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290214.

Gross:1996:NER

[GL96] Shulamith T. Gross and Tze Leung Lai. Nonparametric estimation and regression analysis with left- truncated and rightcensored data. Journal of the American Statistical Association, 91 (435):1166–1180, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291735.

Glass:1991:BRB

[Gla91] Gene V. Glass. Book review: The Future of Meta-Analysis, by Kenneth W. Wachter; Miron L. Straf. Journal of the American Statistical Association, 86(416):1141–1142, December 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290539.

Goldman:1998:CCA

[GLB98] Dana P. Goldman, Arleen Leibowitz, and Joan L. Buchanan. Cost containment and adverse selection in Medicaid HMOs. Journal of the American Statistical Association, 93(441):54–62, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_98/GOLDMAN.HTM;   http://www.jstor.org/
stable/2669602.

Gleser:1992:IAM

[Gle92] Leon Jay Gleser. The importance of assessing measurement reliability in multivariate regression. Journal of the American Statistical Association, 87(419):696–707, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290207.

Gleason:1993:UES

[Gle93] John R. Gleason. Understanding elongation: The scale contaminated normal family. Journal of the American Statistical Association, 88(421):327–337, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2290728.

Giles:1992:OSP

[GLG92] David E. A. Giles, Offer Lieberman, and Judith A. Giles. The optimal size of a preliminary test of linear restrictions in a misspecified regression model. Journal of the American Statistical Association, 87(420):1153–1157, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290654.

Glynn:1993:MIM

[GLR93] Robert J. Glynn, Nan M. Laird, and Donald B. Rubin. Multiple imputation in mixture models for nonignorable nonresponse with follow-ups. Journal of the American Statistical Association, 88 (423):984–993, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290790.

Glymour:1998:BRB

[Gly98] Clark Glymour. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1513–1515, December 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670064.

Garber:1993:NHD

[GM93a] Alan M. Garber and Thomas E. MaCurdy. Nursing home discharges and exhaustion of Medicare benefits. Journal of the American Statistical Association, 88(423):727–736, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290757.

George:1993:VSG

[GM93b] Edward I. George and Robert E. McCulloch. Variable selection via Gibbs sampling. Journal of the American Statistical Association, 88(423):881–889, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290777.

Gomez:1994:EPI

[GM94] Víctor Gómez and Agustín Maravall. Estimation, prediction, and interpolation for nonstationary series with the Kalman filter. Journal of the American Statistical Association, 89(426):611–624, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290864.

Goetghebeur:1996:CIP

[GM96] Els Goetghebeur and Geert Molenberghs. Causal inference in a placebo-controlled clinical trial with binary outcome and ordered compliance. Journal of the American Statistical Association, 91 (435):928–934, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291711.

George:1999:VSF

[GM99] E. I. George and R. E. McCulloch. Variable selection and function estimation in additive nonparametric regression using a databased prior: Comment. Journal of the American Statistical Association, 94(447):798–799, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669992.

Gonzalez-Manteiga:1996:BSS

[GMCM96] W. González-Manteiga, R. Cao, and J. S. Marron. Bootstrap selection of the smoothing parameter in nonparametric hazard rate estimation. Journal of the American Statistical Association, 91 (435):1130–1140, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291732.

Gelfand:1995:MEO

[GMD95] Alan E. Gelfand, Bani K. Mallick, and Dipak K. Dey. Modeling expert opinion arising as a partial probabilistic specification. Journal of the American Statistical Association, 90(430):598–604, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291071.

Gelfand:1997:BBS

[GMP97] Alan E. Gelfand, Bani K. Mallick, and Wolfgang Polasek. Broken biological size relationships: a truncated semiparametric regression approach with measurement error. Journal of the American Statistical Association, 92(439):836–845, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/sept/gelfand.htm;   http://www.jstor.org/
stable/2965548.

Gijbels:1999:EMC

[GMPS99] Irène Gijbels, Enno Mammen, Byeong U. Park, and Léopold Simar. On estimation of monotone and concave frontier functions. Journal of the American Statistical Association, 94(445): 220–228, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2669696.

Geys:1999:PMM

[GMR99] Helena Geys, Geert Molenberghs, and Louise M. Ryan. Pseudolikelihood modeling of multivariate outcomes in developmental toxicology. Journal of the American Statistical Association, 94 (447):734–745, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669986.

Guttorp:1997:COE

[GMS97a] Peter Guttorp, Wendy Meiring, and Paul D. Sampson. Comment: Ozone exposure and population density in Harris County, Texas. Journal of the American Statistical Association, 92(438):405–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CCG^{+}97a].

Guttrop:1997:OEP

[GMS97b] Peter Guttrop, Wendy Meiring, and Paul D. Sampson. Ozone exposure and population density in Harris County, Texas: Comment. Journal of the American Statistical Association, 92(438): 405–408, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2965685. See [CCG^{+}97a].

Greenwood:1999:ISP

[GMWZ99] Michael J. Greenwood, John M. McDowell, Donald M. Waldman, and Steven S. Zahniser. The influence of social programs in source countries on various classes of U.S. immigration. Journal of the American Statistical Association, 94(445):64–74, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669678.

Ghosh:1996:EMI

[GNK96] Malay Ghosh, Narinder Nangia, and Dal Ho Kim. Estimation of median income of four-person families: a Bayesian time series approach. Journal of the American Statistical Association, 91 (436):1423–1431, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291568.

Ghosh:1998:GLM

[GNSC98] Malay Ghosh, Kannan Natarajan, T. W. F. Stroud, and Bradley P. Carlin. Generalized linear models for small area estimation. Journal of the American Statistical Association, 93 (441):273–282, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/GHOSH.HTM;   http:/
/www.jstor.org/stable/2669623.

Goldstein:1993:BRB

[Gol93] Harvey Goldstein. Book review: [untitled]. Journal of the American Statistical Association, 88(421):386–387, March 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290750.

Goldie:1995:BRB

[Gol95] Charles M. Goldie. Book review: [untitled]. Journal of the American Statistical Association, 90(430):806–808, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291106.

Goodman:1991:MMGb

[Goo91a] Leo A. Goodman. Measures, models, and graphical displays in the analysis of cross-classified data: Rejoinder. Journal of the American Statistical Association, 86(416):1124–1138, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290535.

Goodman:1991:MMGa

[Goo91b] Leo A. Goodman. Measures, models, and graphical displays in the analysis of cross-classified data. With comments and a rejoinder by the author. Journal of the American Statistical Association, 86 (416):1085–1111, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290528.

Good:1992:BNB

[Goo92] I. J. Good. The Bayes/non-Bayes compromise: a brief review. Journal of the American Statistical Association, 87(419):597–606, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290192.

Goodall:1994:BRB

[Goo94a] Colin Goodall. Book review: [untitled]. Journal of the American Statistical Association, 89(428):1562–1563, December 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291022.

Goodman:1994:QIQ

[Goo94b] Leo A. Goodman. On quasi-independence and quasi-dependence in contingency tables, with special reference to ordinal triangular contingency tables. Journal of the American Statistical Association, 89(427):1059–1063, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290934.

Goodman:1996:SGM

[Goo96] Leo A. Goodman. A single general method for the analysis of cross-classified data: Reconciliation and synthesis of some methods of Pearson, Yule, and Fisher, and also some methods of correspondence analysis and association analysis. Journal of the American Statistical Association, 91(433):408–428, March 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291421.

Goutis:1997:NEM

[Gou97] Constantinos Goutis. Nonparametric estimation of a mixing density via the kernel method. Journal of the American Statistical Association, 92(440):1445–1450, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/december/GOUTIS.HTM;   http://www.jstor.
org/stable/2965414.

Gumpertz:1992:NRV

[GP92] Marcia L. Gumpertz and Sastry G. Pantula. Nonlinear regression with variance components. Journal of the American Statistical Association, 87(417):201–209, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290470.

Gutierrez-Pena:1995:CPN

[GPS95] E. Gutiérrez-Peña and A. F. M. Smith. Conjugate parameterizations for natural exponential families. Journal of the American Statistical Association, 90(432):1347–1356, December 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291525.

Guo:1992:EMP

[GR92a] Guo Guang and German Rodriguez. Estimating a multivariate proportional hazards model for clustered data using the EM algorithm, with an application to child survival in Guatemala. Journal of the American Statistical Association, 87(420):969–976, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290633.

Gwet:1992:ORA

[GR92b] Jean-Philippe Gwet and Louis-Paul Rivest. Outlier resistant alternatives to the ratio estimator. Journal of the American Statistical Association, 87(420):1174–1182, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290657.

Genest:1993:SIP

[GR93] Christian Genest and Louis-Paul Rivest. Statistical inference procedures for bivariate Archimedean copulas. Journal of the American Statistical Association, 88(423):1034–1043, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290796.

Gatto:1996:GSA

[GR96a] Riccardo Gatto and Elvezio Ronchetti. General saddlepoint approximations of marginal densities and tail probabilities. Journal of the American Statistical Association, 91(434):666–673, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291662.

Givens:1996:LAI

[GR96b] Geof H. Givens and Adrian E. Raftery. Local adaptive importance sampling for multivariate densities with strong nonlinear relationships. Journal of the American Statistical Association, 91 (433):132–141, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291389.

Geiser:1997:NTI

[GR97a] Peter W. Geiser and Ronald H. Randles. A nonparametric test of independence between two vectors. Journal of the American Statistical Association, 92(438):561–??, June 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gieser:1997:NTI

[GR97b] Peter W. Gieser and Ronald H. Randles. A nonparametric test of independence between two vectors. Journal of the American Statistical Association, 92(438):561–567, June 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2965704.

Graybill:1991:BRB

[Gra91] Franklin A. Graybill. Book review: [untitled]. Journal of the American Statistical Association, 86(413):247, March 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289743.

Gray:1992:FMA

[Gra92] Robert J. Gray. Flexible methods for analyzing survival data using splines, with applications to breast cancer prognosis. Journal of the American Statistical Association, 87(420):942–951, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290630.

Gray:1995:TVG

[Gra95] Robert J. Gray. Tests for variation over groups in survival data. Journal of the American Statistical Association, 90(429):198– 203, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291143.

Grambsch:1995:MLI

[GRB^{+}95] Patricia M. Grambsch, Bryan L. Randall, Roberd M. Bostick, John D. Potter, and Thomas A. Louis. Modeling the labeling index distribution: An application of functional data analysis. Journal of the American Statistical Association, 90(431):813–821, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291314.

Greenhouse:1991:BRB

[Gre91] Samuel W. Greenhouse. Book review: [untitled]. Journal of the American Statistical Association, 86(415):827–828, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290426.

Green:1992:SIA

[Gre92] Stephanie Green. Statistical issues arising in AIDS clinical trials: Comment. Journal of the American Statistical Association, 87 (418):571–572, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290293.

Grillenzoni:1990:MTV

[Gri90] Carlo Grillenzoni. Modeling time-varying dynamical systems. Journal of the American Statistical Association, 85(410):499–507, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289790.

Gripenberg:1992:CIP

[Gri92] G. Gripenberg. Confidence intervals for partial rank correlations. Journal of the American Statistical Association, 87(418):546–551, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290289.

Grillenzoni:1994:ORE

[Gri94] Carlo Grillenzoni. Optimal recursive estimation of dynamic models. Journal of the American Statistical Association, 89(427):777– 787, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290903.

Grogger:1990:DEC

[Gro90] Jeffrey Grogger. The deterrent effect of capital punishment: An analysis of daily homicide counts. Journal of the American Statistical Association, 85(410):295–303, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289764.

Gross:1992:BRB

[Gro92] Shulamith T. Gross. Book review: [untitled]. Journal of the American Statistical Association, 87(420):1247–1248, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290678.

Gilks:1998:AMC

[GRS98] Walter R. Gilks, Gareth O. Roberts, and Sujit K. Sahu. Adaptive Markov chain Monte Carlo through regeneration. Journal of the American Statistical Association, 93(443):1045–1054, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_98/sept/gilks.html;   http://www.jstor.
org/stable/2669848.

Grubel:1993:BRB

[Gru93] Rudolf Grubel. Book review: [untitled]. Journal of the American Statistical Association, 88(423):1168–1169, September 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290818.

Grubel:1998:BRB

[Gru98] Rudolf Grubel. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1245–1246, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669886.

Gelfand:1990:SBA

[GS90a] Alan E. Gelfand and Adrian F. M. Smith. Sampling-based approaches to calculating marginal densities. Journal of the American Statistical Association, 85(410):398–409, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289776.

Gordon:1990:MMB

[GS90b] K. Gordon and A. F. M. Smith. Modeling and monitoring biomedical time series. Journal of the American Statistical Association, 85(410):328–337, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289768.

Giles:1991:UEC

[GS91a] D. E. A. Giles and V. K. Srivastava. An unbiased estimator of the covariance matrix of the mixed regression estimator. Journal of the American Statistical Association, 86(414):441–444, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290591.

Green:1991:JST

[GS91b] Edwin J. Green and William E. Strawderman. A James–Stein type estimator for combining unbiased and possibly biased estimators. Journal of the American Statistical Association, 86(416): 1001–1006, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290517.

Ganio:1992:DO

[GS92] Lisa M. Ganio and Daniel W. Schafer. Diagnostics for overdispersion. Journal of the American Statistical Association, 87(419): 795–804, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290217.

Gonul:1993:CPB

[GS93a] Fusun Gonul and Kannan Srinivasan. Consumer purchase behavior in a frequently bought product category: Estimation issues and managerial insights from a hazard function model with heterogeneity. Journal of the American Statistical Association, 88 (424):1219–1227, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291260.

Grummer-Strawn:1993:RAC

[GS93b] Laurence M. Grummer-Strawn. Regression analysis of currentstatus data: An application to breast-feeding. Journal of the American Statistical Association, 88(423):758–765, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290760.

Giani:1994:TSE

[GS94] Guido Giani and Klaus Straßburger. Testing and selecting for equivalence with respect to a control. Journal of the American Statistical Association, 89(425):320–329, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291228.

Gurland:1995:HPF

[GS95] John Gurland and Jayaram Sethuraman. How pooling failure data may reverse increasing failure rates. Journal of the American Statistical Association, 90(432):1416–1423, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291533.

Giani:1997:OPP

[GS97] Guido Giani and Klaus Straßburger. Optimum partition procedures for separating good and bad treatments. Journal of the American Statistical Association, 92(437):291–298, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/march/giani.htm;   http://www.jstor.org/
stable/2291473.

Ge:1998:CHD

[GS98a] Nanxiang Ge and Douglas G. Simpson. Correlation and high dimensional consistency in pattern recognition. Journal of the American Statistical Association, 93(443):995–1006, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_98/sept/ge.html;   http://www.jstor.org/
stable/2669844.

Glickman:1998:SSM

[GS98b] Mark E. Glickman and Hal S. Stern. A state-space model for National Football League scores. Journal of the American Statistical Association, 93(441):25–35, March 1998. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
GLICKMAN.HTM;   http://www.jstor.org/stable/2669599.

Gelfand:1999:IIP

[GS99] Alan E. Gelfand and Sujit K. Sahu. Identifiability, improper priors, and Gibbs sampling for generalized linear models. Journal of the American Statistical Association, 94(445):247–253, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669699.

Gelfand:1992:BAC

[GSL92] Alan E. Gelfand, Adrian F. M. Smith, and Tai-Ming Lee. Bayesian analysis of constrained parameter and truncated data problems using Gibbs sampling. Journal of the American Statistical Association, 87(418):523–532, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2290286.

Guttorp:1991:ESO

[GT91] Peter Guttorp and Mary Lou Thompson. Estimating secondorder parameters of volcanicity from historical data. Journal of the American Statistical Association, 86(415):578–583, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290385.

Geyer:1995:AMC

[GT95] Charles J. Geyer and Elizabeth A. Thompson. Annealing Markov chain Monte Carlo with applications to ancestral inference. Journal of the American Statistical Association, 90(431):909–920, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291325.

Gallant:1998:RPO

[GT98] A. Ronald Gallant and George Tauchen. Reprojecting partially observed systems with application to interest rate diffusions. Journal of the American Statistical Association, 93 (441):10–24, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/GALLANT.HTM;   http://
www.jstor.org/stable/2669598.

Gu:1990:ASS

[Gu90] Chong Gu. Adaptive spline smoothing in non-Gaussian regression models. Journal of the American Statistical Association, 85(411): 801–807, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290018.

Gu:1992:DNR

[Gu92] Chong Gu. Diagnostics for nonparametric regression models with additive terms. Journal of the American Statistical Association, 87(420):1051–1058, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2290642.

Gu:1993:SSD

[Gu93] Chong Gu. Smoothing spline density estimation: a dimensionless automatic algorithm. Journal of the American Statistical Association, 88(422):495–504, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290329.

Guess:1996:BRB

[Gue96] Frank M. Guess. Book review: [untitled]. Journal of the American Statistical Association, 91(433):436, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291431.

Guilbaud:1993:EIA

[Gui93] Olivier Guilbaud. Exact inference about the within-subject variability in 2 × 2 crossover trials. Journal of the American Statistical Association, 88(423):939–946, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290785.

Guo:1993:BRB

[Guo93] Shumei Guo. Book review: [untitled]. Journal of the American Statistical Association, 88(424):1469, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291298.

Gurler:1996:BER

[Gür96] Ülkü Gürler. Bivariate estimation with right-truncated data. Journal of the American Statistical Association, 91(435):1152– 1165, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291734.

Gustafson:1996:LSI

[Gus96] Paul Gustafson. Local sensitivity of inferences to prior marginals. Journal of the American Statistical Association, 91(434):774–781, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291672.

Guthrie:1990:BRBa

[Gut90a] Donald Guthrie. Book review: [untitled]. Journal of the American Statistical Association, 85(409):262–263, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289569.

Guthrie:1990:BRBb

[Gut90b] Donald Guthrie. Book review: [untitled]. Journal of the American Statistical Association, 85(410):606, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289829.

Guttorp:1990:BRB

[Gut90c] Peter Guttorp. Book review: [untitled]. Journal of the American Statistical Association, 85(410):592, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289806.

Guthrie:1991:BRBa

[Gut91a] Donald Guthrie. Book review: [untitled]. Journal of the American Statistical Association, 86(414):549, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290614.

Guthrie:1991:BRBb

[Gut91b] Donald Guthrie. Book review: [untitled]. Journal of the American Statistical Association, 86(416):1151, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290551.

Guthrie:1992:BRB

[Gut92a] Donald Guthrie. Book review: [untitled]. Journal of the American Statistical Association, 87(420):1243, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290671.

Guttorp:1992:BRBa

[Gut92b] Peter Guttorp. Book review: [untitled]. Journal of the American Statistical Association, 87(418):586–587, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290302.

Guttorp:1992:BRBb

[Gut92c] Peter Guttorp. Book review: [untitled]. Journal of the American Statistical Association, 87(419):897, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290231.

Guthrie:1993:BRB

[Gut93a] Donald Guthrie. Book review: [untitled]. Journal of the American Statistical Association, 88(421):388, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290753.

Guttorp:1993:BRB

[Gut93b] Peter Guttorp. Book review: [untitled]. Journal of the American Statistical Association, 88(422):699–700, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290361.

Guthrie:1994:SS

[Gut94a] Donald Guthrie. Statistics in sports. Journal of the American Statistical Association, 89(427):1064–1065, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290935.

Guttorp:1994:ASS

[Gut94b] Peter Guttorp. An approach to statistical spatial-temporal modeling of meteorological fields: Comment. Journal of the American Statistical Association, 89(426):382–384, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290834.

Guthrie:1999:BRB

[Gut99] Donald Guthrie. Book review: [untitled]. Journal of the American Statistical Association, 94(448):1383–1384, December 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669951.

Godambe:1996:OER

[GV96] V. P. Godambe and K. Vijayan. Optimal estimation for responsedependent retrospective sampling. Journal of the American Statistical Association, 91(436):1724–1734, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291602.

Gray:1991:GMA

[GW91] H. L. Gray and Suojin Wang. A general method for approximating tail probabilities. Journal of the American Statistical Association, 86(413):159–166, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289726.

Glidden:1999:RET

[GW99] D. V. Glidden and L. J. Wei. Rank estimation of treatment differences based on repeated measurements subject to dependent censoring. Journal of the American Statistical Association, 94(447): 888–895, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2670004.

Guo:1999:SEA

[GWB99] Wensheng Guo, Yuedong Wang, and Morton B. Brown. A signal extraction approach to modeling hormone time series with pulses and a changing baseline. Journal of the American Statistical Association, 94(447):746–756, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669987.

Gong:1990:CSD

[GWG90] Gail Gong, Alice S. Whittemore, and Stella Grosser. Censored survival data with misclassified covariates: a case study of breastcancer mortality. Journal of the American Statistical Association, 85(409):20–28, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289521.

Gibbons:1998:HSU

[GWG98] Robert D. Gibbons and Virginia Wilcox-Gok. Health service utilization and insurance coverage: a multivariate probit analysis. Journal of the American Statistical Association, 93 (441):63–72, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/GIBBONS.HTM;   http://
www.jstor.org/stable/2669603.

Gu:1993:SAC

[GY93] Minggao Gu and Zhiliang Ying. Sequential analysis for censored regression data. Journal of the American Statistical Association, 88(423):890–898, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290778.

Helland:1994:CPM

[HA94] Inge S. Helland and Trygve Almøy. Comparison of prediction methods when only a few components are relevant. Journal of the American Statistical Association, 89(426):583–591, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290861. See cor-

rection [HA95].

Helland:1995:CCP

[HA95] Inge S. Helland and Trygve Almøy. Correction: “Comparison of prediction methods when only a few components are relevant”. Journal of the American Statistical Association, 90(429): 399, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2291191. See [HA94].

Haas:1990:LMW

[Haa90] Timothy C. Haas. Lognormal and moving window methods of estimating acid deposition. Journal of the American Statistical Association, 85(412):950–963, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289592.

Haas:1995:BRB

[Haa95a] Peter J. Haas. Book review: [untitled]. Journal of the American Statistical Association, 90(430):808–809, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291108.

Haas:1995:LPS

[Haa95b] Timothy C. Haas. Local prediction of a spatio-temporal process with an application to wet sulfate deposition. Journal of the American Statistical Association, 90(432):1189–1199, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291511.

Haas:1996:BRB

[Haa96] Timothy C. Haas. Book review: [untitled]. Journal of the American Statistical Association, 91(435):1375–1376, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291758.

Haberman:1991:MMG

[Hab91] Shelby J. Haberman. Measures, models, and graphical displays in the analysis of cross-classified data: Comment. Journal of the American Statistical Association, 86(416):1121–1123, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290534.

Haberman:1995:BRB

[Hab95a] Shelby J. Haberman. Book review: Statistical Applications Using Fuzzy Sets, by Kenneth G. Manton; Max A. Woodbury; H. Dennis Tolley. Journal of the American Statistical Association, 90(431): 1131–1133, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291362.

Haberman:1995:CML

[Hab95b] Shelby J. Haberman. Computation of maximum likelihood estimates in association models. Journal of the American Statistical Association, 90(432):1438–1446, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291536.

Hagenaars:1997:BRB

[Hag97] Jacques Hagenaars. Book review: [untitled]. Journal of the American Statistical Association, 92(438):797, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965739.

Hahn:1990:ITF

[Hah90] Robert A. Hahn. Interactional troubles in face-to-face survey interviews: Comment. Journal of the American Statistical Association, 85(409):244–246, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289552.

Hall:1994:EBE

[Hal94] Peter Hall. On the erratic behavior of estimators of N in the binomial N, p distribution. Journal of the American Statistical Association, 89(425):344–352, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291231.

Halloran:1997:BRB

[Hal97] M. Elizabeth Halloran. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1640–1642, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2965436.

Hamer:1990:BRB

[Ham90a] Robert M. Hamer. Book review: [untitled]. Journal of the American Statistical Association, 85(409):268–269, March 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289579.

Hamming:1990:BRU

[Ham90b] R. W. Hamming. Book review: [untitled]. Journal of the American Statistical Association, 85(410):592, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289807.

Hamer:1993:BRB

[Ham93] Robert M. Hamer. Book review: [untitled]. Journal of the American Statistical Association, 88(423):1173, September 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290826.

Handcock:1994:ASSb

[Han94] Mark S. Handcock. An approach to statistical spatial-temporal modeling of meteorological fields: Rejoinder. Journal of the American Statistical Association, 89(426):388–390, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290836.

Handcock:1999:CPS

[Han99a] Mark S. Handcock. Comment: Prediction of spatial cumulative distribution functions using subsampling. Journal of the American Statistical Association, 94(445):100–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Handcock:1999:PSC

[Han99b] Mark S. Handcock. Prediction of spatial cumulative distribution functions using subsampling: Comment. Journal of the American Statistical Association, 94(445):100–102, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669683.

Harder:1990:ECV

[Har90a] Melinda H. Harder. Effects of censoring on the validity of confidence intervals. Journal of the American Statistical Association, 85(409):90–96, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289529.

Harris:1990:RDI

[Har90b] Jeffrey E. Harris. Reporting delays and the incidence of AIDS. Journal of the American Statistical Association, 85(412):915–924, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289588.

Harville:1991:IBR

[Har91] David A. Harville. Interpreting blocks and random factors: Comment. Journal of the American Statistical Association, 86(415): 812–815, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290418.

Harder:1992:MDE

[Har92a] Melinda H. Harder. Minimum distance estimators of scale with censored data. Journal of the American Statistical Association, 87(419):832–843, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290222.

Harrington:1992:SIA

[Har92b] Mark Harrington. Statistical issues arising in AIDS clinical trials: Comment. Journal of the American Statistical Association, 87 (418):573–576, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290294.

Harris:1996:BRB

[Har96] Bernard Harris. Book review: [untitled]. Journal of the American Statistical Association, 91(434):914–915, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291696.

Hobert:1997:AFS

[HAS97] James P. Hobert, N. S. Altman, and Carl L. Schofield. Analyses of fish species richness with spatial covariate. Journal of the American Statistical Association, 92(439):846–854, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/sept/hobert.htm;   http://www.jstor.org/
stable/2965549.

Hauptly:1990:BRB

[Hau90a] Denis J. Hauptly. Book review: [untitled]. Journal of the American Statistical Association, 85(410):591, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289804.

Hausch:1990:BRB

[Hau90b] Donald B. Hausch. Book review: [untitled]. Journal of the American Statistical Association, 85(411):911–912, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290051.

Haunsperger:1992:DPS

[Hau92] Deanna B. Haunsperger. Dictionaries of paradoxes for statistical tests on k samples. Journal of the American Statistical Association, 87(417):149–155, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2290463.

Haugh:1997:BRBa

[Hau97a] Larry D. Haugh. Book review: [untitled]. Journal of the American Statistical Association, 92(437):388, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291492.

Haugh:1997:BRBb

[Hau97b] Larry D. Haugh. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1221, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965598.

Have:1995:BRB

[Hav95] Thomas R. Ten Have. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1123–1124, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291352.

Hawkins:1993:AES

[Haw93a] Douglas M. Hawkins. The accuracy of elemental set approximations for regression. Journal of the American Statistical Association, 88(422):580–589, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290339.

Hawkins:1993:BRB

[Haw93b] Douglas M. Hawkins. Book review: [untitled]. Journal of the American Statistical Association, 88(422):695–696, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290355.

Hawkins:1996:BRB

[Haw96] Douglas M. Hawkins. Book review: [untitled]. Journal of the American Statistical Association, 91(433):435–436, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291430.

Hayter:1990:OSS

[Hay90] A. J. Hayter. A one-sided Studentized range test for testing against a simple ordered alternative. Journal of the American Statistical Association, 85(411):778–785, September 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290015.

Hayakawa:1994:CNB

[Hay94] Yu Hayakawa. The construction of new bivariate exponential distributions from a Bayesian perspective. Journal of the American Statistical Association, 89(427):1044–1049, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290932.

Hazeghi:1995:BRB

[Haz95] Kasra Hazeghi. Book review: [untitled]. Journal of the American Statistical Association, 90(429):392–393, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291177.

Hsu:1999:SCI

[HB99] Jason C. Hsu and Roger L. Berger. Stepwise confidence intervals without multiplicity adjustment for dose response and toxicity studies. Journal of the American Statistical Association, 94(446): 468–482, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2670167.

Heiberger:1993:ODA

[HBH93] Richard M. Heiberger, Dulal K. Bhaumik, and Burt Holland. Optimal data augmentation strategies for additive models. Journal of the American Statistical Association, 88(423):926–938, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290784.

Hobert:1996:EIP

[HC96] James P. Hobert and George Casella. The effect of improper priors on Gibbs sampling in hierarchical linear mixed models. Journal of the American Statistical Association, 91(436):1461–1473, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291572.

Hwang:1997:PIA

[HD97] J. T. Gene Hwang and A. Adam Ding. Prediction intervals for artificial neural networks. Journal of the American Statistical Association, 92(438):748–757, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2965723.

Higgins:1997:TSA

[HDG97] Karen M. Higgins, Marie Davidian, and David M. Giltinan. A two-step approach to measurement error in time-dependent covariates in nonlinear mixed-effects models, with application to IGF–I pharmacokinetics. Journal of the American Statistical Association, 92(438):436–448, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2965691.

He:1990:APE

[He90] Kun He. An ancillarity paradox in the estimation of multinomial probabilities. Journal of the American Statistical Association, 85 (411):824–828, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290021.

He:1999:CRD

[He99a] Xuming He. Comment: Regression depth. Journal of the American Statistical Association, 94(446):403–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

He:1999:RDC

[He99b] Xuming He. Regression depth: Comment. Journal of the American Statistical Association, 94(446):403–404, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670156.

Heald:1990:BRB

[Hea90] Mark A. Heald. Book review: [untitled]. Journal of the American Statistical Association, 85(409):266–267, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289575.

Heckman:1996:CIC

[Hec96a] James J. Heckman. Comment: Identification of causal effects using instrumental variables. Journal of the American Statistical Association, 91(434):444–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Heckman:1996:ICE

[Hec96b] James J. Heckman. Identification of causal effects using instrumental variables: Comment. Journal of the American Statistical Association, 91(434):459–462, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291631.

Heiberger:1990:BRB

[Hei90] Richard M. Heiberger. Book review: [untitled]. Journal of the American Statistical Association, 85(409):267–268, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289577.

Heitjan:1991:NMS

[Hei91] Daniel F. Heitjan. Nonlinear modeling of serial immunologic data: a case study. Journal of the American Statistical Association, 86 (416):891–898, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290502.

Heitjan:1995:BAP

[Hei95] Daniel F. Heitjan. Bayesian analysis of a protocol for sampling preserved tumor specimens. Journal of the American Statistical Association, 90(429):38–44, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291127.

Herzig:1991:BRB

[Her91] Abbe H. Herzig. Book review: [untitled]. Journal of the American Statistical Association, 86(414):543, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290606.

Herting:1993:BRB

[Her93] Jerald R. Herting. Book review: [untitled]. Journal of the American Statistical Association, 88(424):1471, December 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291301.

Hertzberg:1997:BRB

[Her97] Vicki Stover Hertzberg. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1222–1223, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2965601.

Herrmann:1998:BRB

[Her98] Douglas Herrmann. Book review: [untitled]. Journal of the American Statistical Association, 93(441):398–399, March 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669637.

Hall:1990:BTD

[HH90] Peter Hall and Jeffrey D. Hart. Bootstrap test for difference between means in nonparametric regression. Journal of the American Statistical Association, 85(412):1039–1049, December 1990.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289600.

Hulting:1991:SBN

[HH91] Frederick L. Hulting and David A. Harville. Some Bayesian and non-Bayesian procedures for the analysis of comparative experiments and for small-area estimation: Computational aspects, frequentist properties, and relationships. Journal of the American Statistical Association, 86(415):557–568, September 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290383.

Hayter:1994:RBS

[HH94] Anthony J. Hayter and Jason C. Hsu. On the relationship between stepwise decision procedures and confidence sets. Journal of the American Statistical Association, 89(425):128–136, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291208.

Horowitz:1996:DSE

[HH96] Joel L. Horowitz and Wolfgang Härdle. Direct semiparametric estimation of single-index models with discrete covariates. Journal of the American Statistical Association, 91(436):1632–1640, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291590.

Hougaard:1992:MSB

[HHH92] Philip Hougaard, Bent Harvald, and Niels V. Holm. Measuring the similarities between the lifetimes of adult Danish twins born between 1881–1930. Journal of the American Statistical Association, 87(417):17–24, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290447.

Hardle:1992:RSP

[HHM92] W. Härdle, P. Hall, and J. S. Marron. Regression smoothing parameters that are not far from their optimum. Journal of the American Statistical Association, 87(417):227–233, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290473.

hardle:1993:CBC

[hHMT93] W. hardle, J. Hart, J. S. Marron, and A. B. Tsybakov. Correction: Bandwidth choice for average derivative estimation. Journal of the American Statistical Association, 88(422):718, June 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290380.

Hall:1997:CMO

[HHS97] Peter Hall, Catherine Huber, and Paul L. Speckman. Covariatematched one-sided tests for the difference between functional means. Journal of the American Statistical Association, 92 (439):1074–1083, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

amstat.org/publications/jasa/abstracts_97/sept/hall_p.
htm;   http://www.jstor.org/stable/2965572.

Higdon:1998:AVM

[Hig98a] David M. Higdon. Auxiliary variable methods for Markov chain Monte Carlo with applications. Journal of the American Statistical Association, 93(442):585–595, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_98/june/
HIGDON.HTM;   http://www.jstor.org/stable/2670110.

Higgins:1998:BRB

[Hig98b] James J. Higgins. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1249, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669891.

Hillmer:1991:BRB

[Hil91] Steven C. Hillmer. Book review: [untitled]. Journal of the American Statistical Association, 86(413):245–246, March 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289741.

Hildebrand:1992:BRB

[Hil92] David K. Hildebrand. Book review: [untitled]. Journal of the American Statistical Association, 87(418):588, June 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290304.

Hill:1993:RSE

[Hil93] Daniel H. Hill. Response and sequencing errors in surveys: a discrete contagious regression analysis. Journal of the American Statistical Association, 88(423):775–781, September 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290762.

Hilton:1995:BRB

[Hil95] Joan Hilton. Book review: [untitled]. Journal of the American Statistical Association, 90(430):800, June 1995. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2291096.

Hipp:1995:BRB

[Hip95] Christina Hipp. Book review: [untitled]. Journal of the American Statistical Association, 90(429):392, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291176.

Hirji:1992:CED

[Hir92] Karim F. Hirji. Computing exact distributions for polytomous response data. Journal of the American Statistical Association, 87(418):487–492, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290281.

Harville:1992:MSE

[HJ92] David A. Harville and Daniel R. Jeske. Mean squared error of estimation or prediction under a general linear model. Journal of the American Statistical Association, 87(419):724–731, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290210.

Hjort:1994:EAN

[Hjo94] Nils Lid Hjort. The exact amount of t-ness that the normal model can tolerate. Journal of the American Statistical Association, 89 (426):665–675, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290869.

Hendricks:1992:HSM

[HK92] Wallace Hendricks and Roger Koenker. Hierarchical spline models for conditional quantiles and the demand for electricity. Journal of the American Statistical Association, 87(417):58–68, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290452.

Harvey:1993:FHE

[HK93] Andrew Harvey and Siem Jan Koopman. Forecasting hourly electricity demand using time-varying splines. Journal of the American Statistical Association, 88(424):1228–1236, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291261.

Hansen:1998:TM

[HKS98] Mark Hansen, Charles Kooperberg, and Sylvain Sardy. Triogram models. Journal of the American Statistical Association, 93(441):101–119, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/HANSEN.HTM;   http:
//www.jstor.org/stable/2669608.

Heitjan:1994:AST

[HL94] Daniel F. Heitjan and J. Richard Landis. Assessing secular trends in blood pressure: a multiple- imputation approach. Journal of the American Statistical Association, 89(427):750–759, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290900.

Hu:1996:ERM

[HL96] X. Joan Hu and Jerald F. Lawless. Estimation of rate and mean functions from truncated recurrent event data. Journal of the American Statistical Association, 91(433):300–310, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291408.

Huffer:1997:ADS

[HL97] Fred W. Huffer and Chien-Tai Lin. Approximating the distribution of the scan statistic using moments of the number of clumps. Journal of the American Statistical Association, 92(440): 1466–1475, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/HUFFER.HTM;
http://www.jstor.org/stable/2965417.

Heagerty:1998:CLA

[HL98] Patrick J. Heagerty and Subhash R. Lele. A composite likelihood approach to binary spatial data. Journal of the American Statistical Association, 93(443):1099–1111, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
sept/heagerty.html;   http://www.jstor.org/stable/2669853.

Hines:1992:DCM

[HLC92] R. J. O’Hara Hines, J. F. Lawless, and E. M. Carter. Diagnostics for a cumulative multinomial generalized linear model, with applications to grouped toxicological mortality data. Journal of the American Statistical Association, 87(420):1059–1069, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290643.

Huffer:1991:WLS

[HM91] Fred W. Huffer and Ian W. McKeague. Weighted least squares estimation for Aalen’s additive risk model. Journal of the American Statistical Association, 86(413):114–129, March 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289721.

Hawkins:1997:HBL

[HM97] Douglas M. Hawkins and Geoffrey J. McLachlan. High-breakdown linear discriminant analysis. Journal of the American Statistical Association, 92(437):136–143, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/march/
hawkins.htm;   http://www.jstor.org/stable/2291457.

Hoover:1993:UED

[HMC^{+}93] D. R. Hoover, A. Munoz, V. Carey, J. M. G. Taylor, M. Van-Raden, J. S. Chmiel, L. Kingsley, and The Multicenter AIDS Cohort Study. Using events from dropouts in nonparametric survival function estimation with application to incubation of AIDS. Journal of the American Statistical Association, 88(421):37–43, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290689.

Hardle:1998:TPV

[HMM98] Wolfgang Härdle, Enno Mammen, and Marlene Müller. Testing parametric versus semiparametric modeling in generalized linear models. Journal of the American Statistical Association, 93 (444):1461–1474, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/dec/hardle.html;
http://www.jstor.org/stable/2670060.

Handcock:1994:KSE

[HMN94] Mark S. Handcock, Kristen Meier, and Douglas Nychka. Kriging and splines: An empirical comparison of their predictive performance in some applications: Comment. Journal of the American Statistical Association, 89(426):401–403, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290838.

Hettmansparger:1997:AIM

[HMO97a] Thomas P. Hettmansparger, Jyrki Mottonen, and Hannu Oja. Affine-invariant multivariate one-sample signed-rank tests. Journal of the American Statistical Association, 92(440):1591– 1600, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/HETTMANS.HTM.

Hettmansperger:1997:AIM

[HMO97b] Thomas P. Hettmansperger, Jyrki Mottonen, and Hannu Oja. Affine-invariant multivariate one-sample signed-rank tests. Journal of the American Statistical Association, 92(440):1591– 1600, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/HETTMANS.HTM;
 http://www.jstor.org/stable/2965430.

Hardle:1992:BCA

[HMT92] W. Härdle, J. S. Marron, and A. B. Tsybakov. Bandwidth choice for average derivative estimation. Journal of the American Statistical Association, 87(417):218–226, March 1992. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2290472.

Hollander:1997:LRB

[HMY97] Myles Hollander, Ian W. McKeague, and Jie Yang. Likelihood ratio — based confidence bands for survival functions. Journal of the American Statistical Association, 92(437):215–226, March 1997.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/march/hollander.htm;   http://www.jstor.
org/stable/2291466.

Horowitz:1992:GMS

[HN92] Joel L. Horowitz and George R. Neumann. A generalized moments specification test of the proportional hazards model. Journal of the American Statistical Association, 87(417):234–240, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290474.

Hocking:1991:IBR

[Hoc91] R. R. Hocking. Interpreting blocks and random factors: Comment. Journal of the American Statistical Association, 86(415): 811, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290417.

Hodges:1991:BRB

[Hod91] James S. Hodges. Book review: [untitled]. Journal of the American Statistical Association, 86(414):547–548, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290611.

Hodges:1994:AUD

[Hod94] James S. Hodges. Analytical use of data from Army training exercises: a case study of tactical reconnaissance. Journal of the American Statistical Association, 89(426):444–451, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290844.

Hoffman:1992:CLB

[Hof92] Lorrie L. Hoffman. Characterizing linear birth and death processes. Journal of the American Statistical Association, 87(420): 1183–1187, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290658.

Hogan:1993:PES

[Hog93] Howard Hogan. The 1990 post-enumeration survey: Operations and results. Journal of the American Statistical Association, 88 (423):1047–1060, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290798.

Holmgren:1995:PMC

[Hol95] Eric B. Holmgren. The P –P plot as a method for comparing treatment effects. Journal of the American Statistical Association, 90(429):360–365, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291161.

Holford:1998:BRB

[Hol98] Theodore R. Holford. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1242, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669879.

Holm:1999:MCS

[Hol99a] Sture Holm. Multiple confidence sets based on stagewise tests. Journal of the American Statistical Association, 94(446):489–495, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670169.

Holmes:1999:BRBa

[Hol99b] Susan Holmes. Book review: [untitled]. Journal of the American Statistical Association, 94(447):981–982, September 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670014.

Holmes:1999:BRBb

[Hol99c] Susan P. Holmes. Book review: [untitled]. Journal of the American Statistical Association, 94(447):989, September 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670026.

Hughes-Oliver:1998:AUS

[HOLDG98] Jacqueline M. Hughes-Oliver, Jye-Chyi Lu, Joseph C. Davis, and Ronald S. Gyurcsik. Achieving uniformity in a semiconductor fabrication process using spatial modeling. Journal of the American Statistical Association, 93(441):36–45, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
HUGHES-O.HTM;   http://www.jstor.org/stable/2669600.

Hong:1999:HTT

[Hon99] Yongmiao Hong. Hypothesis testing in time series via the empirical characteristic function: a generalized spectral density approach. Journal of the American Statistical Association, 94(448): 1201–1220, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2669935.

Hooper:1993:IWL

[Hoo93] Peter M. Hooper. Iterative weighted least squares estimation in heteroscedastic linear models. Journal of the American Statistical Association, 88(421):179–184, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290712.

Hossjer:1994:RBE

[Hös94a] Ola Hössjer. Rank-based estimates in the linear model with high breakdown point. Journal of the American Statistical Association, 89(425):149–158, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291211.

Hughes-Oliver:1994:TSA

[HOS94b] Jacqueline M. Hughes-Oliver and William H. Swallow. A twostage adaptive group-testing procedure for estimating small proportions. Journal of the American Statistical Association, 89 (427):982–993, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290924.

Host:1995:SIE

[HOS95] Gudmund Høst, Henning Omre, and Paul Switzer. Spatial interpolation errors for monitoring data. Journal of the American Statistical Association, 90(431):853–861, September 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291319.

Hossjer:1997:BRB

[Hos97] Ola Hossjer. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1652, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965449.

Houdre:1995:SAC

[Hou95] Christian Houdré. Some applications of covariance identities and inequalities to functions of multivariate normal variables. Journal of the American Statistical Association, 90(431):965–968, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291332.

Hollander:1992:CSG

[HP92] Myles Hollander and Edsel A. Peña. A chi-squared goodness-of-fit test for randomly censored data. Journal of the American Statistical Association, 87(418):458–463, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290277.

Hardwick:1998:SDB

[HPS98] Janis Hardwick, Connie Page, and Quentin F. Stout. Sequentially deciding between two experiments for estimating a common success probability. Journal of the American Statistical Association, 93(444):1502–1511, December 1998. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
dec/hardwick.html;   http://www.jstor.org/stable/2670063.

Heitjan:1990:ICD

[HR90] Daniel F. Heitjan and Donald B. Rubin. Inference from coarse data via multiple imputation with application to age heaping. Journal of the American Statistical Association, 85(410):304–314, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289765.

Heritier:1994:RBI

[HR94] Stéphane Heritier and Elvezio Ronchetti. Robust boundedinfluence tests in general parametric models. Journal of the American Statistical Association, 89(427):897–904, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290914.

Huang:1997:SEP

[HR97] Jian Huang and A. J. Rossini. Sieve estimation for the proportional-odds failure-TIme regression model with interval censoring. Journal of the American Statistical Association, 92(439):960–967, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

amstat.org/publications/jasa/abstracts_97/sept/huang.
htm;   http://www.jstor.org/stable/2965559.

Hale:1996:BRB

[HRF96] Michael D. Hale, Russell L. Reeve, and Carol A. Francisco. Book review: [untitled]. Journal of the American Statistical Association, 91(434):911, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291691.

Heckman:1990:TME

[HRW90] James J. Heckman, Richard Robb, and James R. Walker. Testing the mixture of exponentials hypothesis and estimating the mixing distribution by the method of moments. Journal of the American Statistical Association, 85(410):582–589, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289802.

Hadi:1993:PIM

[HS93a] Ali S. Hadi and Jeffrey S. Simonoff. Procedures for the identification of multiple outliers in linear models. Journal of the American Statistical Association, 88(424):1264–1272, December 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291266.

Hengartner:1993:ABBa

[HS93b] N. Hengartner and T. P. Speed. Assessing between-block heterogeneity within the post-strata of the 1990 post-enumeration survey. Journal of the American Statistical Association, 88(423): 1119–1125, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290806.

Hengartner:1993:ABBb

[HS93c] N. Hengartner and T. P. Speed. Assessing between-block heterogeneity within the post-strata of the 1990 post-enumeration survey: Rejoinder. Journal of the American Statistical Association, 88(423):1128–1129, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290809.

Huggins:1994:VCM

[HS94] Richard M. Huggins and Robert G. Staudte. Variance components models for dependent cell populations. Journal of the American Statistical Association, 89(425):19–29, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291197.

Haas:1998:ENC

[HS98a] Peter J. Haas and Lynne Stokes. Estimating the number of classes in a finite population. Journal of the American Statistical Association, 93(444):1475–1487, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
dec/haas.html;   http://www.jstor.org/stable/2670061.

Hall:1998:EGE

[HS98b] Daniel B. Hall and Thomas A. Severini. Extended generalized estimating equations for clustered data. Journal of the American Statistical Association, 93(444):1365–1375, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
dec/hall.html;   http://www.jstor.org/stable/2670052.

He:1998:MBS

[HS98c] Xuming He and Peide Shi. Monotone B-spline smoothing. Journal of the American Statistical Association, 93(442):643–650, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670115.

He:1998:MSS

[HS98d] Xuming He and Peide Shi. Monotone B-spline smoothing. Journal of the American Statistical Association, 93(442):643–650, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_98/june/HE.HTM.

Hsiang:1990:BRB

[Hsi90] Thomas C. Hsiang. Book review: [untitled]. Journal of the American Statistical Association, 85(410):601, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289822.

Hsiao:1997:ABF

[Hsi97] Chuhsing Kate Hsiao. Approximate Bayes factors when a mode occurs on the boundary. Journal of the American Statistical Association, 92(438):656–663, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2965713.

He:1990:BRT

[HSP90] Xuming He, Douglas G. Simpson, and Stephen L. Portnoy. Breakdown robustness of tests. Journal of the American Statistical Association, 85(410):446–452, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289782.

Hansen:1990:REFa

[HT90a] Morris H. Hansen and Benjamin J. Tepping. Regression estimates in Federal welfare quality control programs. Journal of the American Statistical Association, 85(411):856–864, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290026.

Hansen:1990:REFb

[HT90b] Morris H. Hansen and Benjamin J. Tepping. Regression estimates in Federal Welfare Quality Control Programs: Rejoinder. Journal of the American Statistical Association, 85(411):871–873, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290031.

Hardle:1991:SIR

[HT91] W. Hardle and A. B. Tsybakov. Sliced inverse regression for dimension reduction: Comment. Journal of the American Statistical Association, 86(414):333–335, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290566.

Hall:1997:IMAa

[HT97a] Peter Hall and Berwin A. Turlach. Interpolation methods for adapting to sparse design in nonparametric regression. Journal of the American Statistical Association, 92(438): 466–472, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/june/hall.htm;   http:/
/www.jstor.org/stable/2965694.

Hall:1997:IMAb

[HT97b] Peter Hall and Berwin A. Turlach. Interpolation methods for adapting to sparse design in nonparametric: Regression rejoinder. Journal of the American Statistical Association, 92(438):476, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2965696.

Hall:1997:RIM

[HT97c] Peter Hall and Berwin A. Turlach. Rejoinder: Interpolation methods for adapting to sparse design in nonparametric regression. Journal of the American Statistical Association, 92(438):476–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hsiao:1997:PAL

[HT97d] Cheng Hsiao and A. Kamil Tahmiscioğlu. A panel analysis of liquidity constraints and firm investment. Journal of the American Statistical Association, 92(438):455–465, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965693.

Hastie:1994:FDA

[HTB94] Trevor Hastie, Robert Tibshirani, and Andreas Buja. Flexible discriminant analysis by optimal scoring. Journal of the American Statistical Association, 89(428):1255–1270, December 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290989.

Hughes:1995:OSS

[HTP95] Michael D. Hughes, Simon G. Thompson, and Stuart J. Pocock. Optimal sequential screening guidelines for quantitative risk factors measured with error. Journal of the American Statistical Association, 90(429):19–26, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2291125.

Huang:1997:TGF

[Hua97] Li-Shan Huang. Testing goodness-of-fit based on a roughness measure. Journal of the American Statistical Association, 92(440): 1399–1402, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/HUANG_LS.HTM;
 http://www.jstor.org/stable/2965409.

Hunsberger:1994:SRL

[Hun94a] Sally Hunsberger. Semiparametric regression in likelihood-based models. Journal of the American Statistical Association, 89(428): 1354–1365, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290997.

Hunter:1994:SP

[Hun94b] J. Stuart Hunter. Statistics as a profession. Journal of the American Statistical Association, 89(425):1–6, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291195.

Hadi:1990:AEM

[HW90a] Ali S. Hadi and Martin T. Wells. Assessing the effects of multiple rows on the condition number of a matrix. Journal of the American Statistical Association, 85(411):786–792, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290016.

Heckman:1990:EFD

[HW90b] James J. Heckman and James R. Walker. Estimating fecundability from data on waiting times to first conception. Journal of the American Statistical Association, 85(410):283–294, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289763.

Hall:1991:GMR

[HW91] Peter Hall and Thomas E. Wehrly. A geometrical method for removing edge effects from kernel- type nonparametric regression estimators. Journal of the American Statistical Association, 86 (415):665–672, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290395.

Hart:1992:KRW

[HW92a] Jeffrey D. Hart and Thomas E. Wehrly. Kernel regression when the boundary region is large, with an application to testing the adequacy of polynomial models. Journal of the American Statistical Association, 87(420):1018–1024, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290639.

Hasofer:1992:TEV

[HW92b] A. M. Hasofer and Z. Wang. A test for extreme value domain of attraction. Journal of the American Statistical Association, 87 (417):171–177, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290466.

Handcock:1994:ASSa

[HW94] Mark S. Handcock and James R. Wallis. An approach to statistical spatial-temporal modeling of meteorological fields. With comments and a rejoinder by Handcock. Journal of the American Statistical Association, 89(426):368–378, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290832.

Hamada:1995:ACD

[HW95a] M. Hamada and C. F. J. Wu. Analysis of censored data from fractionated experiments: a Bayesian approach. Journal of the American Statistical Association, 90(430):467–477, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291057.

Huang:1995:EOR

[HW95b] Ying Huang and Mei Cheng Wang. Estimating the occurrence rate for prevalent survival data in competing risks models. Journal of the American Statistical Association, 90(432):1406–1415, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291532.

Hall:1999:MEC

[HWY99] Peter Hall, Rodney C. L. Wolff, and Qiwei Yao. Methods for estimating a conditional distribution function. Journal of the American Statistical Association, 94(445):154–163, March 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669691.

Hart:1998:OSC

[HY98] Jeffrey D. Hart and Seongbaek Yi. One-sided cross-validation. Journal of the American Statistical Association, 93(442):620–631, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_98/june/HART.HTM;   http://www.jstor.org/
stable/2670113.

Hedayat:1997:MIO

[HYP97] A. S. Hedayat, Bo Yan, and John M. Pezzuto. Modeling and identifying optimum designs for fitting dose- response curves based on raw optical density data. Journal of the American Statistical Association, 92(439):1132–1140, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/sept/hedayat.htm;   http://www.jstor.org/
stable/2965578.

Hurvich:1990:FDS

[HZ90] Clifford M. Hurvich and Scott L. Zeger. A frequency domain selection criterion for regression with autocorrelated errors. Journal of the American Statistical Association, 85(411):705–714, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290006.

Heagerty:1996:MRM

[HZ96] Patrick J. Heagerty and Scott L. Zeger. Marginal regression models for clustered ordinal measurements. Journal of the American Statistical Association, 91(435):1024–1036, September 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291722.

Heagerty:1998:LRA

[HZ98] Patrick J. Heagerty and Scott L. Zeger. Lorelogram: a regression approach to exploring dependence in longitudinal categorical responses. Journal of the American Statistical Association, 93 (441):150–162, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/HEAGERTY.HTM;   http://
www.jstor.org/stable/2669612.

Haseman:1990:PRC

[HZS^{+}90] Joseph K. Haseman, Errol Zeiger, Michael D. Shelby, Barry H. Margolin, and Raymond W. Tennant. Predicting rodent carcinogenicity from four in vitro genetic toxicity assays: An evaluation of 114 chemicals studied by the National Toxicology Program. Journal of the American Statistical Association, 85(412): 964–971, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289593.

Ibrahim:1990:IDG

[Ibr90] Joseph G. Ibrahim. Incomplete data in generalized linear models. Journal of the American Statistical Association, 85(411):765–769, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290013.

Iglewicz:1991:BRB

[Igl91] Boris Iglewicz. Book review: [untitled]. Journal of the American Statistical Association, 86(415):833–834, September 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290433.

Iglewicz:1992:BRB

[Igl92] Boris Iglewicz. Book review: [untitled]. Journal of the American Statistical Association, 87(419):904, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290241.

Ibrahim:1991:BAG

[IL91] Joseph G. Ibrahim and Purushottam W. Laud. On Bayesian analysis of generalized linear models using Jeffreys’s prior. Journal of the American Statistical Association, 86(416):981–986, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290514.

Ibrahim:1994:PAA

[IL94] Joseph G. Ibrahim and Purushottam W. Laud. A predictive approach to the analysis of designed experiments. Journal of the American Statistical Association, 89(425):309–319, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291227.

Iman:1995:NPS

[Ima95] Ronald L. Iman. New paradigms for the statistics profession. Journal of the American Statistical Association, 90(429): 1–6, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291123.

Ibrahim:1998:ACS

[IRC98a] Joseph G. Ibrahim, Louise M. Ryan, and Ming-Hui Chen. Applications and case studies — using historical controls to adjust for covariates in trend tests for binary data. Journal of the American Statistical Association, 93(444):1282–1293, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/dec/ibrahim.html.

Ibrahim:1998:UHC

[IRC98b] Joseph G. Ibrahim, Louise M. Ryan, and Ming-Hui Chen. Using historical controls to adjust for covariates in trend tests for binary data. Journal of the American Statistical Association, 93(444): 1282–1293, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2670043.

Inclan:1994:UCS

[IT94] Carla Inclán and George C. Tiao. Use of cumulative sums of squares for retrospective detection of changes of variance. Journal of the American Statistical Association, 89(427):913–923, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290916.

Iyengar:1997:BRBa

[Iye97a] Satish Iyengar. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1220–1221, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965597.

Iyengar:1997:BRBb

[Iye97b] Satish Iyengar. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1651, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965447.

Izenman:1991:RDN

[Ize91] Alan Julian Izenman. Recent developments in nonparametric density estimation. Journal of the American Statistical Association, 86(413):205–224, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289732.

Jacroux:1993:CTR

[Jac93] Mike Jacroux. On the construction of trend-resistant designs for comparing a set of test treatments with a set of controls. Journal of the American Statistical Association, 88(424):1398–1403, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291283.

Jammalamadaka:1998:BRB

[Jam98] S. Rao Jammalamadaka. Book review: [untitled]. Journal of the American Statistical Association, 93(441):404–405, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669647.

Janssen:1995:BRB

[Jan95] Paul Janssen. Book review: [untitled]. Journal of the American Statistical Association, 90(430):804–805, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291103.

Jasso:1992:BRBa

[Jas92a] Guillermina Jasso. Book review: [untitled]. Journal of the American Statistical Association, 87(418):592, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290311.

Jasso:1992:BRBb

[Jas92b] Guillermina Jasso. Book review: [untitled]. Journal of the American Statistical Association, 87(419):904–905, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290242.

Jayasuriya:1996:TPR

[Jay96] Bodhini R. Jayasuriya. Testing for polynomial regression using nonparametric regression techniques. Journal of the American Statistical Association, 91(436):1626–1631, December 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291589.

Ji:1992:BCV

[JB92] Zhang Ji and Dennis D. Boos. Bootstrap critical values for testing homogeneity of covariance matrices. Journal of the American Statistical Association, 87(418):425–??, June 1992. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jae:1991:SCC

[JD91] Won Lee Jae and David L. DeMets. Sequential comparison of changes with repeated measurements data. Journal of the American Statistical Association, 86(415):757–??, September 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jones:1994:VKT

[JDP94] M. C. Jones, S. J. Davies, and B. U. Park. Versions of kernel-type regression estimators. Journal of the American Statistical Association, 89(427):825–832, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290908.

JEB:1998:BRB

[JEB98] JEB. Book review: [untitled]. Journal of the American Statistical Association, 93(442):844, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670148.

Jensen:1991:LDT

[Jen91] Jens Ledet Jensen. A large deviation-type approximation for the “Box Class” of likelihood ratio criteria. Journal of the American Statistical Association, 86(414):437–440, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290590.

Jensen:1995:CLD

[Jen95] Jens Ledet Jensen. Correction: a large deviation-type approximation for the ’box class’ of likelihood ratio criteria. Journal of the American Statistical Association, 90(430):812, June 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291120.

JFM:1998:BRB

[JFM98] JFM. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1529, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670081.

JGB:1999:BRB

[JGB99] JGB. Book review: [untitled]. Journal of the American Statistical Association, 94(446):657, June 1999. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670203.

Jacqmin-Gadda:1995:THG

[JGC95] Hélène Jacqmin-Gadda and Daniel Commenges. Tests of homogeneity for generalized linear models. Journal of the American Statistical Association, 90(432):1237–1246, December 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291514.

Jiang:1998:CEG

[Jia98] Jiming Jiang. Consistent estimators in generalized linear mixed models. Journal of the American Statistical Association, 93 (442):720–729, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/june/JIANG.HTM;   http:
//www.jstor.org/stable/2670122.

Jamshidian:1993:CGA

[JJ93] Mortaza Jamshidian and Robert I. Jennrich. Conjugate gradient acceleration of the EM algorithm. Journal of the American Statistical Association, 88(421):221–228, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290716.

Johnson:1993:EHN

[JK93] Richard Johnson and Jerome Klotz. Estimating hot numbers and testing uniformity for the lottery. Journal of the American Statistical Association, 88(422):662–668, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290349.

Johnson:1996:CIS

[JK96] Darren J. Johnson and K. Krishnamoorthy. Combining independent studies in a calibration problem. Journal of the American Statistical Association, 91(436):1707–1715, December 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291600.

Jones:1996:BSB

[JMS96] M. C. Jones, J. S. Marron, and S. J. Sheather. A brief survey of bandwidth selection for density estimation. Journal of the American Statistical Association, 91(433):401–407, March 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291420.

Jacroux:1997:DCO

[JMS97] Mike Jacroux, Dibyen Majumdar, and Kirti R. Shah. On the determination and construction of optimal block designs in the presence of linear trends. Journal of the American Statistical Association, 92(437):375–382, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/march/
jacroux.htm;   http://www.jstor.org/stable/2291483.

Jewell:1994:NEF

[JMV94] Nicholas P. Jewell, Hina M. Malani, and Eric Vittinghoff. Nonparametric estimation for a form of doubly censored data, with application to two problems in AIDS. Journal of the American Statistical Association, 89(425):7–18, March 1994. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2291196.

Joe:1995:AMN

[Joe95] Harry Joe. Approximations to multivariate normal rectangle probabilities based on conditional expectations. Journal of the American Statistical Association, 90(431):957–964, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291331.

Johnson:1990:MHE

[Joh90] Robert Alan Johnson. Measurement of Hispanic ethnicity in the U.S. census: An evaluation based on latent-class analysis. Journal of the American Statistical Association, 85(409):58–65, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289525.

Johnson:1992:TEM

[Joh92] Valen E. Johnson. A technique for estimating marginal posterior densities in hierarchical models using mixtures of conditional densities. Journal of the American Statistical Association, 87(419): 852–860, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290224.

Johnson:1994:MSA

[Joh94] Valen E. Johnson. A model for segmentation and analysis of noisy images. Journal of the American Statistical Association, 89 (425):230–241, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291219.

Johnson:1995:BRB

[Joh95] Richard Johnson. Book review: [untitled]. Journal of the American Statistical Association, 90(432):1490, December 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291546.

Johnson:1996:BAM

[Joh96a] Valen E. Johnson. On Bayesian analysis of multirater ordinal data: An application to automated essay grading. Journal of the American Statistical Association, 91(433):42–51, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291381.

Johnson:1996:SCM

[Joh96b] Valen E. Johnson. Studying convergence of Markov chain Monte Carlo algorithms using coupled sample paths. Journal of the American Statistical Association, 91(433):154–166, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291391.

Johnson:1998:CEP

[Joh98a] Valen E. Johnson. Comment: Edge preserving smoothers for image processing. Journal of the American Statistical Association, 93(442):541–??, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnson:1998:CRS

[Joh98b] Valen E. Johnson. A coupling-regeneration scheme for diagnosing convergence in Markov chain Monte Carlo algorithms. Journal of the American Statistical Association, 93(441):238–248, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_98/JOHNSON.HTM;   http://www.jstor.org/
stable/2669620.

Johnson:1998:EPS

[Joh98c] Valen E. Johnson. Edge-preserving smoothers for image processing: Comment. Journal of the American Statistical Association, 93(442):541–544, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2670101.

Jones:1990:BRB

[Jon90] Albyn C. Jones. Book review: [untitled]. Journal of the American Statistical Association, 85(412):1176, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289631.

Jones:1996:ISM

[Jon96] Michael P. Jones. Indicator and stratification methods for missing explanatory variables in multiple linear regression. Journal of the American Statistical Association, 91(433):222–230, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291399.

Jorgensen:1996:BRB

[Jor96] Bent Jorgensen. Book review: [untitled]. Journal of the American Statistical Association, 91(433):432–433, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291426.

Jan:1996:ITS

[JR96] Show-Li Jan and Ronald H. Randles. Interdirection tests for simple repeated-measures designs. Journal of the American Statistical Association, 91(436):1611–1618, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291587.

Jensen:1994:AJD

[JS94] D. R. Jensen and Herbert Solomon. Approximations to joint distributions of definite quadratic forms. Journal of the American Statistical Association, 89(426):480–486, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290849.

Jones:1997:CHO

[JS97a] M. C. Jones and D. F. Signorini. A comparison of higher-order bias kernel density estimators. Journal of the American Statistical Association, 92(439):1063–1073, September 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/sept/
jones_mc.htm;   http://www.jstor.org/stable/2965571.

Juster:1997:IQE

[JS97b] F. Thomas Juster and James P. Smith. Improving the quality of economic data: Lessons from the HRS and AHEAD. Journal of the American Statistical Association, 92(440):1268– 1278, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/JUSTER.HTM;
http://www.jstor.org/stable/2965397.

Johnson:1998:MMQ

[JSKK98] Richard A. Johnson, Songyong Sim, Barbara E. Klein, and Ronald Klein. A multivariate MultiSample quantile test for ordered alternatives. Journal of the American Statistical Association, 93(442):807–818, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/june/JOHNSON.HTM;
http://www.jstor.org/stable/2670130.

Jennison:1997:GSA

[JT97] Christopher Jennison and Bruce W. Turnbull. Group-sequential analysis incorporating covariate information. Journal of the American Statistical Association, 92(440):1330–1341, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_97/december/JENNISON.HTM;   http://www.
jstor.org/stable/2965403.

Jiang:1999:SRM

[JTC99] Wenxin Jiang, Bruce W. Turnbull, and Larry C. Clark. Semiparametric regression models for repeated events with random effects and measurement error. Journal of the American Statistical Association, 94(445):111–124, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669687.

Judkins:1996:CVE

[Jud96a] David R. Judkins. Comment: On variance estimation with imputed survey data. Journal of the American Statistical Association, 91(434):499–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Judkins:1996:VEI

[Jud96b] David R. Judkins. On variance estimation with imputed survey data: Comment. Journal of the American Statistical Association, 91(434):507–510, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291638.

Judkins:1998:ACS

[Jud98a] David R. Judkins. Applications and case studies — comment: Not asked and not answered: Multiple imputation for multiple surveys. Journal of the American Statistical Association, 93(443): 861–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Judkins:1998:AAM

[Jud98b] David R. Judkins. Not asked and not answered: Multiple imputation for multiple surveys: Comment. Journal of the American Statistical Association, 93(443):861–864, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669822.

Jung:1996:QLM

[Jun96] Sin-Ho Jung. Quasi-likelihood for median regression models. Journal of the American Statistical Association, 91(433):251– 257, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291402.

Jones:1993:FCA

[JV93] Richard H. Jones and Aldo V. Vecchia. Fitting continuous ARMA models to unequally spaced spatial data. Journal of the American Statistical Association, 88(423):947–954, September 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290786.

Wu:1997:RBS

[jW97] Tiee jian Wu. Root n bandwidth selectors for kernel estimation of density derivatives. Journal of the American Statistical Association, 92(438):536–547, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2965702.

Kohn:1991:SEA

[KA91] Robert Kohn and Craig F. Ansley. A signal extraction approach to the estimation of treatment and control curves. Journal of the American Statistical Association, 86(416):1034–1041, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290522.

Kim:1995:IEI

[KA95] Donguk Kim and Alan Agresti. Improved exact inference about conditional association in three-way contingency tables. Journal of the American Statistical Association, 90(430):632–639, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291076.

Kabanov:1998:BRB

[Kab98] Yuri Kabanov. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1246–1247, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669887.

Kadane:1990:BRB

[Kad90a] Joseph B. Kadane. Book review: [untitled]. Journal of the American Statistical Association, 85(412):1167–1168, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289620.

Kadane:1990:SAA

[Kad90b] Joseph B. Kadane. A statistical analysis of adverse impact of employer decisions. Journal of the American Statistical Association, 85(412):925–933, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289589.

Kadane:1991:BRB

[Kad91] Joseph B. Kadane. Book review: [untitled]. Journal of the American Statistical Association, 86(414):546–547, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290610.

Kadane:1998:ACS

[Kad98a] Joseph B. Kadane. Applications and case studies — comment: Not asked and not answered: Multiple imputation for multiple surveys. Journal of the American Statistical Association, 93(443): 867–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kadane:1998:AAM

[Kad98b] Joseph B. Kadane. Not asked and not answered: Multiple imputation for multiple surveys: Comment. Journal of the American Statistical Association, 93(443):867–868, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669824.

Kadane:1999:BDE

[Kad99] Joseph B. Kadane. Bridging different eras in sports: Comment. Journal of the American Statistical Association, 94(447): 680, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2669975.

Kafadar:1994:BRB

[Kaf94] Karen Kafadar. Book review: [untitled]. Journal of the American Statistical Association, 89(425):355–356, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291235.

Kafadar:1995:BRB

[Kaf95] Karen Kafadar. Book review: [untitled]. Journal of the American Statistical Association, 90(429):395, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291180.

Kafadar:1997:BRBa

[Kaf97a] Karen Kafadar. Book review: [untitled]. Journal of the American Statistical Association, 92(437):388–389, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291493.

Kafadar:1997:BRBb

[Kaf97b] Karen Kafadar. Book review: [untitled]. Journal of the American Statistical Association, 92(438):798, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965742.

Kafadar:1997:BRBc

[Kaf97c] Karen Kafadar. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1651, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965446.

Kafadar:1997:BRBd

[Kaf97d] Karen Kafadar. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1653, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965451.

Kafadar:1997:BRBe

[Kaf97e] Karen Kafadar. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1653–1654, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965452.

Kafadar:1997:BRBf

[Kaf97f] Karen Kafadar. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1654, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965453.

Kafadar:1997:BRBg

[Kaf97g] Karen Kafadar. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1654–1655, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965456.

Kafadar:1997:BRBh

[Kaf97h] Karen Kafadar. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1655, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965457.

Kafadar:1999:BRBa

[Kaf99a] Karen Kafadar. Book review: [untitled]. Journal of the American Statistical Association, 94(445):341, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669712.

Kafadar:1999:BRBb

[Kaf99b] Karen Kafadar. Book review: [untitled]. Journal of the American Statistical Association, 94(447):982–983, September 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670015.

Kaigh:1994:DFT

[Kai94] W. D. Kaigh. Distribution-free two-sample tests based on rank spacings. Journal of the American Statistical Association, 89 (425):159–167, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291212.

Kanazawa:1990:BRB

[Kan90] Yuichiro Kanazawa. Book review: [untitled]. Journal of the American Statistical Association, 85(411):907–908, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290046.

Karr:1990:BRB

[Kar90] Alan F. Karr. Book review: [untitled]. Journal of the American Statistical Association, 85(409):262, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289568.

Kass:1990:BRB

[Kas90] Robert E. Kass. Book review: [untitled]. Journal of the American Statistical Association, 85(410):594–595, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289810.

Kaspi:1997:BRB

[Kas97a] Haya Kaspi. Book review: [untitled]. Journal of the American Statistical Association, 92(438):792–793, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965732.

Kass:1997:BRB

[Kas97b] Robert E. Kass. Book review: Markov Chain Monte Carlo in Practice, by W. R. Gilks; S. Richardson; D. J. Spiegelhalter. Journal of the American Statistical Association, 92(440):1645–1646, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2965438.

Kohn:1991:PCV

[KAT91] Robert Kohn, Craig F. Ansley, and David Tharm. The performance of cross-validation and maximum likelihood estimators of spline smoothing parameters. Journal of the American Statistical Association, 86(416):1042–1050, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290523.

Katzoff:1997:BRBa

[Kat97a] Myron Katzoff. Book review: [untitled]. Journal of the American Statistical Association, 92(438):794–795, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965735.

Katzoff:1997:BRBb

[Kat97b] Myron Katzoff. Book review: [untitled]. Journal of the American Statistical Association, 92(438):795, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965736.

Kawczak:1999:BRB

[Kaw99] Janusz Kawczak. Book review: [untitled]. Journal of the American Statistical Association, 94(448):1385–1386, December 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669955.

KB:1997:BRBa

[KB97a] KB. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1227, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965614.

KB:1997:BRBb

[KB97b] KB. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1227, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965615.

Kano:1993:SIB

[KBB93] Yutaka Kano, Maia Berkane, and Peter M. Bentler. Statistical inference based on pseudo-maximum likelihood estimators in elliptical populations. Journal of the American Statistical Association, 88(421):135–143, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290706.

Kooperberg:1997:PR

[KBS97] Charles Kooperberg, Smarajit Bose, and Charles J. Stone. Polychotomous regression. Journal of the American Statistical Association, 92(437):117–127, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/march/
kooperbe.htm;   http://www.jstor.org/stable/2291455.

Kim:1993:RLR

[KC93] Hyune-Ju Kim and Lujing Cai. Robustness of the likelihood ratio test for a change in simple linear regression. Journal of the American Statistical Association, 88(423):864–871, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290775.

Kosorok:1996:ALO

[KC96] Michael R. Kosorok and Wei-Hsiung Chao. The analysis of longitudinal ordinal response data in continuous time. Journal of the American Statistical Association, 91(434):807–817, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291675.

Kim:1993:SER

[KCHH93] P. T. Kim, E. M. Carter, J. J. Hubert, and K. J. Hand. Shrinkage estimators of relative potency. Journal of the American Statistical Association, 88(422):615–621, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290343.

Keppenne:1994:ASS

[KDG94] Christian L. Keppenne, Michael D. Dettinger, and Michael Ghil. An approach to statistical spatial-temporal modeling of meteorological fields: Comment. Journal of the American Statistical Association, 89(426):384–387, June 1994. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290835.

Kent:1991:SIR

[Ken91] John T. Kent. Sliced inverse regression for dimension reduction: Comment. Journal of the American Statistical Association, 86 (414):336–337, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290567.

Kepler:1997:BRB

[Kep97] Tom Kepler. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1647–1648, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965441.

Kettenring:1997:PAS

[Ket97a] Jon R. Kettenring. Presidential address: Shaping statistics for success in the 21st Century. Journal of the American Statistical Association, 92(440):1229–??, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.amstat.org/publications/jasa/kettenring.html.

Kettenring:1997:SSS

[Ket97b] Jon R. Kettenring. Shaping statistics for success in the 21st century. Journal of the American Statistical Association, 92(440): 1229–1234, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2965390.

Kfadar:1997:BRB

[Kfa97] Karen Kfadar. Book review: [untitled]. Journal of the American Statistical Association, 92(437):387–388, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291490.

Khuri:1994:BRB

[Khu94] Andre I. Khuri. Book review: [untitled]. Journal of the American Statistical Association, 89(425):357–358, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291239.

Kianifard:1990:BRB

[Kia90] Farid Kianifard. Book review: [untitled]. Journal of the American Statistical Association, 85(410):595–596, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289811.

King:1999:FEI

[Kin99] Gary King. The future of ecological inference research: a comment on Freedman et al. Journal of the American Statistical Association, 94(445):352–355, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669734.

Kirk:1998:BRB

[Kir98] Roger E. Kirk. Book review: [untitled]. Journal of the American Statistical Association, 93(442):836–837, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670135.

Kitagawa:1998:SOS

[Kit98] Genshiro Kitagawa. A self-organizing state space model. Journal of the American Statistical Association, 93(443):1203– 1215, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/sept/kitagawa.html;
 http://www.jstor.org/stable/2669862.

KK:1994:BRBa

[KK94a] KK. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1149–1150, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290964.

KK:1994:BRBb

[KK94b] KK. Book review: [untitled]. Journal of the American Statistical Association, 89(428):1569, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291035.

KK:1994:BRBc

[KK94c] KK. Book review: [untitled]. Journal of the American Statistical Association, 89(428):1569, December 1994. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291036.

KK:1995:BRBa

[KK95a] KK. Book review: [untitled]. Journal of the American Statistical Association, 90(430):810, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291114.

KK:1995:BRBb

[KK95b] KK. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1136, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291369.

KK:1995:BRBc

[KK95c] KK. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1136, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291372.

KK:1995:BRBd

[KK95d] KK. Book review: [untitled]. Journal of the American Statistical Association, 90(432):1492–1493, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291552.

KK:1995:BRBe

[KK95e] KK. Book review: [untitled]. Journal of the American Statistical Association, 90(432):1493, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291556.

KK:1996:BRBa

[KK96a] KK. Book review: [untitled]. Journal of the American Statistical Association, 91(434):916, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291706.

KK:1996:BRBb

[KK96b] KK. Book review: [untitled]. Journal of the American Statistical Association, 91(435):1380, September 1996. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291763.

KK:1996:BRBc

[KK96c] KK. Book review: [untitled]. Journal of the American Statistical Association, 91(436):1755, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291621.

KK:1997:BRBa

[KK97a] KK. Book review: [untitled]. Journal of the American Statistical Association, 92(437):389–390, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291499.

KK:1997:BRBb

[KK97b] KK. Book review: [untitled]. Journal of the American Statistical Association, 92(438):799, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965747.

KK:1997:BRBc

[KK97c] KK. Book review: [untitled]. Journal of the American Statistical Association, 92(438):800, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965752.

KK:1997:BRBd

[KK97d] KK. Book review: [untitled]. Journal of the American Statistical Association, 92(438):800, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965755.

KK:1997:BRBe

[KK97e] KK. Book review: [untitled]. Journal of the American Statistical Association, 92(438):800–801, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965757.

KK:1997:BRBf

[KK97f] KK. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1656, December 1997. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965459.

KK:1998:BRBa

[KK98a] KK. Book review: [untitled]. Journal of the American Statistical Association, 93(441):411, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669661.

KK:1998:BRBb

[KK98b] KK. Book review: [untitled]. Journal of the American Statistical Association, 93(441):411–412, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669663.

KK:1998:BRBc

[KK98c] KK. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1250, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669895.

KK:1998:BRBd

[KK98d] KK. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1250, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669896.

KK:1998:BRBe

[KK98e] KK. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1251, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669899.

KK:1998:BRBf

[KK98f] KK. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1251–1252, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669905.

KK:1998:BRBg

[KK98g] KK. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1252, September 1998. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669906.

KK:1998:BRBh

[KK98h] KK. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1252, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669907.

Knight:1998:BCM

[KKT98a] Keith Knight, Rafal Kustra, and Robert Tibshirani. Bayesian CART model search: Comment. Journal of the American Statistical Association, 93(443):950–954, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669834.

Knight:1998:CBC

[KKT98b] Keith Knight, Rafal Kustra, and Robert Tibshirani. Comment: Bayesian CART model search. Journal of the American Statistical Association, 93(443):950–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Keramidas:1990:FTS

[KL90] Elaine M. Keramidas and Jack C. Lee. Forecasting technological substitutions with concurrent short time series. Journal of the American Statistical Association, 85(411):625–632, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289994.

Keyes:1996:GPF

[KL96] Tim K. Keyes and Martin S. Levy. Goodness of prediction fit for multivariate linear models. Journal of the American Statistical Association, 91(433):191–197, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291395.

Kallenberg:1997:DDS

[KL97]  Wilbert C. M. Kallenberg and Teresa Ledwina.  Data-

driven smooth tests when the hypothesis is composite. Journal of the American Statistical Association, 92(439):1094– 1104, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/sept/kallenberg.htm;
http://www.jstor.org/stable/2965574.

Kallenberg:1999:DDR

[KL99a] Wilbert C. M. Kallenberg and Teresa Ledwina. Data-driven rank tests for independence. Journal of the American Statistical Association, 94(445):285–301, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669703.

Kosorok:1999:VFI

[KL99b] Michael R. Kosorok and Chin-Yu Lin. The versatility of functionindexed weighted log-rank statistics. Journal of the American Statistical Association, 94(445):320–332, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669706.

Kluppelberg:1994:BRB

[Klu94] Claudia Kluppelberg. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1148, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290959.

Kong:1994:SIB

[KLW94] Augustine Kong, Jun S. Liu, and Wing Hung Wong. Sequential imputations and Bayesian missing data problems. Journal of the American Statistical Association, 89(425):278–288, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291224.

Kempthorne:1990:UMO

[KM90] Peter J. Kempthorne and Max B. Mendel. Unmasking multivariate outliers and leverage points: Comment. Journal of the American Statistical Association, 85(411):647–648, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289998.

Kao:1991:OSA

[KM91a] Tzu-Cheg Kao and George P. McCabe. Optimal sample allocation for normal discrimination and logistic regression under stratified sampling. Journal of the American Statistical Association, 86 (414):432–436, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290589.

Karlin:1991:SSP

[KM91b] Samuel Karlin and Catherine Macken. Some statistical problems in the assessment of inhomogeneities of DNA sequence data. Journal of the American Statistical Association, 86(413):27–35, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289712.

Kitagawa:1996:DCC

[KM96a] Genshiro Kitagawa and Norio Matsumoto. Detection of coseismic changes of underground water level. Journal of the American Statistical Association, 91(434):521–528, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291644.

Koshevoy:1996:LZM

[KM96b] Gleb Koshevoy and Karl Mosler. The Lorenz zonoid of a multivariate distribution. Journal of the American Statistical Association, 91(434):873–882, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291682.

Kobayashi:1999:TLL

[KM99a] Masahito Kobayashi and Michael McAleer. Tests of linear and logarithmic transformations for integrated processes. Journal of the American Statistical Association, 94(447):860–868, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670001.

Koenker:1999:GFR

[KM99b] Roger Koenker and Jose A. F. Machado. Goodness of fit and related inference processes for quantile regression. Journal of the American Statistical Association, 94(448):1296–1310, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669943.

Kadane:1999:YAP

[KMT99] Joseph B. Kadane, Michael M. Meyer, and John W. Tukey. Yule’s association paradox and ignored stratum heterogeneity in capture-recapture studies. Journal of the American Statistical Association, 94(447):855–859, September 1999. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670000.

KNB:1998:BRB

[KNB98] KNB. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1530, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670086.

Kneip:1998:CSS

[Kne98a] Alois Kneip. Comment: Smoothing spline models for the analysis of nested and crossed samples of curves. Journal of the American Statistical Association, 93(443):984–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kneip:1998:SSM

[Kne98b] Alois Kneip. Smoothing spline models for the analysis of nested and crossed samples of curves: Comment. Journal of the American Statistical Association, 93(443):984–986, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669840.

Kadane:1995:UFA

[KO95] Joseph B. Kadane and Anthony O’Hagan. Using finitely additive probability: Uniform distributions on the natural numbers. Journal of the American Statistical Association, 90(430):626–631, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291075.

Kooperberg:1996:POP

[KO96] Charles Kooperberg and Finbarr O’Sullivan. Predictive oscillation patterns: a synthesis of methods for spatial-temporal decomposition of random fields. Journal of the American Statistical Association, 91(436):1485–1496, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291574.

Koehler:1997:BRB

[Koe97] Kenneth Koehler. Book review: [untitled]. Journal of the American Statistical Association, 92(438):793–794, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965733.

Koenker:1999:DRD

[Koe99a] Roger Koenker. Discussion: Regression depth. Journal of the American Statistical Association, 94(446):405–??, June 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Koenker:1999:RDC

[Koe99b] Roger Koenker. Regression depth: Comment. Journal of the American Statistical Association, 94(446):405–406, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670157.

Kolaczyk:1996:WSA

[Kol96] Eric D. Kolaczyk. A wavelet shrinkage approach to tomographic image reconstruction. Journal of the American Statistical Association, 91(435):1079–1090, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291727.

Kolaczyk:1999:BMM

[Kol99] Eric D. Kolaczyk. Bayesian multiscale models for Poisson processes. Journal of the American Statistical Association, 94(447): 920–933, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2670007.

Koopmans:1991:BRB

[Koo91] Lambert H. Koopmans. Book review: [untitled]. Journal of the American Statistical Association, 86(415):823, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290422.

Koopman:1997:EIK

[Koo97] Siem Jan Koopman. Exact initial Kalman filtering and smoothing for nonstationary time series models. Journal of the American Statistical Association, 92(440):1630–1638, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/december/KOOPMAN.HTM;   http://www.jstor.
org/stable/2965434.

Kott:1993:BRB

[Kot93] Philip S. Kott. Book review: [untitled]. Journal of the American Statistical Association, 88(424):1470–1471, December 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291300.

Kott:1994:NHN

[Kot94] Phillip S. Kott. A note on handling nonresponse in sample surveys. Journal of the American Statistical Association, 89(426): 693–696, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290873.

Koyak:1991:BRB

[Koy91] Robert A. Koyak. Book review: [untitled]. Journal of the American Statistical Association, 86(416):1140–1141, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290538.

Koenker:1990:EMR

[KP90] Roger Koenker and Stephen Portnoy. M estimation of multivariate regressions. Journal of the American Statistical Association, 85(412):1060–1068, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289602.

Kedem:1991:TMR

[KP91] Benjamin Kedem and Harry Pavlopoulos. On the threshold method for rainfall estimation: Choosing the optimal threshold level. Journal of the American Statistical Association, 86(415): 626–633, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290391.

Klein:1992:ABC

[KP92] Ruben Klein and S. James Press. Adaptive Bayesian classification of spatial data. Journal of the American Statistical Association, 87(419):844–851, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290223.

Kateri:1997:AMC

[KP97] Maria Kateri and Takis Papaioannou. Asymmetry models for contingency tables. Journal of the American Statistical Association, 92(439):1124–1131, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

amstat.org/publications/jasa/abstracts_97/sept/kateri.
htm;   http://www.jstor.org/stable/2965577.

Kleffe:1991:OEC

[KPR91] J. Kleffe, N. G. N. Prasad, and J. N. K. Rao. “optimal” estimation of correlated response variance under additive models. Journal of the American Statistical Association, 86(413):144–150, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289723.

Kinderman:1976:CGN

[KR76] A. J. Kinderman and J. G. Ramage. Computer generation of normal random variables. Journal of the American Statistical Association, 71(356):893–896, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2286857. See correction [KR90a].

Kinderman:1990:CCG

[KR90a] A. J. Kinderman and J. G. Ramage. Correction: Computer generation of normal random variables. Journal of the American Statistical Association, 85(409):272, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289582. See [KR76].

Kovar:1990:ITF

[KR90b] Mary Grace Kovar and Patricia Royston. Interactional troubles in face-to-face survey interviews: Comment. Journal of the American Statistical Association, 85(409):246–247, March 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289553.

Krzysztofowicz:1993:SBP

[KR93] Roman Krzysztofowicz and Sharon Reese. Stochastic bifurcation processes and distributions of fractions. Journal of the American Statistical Association, 88(421):345–354, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290730.

Kass:1995:BF

[KR95] Robert E. Kass and Adrian E. Raftery. Bayes factors. Journal of the American Statistical Association, 90(430):773–795, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291091.

Kahn:1996:DRM

[KR96] Michael J. Kahn and Adrian E. Raftery. Discharge rates of Medicare stroke patients to skilled nursing facilities: Bayesian logistic regression with unobserved heterogeneity. Journal of the American Statistical Association, 91(433):29–41, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291380.

Kramer:1990:SPW

[Kra90] Fredrica D. Kramer. Statistics and policy in welfare quality control: a basis for understanding and assessing competing views. Journal of the American Statistical Association, 85(411):850–855, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290025.

Krantz:1999:NHT

[Kra99] David H. Krantz. The null hypothesis testing controversy in psychology. Journal of the American Statistical Association, 94(448): 1372–1381, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2669949.

Krijnen:1994:BRB

[Kri94] Wim P. Krijnen. Book review: [untitled]. Journal of the American Statistical Association, 89(425):361, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291245.

Kroonenberg:1990:BRB

[Kro90] Pieter M. Kroonenberg. Book review: [untitled]. Journal of the American Statistical Association, 85(411):909–910, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290049.

Kvam:1993:EDF

[KS93] Paul H. Kvam and Francisco J. Samaniego. On estimating distribution functions using nomination samples. Journal of the American Statistical Association, 88(424):1317–1322, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291272.

Kvam:1994:NML

[KS94] Paul H. Kvam and Francisco J. Samaniego. Nonparametric maximum likelihood estimation based on ranked set samples. Journal of the American Statistical Association, 89(426):526–537, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290855.

Kianifard:1996:RDA

[KS96] Farid Kianifard and William H. Swallow. A review of the development and application of recursive residuals in linear models. Journal of the American Statistical Association, 91(433):391– 400, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291419.

Kouassi:1997:SAH

[KS97] Djokouri A. Kouassi and Jagbir Singh. A semiparametric approach to hazard estimation with randomly censored observations. Journal of the American Statistical Association, 92(440): 1351–1355, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/KOUASSI.HTM;
http://www.jstor.org/stable/2965405.

Kaiser:1994:SML

[KSJ94] Mark S. Kaiser, Paul L. Speckman, and John R. Jones. Statistical models for limiting nutrient relations in inland waters. Journal of the American Statistical Association, 89(426):410–423, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290841.

Kadane:1996:RFC

[KSS96] Joseph B. Kadane, Mark J. Schervish, and Teddy Seidenfeld. Reasoning to a foregone conclusion. Journal of the American Statistical Association, 91(435):1228–1235, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291741.

Kooperberg:1995:HR

[KST95] Charles Kooperberg, Charles J. Stone, and Young K. Truong. Hazard regression. Journal of the American Statistical Association, 90(429):78–94, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291132.

Kakizawa:1998:DCM

[KST98] Yoshihide Kakizawa, Robert H. Shumway, and Masanobu Taniguchi. Discrimination and clustering for multivariate time series. Journal of the American Statistical Association, 93(441): 328–340, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/KAKIZAWA.HTM;   http://
www.jstor.org/stable/2669629.

Kimeldorf:1992:MMS

[KSW92] George Kimeldorf, Allan R. Sampson, and Lyn R. Whitaker. Min and max scorings for two-sample ordinal data. Journal of the American Statistical Association, 87(417):241–247, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290475.

Kim:1994:TBS

[KT94a] Dong K. Kim and Jeremy M. G. Taylor. Transform-both-sides approach for overdispersed binomial data when N is unobserved. Journal of the American Statistical Association, 89(427):833–845, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290909.

Kolassa:1994:ACI

[KT94b] John E. Kolassa and Martin A. Tanner. Approximate conditional inference in exponential families via the Gibbs sampler. Journal of the American Statistical Association, 89(426):697–702, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290874. See

corrections [KT94d, KT94c].

Kolassa:1994:CAC

[KT94c] John E. Kolassa and Martin A. Tanner. Correction to: “Approximate conditional inference in exponential families via the Gibbs sampler”. Journal of the American Statistical Association, 89(428):1572, ???? 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [KT94b].

Kolassa:1994:CKT

[KT94d] John E. Kolassa and Martin A. Tanner. Correction to Kolassa and Tanner (1994): Approximate Conditional Inference in Exponential Families Via the Gibbs Sampler. Journal of the American Statistical Association, 89(428):1572, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291044. See [KT94b].

Kim:1995:REA

[KT95] Dong K. Kim and Jeremy M. G. Taylor. The restricted EM algorithm for maximum likelihood estimation under linear restrictions on the parameters. Journal of the American Statistical Association, 90(430):708–716, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291083.

Kulasekera:1995:BRB

[Kul95a] K. B. Kulasekera. Book review: [untitled]. Journal of the American Statistical Association, 90(429):397, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291184.

Kulasekera:1995:CRC

[Kul95b] K. B. Kulasekera. Comparison of regression curves using quasiresiduals. Journal of the American Statistical Association, 90 (431):1085–1093, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291346.

Kunert:1990:PTM

[Kun90] Joachim Kunert. On the power of tests for multiple comparison of three normal means. Journal of the American Statistical Association, 85(411):808–812, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290019.

Kupper:1990:BRB

[Kup90] Lawrence L. Kupper. Book review: [untitled]. Journal of the American Statistical Association, 85(411):901, September 1990.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290035.

Kushner:1998:OER

[Kus98] H. B. Kushner. Optimal and efficient repeated measurements designs for uncorrelated observations. Journal of the American Statistical Association, 93(443):1176–1187, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
sept/kushner.html;   http://www.jstor.org/stable/2669860.

Kass:1995:RBT

[KW95] Robert E. Kass and Larry Wasserman. A reference Bayesian test for nested hypotheses and its relationship to the Schwarz criterion. Journal of the American Statistical Association, 90(431):928–934, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291327.

Kass:1996:SPD

[KW96a] Robert E. Kass and Larry Wasserman. The selection of prior distributions by formal rules. Journal of the American Statistical Association, 91(435):1343–1370, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291752.

Kepner:1996:RTT

[KW96b] James L. Kepner and Dennis D. Wackerly. On rank transformation techniques for balanced incomplete repeated-measures designs. Journal of the American Statistical Association, 91(436): 1619–1625, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291588.

Kulasekera:1997:SPS

[KW97] K. B. Kulasekera and J. Wang. Smoothing parameter selection for power optimality in testing of regression curves. Journal of the American Statistical Association, 92(438):500–511, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2965699.

Kuo:1996:BCN

[KY96] Lynn Kuo and Tae Young Yang. Bayesian computation for nonhomogeneous Poisson processes in software reliability. Journal of the American Statistical Association, 91(434):763–773, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291671.

Leger:1993:AIV

[LA93] Christian Léger and Naomi Altman. Assessing influence in variable selection problems. Journal of the American Statistical Association, 88(422):547–556, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2290335.

Lang:1994:SMJ

[LA94] Joseph B. Lang and Alan Agresti. Simultaneously modeling joint and marginal distributions of multivariate categorical responses. Journal of the American Statistical Association, 89(426):625–632, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290865.

Lachenbruch:1990:BRBa

[Lac90a] P. A. Lachenbruch. Book review: [untitled]. Journal of the American Statistical Association, 85(409):267, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289576.

Lachenbruch:1990:BRBb

[Lac90b] Peter A. Lachenbruch. Book review: [untitled]. Journal of the American Statistical Association, 85(410):593–594, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289809.

Lachenbruch:1990:BRBc

[Lac90c] Peter A. Lachenbruch. Book review: [untitled]. Journal of the American Statistical Association, 85(410):597–598, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289815.

Lachenbruch:1990:BRBd

[Lac90d] Peter A. Lachenbruch. Book review: [untitled]. Journal of the American Statistical Association, 85(412):1179, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289636.

Lachenbruch:1991:BRB

[Lac91] P. A. Lachenbruch. Book review: [untitled]. Journal of the American Statistical Association, 86(416):1149, December 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290547.

Lachenbruch:1992:BRB

[Lac92] Peter A. Lachenbruch. Book review: [untitled]. Journal of the American Statistical Association, 87(418):591, June 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290309.

Lachenbruch:1993:BRBa

[Lac93a] Peter A. Lachenbruch. Book review: [untitled]. Journal of the American Statistical Association, 88(421):380, March 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290742.

Lachenbruch:1993:BRBb

[Lac93b] Peter A. Lachenbruch. Book review: [untitled]. Journal of the American Statistical Association, 88(421):380–381, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290743.

Lachenbruch:1993:BRBc

[Lac93c] Peter A. Lachenbruch. Book review: [untitled]. Journal of the American Statistical Association, 88(422):707–708, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290372.

Lamotte:1997:BRB

[Lam97] Lynn Roy Lamotte. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1219, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965595.

Lange:1993:BRB

[Lan93] Nicholas Lange. Book review: [untitled]. Journal of the American Statistical Association, 88(424):1473, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291304.

Lange:1995:BRBa

[Lan95a] Nicholas Lange. Book review: [untitled]. Journal of the American Statistical Association, 90(430):801, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291098.

Lange:1995:BRBb

[Lan95b] Nicholas Lange. Book review: [untitled]. Journal of the American Statistical Association, 90(432):1489, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291544.

Lang:1996:PGF

[Lan96] Joseph B. Lang. On the partitioning of goodness-of-fit statistics for multivariate categorical response models. Journal of the American Statistical Association, 91(435):1017–1023, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291721.

Lange:1997:BRB

[Lan97] Nicholas Lange. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1642–1645, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965437.

Lange:1998:CSS

[Lan98a] Nicholas Lange. Comment: Smoothing spline models for the analysis of nested and crossed samples of curves. Journal of the American Statistical Association, 93(443):986–??, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lange:1998:SSM

[Lan98b] Nicholas Lange. Smoothing spline models for the analysis of nested and crossed samples of curves: Comment. Journal of the American Statistical Association, 93(443):986–988, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669841.

LaRiccia:1991:SGF

[LaR91a] V. N. LaRiccia. Smooth goodness-of-fit tests: a quantile function approach. Journal of the American Statistical Association, 86 (414):427–431, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290588.

Larntz:1991:BRB

[Lar91b] Kinley Larntz. Book review: [untitled]. Journal of the American Statistical Association, 86(415):835, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290436.

Laslett:1994:KSEb

[Las94a] Geoffrey M. Laslett. Kriging and splines: An empirical comparison of their predictive performance in some applications: Rejoinder. Journal of the American Statistical Association, 89(426): 406–409, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290840.

Laslett:1994:KSEa

[Las94b] Geoffrey M. Laslett. Kriging and splines: an empirical comparison of their predictive performance in some applications. With comments and a rejoinder by the author. Journal of the American Statistical Association, 89(426):391–400, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290837.

Lauritzen:1992:BRB

[Lau92a] Steffen L. Lauritzen. Book review: [untitled]. Journal of the American Statistical Association, 87(417):251–252, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290481.

Lauritzen:1992:PPM

[Lau92b] Steffen L. Lauritzen. Propagation of probabilities, means, and variances in mixed graphical association models. Journal of the American Statistical Association, 87(420):1098–1108, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290647.

LaVange:1990:BRB

[LaV90] Lisa M. LaVange. Book review: [untitled]. Journal of the American Statistical Association, 85(409):268, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289578.

Lavine:1991:ARB

[Lav91a] Michael Lavine. An approach to robust Bayesian analysis for multidimensional parameter spaces. Journal of the American Statistical Association, 86(414):400–403, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290584.

Lavine:1991:PEI

[Lav91b] Michael Lavine. Problems in extrapolation illustrated with Space Shuttle O-ring data. Journal of the American Statistical Association, 86(416):919–921, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290505.

Lavine:1991:SBS

[Lav91c] Michael Lavine. Sensitivity in Bayesian statistics: The prior and the likelihood. Journal of the American Statistical Association, 86(414):396–399, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290583.

Lindsay:1993:MNM

[LB93] Bruce G. Lindsay and Prasanta Basak. Multivariate normal mixtures: a fast consistent method of moments. Journal of the American Statistical Association, 88(422):468–476, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290326.

Lewis:1994:GSC

[LB94a] Roger J. Lewis and Donald A. Berry. Group sequential clinical trials: a classical evaluation of Bayesian decision-theoretic designs. Journal of the American Statistical Association, 89(428): 1528–1534, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291016.

Li:1994:REM

[LB94b] Zhaohai Li and Colin B. Begg. Random effects models for combining results from controlled and uncontrolled studies in a metaanalysis. Journal of the American Statistical Association, 89(428): 1523–1527, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291015.

Lindstrom:1994:CLB

[LB94c] Mary J. Lindstrom and Douglas M. Bates. Correction to Lindstrom and Bates (1988): Newton–Raphson and EM algorithms for linear mixed effects models for repeated measures data. Journal of the American Statistical Association, 89(428):1572, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291042.

Lindstrom:1994:CNR

[LB94d] Mary J. Lindstrom and Douglas M. Bates. Correction to: “Newton–Raphson and EM algorithms for linear mixed-effects models for repeated measures-data” [J. Amer. Statist. Assoc. 83 (1988), no. 404, 1014–1022]. Journal of the American Statistical Association, 89(428):1572, ???? 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [LB94e].

Lindstrom:1994:NRE

[LB94e] Mary J. Lindstrom and Douglas M. Bates. Newton–Raphson and EM algorithms for linear mixed-effects models for repeated measures-data. Journal of the American Statistical Association, 89(404):1014–1022, ???? 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [LB94d].

Lin:1996:BCG

[LB96] Xihong Lin and Norman E. Breslow. Bias correction in generalized linear mixed models with multiple components of dispersion. Journal of the American Statistical Association, 91(435): 1007–1016, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291720.

Lee:1992:MFUb

[LC92] Ronald D. Lee and Lawrence R. Carter. Modeling and forecasting U.S. mortality: Rejoinder. Journal of the American Statistical Association, 87(419):674–675, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290204.

LeBlanc:1993:STG

[LC93] Michael LeBlanc and John Crowley. Survival trees by goodness of split. Journal of the American Statistical Association, 88(422): 457–467, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290325.

LeBlanc:1995:SRF

[LC95a] Michael LeBlanc and John Crowley. Semiparametric regression functionals. Journal of the American Statistical Association, 90 (429):95–105, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291133.

Liu:1995:BDS

[LC95b] Jun S. Liu and Rong Chen. Blind deconvolution via sequential imputations. Journal of the American Statistical Association, 90 (430):567–576, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291068.

Liu:1998:SMC

[LC98a] Jun S. Liu and Rong Chen. Sequential Monte Carlo methods for dynamic systems. Journal of the American Statistical Association, 93(443):1032–1044, September 1998. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
sept/liu2.html;   http://www.jstor.org/stable/2669847.

Llu:1998:SMC

[LC98b] Jun S. Llu and Rong Chen. Sequential Monte Carlo methods for dynamic systems. Journal of the American Statistical Association, 93(443):1032–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lindsay:1991:SER

[LCG91] Bruce Lindsay, Clifford C. Clogg, and John Grego. Semiparametric estimation in the Rasch model and related exponential response models, including a simple latent class model for item analysis. Journal of the American Statistical Association, 86(413): 96–107, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289719.

Lange:1992:HBMb

[LCG92a] Nicholas Lange, Bradley P. Carlin, and Alan E. Gelfand. Hierarchical Bayes models for the progression of HIV infection using longitudinal CD4 T-cell numbers: Rejoinder. Journal of the American Statistical Association, 87(419):631–632, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290197.

Lange:1992:HBMa

[LCG^{+}92b] Nicholas Lange, Bradley P. Carlin, Alan E. Gelfand, Thomas A. Louis, Jeremy M. G. Taylor, Mark R. Segal, and Bradley P. Carlin. Hierarchical Bayes models for the progression of HIV infection using longitudinal CD4 T-cell numbers. Journal of the American Statistical Association, 87(419):615–626, September 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290194.

Lambert:1997:ER

[LCK97] Diane Lambert, George Casella, and Karen Kafadar. Editors’ report for 1996. Journal of the American Statistical Association, 92 (438):391, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2965683.

Lee:1992:MFUa

[LCMA92] Ronald D. Lee, Lawrence R. Carter, Robert McNown, and Juha M. Alho. Modeling and forecasting U.S. mortality. Journal of the American Statistical Association, 87(419):659–671, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290201.

Liu:1998:RCS

[LCW98] Jun S. Liu, Rong Chen, and Wing Hung Wong. Rejection control and sequential importance sampling. Journal of the American Statistical Association, 93(443):1022–1031, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_98/sept/liu1.html;   http://www.jstor.
org/stable/2669846.

Linton:1997:ATA

[LCWH97] Oliver B. Linton, Rong Chen, Naiysin Wang, and Wolfgang Hardle. An analysis of transformations for additive nonparametric regression. Journal of the American Statistical Association, 92(440): 1512–1521, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/LINTON.HTM;
http://www.jstor.org/stable/2965422.

Lee:1991:SCC

[LD91] Jae Won Lee and David L. DeMets. Sequential comparison of changes with repeated measurements data. Journal of the American Statistical Association, 86(415):757–762, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290409.

Lee:1992:SRT

[LD92] Jae Won Lee and David L. DeMets. Sequential rank tests with repeated measurements in clinical trials. Journal of the American Statistical Association, 87(417):136–142, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290461.

LD:1996:BRBa

[LD96a] LD. Book review: [untitled]. Journal of the American Statistical Association, 91(434):915, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291699.

LD:1996:BRBb

[LD96b] LD. Book review: [untitled]. Journal of the American Statistical Association, 91(434):916, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291703.

LD:1997:BRBa

[LD97a] LD. Book review: [untitled]. Journal of the American Statistical Association, 92(438):799, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965748.

LD:1997:BRBb

[LD97b] LD. Book review: [untitled]. Journal of the American Statistical Association, 92(438):799–800, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965751.

LD:1998:BRBa

[LD98a] LD. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1250, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669894.

LD:1998:BRBb

[LD98b] LD. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1251, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669902.

LDH:1999:BRB

[LDH99] LDH. Book review: [untitled]. Journal of the American Statistical Association, 94(447):990, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670032.

Ledwina:1994:DDV

[Led94] Teresa Ledwina. Data-driven version of Neyman’s smooth test of fit. Journal of the American Statistical Association, 89(427): 1000–1005, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290926.

Lee:1993:BRB

[Lee93] Martin L. Lee. Book review: [untitled]. Journal of the American Statistical Association, 88(423):1169, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290819.

Lee:1996:EMD

[Lee96a] Chu-In Charles Lee. On estimation for monotone dose-response curves. Journal of the American Statistical Association, 91(435): 1110–1119, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291730.

Lee:1996:BRB

[Lee96b] Mei-Ling Ting Lee. Book review: [untitled]. Journal of the American Statistical Association, 91(435):1377–1378, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291760.

Lehmann:1993:FNP

[Leh93] E. L. Lehmann. The Fisher, Neyman–Pearson theories of testing hypotheses: One theory or two? Journal of the American Statistical Association, 88(424):1242–1249, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291263.

Lehmann:1997:BRB

[Leh97] Erich L. Lehmann. Book review: [untitled]. Journal of the American Statistical Association, 92(438):789, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965729.

Lehmann:1998:BRB

[Leh98] Erich L. Lehmann. Book review: [untitled]. Journal of the American Statistical Association, 93(441):398, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669636.

Lele:1994:EFC

[Lel94] Subhash Lele. Estimating functions in chaotic systems. Journal of the American Statistical Association, 89(426):512–516, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290853.

Lele:1999:BRB

[Lel99] Subhash Lele. Book review: [untitled]. Journal of the American Statistical Association, 94(445):340–341, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669711.

Lenth:1998:BRB

[Len98] Russell V. Lenth. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1240, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669875.

Lenth:1999:BRBa

[Len99a] Russell V. Lenth. Book review: [untitled]. Journal of the American Statistical Association, 94(446):653–654, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670194.

Lenth:1999:BRBb

[Len99b] Russell V. Lenth. Book review: [untitled]. Journal of the American Statistical Association, 94(446):656, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670199.

Lenth:1999:BRBc

[Len99c] Russell V. Lenth. Book review: [untitled]. Journal of the American Statistical Association, 94(447):983–984, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670017.

Lenth:1999:BRBd

[Len99d] Russell V. Lenth. Book review: [untitled]. Journal of the American Statistical Association, 94(448):1384, December 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669953.

Leonard:1992:BRB

[Leo92a] Tom Leonard. Book review: [untitled]. Journal of the American Statistical Association, 87(420):1250, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290681.

Leone:1992:BRB

[Leo92b] Fred C. Leone. Book review: [untitled]. Journal of the American Statistical Association, 87(420):1239, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290666.

Levin:1994:BRB

[Lev94] Andrew Levin. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1143–1144, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290953.

Liu:1993:TAO

[LGWC93] P. Y. Liu, Stephanie Green, Michael Wolf, and John Crowley. Testing against ordered alternatives for censored survival data. Journal of the American Statistical Association, 88(421):153– 160, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290708.

Lin:1991:SAW

[LH91] Tsung-Hua Lin and David A. Harville. Some alternatives to Wald’s confidence interval and test. Journal of the American Statistical Association, 86(413):179–187, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289729.

LH:1996:BRB

[LH96] LH. Book review: [untitled]. Journal of the American Statistical Association, 91(436):1755, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291620.

LH:1997:BRBa

[LH97a] LH. Book review: [untitled]. Journal of the American Statistical Association, 92(437):389, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291496.

LH:1997:BRBb

[LH97b] LH. Book review: [untitled]. Journal of the American Statistical Association, 92(437):389, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291497.

LH:1997:BRBc

[LH97c] LH. Book review: [untitled]. Journal of the American Statistical Association, 92(438):799, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965744.

LH:1997:BRBd

[LH97d] LH. Book review: [untitled]. Journal of the American Statistical Association, 92(438):800, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965753.

LH:1997:BRBe

[LH97e] LH. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1228, September 1997. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965620.

Little:1994:ER

[LHA94] Roderick J. A. Little, Myles Hollander, and Alan Agresti. Editor’s report for 1993. Journal of the American Statistical Association, 89(426):367, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290831.

Lambert:1996:ER

[LHA96] Diane Lambert, Myles Hollander, and Alan Agresti. Editors’ report for 1995. Journal of the American Statistical Association, 91 (434):443, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291628.

Little:1995:ER

[LHAL95] Roderick J. A. Little, Myles Hollander, Alan Agresti, and Diane Lambert. Editor’s report for 1994. Journal of the American Statistical Association, 90(430):401, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291047.

Li:1991:SIRb

[Li91a] Ker-Chau Li. Sliced inverse regression for dimension reduction: Rejoinder. Journal of the American Statistical Association, 86 (414):337–342, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290568.

Li:1991:SIRa

[Li91b] Ker-Chau Li. Sliced inverse regression for dimension reduction. With discussion and a rejoinder by the author. Journal of the American Statistical Association, 86(414):316–327, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290563.

Li:1992:PHD

[Li92] Ker-Chau Li. On principal Hessian directions for data visualization and dimension reduction: Another application of Stein’s lemma. Journal of the American Statistical Association, 87(420): 1025–1039, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290640.

Li:1995:NLR

[Li95] Gang Li. Nonparametric likelihood ratio estimation of probabilities for truncated data. Journal of the American Statistical Association, 90(431):997–1003, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291336.

Li:1996:DTS

[Li96] Ta-Hsin Li. Discrimination of time series by parametric filtering. Journal of the American Statistical Association, 91(433): 284–293, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291406.

Lancaster:1998:PDS

[LI98a] Tony Lancaster and Orna Intrator. Panel data with survival: Hospitalization of HIV-positive patients. Journal of the American Statistical Association, 93(441):46–53, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
LANCASTE.HTM;   http://www.jstor.org/stable/2669601.

Li:1998:CPH

[Li98b] Ker-Chau Li. Comment: Principle Hessian directions revisited. Journal of the American Statistical Association, 93(441):94–??, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:1998:PHD

[Li98c] Ker-Chau Li. Principal Hessian directions revisited: Comment. Journal of the American Statistical Association, 93(441): 94–97, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2669606.

Li:1998:BRB

[Li98d] Ta-Hsin Li. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1525–1526, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670072.

Li:1999:BRB

[Li99] Ker-Chau Li. Book review: [untitled]. Journal of the American Statistical Association, 94(447):980, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670012.

Lieberman:1994:SAD

[Lie94] Offer Lieberman. Saddlepoint approximation for the distribution of a ratio of quadratic forms in normal variables. Journal of the American Statistical Association, 89(427):924–928, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290917.

Light:1994:BRB

[Lig94] Richard J. Light. Book review: [untitled]. Journal of the American Statistical Association, 89(428):1560–1562, December 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291021.

Liggett:1998:BRB

[Lig98] Walter Liggett. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1527–1528, December 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670077.

Lin:1991:GFA

[Lin91] D. Y. Lin. Goodness-of-fit analysis for the Cox regression model based on a class of parameter estimators. Journal of the American Statistical Association, 86(415):725–728, September 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290404.

Lind:1994:BRB

[Lin94] R. Bruce Lind. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1147, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290957.

Lin:1998:BRB

[Lin98] Dennis K. J. Lin. Book review: [untitled]. Journal of the American Statistical Association, 93(441):401–402, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669641.

Lindsay:1999:BRB

[Lin99a] Bruce G. Lindsay. Book review: [untitled]. Journal of the American Statistical Association, 94(446):646, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670184.

Linton:1999:BRB

[Lin99b] Oliver B. Linton. Book review: [untitled]. Journal of the American Statistical Association, 94(448):1388–1389, December 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669961.

Lipsitz:1995:BRB

[Lip95] Stuart R. Lipsitz. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1124–1125, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291353.

Little:1992:RMR

[Lit92] Roderick J. A. Little. Regression with missing X’s: a review. Journal of the American Statistical Association, 87(420):1227– 1237, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290664.

Little:1993:PSM

[Lit93a] R. J. A. Little. Post-stratification: a Modeler’s perspective. Journal of the American Statistical Association, 88(423):1001–1012, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290792.

Little:1993:HLR

[Lit93b] Roderick J. A. Little. Hierarchical logistic regression models for imputation of unresolved enumeration status in undercount estimation: Comment. Journal of the American Statistical Association, 88(423):1159–1161, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290813.

Little:1993:PMM

[Lit93c] Roderick J. A. Little. Pattern-mixture models for multivariate incomplete data. Journal of the American Statistical Association, 88(421):125–134, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290705.

Littell:1994:BRB

[Lit94] Ramon Littell. Book review: [untitled]. Journal of the American Statistical Association, 89(425):361–362, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291246.

Little:1995:MDM

[Lit95] Roderick J. A. Little. Modeling the drop-out mechanism in repeated-measures studies. Journal of the American Statistical Association, 90(431):1112–1121, September 1995. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2291350.

Littell:1996:BRB

[Lit96] Ramon C. Littell. Book review: [untitled]. Journal of the American Statistical Association, 91(434):912, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291693.

Liu:1994:CGS

[Liu94] Jun S. Liu. The collapsed Gibbs sampler in Bayesian computations with applications to a gene regulation problem. Journal of the American Statistical Association, 89(427):958–966, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290921.

Liu:1995:CCM

[Liu95] Regina Y. Liu. Control charts for multivariate processes. Journal of the American Statistical Association, 90(432):1380–1387, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291529.

Liu:1996:BRM

[Liu96] Chuanhai Liu. Bayesian robust multivariate linear regression with incomplete data. Journal of the American Statistical Association, 91(435):1219–1227, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2291740.

Liu:1997:BRB

[Liu97] Chuanhai Liu. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1223, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965603.

Lipsitz:1999:WEE

[LIZ99] Stuart R. Lipsitz, Joseph G. Ibrahim, and Lue Ping Zhao. A weighted estimating equation for missing covariate data with properties similar to maximum likelihood. Journal of the American Statistical Association, 94(448):1147–1160, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669931.

Lou:1993:COG

[LJK93] Sheldon Lou, Jiong Jiang, and Kenneth Keng. Clustering objects generated by linear regression models. Journal of the American Statistical Association, 88(424):1356–1362, December 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291277.

Ljungovist:1993:UAM

[Lju93a] Lars Ljungovist. A unified approach to measures of privacy in randomized response models: a utilitarian perspective. Journal of the American Statistical Association, 88(421):97–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ljungqvist:1993:UAM

[Lju93b] Lars Ljungqvist. A unified approach to measures of privacy in randomized response models: a utilitarian perspective. Journal of the American Statistical Association, 88(421):97–103, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290702.

Ljung:1995:BRB

[Lju95] Greta M. Ljung. Book review: [untitled]. Journal of the American Statistical Association, 90(429):394–395, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291179.

Langholz:1991:TDH

[LK91] Bryan Langholz and Richard A. Kronmal. Tests of distributional hypotheses with nuisance parameters using Fourier series methods. Journal of the American Statistical Association, 86(416): 1077–1084, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290527.

Lesaffre:1992:EUM

[LK92a] Emmanuel Lesaffre and Heinz Kaufmann. Existence and uniqueness of the maximum likelihood estimator for a multivariate probit model. Journal of the American Statistical Association, 87(419): 805–811, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290218.

Lesperance:1992:ACN

[LK92b] Mary L. Lesperance and John D. Kalbfleisch. An algorithm for computing the nonparametric MLE of a mixing distribution. Journal of the American Statistical Association, 87(417): 120–126, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290459.

Long:1995:FBD

[LK95] Dou Long and Roman Krzysztofowicz. A family of bivariate densities constructed from marginals. Journal of the American Statistical Association, 90(430):739–746, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291086.

Lam:1997:MLA

[LK97] K. F. Lam and Anthony Y. C. Kuk. A marginal likelihood approach to estimation in frailty models. Journal of the American Statistical Association, 92(439):985–990, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/sept/lam.htm;   http://www.jstor.org/stable/
2965562.

Lu:1997:SIM

[LKC97] Jiandong Lu, Daijin Ko, and Ted Chang. The standardized influence matrix and its applications. Journal of the American Statistical Association, 92(440):1572–1580, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965428.

Lahiri:1999:PSCa

[LKCH99a] Soumendra N. Lahiri, Mark S. Kaiser, Noel Cressie, and Nan-Jung Hsu. Prediction of spatial cumulative distribution functions using subsampling. Journal of the American Statistical Association, 94 (445):86–97, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2669680.

Lahiri:1999:PSCb

[LKCH99b] Soumendra N. Lahiri, Mark S. Kaiser, Noel Cressie, and Nan-Jung Hsu. Prediction of spatial cumulative distribution functions using subsampling: Rejoinder. Journal of the American Statistical Association, 94(445):107–110, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669686.

Lahiri:1999:RPS

[LKCH99c] Soumendra N. Lahiri, Mark S. Kaiser, Noel Cressie, and Nan-Jung Hsu. Rejoinder: Prediction of spatial cumulative distribution functions using subsampling. Journal of the American Statistical Association, 94(445):107–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lancaster:1998:RCS

[LKM98] Vicki A. Lancaster and Sallie Keller-McNulty. A review of composite sampling methods. Journal of the American Statistical Association, 93(443):1216–1230, September 1998. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2669863.

Lin:1997:ANC

[LL97a] Xihong Lin and Roderick J. Little. Analysis of nonrandomly censored ordered categorical longitudinal data from analgesic trials: Comment. Journal of the American Statistical Association, 92 (440):1245–1247, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2965392.

Lin:1997:CAN

[LL97b] Xihong Lin and Roderick J. Little. Comment: Analysis of nonrandomly censored ordered categorical longitudinal data from analgesic trials. Journal of the American Statistical Association, 92 (440):1245–??, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ling:1997:FIA

[LL97c] Shiqing Ling and W. K. Li. On fractionally integrated autoregressive moving-average time series models with conditional heteroscedasticity. Journal of the American Statistical Association, 92(439):1184–1194, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/sept/
ling.htm;   http://www.jstor.org/stable/2965585.

Lee:1993:RCP

[LLE93] J. Jack Lee, Ker-Chau Li, and Robert M. Elashoff. On recensoring for censored paired data. Journal of the American Statistical Association, 88(421):104–118, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290703.

Lloyd:1990:CID

[Llo90] Christopher J. Lloyd. Confidence intervals from the difference between two correlated proportions. Journal of the American Statistical Association, 85(412):1154–1158, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289615.

Lloyd:1998:USR

[Llo98] Chris J. Lloyd. Using smoothed receiver operating characteristic curves to summarize and compare diagnostic systems. Journal of the American Statistical Association, 93(444):1356– 1364, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/dec/lloyd.html;   http:
//www.jstor.org/stable/2670051.

Legler:1995:EPT

[LLR95] Julie M. Legler, Myrto Lefkopoulou, and Louise M. Ryan. Efficiency and power of tests for multiple binary outcomes. Journal of the American Statistical Association, 90(430):680–693, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291081.

Lye:1993:REN

[LM93] Jenny N. Lye and Vance L. Martin. Robust estimation, nonnormalities, and generalized exponential distributions. Journal of the American Statistical Association, 88(421):261–267, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290721.

Le:1996:MFS

[LMR96] Nhu D. Le, R. Douglas Martin, and Adrian E. Raftery. Modeling flat stretches, bursts, and outliers in time series using mixture transition distribution models. Journal of the American Statistical Association, 91(436):1504–1515, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291576.

Lang:1999:AMM

[LMS99] Joseph B. Lang, John W. McDonald, and Peter W. F. Smith. Association-marginal modeling of multivariate categorical responses: a maximum likelihood approach. Journal of the American Statistical Association, 94(448):1161–1171, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669932.

Liu:1995:BMM

[LNL95] Jun S. Liu, Andrew F. Neuwald, and Charles E. Lawrence. Bayesian models for multiple local sequence alignment and Gibbs sampling strategies. Journal of the American Statistical Association, 90(432):1156–1170, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291508.

Liu:1999:MSB

[LNL99] Jun S. Liu, Andrew F. Neuwald, and Charles E. Lawrence. Markovian structures in biological sequence alignments. Journal of the American Statistical Association, 94(445):1–15, March 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669673.

Lock:1990:BRB

[Loc90] Robin H. Lock. Book review: [untitled]. Journal of the American Statistical Association, 85(411):912, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290052.

Lohr:1996:BRB

[Loh96] Sharon L. Lohr. Book review: [untitled]. Journal of the American Statistical Association, 91(435):1374–1375, September 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291757.

Longford:1991:BRB

[Lon91] Nicholas T. Longford. Book review: [untitled]. Journal of the American Statistical Association, 86(413):245, March 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289740.

Lorenz:1992:BRB

[Lor92] Frederick O. Lorenz. Book review: [untitled]. Journal of the American Statistical Association, 87(419):902–903, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290239.

Louis:1992:HBM

[Lou92a] Thomas A. Louis. Hierarchical Bayes models for the progression of HIV infection using longitudinal CD4 T-cell numbers: Comment. Journal of the American Statistical Association, 87(419):626–628, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290195.

Louis:1992:SIA

[Lou92b] Thomas A. Louis. Statistical issues arising in AIDS clinical trials: Comment. Journal of the American Statistical Association, 87 (418):576–579, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290295.

Lou:1996:RLR

[Lou96a] W. Y. Wendy Lou. On runs and longest run tests: a method of finite Markov chain imbedding. Journal of the American Statistical Association, 91(436):1595–1601, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291585.

Louis:1996:BRB

[Lou96b] Thomas A. Louis. Book review: [untitled]. Journal of the American Statistical Association, 91(433):430–431, March 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291424.

Loughin:1999:BRB

[Lou99] Thomas M. Loughin. Book review: [untitled]. Journal of the American Statistical Association, 94(445):334–335, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669707.

Laird:1999:AND

[LP99] Nan M. Laird and Donna K. Pauler. Adjusting for nonignorable drop-out using semiparametric nonresponse models: Comment. Journal of the American Statistical Association, 94(448):1133– 1135, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2669929.

Lambert:1991:NDL

[LPT91] Diane Lambert, Bruce Peterson, and Irma Terpenning. Nondetects, detection limits, and the probability of detection. Journal of the American Statistical Association, 86(414):266–277, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290558.

Li:1997:SLR

[LQT97] Gang Li, Jing Qin, and Ram C. Tiwari. Semiparametric likelihood ratio — based inferences for truncated data. Journal of the American Statistical Association, 92(437):236– 245, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/march/li.htm;   http://
www.jstor.org/stable/2291468.

Lindsay:1992:RDM

[LR92] Bruce G. Lindsay and Kathryn Roeder. Residual diagnostics for mixture models. Journal of the American Statistical Association, 87(419):785–794, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290216.

Lahiri:1995:REM

[LR95a] P. Lahiri and J. N. K. Rao. Robust estimation of mean squared error of small area estimators. Journal of the American Statistical Association, 90(430):758–766, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291089.

Lambert:1995:ODG

[LR95b] Diane Lambert and Kathryn Roeder. Overdispersion diagnostics for generalized linear models. Journal of the American Statistical Association, 90(432):1225–1236, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291513.

Lawrence:1996:LIP

[LR96] Charles Lawrence and Andrew Reilly. Likelihood inference for permuted data with application to gene regulation. Journal of the American Statistical Association, 91(433):76–85, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291384.

Legler:1997:LVM

[LR97a] Julie M. Legler and Louise M. Ryan. Latent variable models for teratogenesis using multiple binary outcomes. Journal of the American Statistical Association, 92(437):13–20, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291445.

Lewis:1997:MLR

[LR97b] Peter A. W. Lewis and Bonnie K. Ray. Modeling long-range dependence, nonlinearity, and periodic phenomena in sea surface temperatures using TSMARS. Journal of the American Statistical Association, 92(439):881–893, September 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/sept/
lewis_aw.htm;   http://www.jstor.org/stable/2965552.

Lewis:1997:EBF

[LR97c] Steven M. Lewis and Adrian E. Raftery. Estimating Bayes factors via posterior simulation with the Laplace-Metropolis estimator. Journal of the American Statistical Association, 92(438):648–655, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2965712.

Little:1999:AND

[LR99] Roderick J. Little and Donald B. Rubin. Adjusting for nonignorable drop-out using semiparametric nonresponse models: Comment. Journal of the American Statistical Association, 94(448): 1130–1132, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2669928.

Le:1996:RBM

[LRM96] Nhu D. Le, Adrian E. Raftery, and R. Douglas Martin. Robust Bayesian model selection for autoregressive processes with additive outliers. Journal of the American Statistical Association, 91 (433):123–131, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291388.

Li:1991:LSS

[LRR91] K. H. Li, T. E. Raghunathan, and D. B. Rubin. Large-sample significance levels from multiply imputed data using moment-based statistics and an F reference distribution. Journal of the American Statistical Association, 86(416):1065–1073, December 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290525.

Lewis:1991:NMT

[LS91a] P. A. W. Lewis and J. G. Stevens. Nonlinear modeling of time series using multivariate adaptive regression splines (MARS). Journal of the American Statistical Association, 86(416):864–877, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290499.

Lindley:1991:ENR

[LS91b] Dennis V. Lindley and Nozer D. Singpurwalla. On the evidence needed to reach agreed action between adversaries, with application to acceptance sampling. Journal of the American Statistical Association, 86(416):933–937, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290509.

Liu:1993:QIB

[LS93] Regina Y. Liu and Kesar Singh. A quality index based on data depth and multivariate rank tests. Journal of the American Statistical Association, 88(421):252–260, March 1993. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2290720.

Lee:1995:ELN

[LS95] Lung Fei Lee and Jungsywan H. Sepanski. Estimation of linear and nonlinear errors-in-variables models using validation data. Journal of the American Statistical Association, 90(429):130– 140, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291136.

Liu:1997:NLV

[LS97] Regina Y. Liu and Kesar Singh. Notions of limiting P values based on data depth and bootstrap. Journal of the American Statistical Association, 92(437):266–277, March 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/march/
liu.htm;   http://www.jstor.org/stable/2291471.

Liu:1999:DRD

[LS99a] Regina Liu and Kesar Singh. Discussion: Regression depth. Journal of the American Statistical Association, 94(446):407–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:1999:RDC

[LS99b] Regina Y. Liu and Kesar Singh. Regression depth: Comment. Journal of the American Statistical Association, 94(446):407–409, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670158.

LeBlanc:1994:APS

[LT94a] Michael LeBlanc and Robert Tibshirani. Adaptive principal surfaces. Journal of the American Statistical Association, 89(425): 53–64, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291200.

Lee:1994:SPF

[LT94b] Ronald D. Lee and Shripad Tuljapurkar. Stochastic population forecasts for the United States: Beyond high, medium, and low. Journal of the American Statistical Association, 89(428):1175– 1189, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290980.

LeBlanc:1996:CER

[LT96a] Michael LeBlanc and Robert Tibshirani. Combining estimates in regression and classification. Journal of the American Statistical Association, 91(436):1641–1650, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291591.

Liu:1996:CCD

[LT96b] Regina Y. Liu and Jen Tang. Control charts for dependent and independent measurements based on bootstrap methods. Journal of the American Statistical Association, 91(436):1694–1700, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291598.

Lambert:1997:NML

[LT97] Diane Lambert and Luke Tierney. Nonparametric maximum likelihood estimation from samples with irrelevant data and verification bias. Journal of the American Statistical Association, 92 (439):937–944, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/sept/lambert.htm;
http://www.jstor.org/stable/2965557.

Li:1998:UAI

[LT98] Hong Li and Ruey S. Tsay. A unified approach to identifying multivariate time series models. Journal of the American Statistical Association, 93(442):770–782, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.amstat.org/publications/jasa/abstracts_98/june/LI.
HTM;   http://www.jstor.org/stable/2670127.

Laska:1992:THA

[LTM92] Eugene M. Laska, Dei-In Tang, and Morris J. Meisner. Testing hypotheses about an identified treatment when there are multiple endpoints. Journal of the American Statistical Association, 87 (419):825–831, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290221.

Litvak:1994:SPD

[LTP94] Eugene Litvak, Xin M. Tu, and Marcello Pagano. Screening for the presence of a disease by pooling sera samples. Journal of the American Statistical Association, 89(426):424–434, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290842.

Li:1996:QCF

[LTW96] Gang Li, Ram C. Tiwari, and Martin T. Wells. Quantile comparison functions in two-sample problems, with application to comparisons of diagnostic markers. Journal of the American Statistical Association, 91(434):689–698, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291664.

Lucas:1993:WCN

[Luc93a] T. W. Lucas. When is conflict normal? Journal of the American Statistical Association, 88(424):1433–1437, December 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291288.

Luceno:1993:FAR

[Luc93b] Alberto Luceno. A fast algorithm for the repeated evaluation of the likelihood of a general linear process for long series. Journal of the American Statistical Association, 88(421):229–236, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290717.

Luckey:1998:BRB

[Luc98] Dennis W. Luckey. Book review: [untitled]. Journal of the American Statistical Association, 93(441):409, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669654.

Lund:1997:BRBa

[Lun97a] Robert Lund. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1224–1225, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965605.

Lund:1997:BRBb

[Lun97b] Robert Lund. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1225, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965606.

Lund:1998:BRBb

[Lun98a] Robert Lund. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1242–1243, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669880.

Lund:1998:BRBa

[Lun98b] Robert B. Lund. Book review: [untitled]. Journal of the American Statistical Association, 93(442):842, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670143.

Lund:1998:BRBc

[Lun98c] Robert B. Lund. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1528, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670078.

Lund:1998:BRBd

[Lun98d] Robert B. Lund. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1528–1529, December 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670079.

Lund:1999:BRB

[Lun99] Robert B. Lund. Book review: [untitled]. Journal of the American Statistical Association, 94(446):654–655, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670196.

LuValle:1991:BRB

[LuV91] Michael LuValle. Book review: [untitled]. Journal of the American Statistical Association, 86(415):824–825, September 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290424.

Lian:1990:ESE

[LW90a] Kung-Yee Lian and Myron A. Waclawiw. Extension of the Stein estimating procedure through the use of estimating functions. Journal of the American Statistical Association, 85(410):435–440, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289780.

Liang:1990:ESE

[LW90b] Kung-Yee Liang and Myron A. Waclawiw. Extension of the Stein estimating procedure through the use of estimating functions. Journal of the American Statistical Association, 85(410):435–440, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Little:1991:MCT

[LW91] Roderick J. A. Little and Mei-Miau Wu. Models for contingency tables with known margins when target and sampled populations differ. Journal of the American Statistical Association, 86(413): 87–95, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289718.

Lin:1992:LRA

[LW92] J. S. Lin and L. J. Wei. Linear regression analysis for multivariate failure time observations. Journal of the American Statistical Association, 87(420):1091–1097, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290646.

Luo:1997:HAS

[LW97] Zhen Luo and Grace Wahba. Hybrid adaptive splines. Journal of the American Statistical Association, 92(437):107– 116, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/march/luo.htm;   http:/
/www.jstor.org/stable/2291454.

Liu:1999:PED

[LW99] Jun S. Liu and Ying Nian Wu. Parameter expansion for data augmentation. Journal of the American Statistical Association, 94 (448):1264–1274, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669940.

Little:1993:ER

[LWG93] Roderick J. A. Little, Edward J. Wegman, and Donald Guthrie. Editors’ report for 1992. Journal of the American Statistical Association, 88(422):389, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290316.

Lu:1994:ISF

[LWT94] Henry H. S. Lu, Martin T. Wells, and Ram C. Tiwari. Inference for shift functions in the two-sample problem with rightcensored data: With applications. Journal of the American Statistical Association, 89(427):1017–1026, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290929.

Lavine:1991:BIS

[LWW91] Michael Lavine, Larry Wasserman, and Robert L. Wolpert. Bayesian inference with specified prior marginals. Journal of the American Statistical Association, 86(416):964–971, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290512.

Lee:1993:LRA

[LWY93] Eric W. Lee, L. J. Wei, and Z. Ying. Linear regression analysis for highly stratified failure time data. Journal of the American Statistical Association, 88(422):557–565, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290336.

Lin:1993:CRI

[LY93] D. Y. Lin and Z. Ying. Cox regression with incomplete covariate measurements. Journal of the American Statistical Association, 88(424):1341–1349, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291275.

Lyerla:1997:BRB

[Lye97] Rob Lyerla. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1223, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965602.

Lynch:1998:BRB

[Lyn98] James Lynch. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1244, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669883.

Lin:1999:GTS

[LYY99] D. Y. Lin, Q. Yao, and Zhiliang Ying. A general theory on stochastic curtailment for censored survival data. Journal of the American Statistical Association, 94(446):510–521, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670171.

Lee:1998:ACS

[LZ98a] Sandra J. Lee and Marvin Zelen. Applications and case studies — scheduling periodic examinations for the early detection of disease: Applications to breast cancer. Journal of the American Statistical Association, 93(444):1271–1281, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/dec/lee.html.

Lee:1998:SPE

[LZ98b] Sandra J. Lee and Marvin Zelen. Scheduling periodic examinations for the early detection of disease: Applications to breast cancer. Journal of the American Statistical Association, 93(444): 1271–1281, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2670042.

Marshall:1995:BRBb

[MA95] Guillermo Marshall and Renato Allende. Book review: [untitled]. Journal of the American Statistical Association, 90(432):1491– 1492, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291549.

Machler:1994:KSE

[Mac94] Martin B. Machler. Kriging and splines: An empirical comparison of their predictive performance in some applications: Comment. Journal of the American Statistical Association, 89(426):403–405, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290839.

Madura:1993:BRB

[Mad93] Jeff Madura. Book review: [untitled]. Journal of the American Statistical Association, 88(422):704, June 1993. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2290366.

Mallows:1995:BRB

[Mal95] Colin Mallows. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1133, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291363.

Manski:1990:UID

[Man90] Charles F. Manski. The use of intentions data to predict behavior: a best-case analysis. Journal of the American Statistical Association, 85(412):934–940, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289590.

Mann:1994:MWE

[Man94] Nancy R. Mann. In memoriam: W. Edwards Deming 1900–1993. Journal of the American Statistical Association, 89(426):365–366, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290830.

Marden:1990:BRB

[Mar90a] John I. Marden. Book review: [untitled]. Journal of the American Statistical Association, 85(410):598–599, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289817.

Martin:1990:BRB

[Mar90b] Michael A. Martin. Book review: [untitled]. Journal of the American Statistical Association, 85(409):264–265, March 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289572.

Martin:1990:BIC

[Mar90c] Michael A. Martin. On bootstrap iteration for coverage correction in confidence intervals. Journal of the American Statistical Association, 85(412):1105–1118, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289608.

Marden:1992:BRB

[Mar92a] John I. Marden. Book review: [untitled]. Journal of the American Statistical Association, 87(417):249–250, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290477.

Marden:1992:UNO

[Mar92b] John I. Marden. Use of nested orthogonal contrasts in analyzing rank data. Journal of the American Statistical Association, 87 (418):307–318, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290261.

Marx:1992:BRB

[Mar92c] Brian D. Marx. Book review: [untitled]. Journal of the American Statistical Association, 87(420):1242, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290670.

Marasinghe:1994:BRB

[Mar94] Mervyn G. Marasinghe. Book review: [untitled]. Journal of the American Statistical Association, 89(426):720, June 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290890.

Marshall:1995:BRBa

[Mar95a] Pablo Marshall. Book review: [untitled]. Journal of the American Statistical Association, 90(429):393–394, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291178.

Martin:1995:BRB

[Mar95b] Frank B. Martin. Book review: [untitled]. Journal of the American Statistical Association, 90(430):802–803, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291100.

Marron:1998:BRB

[Mar98a] J. S. Marron. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1232, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669864.

Martin:1998:BRB

[Mar98b] Nadine W. Martin. Book review: [untitled]. Journal of the American Statistical Association, 93(441):409–410, March 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669655.

Martinsek:1998:BRB

[Mar98c] Adam T. Martinsek. Book review: [untitled]. Journal of the American Statistical Association, 93(441):405–406, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669648.

Marx:1998:BRB

[Mar98d] Brian D. Marx. Book review: [untitled]. Journal of the American Statistical Association, 93(441):399–400, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669639.

Mason:1992:BRB

[Mas92] Robert L. Mason. Book review: [untitled]. Journal of the American Statistical Association, 87(418):591–592, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290310.

Mason:1998:BRB

[Mas98] Thomas J. Mason. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1232–1233, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669865.

Matthews:1993:BRB

[Mat93] Peter Matthews. Book review: [untitled]. Journal of the American Statistical Association, 88(422):699, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290360.

Matis:1994:BRB

[Mat94] James H. Matis. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1146–1147, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290956.

Mathar:1999:BRB

[Mat99] Rudolf Mathar. Book review: [untitled]. Journal of the American Statistical Association, 94(448):1387, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669957.

Mauricio:1995:EML

[Mau95] José Alberto Mauricio. Exact maximum likelihood estimation of stationary vector ARMA models. Journal of the American Statistical Association, 90(429):282–291, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291152.

Markatou:1998:WLE

[MBL98] Marianthi Markatou, Ayanendranath Basu, and Bruce G. Lindsay. Weighted likelihood equations with bootstrap root search. Journal of the American Statistical Association, 93 (442):740–750, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/june/MARKATOU.HTM;
 http://www.jstor.org/stable/2670124.

Magee:1991:CKS

[MBR91] Lonnie Magee, John B. Burbidge, and A. Leslie Robb. Computing kernel-smoothed conditional quantiles from many observations. Journal of the American Statistical Association, 86(415):673–677, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290396.

Moreno:1998:ILP

[MBR98] Elias Moreno, Francesco Bertolino, and Walter Racugno. An intrinsic limiting procedure for model selection and hypotheses testing. Journal of the American Statistical Association, 93 (444):1451–1460, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/dec/moreno.html;
http://www.jstor.org/stable/2670059.

McCulloch:1994:MLV

[McC94] Charles E. McCulloch. Maximum likelihood variance components estimation for binary data. Journal of the American Statistical Association, 89(425):330–335, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291229.

McCulloch:1997:MLA

[McC97] Charles E. McCulloch. Maximum likelihood algorithms for generalized linear mixed models. Journal of the American Statistical Association, 92(437):162–170, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/march/
mcculloch.htm;   http://www.jstor.org/stable/2291460.

McCulloch:1998:BRB

[McC98a] Charles E. McCulloch. Book review: [untitled]. Journal of the American Statistical Association, 93(441):403–404, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669645.

McCulloch:1998:BIN

[McC98b] Robert McCulloch. Bayesian inference on network traffic using link count data: Comment. Journal of the American Statistical Association, 93(442):575, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2670107.

McCulloch:1998:CBI

[McC98c] Robert McCulloch. Comment: Bayesian inference on network traffic using link count data. Journal of the American Statistical Association, 93(442):575–??, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McDonald:1990:BRB

[McD90] Roderick P. McDonald. Book review: [untitled]. Journal of the American Statistical Association, 85(412):1175–1176, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289630.

McDonald:1992:BRB

[McD92] Roderick P. McDonald. Book review: [untitled]. Journal of the American Statistical Association, 87(419):908, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290247.

McDonald:1997:BRB

[McD97] Roderick P. McDonald. Book review: [untitled]. Journal of the American Statistical Association, 92(437):386, March 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291488.

McGarity:1990:BRB

[McG90] Arthur E. McGarity. Book review: [untitled]. Journal of the American Statistical Association, 85(409):265, March 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289573.

McKenzie:1992:BRB

[McK92] John D. McKenzie, Jr. Book review: [untitled]. Journal of the American Statistical Association, 87(420):1241–1242, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290669.

McKenzie:1995:BRB

[McK95] Sandra McKenzie. Book review: [untitled]. Journal of the American Statistical Association, 90(432):1491, December 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291548.

McNown:1992:MFU

[McN92] Robert McNown. Modeling and forecasting U.S. mortality: Comment. Journal of the American Statistical Association, 87(419): 671–672, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290202.

McNeil:1997:BRB

[McN97] Alexander J. McNeil. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1654, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2965455.

McNeil:1998:BRBa

[McN98a] Alexander J. McNeil. Book review: [untitled]. Journal of the American Statistical Association, 93(442):839–840, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670140.

McNeil:1998:BRBb

[McN98b] Alexander J. McNeil. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1516–1518, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670066.

McWilliams:1990:DFT

[McW90] Thomas P. McWilliams. A distribution-free test for symmetry based on a runs statistic. Journal of the American Statistical Association, 85(412):1130–1133, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289611.

McCann:1996:PLI

[ME96] Melinda McCann and Don Edwards. A path length inequality for the multivariate-t distribution, with applications to multiple comparisons. Journal of the American Statistical Association, 91 (433):211–216, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291397.

Mee:1990:CIP

[Mee90] Robert W. Mee. Confidence intervals for probabilities and tolerance regions based on a generalization of the Mann–Whitney statistic. Journal of the American Statistical Association, 85(411): 793–800, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290017.

Mee:1995:BRB

[Mee95] Robert W. Mee. Book review: [untitled]. Journal of the American Statistical Association, 90(429):397, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291183.

McCaffrey:1992:ELE

[MEGN92] Daniel F. McCaffrey, Stephen Ellner, A. Ronald Gallant, and Douglas W. Nychka. Estimating the Lyapunov exponent of a chaotic system with nonparametric regression. Journal of the American Statistical Association, 87(419):682–695, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290206.

Meier:1991:CEV

[Mei91] Paul Meier. Compliance as an explanatory variable in clinical trials: Comment. Journal of the American Statistical Association, 86(413):19–22, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289709.

Meilijson:1994:CRC

[Mei94] Isaac Meilijson. Competing risks on coherent reliability systems: Estimation in the parametric case. Journal of the American Statistical Association, 89(428):1459–1464, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291007.

Melnick:1990:BRB

[Mel90] Daniel Melnick. Book review: [untitled]. Journal of the American Statistical Association, 85(410):602–603, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289824.

Mellen:1999:MET

[Mel99] Beverly G. Mellen. Modeling epidemiologic typing data and likelihood inference for disease spread. Journal of the American Statistical Association, 94(448):1015–1024, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669915.

Merrill:1994:PMSa

[Mer94a] Samuel Merrill, III. A probabilistic model for the spatial distribution of party support in multiparty electorates. Journal of the American Statistical Association, 89(428):1190–1197, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290981.

Merrill:1994:PMSb

[Mer94b] Samuel Merrill, III. A probabilistic model for the spatial distribution of party support in multiparty electorates: Rejoinder. Journal of the American Statistical Association, 89(428):1199, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290983.

Merrill:1994:RPM

[Mer94c] Samuel Merrill, III. Rejoinder: a probabilistic model for the spatial distribution of party support in multiparty electorates. Journal of the American Statistical Association, 89(428):1199–??, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Meulman:1995:BRB

[Meu95] Jacqueline J. Meulman. Book review: [untitled]. Journal of the American Statistical Association, 90(429):395–396, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291181.

Meyer:1999:BRB

[Mey99] Mary C. Meyer. Book review: [untitled]. Journal of the American Statistical Association, 94(447):988, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670024.

Monahan:1997:SRI

[MG97a] John Monahan and Alan Genz. Spherical-radial integration rules for Bayesian computation. Journal of the American Statistical Association, 92(438):664–674, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2965714.

Mukhopadhyay:1997:DPM

[MG97b] Saurabh Mukhopadhyay and Alan E. Gelfand. Dirichlet process mixed generalized linear models. Journal of the American Statistical Association, 92(438):633–639, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965710.

Markatou:1990:RBI

[MH90] Marianthi Markatou and Thomas P. Hettmansperger. Robust bounded-influence tests in linear models. Journal of the American Statistical Association, 85(409):187–190, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289543.

Markatou:1994:BIH

[MH94a] Marianthi Markatou and Xuming He. Bounded influence and high breakdown point testing procedures in linear models. Journal of the American Statistical Association, 89(426):543–549, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290857.

McArdle:1994:LML

[MH94b] John J. McArdle and Fumiaki Hamagami. Logit and multilevel logit modeling of college graduation for 1984–1985 freshman student-athletes. Journal of the American Statistical Association, 89(427):1107–1123, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290941.

Miwa:1999:CAO

[MH99] Tetsuhisa Miwa and Tony Hayter. Combining the advantages of one-sided and two-sided test procedures for comparing several treatment effects. Journal of the American Statistical Association, 94(445):302–307, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2669704.

Mickey:1997:BRB

[Mic97] Ruth M. Mickey. Book review: [untitled]. Journal of the American Statistical Association, 92(437):384, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291485.

Mickey:1998:BRBb

[Mic98a] Ruth Mickey. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1239, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669873.

Mickey:1998:BRBa

[Mic98b] Ruth M. Mickey. Book review: [untitled]. Journal of the American Statistical Association, 93(442):839, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670139.

Mikosch:1993:BRB

[Mik93] Thomas Mikosch. Book review: [untitled]. Journal of the American Statistical Association, 88(422):702, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290363.

Mikosch:1994:BRB

[Mik94] Thomas Mikosch. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1147–1148, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290958.

Mikosch:1995:BRBa

[Mik95a] Thomas Mikosch. Book review: [untitled]. Journal of the American Statistical Association, 90(430):805–806, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291104.

Mikosch:1995:BRBb

[Mik95b] Thomas Mikosch. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1133–1134, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291364.

Mikosch:1997:BRB

[Mik97] Thomas Mikosch. Book review: [untitled]. Journal of the American Statistical Association, 92(438):794, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965734.

Mikosch:1998:BRB

[Mik98] Thomas Mikosch. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1243, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669881.

Millar:1992:ESS

[Mil92] Russell B. Millar. Estimating the size-selectivity of fishing gear by conditioning on the total catch. Journal of the American Statistical Association, 87(420):962–968, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290632.

Minkin:1993:EDC

[Min93] Salomon Minkin. Experimental design for clonogenic assays in chemotherapy. Journal of the American Statistical Association, 88(422):410–420, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290319.

Minnotte:1998:AHO

[Min98] Michael C. Minnotte. Achieving higher-order convergence rates for density estimation with binned data. Journal of the American Statistical Association, 93(442):663–672, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
june/MINNOTTE.HTM;   http://www.jstor.org/stable/2670117.

Mittal:1991:HSS

[Mit91] Yashaswini Mittal. Homogeneity of subpopulations and Simpson’s paradox. Journal of the American Statistical Association, 86(413): 167–172, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289727.

Mitchell:1994:BRB

[Mit94] Ann F. S. Mitchell. Book review: [untitled]. Journal of the American Statistical Association, 89(426):719, June 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290887.

Mason:1990:CLE

[MKSB90] Robert L. Mason, Jerome P. Keating, Pranab K. Sen, and Neil W. Blaylock, Jr. Comparison of linear estimators using Pitman’s measure of closeness. Journal of the American Statistical Association, 85(410):579–581, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289801.

Molenberghs:1994:MMC

[ML94] Geert Molenberghs and Emmanuel Lesaffre. Marginal modeling of correlated ordinal data using a multivariate Plackett distribution. Journal of the American Statistical Association, 89(426):633–644, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290866.

Mesenbrink:1994:COW

[MLMT94] Peter Mesenbrink, Jye-Chyi Lu, Richard McKenzie, and Javad Taheri. Characterization and optimization of a wave-soldering process. Journal of the American Statistical Association, 89(428): 1209–1217, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290985.

Muller:1992:PCG

[MLRR92] Keith E. Muller, Lisa M. La Vange, Sharon Landesman Ramey, and Craig T. Ramey. Power calculations for general linear multivariate models including repeated measures applications. Journal of the American Statistical Association, 87(420):1209–1226, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290663.

McDermott:1993:SAT

[MM93] Michael P. McDermott and Govind S. Mudholkar. A simple approach to testing homogeneity of order-constrained means. Journal of the American Statistical Association, 88(424):1371–1379, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291279.

Marden:1995:RTM

[MM95] John I. Marden and Maria Emerita T. Muyot. Rank tests for main and interaction effects in analysis of variance. Journal of the American Statistical Association, 90(432):1388–1398, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291530.

Meier:1993:NER

[MN93] Kristen Meier and Douglas Nychka. Nonparametric estimation of rate equations for nutrient uptake. Journal of the American Statistical Association, 88(422):602–614, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290342.

Miller:1999:IMP

[MNS99] Forrest R. Miller, James W. Neill, and Brian W. Sherfey. Implementation of a maximin power clustering criterion to select near replicates for regression lack-of-fit tests. Journal of the American Statistical Association, 94(446):610–620, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670181.

Maitra:1998:VAP

[MO98] Ranjan Maitra and Finbarr O’Sullivan. Variability assessment in positron emission tomography and related generalized deconvolution models. Journal of the American Statistical Association, 93 (444):1340–1355, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/dec/maitra.html;
http://www.jstor.org/stable/2670050.

Model:1993:EMD

[Mod93] Karyn E. Model. The effect of marijuana decriminalization on hospital emergency Room drug episodes: 1975–1978. Journal of the American Statistical Association, 88(423):737–747, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290758.

Modarres:1995:BRB

[Mod95] Reza Modarres. Book review: [untitled]. Journal of the American Statistical Association, 90(429):384–385, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291167.

Moffitt:1996:CIC

[Mof96a] Robert A. Moffitt. Comment: Identification of causal effects using instrumental variables. Journal of the American Statistical Association, 91(434):444–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Moffitt:1996:ICE

[Mof96b] Robert A. Moffitt. Identification of causal effects using instrumental variables: Comment. Journal of the American Statistical Association, 91(434):462–465, June 1996. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291632.

Mojirsheibani:1999:CCD

[Moj99] Majid Mojirsheibani. Combining classifiers via discretization. Journal of the American Statistical Association, 94(446):600–609, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670180.

Moore:1991:BRB

[Moo91] Dirk F. Moore. Book review: [untitled]. Journal of the American Statistical Association, 86(416):1142–1143, December 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290541.

Moore:1998:PAS

[Moo98a] David S. Moore. Presidential address — statistics among the liberal arts. Journal of the American Statistical Association, 93 (444):1253–??, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/moore.html.

Moore:1998:SAL

[Moo98b] David S. Moore. Statistics among the liberal arts. Journal of the American Statistical Association, 93(444):1253–1259, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670040.

Moore:1999:BRB

[Moo99] Dirk F. Moore. Book review: [untitled]. Journal of the American Statistical Association, 94(445):348, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669721.

Morgan:1990:BRB

[Mor90] Timothy M. Morgan. Book review: [untitled]. Journal of the American Statistical Association, 85(412):1172, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289626.

Morettin:1991:WFA

[Mor91a] Pedro A. Morettin. Walsh–Fourier analysis and its statistical applications: Comment. Journal of the American Statistical Association, 86(414):482–483, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290598.

Morganstein:1991:BRB

[Mor91b] David Morganstein. Book review: [untitled]. Journal of the American Statistical Association, 86(414):550–551, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290617.

Morris:1993:BRBa

[Mor93a] John D. Morris. Book review: [untitled]. Journal of the American Statistical Association, 88(421):387–388, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290752.

Morris:1993:BRBb

[Mor93b] John D. Morris. Book review: [untitled]. Journal of the American Statistical Association, 88(424):1472–1473, December 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291303.

Morris:1996:CEB

[Mor96a] Carl N. Morris. Comment: Empirical Bayes methods for combining likelihoods. Journal of the American Statistical Association, 91(434):538–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Morris:1996:EBM

[Mor96b] Carl N. Morris. Empirical Bayes methods for combining likelihoods: Comment. Journal of the American Statistical Association, 91(434):555–558, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291649.

Mouhab:1995:BRB

[Mou95] Abderrahamane Mouhab. Book review: [untitled]. Journal of the American Statistical Association, 90(432):1489–1490, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291545.

Moynihan:1999:DDP

[Moy99] Daniel Patrick Moynihan. Data and dogma in public policy. Journal of the American Statistical Association, 94(446):359–364, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/moynihan.pdf;   http://www.jstor.org/stable/2670152.

Muller:1995:ODC

[MP95] Peter Müller and Giovanni Parmigiani. Optimal design via curve fitting of Monte Carlo experiments. Journal of the American Statistical Association, 90(432):1322–1330, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291522.

Mohanty:1996:EGP

[MP96] R. Mohanty and M. Pourahmadi. Estimation of the generalized prediction error variance of a multiple time series. Journal of the American Statistical Association, 91(433):294–299, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291407.

Munk:1999:ECR

[MP99] Axel Munk and Rafael Pfluger. 1 − α equivariant confidence rules for convex alternatives are α/2-level tests — with applications to the multivariate assessment of bioequivalence. Journal of the American Statistical Association, 94(448):1311–1319, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669944.

Morrell:1995:EUT

[MPCB95] Christopher H. Morrell, Jay D. Pearson, H. Ballentine Carter, and Larry J. Brant. Estimating unknown transition times using a piecewise nonlinear mixed-effects model in men with prostate cancer. Journal of the American Statistical Association, 90(429): 45–53, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291128.

Mehta:1998:AEI

[MPS98a] Cyrus R. Mehta, Nitin R. Patel, and Pralay Senchaudhuri. Approximately exact inference for the common odds ration in several 2×2 tables: Comment. Journal of the American Statistical Association, 93(444):1313–1316, December 1998. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670047.

Mehta:1998:CAE

[MPS98b] Cyrus R. Mehta, Nitin R. Patel, and Pralay Senchaudhuri. Comment: Approximately exact inference for the common odds ratio in several 2×2 tables. Journal of the American Statistical Association, 93(444):1313–??, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Meng:1991:UEO

[MR91] Xiao-Li Meng and Donald B. Rubin. Using EM to obtain asymptotic variance-covariance matrices: The SEM algorithm. Journal of the American Statistical Association, 86(416):899–909, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290503.

Madigan:1994:MSA

[MR94] David Madigan and Adrian E. Raftery. Model selection and accounting for model uncertainty in graphical models using Occam’s Window. Journal of the American Statistical Association, 89 (428):1535–1546, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291017.

Moller:1997:BPM

[MR97a] Peter Möller and Gary Rosner. A Bayesian population model with hierarchical mixture priors applied to blood count data. Journal of the American Statistical Association, 92(440):1279– 1292, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/MULLER.HTM.

Muller:1997:BPM

[MR97b] Peter Muller and Gary L. Rosner. A Bayesian population model with hierarchical mixture priors applied to blood count data. Journal of the American Statistical Association, 92(440):1279– 1292, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/MULLER.HTM;
http://www.jstor.org/stable/2965398.

Murphy:1997:MLE

[MRvdV97] S. A. Murphy, A. J. Rossini, and A. W. van der Vaart. Maximum likelihood estimation in the proportional odds model. Journal of the American Statistical Association, 92(439):968–976, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_97/sept/murphy.htm;   http://www.jstor.
org/stable/2965560.

Mehta:1990:BRB

[MS90] J. S. Mehta and P. A. V. B. Swamy. Book review: [untitled]. Journal of the American Statistical Association, 85(409):257, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289559.

Muller:1991:EME

[MS91a] D. W. Müller and G. Sawitzki. Excess mass estimates and tests for multimodality. Journal of the American Statistical Association, 86(415):738–746, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2290406. See acknowledgement of priority [MS95a].

Mulry:1991:TEPa

[MS91b] Mary H. Mulry and Bruce D. Spencer. Total error in PES estimates of population. Journal of the American Statistical Association, 86(416):839–855, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290495.

Mulry:1991:TEPb

[MS91c] Mary H. Mulry and Bruce D. Spencer. Total error in PES estimates of population: Rejoinder. Journal of the American Statistical Association, 86(416):861–863, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290498.

Mathew:1992:EOT

[MS92] Thomas Mathew and Bimal Kumar Sinha. Exact and optimum tests in unbalanced split-plot designs under mixed and random models. Journal of the American Statistical Association, 87(417): 192–200, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290469.

Mulry:1993:ACU

[MS93] Mary H. Mulry and Bruce D. Spencer. Accuracy of the 1990 census and undercount adjustments. Journal of the American Statistical Association, 88(423):1080–1091, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290803.

Mukarjee:1994:FNE

[MS94] Hari Mukarjee and Steven Stern. Feasible nonparametric estimation of multiargument monotone functions. Journal of the American Statistical Association, 89(425):77–80, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291202.

Muller:1995:APE

[MS95a] D. W. Müller and G. Sawitzki. Acknowledgement of priority: “Excess mass estimates and tests for multimodality” [J. Amer. Statist. Assoc. 86 (1991), no. 415, 738–746, MR 92i:62091]. Journal of the American Statistical Association, 90(429):399, ???? 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [MS91a].

Muller:1995:CEM

[MS95b] D. W. Muller and G. Sawitzki. Corrections: Excess mass estimates and tests for multimodality. Journal of the American Statistical Association, 90(429):399, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291192.

Meng:1996:FFI

[MS96] Xiao-Li Meng and Stephen Schilling. Fitting full-information item factor models and an empirical investigation of bridge sampling. Journal of the American Statistical Association, 91(435): 1254–1267, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291744.

McCabe:1998:EPS

[MS98a] Brendan P. M. McCabe and Richard J. Smith. Errata: “The power of some tests for difference stationarity under local heteroscedastic integration” [J. Amer. Statist. Assoc. 93 (1998), no.

442, 751–761; 1 631 382]. Journal of the American Statistical Association, 93(443):1252, 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [MS98b].

McCabe:1998:PST

[MS98b] Brendan P. M. McCabe and Richard J. Smith. The power of some tests for difference stationarity under local heteroscedastic integration. Journal of the American Statistical Association, 93(442):751–761, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

amstat.org/publications/jasa/abstracts_98/june/MCCABE.
HTM;   http://www.jstor.org/stable/2670125. See errata

[MS98a].

MacEachern:1999:VSF

[MS99a] Steven N. MacEachern and Xiaotong Shen. Variable selection and function estimation in additive nonparametric regression using a data-based prior: Comment. Journal of the American Statistical Association, 94(447):799–802, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669993.

McKean:1999:DRD

[MS99b] Joseph W. McKean and Simon J. Sheather. Discussion: Regression depth. Journal of the American Statistical Association, 94 (446):411–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McKean:1999:RDC

[MS99c] Joseph W. McKean and Simon J. Sheather. Regression depth: Comment. Journal of the American Statistical Association, 94 (446):411–415, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2670160.

McKean:1990:RDR

[MSH90] Joseph W. McKean, Simon J. Sheather, and Thomas P. Hettmansperger. Regression diagnostics for rank-based methods. Journal of the American Statistical Association, 85(412): 1018–1028, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289598.

McKean:1993:UIR

[MSH93] Joseph W. McKean, Simon J. Sheather, and Thomas P. Hettmansperger. The use and interpretation of residuals based on robust estimation. Journal of the American Statistical Association, 88(424):1254–1263, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291265.

Mudholkar:1996:GWD

[MSK96] Govind S. Mudholkar, Deo Kumar Srivastava, and Georgia D. Kollia. A generalization of the Weibull distribution with application to the analysis of survival data. Journal of the American Statistical Association, 91(436):1575–1583, December 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291583.

Murry:1993:EAD

[MSL93] John P. Murry, Antonie Stam, and John L. Lastovicka. Evaluating an anti-drinking and driving advertising campaign with a sample survey and time series intervention analysis. Journal of the American Statistical Association, 88(421):50–56, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290691.

Malec:1997:SAI

[MSML97] Donald Malec, J. Sedransk, Christopher L. Moriarity, and Felicia B. LeClere. Small area inference for binary variables in the National Health Interview Survey. Journal of the American Statistical Association, 92(439):815–826, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_97/
sept/malec.htm;   http://www.jstor.org/stable/2965546.

Manski:1992:AEE

[MSMP92] Charles F. Manski, Gary D. Sandefur, Sara Mclanahan, and Daniel Powers. Alternative estimates of the effect of family structure during adolescence on high school graduation. Journal of the American Statistical Association, 87(417):25–37, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290448.

Muller:1997:SSG

[MST97] Hans-Georg Müller, Ulrich Stadtmüller, and Farzaneh Tabnak. Spatial smoothing of geographically aggregated data, with application to the construction of incidence maps. Journal of the American Statistical Association, 92(437):61–71, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291450.

Mathew:1993:SSP

[MSZ93] Thomas Mathew, Bimal Kumar Sinha, and Leping Zhou. Some statistical procedures for combining independent tests. Journal of the American Statistical Association, 88(423):912–919, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290781.

McCulloch:1993:BIP

[MT93] Robert E. McCulloch and Ruey S. Tsay. Bayesian inference and prediction for mean and variance shifts in autoregressive time series. Journal of the American Statistical Association, 88(423): 968–978, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290788.

Marron:1995:VEC

[MT95a] J. S. Marron and A. B. Tsybakov. Visual error criteria for qualitative smoothing. Journal of the American Statistical Association, 90(430):499–507, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291060.

Mukerjee:1995:ORI

[MT95b] Hari Mukerjee and Renjin Tu. Order-restricted inferences in linear regression. Journal of the American Statistical Association, 90 (430):717–728, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291084.

Meer:1998:CEP

[MT98a] Peter Meer and David Tyler. Comment: Edge preserving smoothers for image processing. Journal of the American Statistical Association, 93(442):548–??, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Meer:1998:EPS

[MT98b] Peter Meer and David Tyler. Edge-preserving smoothers for image processing: Comment. Journal of the American Statistical Association, 93(442):548–553, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2670103.

MTW:1996:BRBa

[MTW96a] MTW. Book review: [untitled]. Journal of the American Statistical Association, 91(435):1380, September 1996. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2291765.

MTW:1996:BRBb

[MTW96b] MTW. Book review: [untitled]. Journal of the American Statistical Association, 91(436):1754, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291615.

MTW:1996:BRBc

[MTW96c] MTW. Book review: [untitled]. Journal of the American Statistical Association, 91(436):1754, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291617.

MTW:1996:BRBd

[MTW96d] MTW. Book review: [untitled]. Journal of the American Statistical Association, 91(436):1754–1755, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291619.

MTW:1997:BRBa

[MTW97a] MTW. Book review: [untitled]. Journal of the American Statistical Association, 92(438):799, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965746.

MTW:1997:BRBb

[MTW97b] MTW. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1655, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965458.

MTW:1997:BRBc

[MTW97c] MTW. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1656, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965461.

MTW:1998:BRBa

[MTW98a] MTW. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1251, September 1998. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2669900.

MTW:1998:BRBb

[MTW98b] MTW. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1251, September 1998. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2669901.

MTW:1998:BRBc

[MTW98c] MTW. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1251, September 1998. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2669903.

MTW:1998:BRBd

[MTW98d] MTW. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1251, September 1998. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2669904.

MTW:1998:BRBe

[MTW98e] MTW. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1529, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670083.

MTW:1998:BRBf

[MTW98f] MTW. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1529–1530, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670084.

MTW:1999:BRBa

[MTW99a] MTW. Book review: [untitled]. Journal of the American Statistical Association, 94(445):350–351, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669729.

MTW:1999:BRBb

[MTW99b] MTW. Book review: [untitled]. Journal of the American Statistical Association, 94(445):351, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669731.

MTW:1999:BRBc

[MTW99c] MTW. Book review: [untitled]. Journal of the American Statistical Association, 94(445):351, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669732.

MTW:1999:BRBd

[MTW99d] MTW. Book review: [untitled]. Journal of the American Statistical Association, 94(445):351, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669733.

MTW:1999:BRBe

[MTW99e] MTW. Book review: [untitled]. Journal of the American Statistical Association, 94(446):656, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670200.

MTW:1999:BRBf

[MTW99f] MTW. Book review: [untitled]. Journal of the American Statistical Association, 94(446):658, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670210.

MTW:1999:BRBg

[MTW99g] MTW. Book review: [untitled]. Journal of the American Statistical Association, 94(446):658, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670211.

MTW:1999:BRBh

[MTW99h] MTW. Book review: [untitled]. Journal of the American Statistical Association, 94(447):990, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670033.

MTW:1999:BRBi

[MTW99i] MTW. Book review: [untitled]. Journal of the American Statistical Association, 94(447):991, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670034.

MTW:1999:BRBj

[MTW99j] MTW. Book review: [untitled]. Journal of the American Statistical Association, 94(447):991, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670035.

MTW:1999:BRBk

[MTW99k] MTW. Book review: [untitled]. Journal of the American Statistical Association, 94(447):991, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670036.

MTW:1999:BRBl

[MTW99l] MTW. Book review: [untitled]. Journal of the American Statistical Association, 94(448):1389–1390, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669965.

MTW:1999:BRBm

[MTW99m] MTW. Book review: [untitled]. Journal of the American Statistical Association, 94(448):1390, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669966.

MTW:1999:BRBn

[MTW99n] MTW. Book review: [untitled]. Journal of the American Statistical Association, 94(448):1390, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669967.

Mykland:1995:RMC

[MTY95] Per Mykland, Luke Tierney, and Bin Yu. Regeneration in Markov chain samplers. Journal of the American Statistical Association, 90(429):233–241, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291148.

Mukerjee:1990:BRB

[Muk90] Hari Mukerjee. Book review: [untitled]. Journal of the American Statistical Association, 85(409):260–261, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289565.

Mukerjee:1992:BRB

[Muk92] Hari Mukerjee. Book review: [untitled]. Journal of the American Statistical Association, 87(417):254, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290484.

Mukerjee:1996:ESF

[Muk96] Hari Mukerjee. Estimation of survival functions under uniform stochastic ordering. Journal of the American Statistical Association, 91(436):1684–1689, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291596.

Muller:1991:BRB

[Mul91] Hans-Georg Muller. Book review: [untitled]. Journal of the American Statistical Association, 86(415):834, September 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290434.

Muller:1993:BRB

[Mul93] Keith E. Muller. Book review: [untitled]. Journal of the American Statistical Association, 88(421):379–380, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290741.

Muller:1998:BRB

[Mul98] Hans-Georg Muller. Book review: [untitled]. Journal of the American Statistical Association, 93(442):835–836, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670134.

Muller:1999:VSF

[Mul99] Peter Muller. Variable selection and function estimation in additive nonparametric regression using a data-based prior: Comment. Journal of the American Statistical Association, 94(447): 803, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2669994.

Munk:1996:EIT

[Mun96] Axel Munk. Equivalence and interval testing for Lehmann’s alternative. Journal of the American Statistical Association, 91(435): 1187–1196, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291737.

Murray:1990:BRB

[Mur90] Leigh Murray. Book review: [untitled]. Journal of the American Statistical Association, 85(412):1170–1171, December 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289624.

Murtagh:1991:BRB

[Mur91] Fionn Murtagh. Book review: [untitled]. Journal of the American Statistical Association, 86(413):252, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289752.

Murphy:1995:LRB

[Mur95] S. A. Murphy. Likelihood ratio-based confidence intervals in survival analysis. Journal of the American Statistical Association, 90 (432):1399–1405, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291531.

Murphy:1997:BRB

[Mur97] James R. Murphy. Book review: [untitled]. Journal of the American Statistical Association, 92(438):797–798, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965741.

Meeden:1991:NBA

[MV91] Glen Meeden and Stephen Vardeman. A noninformative Bayesian approach to interval estimation in finite population sampling. Journal of the American Statistical Association, 86(416):972–980, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290513.

MW:1995:BRBa

[MW95a] MW. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1135, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291367.

MW:1995:BRBb

[MW95b] MW. Book review: [untitled]. Journal of the American Statistical Association, 90(432):1493, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291558.

MW:1995:BRBc

[MW95c] MW. Book review: [untitled]. Journal of the American Statistical Association, 90(432):1493–1494, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291559.

MW:1996:BRBa

[MW96a] MW. Book review: [untitled]. Journal of the American Statistical Association, 91(433):441, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291438.

MW:1996:BRBb

[MW96b] MW. Book review: [untitled]. Journal of the American Statistical Association, 91(433):441, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291441.

MW:1996:BRBc

[MW96c] MW. Book review: [untitled]. Journal of the American Statistical Association, 91(434):915, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291700.

MW:1996:BRBd

[MW96d] MW. Book review: [untitled]. Journal of the American Statistical Association, 91(434):916, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291705.

MW:1997:BRBa

[MW97a] MW. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1226–1227, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965611.

MW:1997:BRBb

[MW97b] MW. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1227, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965612.

MW:1997:BRBc

[MW97c] MW. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1227, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965616.

MW:1997:BRBd

[MW97d] MW. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1227, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965617.

MW:1997:BRBe

[MW97e] MW. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1227, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965618.

Marais:1998:COVa

[MW98a] M. Laurentius Marais and William E. Wecker. Correcting for omitted-variables and measurement-error bias in regression with an application to the effect of lead on IQ. Journal of the American Statistical Association, 93(442):494–504, June 1998.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/june/MARAIS.HTM;   http://www.jstor.org/
stable/2670095.

Marais:1998:COVb

[MW98b] M. Laurentius Marais and William E. Wecker. Correcting for omitted-variables and measurement-error bias in regression with an application to the effect of lead on IQ: Rejoinder. Journal of the American Statistical Association, 93(442):515–517, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670098.

Marais:1998:RCO

[MW98c] M. Laurentius Marais and William E. Wecker. Rejoinder: Correcting for omitted-variables and measurement-error bias in regression with an application to the effect of lead on IQ. Journal of the American Statistical Association, 93(442):515–??, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

MW:1998:BRBa

[MW98d] MW. Book review: [untitled]. Journal of the American Statistical Association, 93(441):412, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669666.

MW:1998:BRBb

[MW98e] MW. Book review: [untitled]. Journal of the American Statistical Association, 93(441):412, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669667.

Maronna:1991:BPS

[MY91] Ricardo A. Maronna and Víctor J. Yohai. The breakdown point of simultaneous general M estimates of regression and scale. Journal of the American Statistical Association, 86(415):699–703, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290400.

Maronna:1995:BSD

[MY95] Ricardo A. Maronna and Víctor J. Yohai. The behavior of the Stahel-Donoho robust multivariate estimator. Journal of the American Statistical Association, 90(429):330–341, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291158.

Muller:1991:TDE

[MZ91] Hans-Georg Muller and Hong Zhou. Transformations in density estimation: Comment. Journal of the American Statistical Association, 86(414):356–358, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2290573.

Maller:1994:TSF

[MZ94] R. A. Maller and S. Zhou. Testing for sufficient follow-up and outliers in survival data. Journal of the American Statistical Association, 89(428):1499–1506, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291012.

Magder:1996:SNE

[MZ96] Laurence S. Magder and Scott L. Zeger. A smooth nonparametric estimate of a mixing distribution using mixtures of Gaussians. Journal of the American Statistical Association, 91(435): 1141–1151, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291733.

Mathew:1997:MUC

[MZ97] Thomas Mathew and Wenxing Zha. Multiple use confidence regions in multivariate calibration. Journal of the American Statistical Association, 92(439):1141–1150, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/sept/mathew.htm;   http://www.jstor.org/
stable/2965579.

Montgomery:1998:FUU

[MZTT98] Alan L. Montgomery, Victor Zarnowitz, Ruey S. Tsay, and George C. Tiao. Forecasting the U.S. unemployment rate. Journal of the American Statistical Association, 93(442):478–493, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_98/june/MONTGOME.HTM;   http://www.jstor.
org/stable/2670094.

Navidi:1999:CDP

[NA99] William Navidi and Norman Arnheim. Combining data from polymerase chain reaction DNA typing experiments: Applications to sperm typing data. Journal of the American Statistical Association, 94(447):726–733, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669985.

Nachtsheim:1991:BRB

[Nac91] Christopher J. Nachtsheim. Book review: [untitled]. Journal of the American Statistical Association, 86(414):553, June 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290621.

Naiman:1991:PIA

[Nai91] Daniel Q. Naiman. On the problem of interactions in the analysis of variance: Comment. Journal of the American Statistical Association, 86(414):369–372, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290578.

Nakamura:1992:BRB

[Nak92] Alice Nakamura. Book review: [untitled]. Journal of the American Statistical Association, 87(420):1240–1241, December 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290668.

Naranjo:1993:BRB

[Nar93] Joshua D. Naranjo. Book review: [untitled]. Journal of the American Statistical Association, 88(424):1468–1469, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291297.

Ng:1999:AST

[NC99] Edmund T. M. Ng and Richard J. Cook. Adjusted score tests of homogeneity for Poisson processes. Journal of the American Statistical Association, 94(445):308–319, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669705.

Newton:1996:BIS

[NCC96] Michael A. Newton, Claudia Czado, and Rick Chappell. Bayesian inference for semiparametric binary regression. Journal of the American Statistical Association, 91(433):142–153, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291390.

Nusser:1996:STA

[NCDF96] S. M. Nusser, A. L. Carriquiry, K. W. Dodd, and W. A. Fuller. A semiparametric transformation approach to estimating usual daily intake distributions. Journal of the American Statistical Association, 91(436):1440–1449, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291570.

Neill:1997:BRB

[Nei97] James W. Neill. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1652–1653, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965450.

Nelson:1992:BRB

[Nel92] Charles R. Nelson. Book review: [untitled]. Journal of the American Statistical Association, 87(419):905, September 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290243.

Nelson:1997:BRB

[Nel97] Paul I. Nelson. Book review: [untitled]. Journal of the American Statistical Association, 92(438):795–796, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965737.

Nelson:1998:BRBa

[Nel98a] Paul I. Nelson. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1233–1234, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669867.

Nelson:1998:BRBb

[Nel98b] Paul I. Nelson. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1234–1235, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669868.

Neter:1990:REF

[Net90] John Neter. Regression estimates in Federal Welfare Quality Control Programs: Comment. Journal of the American Statistical Association, 85(411):864–865, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290027.

Neuberg:1992:BRB

[Neu92] Leland Gerson Neuberg. Book review: [untitled]. Journal of the American Statistical Association, 87(420):1246–1247, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290677.

Neuhaus:1998:EEO

[Neu98] John M. Neuhaus. Estimation efficiency with omitted covariates in generalized linear models. Journal of the American Statistical Association, 93(443):1124–1129, September 1998. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_98/sept/
neuhaus.html;   http://www.jstor.org/stable/2669855.

Neuhaus:1999:ECE

[Neu99] John M. Neuhaus. Estimation of the causal effect of a timevarying exposure on the marginal mean of a repeated binary outcome: Comment. Journal of the American Statistical Association, 94(447):701–702, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669979.

Newton:1990:BRB

[New90] H. Joseph Newton. Book review: [untitled]. Journal of the American Statistical Association, 85(412):1178–1179, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289635.

Newbold:1992:BRB

[New92] Paul Newbold. Book review: [untitled]. Journal of the American Statistical Association, 87(420):1243–1244, December 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290672.

Newcombe:1992:UNLa

[NFL92a] Howard B. Newcombe, Martha E. Fair, and Pierre Lalonde. The use of names for linking personal records. Journal of the American Statistical Association, 87(420):1193–1204, December 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290660.

Newcombe:1992:UNLb

[NFL92b] Howard B. Newcombe, Martha E. Fair, and Pierre Lalonde. The use of names for linking personal records: Rejoinder. Journal of the American Statistical Association, 87(420):1207–1208, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290662.

Newton:1994:BRM

[NG94] Michael A. Newton and Charles J. Geyer. Bootstrap recycling: a Monte Carlo alternative to the nested bootstrap. Journal of the American Statistical Association, 89(427):905–912, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290915.

Newton:1995:SME

[NGC^{+}95] Michael A. Newton, Peter Guttorp, Sandra Catlin, Renato Assuncao, and Janis L. Abkowitz. Stochastic modeling of early hematopoiesis. Journal of the American Statistical Association, 90(432):1146–1155, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2291507.

Normand:1997:SMP

[NGG97] Sharon-Lise T. Normand, Mark E. Glickman, and Constantine A. Gatsonis. Statistical methods for profiling providers of medical care: Issues and applications. Journal of the American Statistical Association, 92(439):803–814, September 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2965545.

Nychka:1995:NRA

[NGH^{+}95] Doug Nychka, Gerry Gray, Perry Haaland, David Martin, and Michael O’Connell. A nonparametric regression approach to syringe grading for quality improvement. Journal of the American Statistical Association, 90(432):1171–1178, December 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291509.

Niu:1996:NAM

[Niu96] Xu-Feng Niu. Nonlinear additive models for environmental time series, with applications to ground-level ozone data analysis. Journal of the American Statistical Association, 91(435):1310–1321, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291749.

NL:1994:BRBa

[NL94a] NL. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1149, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290961.

NL:1994:BRBb

[NL94b] NL. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1149, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290962.

NL:1994:BRBc

[NL94c] NL. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1149, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290963.

NL:1995:BRB

[NL95] NL. Book review: [untitled]. Journal of the American Statistical Association, 90(430):810, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291112.

Neerchal:1998:LCR

[NM98] Nagaraj K. Neerchal and Jorge G. Morel. Large cluster results for TWO parametric multinomial extra variation models. Journal of the American Statistical Association, 93(443): 1078–1087, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/sept/neerchal.html;
 http://www.jstor.org/stable/2669851.

Naik-Nimbalkar:1995:FSE

[NNR95] U. V. Naik-Nimbalkar and M. B. Rajarshi. Filtering and smoothing via estimating functions. Journal of the American Statistical Association, 90(429):301–306, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291154.

Nolan:1992:BRB

[Nol92] Deborah Nolan. Book review: [untitled]. Journal of the American Statistical Association, 87(417):251, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290480.

Norwood:1990:SPP

[Nor90] Janet L. Norwood. Statistics and public policy: Reflections of a changing world. Journal of the American Statistical Association, 85(409):1–5, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289519.

Notz:1990:BRB

[Not90] William I. Notz. Book review: [untitled]. Journal of the American Statistical Association, 85(412):1179–1180, December 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289637.

Notz:1994:BRB

[Not94] William I. Notz. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1139–1140, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290947.

Nowak:1992:IMD

[Now92] Eugen Nowak. Identifiability in multivariate dynamic linear errors-in- variables models. Journal of the American Statistical Association, 87(419):714–723, September 1992. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2290209.

Ng:1995:URT

[NP95] Serena Ng and Pierre Perron. Unit root tests in ARMA models with data-dependent methods for the selection of the truncation lag. Journal of the American Statistical Association, 90(429):268– 281, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291151.

Nandram:1993:EBE

[NS93] B. Nandram and J. Sedransk. Empirical Bayes estimation for the finite population mean on the current occasion. Journal of the American Statistical Association, 88(423):994–1000, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290791.

Nandram:1997:ORB

[NSS97] Balgobin Nandram, J. Sedransk, and Stephen J. Smith. Orderrestricted Bayesian estimation of the age composition of a population of Atlantic cod. Journal of the American Statistical Association, 92(437):33–40, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/march/
nandram.htm;   http://www.jstor.org/stable/2291447.

Normand:1992:PUB

[NT92] Sharon-Lise Normand and David Tritchler. Parameter updating in a Bayes network. Journal of the American Statistical Association, 87(420):1109–1115, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290648.

Niu:1995:MSO

[NT95] Xu Feng Niu and George C. Tiao. Modeling satellite ozone data. Journal of the American Statistical Association, 90(431):969–983, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291333.

Naus:1997:DSS

[NW97] Joseph I. Naus and Dan Wartenberg. A double-scan statistic for clusters of two types of events. Journal of the American Statistical Association, 92(439):1105–1113, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_97/
sept/naus.htm;   http://www.jstor.org/stable/2965575.

Nychka:1991:CRA

[Nyc91] Douglas Nychka. Choosing a range for the amount of smoothing in nonparametric regression. Journal of the American Statistical Association, 86(415):653–664, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290394.

Oh:1993:IMF

[OB93] Man-Suk Oh and James O. Berger. Integration of multimodal functions by Monte Carlo importance sampling. Journal of the American Statistical Association, 88(422):450–456, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290324.

ONeill:1992:SIA

[OBW92] Robert T. O’Neill, Paul Beninger, and Randolf Wykoff. Statistical issues arising in AIDS clinical trials: Comment. Journal of the American Statistical Association, 87(418):579–581, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290296.

OConnell:1999:BRB

[O’C99] Ann A. O’Connell. Book review: Modern Multidimensional Scaling: Theory and Applications, by Ingwer Borg; Patrick Groenen. Journal of the American Statistical Association, 94(445): 338–339, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2669710.

Oden:1991:AEM

[Ode91] Neal L. Oden. Allocation of effort in Monte Carlo simulation for power of permutation tests. Journal of the American Statistical Association, 86(416):1074–1076, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290526.

Oehlert:1993:BRB

[Oeh93a] Gary W. Oehlert. Book review: [untitled]. Journal of the American Statistical Association, 88(422):710, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290375.

Oehlert:1993:RTS

[Oeh93b] Gary W. Oehlert. Regional trends in sulfate wet deposition. Journal of the American Statistical Association, 88(422):390–399, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290317.

Ogunnaike:1999:BRB

[Ogu99] Babatunde A. Ogunnaike. Book review: [untitled]. Journal of the American Statistical Association, 94(446):651–652, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670190.

OHagan:1990:OCL

[O’H90] A. O’Hagan. Outliers and credence for location parameter inference. Journal of the American Statistical Association, 85(409): 172–176, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289540.

Olive:1999:DRD

[OH99a] David Olive and Douglas M. Hawkins. Discussion: Regression depth. Journal of the American Statistical Association, 94(446): 416–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Olive:1999:RDC

[OH99b] David Olive and Douglas M. Hawkins. Regression depth: Comment. Journal of the American Statistical Association, 94(446): 416–417, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2670161.

OlusegunGeorge:1996:TIT

[OK96] E. Olusegun George and Ralph L. Kodell. Tests of independence, treatment heterogeneity, and dose-related trend with exchangeable binary data. Journal of the American Statistical Association, 91(436):1602–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Oliveira:1997:BPT

[OKS97a] Victor De Oliveira, Benjamin Kedem, and David A. Short. Bayesian prediction of transformed Gaussian random fields. Journal of the American Statistical Association, 92(440):1422–1433, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2965412.

Ord:1997:EPC

[OKS97b] J. K. Ord, A. B. Koehler, and R. D. Snyder. Estimation and prediction for a class of dynamic nonlinear statistical models. Journal of the American Statistical Association, 92(440): 1621–1629, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/december/ORD.HTM;
http://www.jstor.org/stable/2965433.

Olkin:1991:BRB

[Olk91] Ingram Olkin. Book review: [untitled]. Journal of the American Statistical Association, 86(416):1143–1144, December 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290543.

Olkin:1992:BRBa

[Olk92a] Ingram Olkin. Book review: [untitled]. Journal of the American Statistical Association, 87(420):1244–1245, December 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290674.

Olkin:1992:BRBb

[Olk92b] Ingram Olkin. Book review: [untitled]. Journal of the American Statistical Association, 87(420):1245, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290675.

Olkin:1993:BRB

[Olk93] Ingram Olkin. Book review: [untitled]. Journal of the American Statistical Association, 88(422):702–703, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290364.

Olkin:1997:BRB

[Olk97] Ingram Olkin. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1654, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965454.

Olkin:1999:BRB

[Olk99] Ingram Olkin. Book review: [untitled]. Journal of the American Statistical Association, 94(446):652, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670191.

Olsen:1990:BRB

[Ols90] Randall J. Olsen. Book review: [untitled]. Journal of the American Statistical Association, 85(409):269–270, March 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289580.

Oakes:1993:UCM

[OMF^{+}93] D. Oakes, A. J. Moss, J. L. Fleiss, J. T. Bigger, Jr., T. Therneau, S. W. Eberly, M. P. McDermott, A. Manatunga, E. Carleen, and J. Benhorin. Use of compliance measures in an analysis of the effect of diltiazem on mortality and reinfarction after myocardial infarction. Journal of the American Statistical Association, 88 (421):44–49, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290690.

Omori:1995:BRB

[Omo95] Yasuhiro Omori. Book review: [untitled]. Journal of the American Statistical Association, 90(432):1490–1491, December 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291547.

ONeill:1994:BEE

[O’N94] Terence J. O’Neill. The bias of estimating equations with application to the error RATE of logistic discrimination. Journal of the American Statistical Association, 89(428):1492–1498, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291011.

Olshen:1997:CDA

[OO97] Adam B. Olshen and Finbarr O’Sullivan. Camouflaged deconvolution with application to blood curve modeling in FDG PET studies. Journal of the American Statistical Association, 92(440): 1293–1303, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/OLSHEN.HTM;
http://www.jstor.org/stable/2965399.

OSullivan:1996:BSI

[OP96] Finbarr O’Sullivan and Yudi Pawitan. Bandwidth selection for indirect density estimation based on corrupted histogram data. Journal of the American Statistical Association, 91(434):610–626, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291657.

ORourke:1991:BRB

[O’R91] Diane O’Rourke. Book review: [untitled]. Journal of the American Statistical Association, 86(413):252, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289751.

Osius:1992:NGF

[OR92] Gerhard Osius and Dieter Rojek. Normal goodness-of-fit tests for multinomial models with large degrees of freedom. Journal of the American Statistical Association, 87(420):1145–1152, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290653.

Olsson:1996:QER

[OR96] J. Olsson and H. Rootzén. Quantile estimation from repeated measurements. Journal of the American Statistical Association, 91(436):1560–1565, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2291581.

Opsomer:1998:FAB

[OR98] Jean D. Opsomer and David Ruppert. A fully automated bandwidth selection method for fitting additive models. Journal of the American Statistical Association, 93(442):605–619, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/

abstracts_98/june/OPSOMER.HTM;   http://www.jstor.org/
stable/2670112.

OSullivan:1990:IAT

[O’S90] Finbarr O’Sullivan. An iterative approach to two-dimensional Laplacian smoothing with application to image restoration. Journal of the American Statistical Association, 85(409):213–219, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289547.

OSullivan:1991:DLS

[O’S91] Finbarr O’Sullivan. Discretized Laplacian smoothing by Fourier methods. Journal of the American Statistical Association, 86 (415):634–642, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290392.

OSullivan:1993:MEM

[O’S93] Finbarr O’Sullivan. Mixture estimation from multichannel image data. Journal of the American Statistical Association, 88(421): 209–220, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290715.

Olkin:1995:CAE

[OV95] Ingram Olkin and Marlos Viana. Correlation analysis of extreme observations from a multivariate normal distribution. Journal of the American Statistical Association, 90(432):1373–1379, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291528.

Overton:1991:BRB

[Ove91] W. Scott Overton. Book review: [untitled]. Journal of the American Statistical Association, 86(416):1149–1150, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290549.

Overton:1994:BRB

[Ove94] W. Scott Overton. Book review: [untitled]. Journal of the American Statistical Association, 89(426):717–718, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290885.

Owen:1992:BRB

[Owe92] William B. Owen. Book review: [untitled]. Journal of the American Statistical Association, 87(419):897, September 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290232.

Owen:1994:CCL

[Owe94] Art B. Owen. Controlling correlations in Latin hypercube samples. Journal of the American Statistical Association, 89(428): 1517–1522, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291014.

Owen:1995:NLC

[Owe95] Art B. Owen. Nonparametric likelihood confidence bands for a distribution function. Journal of the American Statistical Association, 90(430):516–521, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291062.

Padgett:1993:BRB

[Pad93] W. J. Padgett. Book review: [untitled]. Journal of the American Statistical Association, 88(421):385–386, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290749.

Page:1990:APC

[Pag90] Connie Faye Page. Allocation proportional to coefficients of variation when estimating the product of parameters. Journal of the American Statistical Association, 85(412):1134–1139, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289612.

Paik:1997:GEE

[Pai97] Myunghee Cho Paik. The generalized estimating equation approach when data are not missing completely at random. Journal of the American Statistical Association, 92(440):1320– 1329, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/december/PAIK.HTM;
 http://www.jstor.org/stable/2965402.

PALE:1996:BRBa

[PAL96a] PALE. Book review: [untitled]. Journal of the American Statistical Association, 91(436):1754, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291616.

PALE:1996:BRBb

[PAL96b] PALE. Book review: [untitled]. Journal of the American Statistical Association, 91(436):1754, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291618.

Pan:1997:CSC

[Pan97] Guohua Pan. Confidence subset containing the unknown peaks of an umbrella ordering. Journal of the American Statistical Association, 92(437):307–314, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/march/
pan.htm;   http://www.jstor.org/stable/2291475.

Paparoditis:1996:FDB

[Pap96] Efstathios Paparoditis. A frequency domain bootstrap-based method for checking the fit of a transfer function model. Journal of the American Statistical Association, 91(436):1535–1550, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291579.

Parmigiani:1991:BRB

[Par91] Giovanni Parmigiani. Book review: [untitled]. Journal of the American Statistical Association, 86(415):825–827, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290425.

Park:1993:CVC

[Par93a] Byeong U. Park. A cross-validatory choice of smoothing parameter in adaptive location estimation. Journal of the American Statistical Association, 88(423):848–854, September 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290773.

Parmigiani:1993:OSA

[Par93b] Giovanni Parmigiani. On optimal screening ages. Journal of the American Statistical Association, 88(422):622–628, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290344.

Park:1996:FIO

[Par96] Sangun Park. Fisher information in order statistics. Journal of the American Statistical Association, 91(433):385–390, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291418.

Passel:1993:EPC

[Pas93] Jeffrey S. Passel. Estimation of population coverage in the 1990 United States census based on demographic analysis: Comment. Journal of the American Statistical Association, 88(423):1074– 1077, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290801.

Paulauskas:1996:BRB

[Pau96] V. Paulauskas. Book review: [untitled]. Journal of the American Statistical Association, 91(434):915, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291698.

Picard:1992:CSV

[PB92] Richard Picard and Maurice Bryson. Calibrated seismic verification of the threshold test ban treaty. Journal of the American Statistical Association, 87(418):293–299, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290259.

Park:1994:MCD

[PB94] Taesung Park and Morton B. Brown. Models for categorical data with nonignorable nonresponse. Journal of the American Statistical Association, 89(425):44–52, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291199.

Paul:1998:ATW

[PB98] S. R. Paul and T. Banerjee. Analysis of two-way layout of count data involving multiple counts in each cell. Journal of the American Statistical Association, 93(444):1419– 1429, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/dec/paul.html;   http:/
/www.jstor.org/stable/2670056.

Pepe:1993:SGD

[PC93] Margaret Sullivan Pepe and Jianwen Cai. Some graphical displays and marginal regression analyses for recurrent failure times and time dependent covariates. Journal of the American Statistical Association, 88(423):811–820, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290770.

Peddada:1996:BCR

[PC96] Shyamal Das Peddada and Ted Chang. Bootstrap confidence region estimation of the motion of rigid bodies. Journal of the American Statistical Association, 91(433):231–241, March 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291400.

Pepe:1997:MPC

[PC97] Margaret Sullivan Pepe and David Couper. Modeling partly conditional means with longitudinal data. Journal of the American Statistical Association, 92(439):991–998, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/sept/sullivan.htm;   http://www.jstor.org/
stable/2965563.

PD:1999:BRB

[PD99] PD. Book review: [untitled]. Journal of the American Statistical Association, 94(447):989, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670027.

PE:1994:BRBa

[PE94a] PE. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1150, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290965.

PE:1994:BRBb

[PE94b] PE. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1150, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290966.

PE:1994:BRBc

[PE94c] PE. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1150, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290967.

PE:1994:BRBd

[PE94d] PE. Book review: [untitled]. Journal of the American Statistical Association, 89(428):1570, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291040.

PE:1995:BRBa

[PE95a] PE. Book review: [untitled]. Journal of the American Statistical Association, 90(429):398, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291190.

PE:1995:BRBb

[PE95b] PE. Book review: [untitled]. Journal of the American Statistical Association, 90(430):810, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291113.

PE:1995:BRBc

[PE95c] PE. Book review: [untitled]. Journal of the American Statistical Association, 90(430):810, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291116.

PE:1995:BRBd

[PE95d] PE. Book review: [untitled]. Journal of the American Statistical Association, 90(430):810, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291117.

PE:1995:BRBe

[PE95e] PE. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1136, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291373.

PE:1995:BRBf

[PE95f] PE. Book review: [untitled]. Journal of the American Statistical Association, 90(432):1493, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291553.

PE:1995:BRBg

[PE95g] PE. Book review: [untitled]. Journal of the American Statistical Association, 90(432):1493, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291554.

PE:1995:BRBh

[PE95h] PE. Book review: [untitled]. Journal of the American Statistical Association, 90(432):1493, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291555.

PE:1996:BRBa

[PE96a] PE. Book review: [untitled]. Journal of the American Statistical Association, 91(433):441, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291440.

PE:1996:BRBb

[PE96b] PE. Book review: [untitled]. Journal of the American Statistical Association, 91(434):916, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291704.

PE:1997:BRBa

[PE97a] PE. Book review: [untitled]. Journal of the American Statistical Association, 92(437):389, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291498.

PE:1997:BRBb

[PE97b] PE. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1226, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965610.

PE:1998:BRBa

[PE98a] PE. Book review: [untitled]. Journal of the American Statistical Association, 93(441):411, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669659.

PE:1998:BRBb

[PE98b] PE. Book review: [untitled]. Journal of the American Statistical Association, 93(441):411, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669660.

PE:1998:BRBc

[PE98c] PE. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1251, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669897.

PE:1998:BRBd

[PE98d] PE. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1529, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670082.

PE:1999:BRB

[PE99] PE. Book review: [untitled]. Journal of the American Statistical Association, 94(446):658, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670209.

Pearson:1991:SSA

[Pea91] Robert W. Pearson. Social statistics and an American urban underclass: Improving the knowledge base for social policy in the 1990s. Journal of the American Statistical Association, 86(414): 504–512, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290602.

Peddada:1993:PCR

[Ped93] Shyamal Das Peddada. Is Pitman closeness a reasonable criterion?: Comment. Journal of the American Statistical Association, 88(421):67–69, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290694.

Peddada:1994:BRB

[Ped94] Shyamal D. Peddada. Book review: [untitled]. Journal of the American Statistical Association, 89(426):711, June 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290877.

Pennello:1997:RMC

[Pen97] Gene Pennello. The k-ratio multiple comparisons Bayes rule for the balanced two-way design. Journal of the American Statistical Association, 92(438):675–684, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965715.

Pena:1998:SGF

[Peñ98] Edsel A. Peña. Smooth goodness-of-fit tests for the baseline hazard in Cox’s proportional hazards model. Journal of the American Statistical Association, 93(442):673–692, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
june/PENA.HTM;   http://www.jstor.org/stable/2670118.

Pepe:1991:IED

[Pep91] Margaret Sullivan Pepe. Inference for events with dependent risks in multiple endpoint studies. Journal of the American Statistical Association, 86(415):770–778, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290411.

Perez:1990:BRB

[Per90] Lisandro Perez. Book review: [untitled]. Journal of the American Statistical Association, 85(411):902–903, September 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290037.

Percival:1996:BRB

[Per96] Donald B. Percival. Book review: [untitled]. Journal of the American Statistical Association, 91(435):1378–1379, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291761.

Peruggia:1997:VCD

[Per97] Mario Peruggia. On the variability of case-deletion importance sampling weights in the Bayesian linear model. Journal of the American Statistical Association, 92(437):199– 207, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/march/peruggia.htm;
 http://www.jstor.org/stable/2291464.

Peskun:1993:NCI

[Pes93] Peter H. Peskun. A new confidence interval method based on the normal approximation for the difference of two binomial probabilities. Journal of the American Statistical Association, 88(422): 656–661, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290348.

Pettit:1992:BFO

[Pet92] Lawrence I. Pettit. Bayes factors for outlier models using the device of imaginary observations. Journal of the American Statistical Association, 87(418):541–545, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290288.

Pepe:1991:NMD

[PF91] Margaret Sullivan Pepe and Thomas R. Fleming. A nonparametric method for dealing with mismeasured covariate data. Journal of the American Statistical Association, 86(413):108–113, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289720.

Proschan:1997:RTC

[PF97] Michael A. Proschan and Dean A. Follmann. A restricted test of circadian rhythm. Journal of the American Statistical Association, 92(438):717–724, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2965720.

Pfeifer:1999:BRB

[Pfe99] Dietmar Pfeifer. Book review: [untitled]. Journal of the American Statistical Association, 94(447):983, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670016.

Porter-Hudak:1990:ASF

[PH90a] Susan Porter-Hudak. An application of the seasonal fractionally differenced model to the monetary aggregates. Journal of the American Statistical Association, 85(410):338–344, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289769.

Puma:1990:FSP

[PH90b] Michael J. Puma and David C. Hoaglin. Food stamp payment error rates: Can state-specific performance standards be developed? Journal of the American Statistical Association, 85(411): 891–899, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290033.

Pulkstenis:1998:MAB

[PHL98] Erik P. Pulkstenis, Thomas R. Ten Have, and J. Richard Landis. Model for the analysis of binary longitudinal pain data subject to informative dropout through remedication. Journal of the American Statistical Association, 93(442):438–450, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
june/PULKSTEN.HTM;   http://www.jstor.org/stable/2670091.

Presnell:1994:TMR

[PHS94] Brett Presnell, Myles Hollander, and Jayaram Sethuraman. Testing the minimal repair assumption in an imperfect repair model. Journal of the American Statistical Association, 89(425):289– 297, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291225.

Pickle:1997:BRB

[Pic97] Linda Williams Pickle. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1221–1222, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2965599.

Pierce:1993:BRB

[Pie93] David A. Pierce. Book review: [untitled]. Journal of the American Statistical Association, 88(423):1170–1171, September 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290821.

Pinheiro:1997:BRB

[Pin97] Jose C. Pinheiro. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1219–1220, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965596.

Pitts:1999:BRBa

[Pit99a] Susan M. Pitts. Book review: [untitled]. Journal of the American Statistical Association, 94(445):348, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669722.

Pitts:1999:BRBb

[Pit99b] Susan M. Pitts. Book review: [untitled]. Journal of the American Statistical Association, 94(446):655, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670197.

Poli:1994:NNM

[PJ94] I. Poli and R. D. Jones. A neural net model for prediction. Journal of the American Statistical Association, 89(425):117–121, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291206.

Peng:1996:BIM

[PJT96] Fengchun Peng, Robert A. Jacobs, and Martin A. Tanner. Bayesian inference in mixtures-of-experts and hierarchical mixtures-of-experts models with an application to speech recognition. Journal of the American Statistical Association, 91(435): 953–960, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291714.

Plewes:1991:BRB

[Ple91] Thomas J. Plewes. Book review: [untitled]. Journal of the American Statistical Association, 86(414):552, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290619.

Park:1990:CDD

[PM90] Byeong U. Park and J. S. Marron. Comparison of data-driven bandwidth selectors. Journal of the American Statistical Association, 85(409):66–72, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289526.

Percival:1997:ASC

[PM97] Donald B. Percival and Harold O. Mofjeld. Analysis of subtidal coastal sea level fluctuations using wavelets. Journal of the American Statistical Association, 92(439):868–880, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/sept/percival.htm;   http://www.jstor.org/
stable/2965551.

Portnoy:1999:DRD

[PM99a] Stephen Portnoy and Ivan Mizera. Discussion: Regression depth. Journal of the American Statistical Association, 94(446):417–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Portnoy:1999:RDC

[PM99b] Stephen Portnoy and Ivan Mizera. Regression depth: Comment. Journal of the American Statistical Association, 94(446):417–419, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670162.

Presnell:1998:PML

[PML98] Brett Presnell, Scott P. Morrison, and Ramon C. Littell. Projected multivariate linear models for directional data. Journal of the American Statistical Association, 93(443):1068– 1077, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/sept/presnell.html;
 http://www.jstor.org/stable/2669850.

Potthoff:1993:CNB

[PMW93] Richard F. Potthoff, Kenneth G. Manton, and Max A. Woodbury. Correcting for nonavailability bias in surveys by weighting based on number of callbacks. Journal of the American Statistical Association, 88(424):1197–1207, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291258.

Pawitan:1994:NSD

[PO94] Yudi Pawitan and Finbarr O’Sullivan. Nonparametric spectral density estimation using penalized Whittle likelihood. Journal of the American Statistical Association, 89(426):600–610, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290863.

Podgor:1994:BRB

[Pod94] Marvin J. Podgor. Book review: [untitled]. Journal of the American Statistical Association, 89(426):719, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290888.

Priebe:1997:SSS

[POH97] Carey E. Priebe, Tim Olson, and Dennis M. Healy, Jr. A spatial scan statistic for stochastic scan partitions. Journal of the American Statistical Association, 92(440):1476–1484, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/december/PRIEBE.HTM;   http://www.jstor.
org/stable/2965418.

Polson:1990:BRB

[Pol90] Nicholas G. Polson. Book review: [untitled]. Journal of the American Statistical Association, 85(412):1167, December 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289619.

Pollock:1991:MCR

[Pol91] Kenneth H. Pollock. Modeling capture, recapture, and removal statistics for estimation of demographic parameters for fish and wildlife populations: Past, present, and future. Journal of the American Statistical Association, 86(413):225–238, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289733.

Portnoy:1997:CIM

[Por97a] Stephen Portnoy. Comment: Interpolation methods for adapting to sparse design in nonparametric regression. Journal of the American Statistical Association, 92(438):473–??, June 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Portnoy:1997:IMA

[Por97b] Stephen Portnoy. Interpolation methods for adapting to sparse design in nonparametric regression: Comment. Journal of the American Statistical Association, 92(438):473–475, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965695.

Potscher:1996:BRB

[Pot96] Benedikt M. Potscher. Book review: [untitled]. Journal of the American Statistical Association, 91(433):439–440, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291435.

Peters:1990:MSR

[PR90a] Dawn Peters and Ronald H. Randles. A multivariate signed-rank test for the one-sample location problem. Journal of the American Statistical Association, 85(410):552–557, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289797.

Prasad:1990:EMS

[PR90b] N. G. N. Prasad and J. N. K. Rao. The estimation of the mean squared error of small-area estimators. Journal of the American Statistical Association, 85(409):163–171, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289539.

Pukelsheim:1993:EDM

[PR93] Friedrich Pukelsheim and James L. Rosenberger. Experimental designs for model discrimination. Journal of the American Statistical Association, 88(422):642–649, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290346.

Politis:1994:SB

[PR94] Dimitris N. Politis and Joseph P. Romano. The stationary bootstrap. Journal of the American Statistical Association, 89(428):

1303–1313, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290993.

Prewitt:1992:BRB

[Pre92] Kathryn Prewitt. Book review: [untitled]. Journal of the American Statistical Association, 87(418):586, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290301.

Pregibon:1994:BRB

[Pre94] Daryl Pregibon. Book review: [untitled]. Journal of the American Statistical Association, 89(425):360, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291243.

Presnell:1996:BRB

[Pre96a] Brett Presnell. Book review: [untitled]. Journal of the American Statistical Association, 91(433):433–434, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291428.

Press:1996:DNA

[Pre96b] S. James Press. The directional neighborhoods approach to contextual classification of images from noisy data. Journal of the American Statistical Association, 91(435):1091–1100, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291728.

Priebe:1994:AM

[Pri94] Carey E. Priebe. Adaptive mixtures. Journal of the American Statistical Association, 89(427):796–806, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290905.

Priebe:1996:NAU

[Pri96] Carey E. Priebe. Nonhomogeneity analysis using borrowed strength. Journal of the American Statistical Association, 91 (436):1497–1503, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291575.

Pruitt:1993:IBS

[Pru93] Ronald C. Pruitt. Identifiability of bivariate survival curves from censored data. Journal of the American Statistical Association, 88(422):573–579, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290338.

Pawitan:1993:MDM

[PS93] Yudi Pawitan and Steve Self. Modeling disease market processes in AIDS. Journal of the American Statistical Association, 88(423): 719–726, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290756.

Park:1994:ESE

[PS94a] B. U. Park and L. Simar. Efficient semiparametric estimation in a stochastic frontier model. Journal of the American Statistical Association, 89(427):929–936, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290918.

Phillips:1994:BFH

[PS94b] D. B. Phillips and A. F. M. Smith. Bayesian faces via hierarchical template modeling. Journal of the American Statistical Association, 89(428):1151–1163, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290978.

Peruggia:1996:BAT

[PS96] Mario Peruggia and Thomas Santner. Bayesian analysis of time evolution of earthquakes. Journal of the American Statistical Association, 91(435):1209–1218, September 1996. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2291739.

Pitt:1999:FSA

[PS99] Michael K. Pitt and Neil Shephard. Filtering via simulation: Auxiliary particle filters. Journal of the American Statistical Association, 94(446):590–599, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2670179.

Pericchi:1993:PCR

[PSS93] L. R. Pericchi, B. Sansó, and A. F. M. Smith. Posterior cumulant relationships in Bayesian inference involving the exponential family. Journal of the American Statistical Association, 88(424): 1419–1426, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291286.

Pesaran:1999:PMG

[PSS99] M. Hashem Pesaran, Yongcheol Shin, and Ron P. Smith. Pooled mean group estimation of dynamic heterogeneous panels. Journal of the American Statistical Association, 94(446):621–634, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670182.

Pierce:1992:EVP

[PSVS92] Donald A. Pierce, Daniel O. Stram, Michael Vaeth, and Daniel W. Schafer. The errors-in-variables problem: Considerations provided by radiation dose-response analyses of the A-bomb survivor data. Journal of the American Statistical Association, 87(418): 351–359, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290265.

Press:1990:REF

[PT90] S. James Press and Judith M. Tanur. Regression estimates in Federal Welfare Quality Control Programs: Comment. Journal of the American Statistical Association, 85(411):866–867, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290028.

Puhalskii:1998:BRB

[Puh98] Anatolii Puhalskii. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1245, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669885.

Pan:1996:CGU

[PW96] Guohua Pan and Douglas A. Wolfe. Comparing groups with umbrella orderings. Journal of the American Statistical Association, 91(433):311–317, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291409.

Pauler:1999:BFA

[PWK99] Donna K. Pauler, Jonathan C. Wakefield, and Robert E. Kass. Bayes factors and approximations for variance component models. Journal of the American Statistical Association, 94(448):1242– 1253, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2669938.

Potthoff:1992:ESS

[PWM92] Richard F. Potthoff, Max A. Woodbury, and Kenneth G. Manton. “equivalent sample size” and “equivalent degrees of freedom” refinements for inference using survey weights under super- population models. Journal of the American Statistical Association, 87(418):383–396, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290269.

Pena:1999:FPO

[PY99] Daniel Peña and Victor Yohai. A fast procedure for outlier diagnostics in large regression problems. Journal of the American Statistical Association, 94(446):434–445, June 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/pena.pdf;
 http://www.jstor.org/stable/2670164.

Qu:1998:ACS

[QH98a] Yinsheng Qu and Alula Hadgu. Applications and case studies — a model for evaluating sensitivity and specificity for correlated diagnostic tests in efficacy studies with an imperfect reference test. Journal of the American Statistical Association, 93(443): 920–928, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/sept/qu.html.

Qu:1998:MES

[QH98b] Yinsheng Qu and Alula Hadgu. A model for evaluating sensitivity and specificity for correlated diagnostic tests in efficacy studies with an imperfect reference test. Journal of the American Statistical Association, 93(443):920–928, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669830.

Quintana:1998:AOD

[QN98] Fernando A. Quintana and Michael A. Newton. Assessing the order of dependence for partially exchangeable binary data. Journal of the American Statistical Association, 93(441): 194–202, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/QUINTANA.HTM;   http://
www.jstor.org/stable/2669616.

Qian:1996:IAP

[QTC96] W. Qian, D. M. Titterington, and J. N. Chapman. An image analysis problem in electron microscopy. Journal of the American Statistical Association, 91(435):944–952, September 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291713.

Quade:1990:BRB

[Qua90] Dana Quade. Book review: [untitled]. Journal of the American Statistical Association, 85(412):1180, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289638.

Quintana:1998:NBA

[Qui98a] Fernando A. Quintana. Nonparametric Bayesian analysis for assessing homogeneity in k × l contingency tables with fixed right margin totals. Journal of the American Statistical Association, 93 (443):1140–1149, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/sept/quintana.html;
 http://www.jstor.org/stable/2669857.

Quintos:1998:FMV

[Qui98b] Carmela E. Quintos. Fully modified vector autoregressive inference in partially nonstationary models. Journal of the American Statistical Association, 93(442):783–795, June 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/june/QUINTOS.HTM;   http://www.jstor.org/
stable/2670128.

Rade:1995:BRB

[Rad95] Lennart Rade. Book review: [untitled]. Journal of the American Statistical Association, 90(429):391–392, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291175.

Raghunathan:1993:QEB

[Rag93] T. E. Raghunathan. A quasi-empirical Bayes method for small area estimation. Journal of the American Statistical Association, 88(424):1444–1448, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2291290.

Robinson:1993:EPCa

[RAGW93a] J. Gregory Robinson, Bashir Ahmed, Prithwis Das Gupta, and Karen A. Woodrow. Estimation of population coverage in the 1990 United States census based on demographic analysis. Journal of the American Statistical Association, 88(423):1061–1071, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290799.

Robinson:1993:EPCb

[RAGW93b] J. Gregory Robinson, Bashir Ahmed, Prithwis Das Gupta, and Karen A. Woodrow. Estimation of population coverage in the 1990 United States census based on demographic analysis: Rejoinder. Journal of the American Statistical Association, 88(423): 1077–1079, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290802.

RAJ:1996:BRB

[RAJ96] RAJ. Book review: [untitled]. Journal of the American Statistical Association, 91(434):915–916, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291702.

Ramsey:1993:BRB

[Ram93] Fred L. Ramsey. Book review: [untitled]. Journal of the American Statistical Association, 88(422):704, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290367.

Rao:1993:PCR

[Rao93] C. R. Rao. Is Pitman closeness a reasonable criterion?: Comment. Journal of the American Statistical Association, 88(421):

69–70, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290695.

Rao:1996:VEIa

[Rao96a] J. N. K. Rao. On variance estimation with imputed survey data. Journal of the American Statistical Association, 91(434):499–506, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291637.

Rao:1996:VEIb

[Rao96b] J. N. K. Rao. On variance estimation with imputed survey data: Rejoinder. Journal of the American Statistical Association, 91 (434):519–520, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291643.

Rao:1996:RVE

[Rao96c] J. N. K. Rao. Rejoinder: On variance estimation with imputed survey data. Journal of the American Statistical Association, 91 (434):499–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ratnaparkhi:1994:BRB

[Rat94] Makarand V. Ratnaparkhi. Book review: [untitled]. Journal of the American Statistical Association, 89(425):363, March 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291248.

Rathbun:1998:BRB

[Rat98] Stephen L. Rathbun. Book review: [untitled]. Journal of the American Statistical Association, 93(441):402–403, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669643.

Raudenbush:1991:BRB

[Rau91] Stephen W. Raudenbush. Book review: [untitled]. Journal of the American Statistical Association, 86(413):242–244, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289738.

Ravishanker:1997:BRB

[Rav97] Nalini Ravishanker. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1646–1647, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2965439.

Raz:1990:CTN

[Raz90a] Jonathan Raz. Correction: Testing for no effect when estimating a smooth function by nonparametric regression: a randomization approach. Journal of the American Statistical Association, 85 (412):1182, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289641.

Raz:1990:TNE

[Raz90b] Jonathan Raz. Testing for no effect when estimating a smooth function by nonparametric regression: a randomization approach. Journal of the American Statistical Association, 85(409):132– 138, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289535.

Rousseeuw:1990:RRA

[RB90] Peter J. Rousseeuw and Gilbert W. Bassett, Jr. The remedian: a robust averaging method for large data sets. Journal of the American Statistical Association, 85(409):97–104, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289530.

Rodgers:1993:ESR

[RBD93] Willard L. Rodgers, Charles Brown, and Greg J. Duncan. Errors in survey reports of earnings, hours worked, and hourly wages. Journal of the American Statistical Association, 88(424):1208– 1218, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291259.

RBD:1996:BRB

[RBD96] RBD. Book review: [untitled]. Journal of the American Statistical Association, 91(434):915, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291701.

RBL:1997:BRBa

[RBL97a] RBL. Book review: [untitled]. Journal of the American Statistical Association, 92(437):390, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291500.

RBL:1997:BRBb

[RBL97b] RBL. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1228, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965619.

RBL:1999:BRBa

[RBL99a] RBL. Book review: [untitled]. Journal of the American Statistical Association, 94(447):989–990, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670028.

RBL:1999:BRBb

[RBL99b] RBL. Book review: [untitled]. Journal of the American Statistical Association, 94(447):990, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670030.

RBL:1999:BRBc

[RBL99c] RBL. Book review: [untitled]. Journal of the American Statistical Association, 94(448):1389, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669963.

Rousseeuw:1993:AMA

[RC93] Peter J. Rousseeuw and Christophe Croux. Alternatives to the median absolute deviation. Journal of the American Statistical Association, 88(424):1273–1283, December 1993. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2291267.

Rathbun:1994:STS

[RC94a] Stephen L. Rathbun and Noel Cressie. A space-time survival point process for a longleaf pine forest in Southern Georgia. Journal of the American Statistical Association, 89(428):1164–1174, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290979.

Rathburn:1994:STS

[RC94b] Stephen L. Rathburn and Noel Cressie. A space-time survival point process for a longleaf pine forest in Southern Georgia. Journal of the American Statistical Association, 89(428):1164–??, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

RC:1999:BRB

[RC99] RC. Book review: [untitled]. Journal of the American Statistical Association, 94(446):657, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670204.

Ridder:1994:TEE

[RCH94] G. Ridder, J. S. Cramer, and P. Hopstaken. Down to ten: Estimating the effect of a red card in soccer. Journal of the American Statistical Association, 89(427):1124–1127, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290942.

Roeder:1996:SMA

[RCL96] Kathryn Roeder, Raymond J. Carroll, and Bruce G. Lindsay. A semiparametric mixture approach to case-control studies with errors in covariables. Journal of the American Statistical Association, 91(434):722–732, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291667.

Reeves:1998:BRB

[Ree98] Jaxk H. Reeves. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1235–1236, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669869.

Reid:1990:BRB

[Rei90] N. Reid. Book review: [untitled]. Journal of the American Statistical Association, 85(409):260, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289564.

Reid:1996:CEB

[Rei96a] N. Reid. Comment: Empirical Bayes methods for combining likelihoods. Journal of the American Statistical Association, 91(434):

538–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Reid:1996:EBM

[Rei96b] N. Reid. Empirical Bayes methods for combining likelihoods: Comment. Journal of the American Statistical Association, 91 (434):559, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291650.

Rubin:1999:ECE

[RF99] Donald B. Rubin and Constantine E. Frangakis. Estimation of the causal effect of a time-varying exposure on the marginal mean of a repeated binary outcome: Comment. Journal of the American Statistical Association, 94(447):702–704, September 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669980.

Ronchetti:1997:RLM

[RFB97] Elvezio Ronchetti, Christopher Field, and Wade Blanchard. Robust linear model selection by cross-validation. Journal of the American Statistical Association, 92(439):1017– 1023, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/sept/ronchetti.htm;
 http://www.jstor.org/stable/2965566.

Ritov:1993:ACT

[RG93] Ya’acov Ritov and Zvi Gilula. Analysis of contingency tables by correspondence models subject to order constraints. Journal of the American Statistical Association, 88(424):1380–1387, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291280.

Robins:1994:ADR

[RG94] James M. Robins and Sander Greenland. Adjusting for differential rates of prophylaxis therapy for PCP in high- versus low-dose AZT treatment arms in an AIDS randomized trial. Journal of the American Statistical Association, 89(427):737–749, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290899.

Raghunathan:1995:SQS

[RG95] Trivellore E. Raghunathan and James E. Grizzle. A split questionnaire survey design. Journal of the American Statistical Association, 90(429):54–63, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291129.

Robins:1996:CIC

[RG96a] James M. Robins and Sander Greenland. Comment: Identification of causal effects using instrumental variables. Journal of the American Statistical Association, 91(434):444–??, ???? 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Robins:1996:ICE

[RG96b] James M. Robins and Sander Greenland. Identification of causal effects using instrumental variables: Comment. Journal of the American Statistical Association, 91(434):456–458, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291630.

Robins:1999:ECEb

[RGH99a] J. M. Robins, S. Greenland, and F. C. Hu. Estimation of the causal effect of a time-varying exposure on the marginal mean of a repeated binary outcome: Rejoinder. Journal of the American Statistical Association, 94(447):708–712, September 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669983.

Robins:1999:ECEa

[RGH99b] James M. Robins, Sander Greenland, and Fu-Chang Hu. Estimation of the causal effect of a time-varying exposure on the marginal mean of a repeated binary outcome. Journal of the American Statistical Association, 94(447):687–700, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669978.

Raftery:1995:IDPa

[RGZ95a] Adrian E. Raftery, Geof H. Givens, and Judith E. Zeh. Inference from a deterministic population dynamics model for bowhead whales. Journal of the American Statistical Association, 90 (430):402–416, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291048.

Raftery:1995:IDPb

[RGZ95b] Adrian E. Raftery, Geof H. Givens, and Judith E. Zeh. Inference from a deterministic population dynamics model for bowhead whales: Rejoinder. Journal of the American Statistical Association, 90(430):427–430, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291053.

Raftery:1995:RID

[RGZ95c] Adrian E. Raftery, Geof H. Givens, and Judith E. Zeh. Rejoinder: Inference from a deterministic population dynamics model for bowhead whales. Journal of the American Statistical Association, 90(430):427–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Robert:1996:MLE

[RH96] Christian P. Robert and J. T. Gene Hwang. Maximum likelihood estimation under order restrictions by the prior feedback method. Journal of the American Statistical Association, 91(433): 167–172, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291392.

Rousseeuw:1999:RD

[RH99] Peter J. Rousseeuw and Mia Hubert. Regression depth. Journal of the American Statistical Association, 94(446):388–402, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
rousseeuw.pdf;   http://www.jstor.org/stable/2670155.

Robert:1993:PCRb

[RHS93a] Christian P. Robert, J. T. Gene Hwang, and William E. Strawderman. Is Pitman closeness a reasonable criterion?: Rejoinder. Journal of the American Statistical Association, 88(421):74–76, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290698.

Robert:1993:PCRa

[RHS93b] Christian P. Robert, J. T. Gene Hwang, and William E. Strawderman. Is Pitman closeness a reasonable criterion? With discussion and a reply by the authors. Journal of the American Statistical Association, 88(421):57–63, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290692.

Rousseeuw:1999:RRD

[RHV99] Peter J. Rousseeuw, Mia Hubert, and Stefan Van Aelst. Rejoinder: Regression depth. Journal of the American Statistical Association, 94(446):419–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rice:1990:BRB

[Ric90] John Rice. Book review: [untitled]. Journal of the American Statistical Association, 85(409):261, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289566.

Richardson:1997:BIE

[Ric97a] Alice M. Richardson. Bounded influence estimation in the mixed linear model. Journal of the American Statistical Association, 92 (437):154–161, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/march/richardson.htm;
http://www.jstor.org/stable/2291459.

Richardson:1997:BRB

[Ric97b] Thomas Richardson. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1215–1216, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2965591.

Richardson:1998:BRB

[Ric98] Mary Richardson. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1249–1250, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669892.

Rigby:1997:BDB

[Rig97] Robert A. Rigby. Bayesian discrimination between TWO multivariate normal populations with equal covariance matrices. Journal of the American Statistical Association, 92(439):1151– 1154, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/sept/rigby.htm;   http:
//www.jstor.org/stable/2965580.

Rivest:1998:SLM

[Riv98] Louis-Paul Rivest. Some linear models for estimating the motion of rigid bodies with applications to geometric quality assurance. Journal of the American Statistical Association, 93 (442):632–642, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/june/RIVEST.HTM;
http://www.jstor.org/stable/2670114.

Rizzo:1992:CCE

[Riz92] Louis Rizzo. Conditionally consistent estimators using only probabilities of selection in complex sample surveys. Journal of the American Statistical Association, 87(420):1166–1173, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290656.

RJ:1995:BRB

[RJ95] RJ. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1136, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291374.

RJ:1996:BRB

[RJ96] RJ. Book review: [untitled]. Journal of the American Statistical Association, 91(433):440–441, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291436.

Rosenbaum:1990:STS

[RK90] Paul R. Rosenbaum and Abba M. Krieger. Sensitivity of twosample permutation inferences in observational studies. Journal of the American Statistical Association, 85(410):493–498, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289789.

Ronning:1996:EEO

[RK96] Gerd Ronning and Martin Kukuk. Efficient estimation of ordered probit models. Journal of the American Statistical Association, 91 (435):1120–1129, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291731.

RL:1996:BRB

[RL96] RL. Book review: [untitled]. Journal of the American Statistical Association, 91(436):1754, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291614.

RL:1998:BRBa

[RL98a] RL. Book review: [untitled]. Journal of the American Statistical Association, 93(441):411, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669657.

RL:1998:BRBb

[RL98b] RL. Book review: [untitled]. Journal of the American Statistical Association, 93(441):411, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669658.

RL:1998:BRBc

[RL98c] RL. Book review: [untitled]. Journal of the American Statistical Association, 93(441):411, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669662.

Roeder:1999:MUL

[RLN99] Kathryn Roeder, Kevin G. Lynch, and Daniel S. Nagin. Modeling uncertainty in latent class membership: a case study in criminology. Journal of the American Statistical Association, 94(447): 766–776, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2669989.

RLS:1999:BRBa

[RLS99a] RLS. Book review: [untitled]. Journal of the American Statistical Association, 94(446):657, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670201.

RLS:1999:BRBb

[RLS99b] RLS. Book review: [untitled]. Journal of the American Statistical Association, 94(446):658, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670207.

Raftery:1997:BMA

[RMH97] Adrian E. Raftery, David Madigan, and Jennifer A. Hoeting. Bayesian model averaging for linear regression models. Journal of the American Statistical Association, 92(437): 179–191, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/march/raftery.htm;
 http://www.jstor.org/stable/2291462.

Rubin:1990:MRT

[RMS90] Gail Rubin, Charles E. McCulloch, and Michael A. Shapiro. Multinomial runs tests to detect clustering in constrained free recall. Journal of the American Statistical Association, 85(410): 315–320, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289766.

Reichardt:1997:BRB

[RN97a] Charles S. Reichardt and Jenny A. Novotny. Book review: [untitled]. Journal of the American Statistical Association, 92(438): 796–797, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2965738.

Renssen:1997:AEC

[RN97b] Robbert H. Renssen and Nico J. Nieuwenbroek. Aligning estimates for common variables in two or more sample surveys. Journal of the American Statistical Association, 92(437): 368–374, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/march/renssen.htm;
 http://www.jstor.org/stable/2291482.

Robinson:1990:BRB

[Rob90] James M. Robinson. Book review: [untitled]. Journal of the American Statistical Association, 85(411):908, September 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290047.

Robinson:1991:BRB

[Rob91] Jeff Robinson. Book review: [untitled]. Journal of the American Statistical Association, 86(414):548, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290612.

Robinson:1994:ETN

[Rob94] P. M. Robinson. Efficient tests of nonstationary hypotheses. Journal of the American Statistical Association, 89(428):1420–1437, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291004.

Rockwell:1991:SSP

[Roc91] Richard C. Rockwell. Social statistics and public policy for the 1990s: [introduction]. Journal of the American Statistical Association, 86(414):486, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290600.

Rochon:1992:ACS

[Roc92] James Rochon. ARMA covariance structures with time heteroscedasticity for repeated measures experiments. Journal of the American Statistical Association, 87(419):777–784, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290215.

Rocke:1993:BRB

[Roc93] David M. Rocke. Book review: [untitled]. Journal of the American Statistical Association, 88(422):710–711, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290376.

Rochon:1995:SIT

[Roc95] James Rochon. Supplementing the intent-to-treat analysis: Accounting for covariates observed postrandomization in clinical trials. Journal of the American Statistical Association, 90(429):292– 300, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291153.

Rodgers:1990:BRB

[Rod90] Joseph Lee Rodgers. Book review: [untitled]. Journal of the American Statistical Association, 85(412):1177–1178, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289633.

Roeder:1990:DEC

[Roe90] Kathryn Roeder. Density estimation with confidence sets exemplified by superclusters and voids in the galaxies. Journal of the American Statistical Association, 85(411):617–624, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289993.

Roecker:1992:BRB

[Roe92] Ellen B. Roecker. Book review: [untitled]. Journal of the American Statistical Association, 87(417):250, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290478.

Roeder:1994:GTD

[Roe94] Kathryn Roeder. A graphical technique for determining the number of components in a mixture of normals. Journal of the American Statistical Association, 89(426):487–495, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290850.

Rojo:1996:TCP

[Roj96] Javier Rojo. On tail categorization of probability laws. Journal of the American Statistical Association, 91(433):378–384, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291417.

Rolph:1990:QCW

[Rol90] John E. Rolph. Quality control of welfare programs: [introduction]. Journal of the American Statistical Association, 85(411): 848–849, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290024.

Rolle:1994:BNI

[Rol94] Jean-Daniel Rolle. Best nonnegative invariant partially orthogonal quadratic estimation in normal regression. Journal of the American Statistical Association, 89(428):1378–1385, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291000.

Romano:1990:BRT

[Rom90] Joseph P. Romano. On the behavior of randomization tests without a group invariance assumption. Journal of the American Statistical Association, 85(411):686–692, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290003.

Rosenbaum:1993:HLP

[Ros93a] Paul R. Rosenbaum. Hodges-Lehmann point estimates of treatment effect in observational studies. Journal of the American Statistical Association, 88(424):1250–1253, December 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291264.

Rosenbaum:1993:SLT

[Ros93b] Paul R. Rosenbaum. Sampling the leaves of a tree with equal probabilities. Journal of the American Statistical Association, 88 (424):1455–1457, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291292.

Rosalsky:1994:BRB

[Ros94] Andrew Rosalsky. Book review: [untitled]. Journal of the American Statistical Association, 89(428):1566–1567, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291029.

Rosenbaum:1995:QNS

[Ros95a] Paul R. Rosenbaum. Quantiles in nonrandom samples and observational studies. Journal of the American Statistical Association, 90(432):1424–1431, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291534.

Rosenthal:1995:CMC

[Ros95b] Jeffrey Rosenthal. Correction: “Minorization conditions and convergence rates for Markov chain Monte Carlo”. Journal of the American Statistical Association, 90(431):1136, September 1995.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291375. See

[Ros95c].

Rosenthal:1995:MCC

[Ros95c] Jeffrey S. Rosenthal. Minorization conditions and convergence rates for Markov chain Monte Carlo. Journal of the American Statistical Association, 90(430):558–566, June 1995. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2291067. See correction [Ros95b].

Rosenbaum:1996:CIC

[Ros96a] Paul R. Rosenbaum. Comment: Identification of causal effects using instrumental variables. Journal of the American Statistical Association, 91(434):444–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosenbaum:1996:ICE

[Ros96b] Paul R. Rosenbaum. Identification of causal effects using instrumental variables: Comment. Journal of the American Statistical Association, 91(434):465–468, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291633.

Rosenberger:1997:BRB

[Ros97a] William F. Rosenberger. Book review: [untitled]. Journal of the American Statistical Association, 92(437):384–385, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291486.

Rossman:1997:BRB

[Ros97b] Allan J. Rossman. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1226, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965608.

Rosalsky:1999:BRB

[Ros99] Andrew Rosalsky. Book review: [untitled]. Journal of the American Statistical Association, 94(448):1387–1388, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669958.

Runger:1996:MCM

[RP96] George C. Runger and Sharad S. Prabhu. A Markov chain model for the multivariate exponentially weighted moving averages control chart. Journal of the American Statistical Association, 91 (436):1701–1706, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291599.

Robins:1995:SEM

[RR95] James M. Robins and Andrea Rotnitzky. Semiparametric efficiency in multivariate regression models with missing data. Journal of the American Statistical Association, 90(429):122– 129, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291135.

Rotnitzky:1998:SRR

[RRS98] Andrea Rotnitzky, James M. Robins, and Daniel O. Scharfstein. Semiparametric regression for repeated outcomes with non-ignorable non-response. Journal of the American Statistical Association, 93(444):1321–1339, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
dec/rotnitzky.html;   http://www.jstor.org/stable/2670049.

Robins:1994:ERC

[RRZ94] James M. Robins, Andrea Rotnitzky, and Lue Ping Zhao. Estimation of regression coefficients when some regressors are not always observed. Journal of the American Statistical Association, 89(427):846–866, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290910.

Robins:1995:ASR

[RRZ95] James M. Robins, Andrea Rotnitzky, and Lue Ping Zhao. Analysis of semiparametric regression models for repeated outcomes in the presence of missing data. Journal of the American Statistical Association, 90(429):106–121, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291134.

Reneau:1990:ESC

[RS90a] D. M. Reneau and F. J. Samaniego. Estimating the survival curve when new is better than used of a specified age. Journal of the American Statistical Association, 85(409):123–131, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289534.

Ruppert:1990:UMO

[RS90b] David Ruppert and Douglas G. Simpson. Unmasking multivariate outliers and leverage points: Comment. Journal of the American Statistical Association, 85(411):644–646, September 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289997.

Rojo:1993:ESC

[RS93] Javier Rojo and Francisco J. Samaniego. On estimating a survival curve subject to a uniform stochastic ordering constraint. Journal of the American Statistical Association, 88(422):566–572, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290337.

Rayens:1994:DPG

[RS94a] William S. Rayens and Cidambi Srinivasan. Dependence properties of generalized Liouville distributions on the simplex. Journal of the American Statistical Association, 89(428):1465–1470, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291008.

Ronchetti:1994:RVM

[RS94b] Elvezio Ronchetti and Robert G. Staudte. A robust version of Mallow’s C_{p}. Journal of the American Statistical Association, 89 (426):550–559, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290858.

RS:1995:BRBa

[RS95a] RS. Book review: [untitled]. Journal of the American Statistical Association, 90(429):398, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291186.

RS:1995:BRBb

[RS95b] RS. Book review: [untitled]. Journal of the American Statistical Association, 90(430):809, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291109.

RS:1995:BRBc

[RS95c] RS. Book review: [untitled]. Journal of the American Statistical Association, 90(430):809, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291111.

RS:1995:BRBd

[RS95d] RS. Book review: [untitled]. Journal of the American Statistical Association, 90(430):810, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291115.

RS:1995:BRBe

[RS95e] RS. Book review: [untitled]. Journal of the American Statistical Association, 90(432):1492, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291551.

RS:1995:BRBf

[RS95f] RS. Book review: [untitled]. Journal of the American Statistical Association, 90(432):1493, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291557.

Rao:1996:BHS

[RS96a] J. N. K. Rao and J. Shao. On balanced half-sample variance estimation in stratified random sampling. Journal of the American Statistical Association, 91(433):343–348, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291413.

RS:1996:BRB

[RS96b] RS. Book review: [untitled]. Journal of the American Statistical Association, 91(433):441, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291437.

Ramsay:1998:CSS

[RS98a] James O. Ramsay and Bernard W. Silverman. Comment: Smoothing spline models for the analysis of nested and crossed samples of curves. Journal of the American Statistical Association, 93(443):988–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ramsay:1998:SSM

[RS98b] James O. Ramsay and Bernard W. Silverman. Smoothing spline models for the analysis of nested and crossed samples of curves: Comment. Journal of the American Statistical Association, 93 (443):988–991, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669842.

Rubin:1995:HDK

[RSV95] Donald B. Rubin, Hal S. Stern, and Vasja Vehovar. Handling “don’t know” survey responses: The case of the Slovenian plebiscite. Journal of the American Statistical Association, 90 (431):822–828, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291315.

Ruppert:1995:EBS

[RSW95] D. Ruppert, S. J. Sheather, and M. P. Wand. An effective bandwidth selector for local least squares regression. Journal of the American Statistical Association, 90(432):1257–1270, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291516.

Rao:1993:GLS

[RSY93] J. N. K. Rao, Brajendra C. Sutradhar, and Kim Yue. Generalized least squares F test in regression analysis with two-stage cluster samples. Journal of the American Statistical Association, 88(424): 1388–1391, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291281.

Ritter:1992:FGS

[RT92] Christian Ritter and Martin A. Tanner. Facilitating the Gibbs sampler: The Gibbs stopper and the Griddy-Gibbs sampler. Journal of the American Statistical Association, 87(419):861–868, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290225.

Romano:1996:IAU

[RT96a] Joseph P. Romano and Lori A. Thombs. Inference for autocorrelations under weak assumptions. Journal of the American Statistical Association, 91(434):590–600, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291655.

Rossini:1996:SPO

[RT96b] A. J. Rossini and A. A. Tsiatis. A semiparametric proportional odds regression model for the analysis of current status data. Journal of the American Statistical Association, 91(434):713–721, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291666.

Raz:1992:FDE

[RTF92] Jonathan Raz, Bruce Turetsky, and George Fein. Frequency domain estimation of the parameters of human brain electrical dipoles. Journal of the American Statistical Association, 87(417): 69–77, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290453.

Ruberg:1990:CCI

[Rub90] Stephen J. Ruberg. Correction: Contrasts for identifying the minimum effective dose. Journal of the American Statistical Association, 85(410):608, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289833.

Rubin:1991:CEV

[Rub91] Donald B. Rubin. Compliance as an explanatory variable in clinical trials: Comment: Dose-response estimands. Journal of the American Statistical Association, 86(413):22–24, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289710.

Rubin:1994:MDI

[Rub94] Donald B. Rubin. Missing data, imputation, and the bootstrap: Comment. Journal of the American Statistical Association, 89 (426):475–478, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290847.

Rubin:1996:MIA

[Rub96a] Donald B. Rubin. Multiple imputation after 18+ years. Journal of the American Statistical Association, 91(434):473–489, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291635.

Rubin:1996:VEI

[Rub96b] Donald B. Rubin. On variance estimation with imputed survey data: Rejoinder. Journal of the American Statistical Association, 91(434):515–517, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291641.

Rubin:1996:RVE

[Rub96c] Donald B. Rubin. Rejoinder: On variance estimation with imputed survey data. Journal of the American Statistical Association, 91(434):499–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rudemo:1991:TDE

[Rud91] Mats Rudemo. Transformations in density estimation: Comment. Journal of the American Statistical Association, 86(414):353–354, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290570.

Rue:1995:NLF

[Rue95] Håvard Rue. New loss functions in Bayesian imaging. Journal of the American Statistical Association, 90(431):900–908, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291324.

Rukhin:1993:BET

[Ruk93] Andrew L. Rukhin. Bahadur efficiency of tests of separate hypotheses and adaptive test statistics. Journal of the American Statistical Association, 88(421):161–165, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290709.

Rukhin:1998:BRB

[Ruk98] Andrew L. Rukhin. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1525, December 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670071.

Runger:1999:BRB

[Run99] George C. Runger. Book review: [untitled]. Journal of the American Statistical Association, 94(446):649–650, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670188.

Ruppert:1997:EBB

[Rup97] David Ruppert. Empirical-bias bandwidths for local polynomial nonparametric regression and density estimation. Journal of the American Statistical Association, 92(439):1049– 1062, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/sept/ruppert.htm;
http://www.jstor.org/stable/2965570.

Rutter:1991:BRB

[Rut91] Carolyn M. Rutter. Book review: [untitled]. Journal of the American Statistical Association, 86(414):548–549, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290613.

Rousseeuw:1990:RUM

[Rv90a] Peter J. Rousseeuw and Bert C. van Zomeren. Rejoinder: Unmasking multivariate outliers and leverage points. Journal of the American Statistical Association, 85(411):648–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rousseeuw:1990:UMOa

[Rv90b] Peter J. Rousseeuw and Bert C. van Zomeren. Unmasking multivariate outliers and leverage points. Journal of the American Statistical Association, 85(411):633–639, September 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289995.

Rust:1990:SDC

[RV90c] Philip F. Rust and Eberhard O. Voit. Statistical densities, cumulatives, quantiles, and power obtained by S-system differential equations. Journal of the American Statistical Association, 85 (410):572–578, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289800.

Rukhin:1998:ECM

[RV98] Andrew L. Rukhin and Mark G. Vangel. Estimation of a common mean and weighted means statistics. Journal of the American Statistical Association, 93(441):303–308, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
RUKHIN.HTM;   http://www.jstor.org/stable/2669626.

Rousseeuw:1999:RDR

[RVH99] Peter J. Rousseeuw, Stefan Van Aelst, and Mia Hubert. Regression depth: Rejoinder. Journal of the American Statistical Association, 94(446):419–433, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2670163.

RVL:1997:BRBa

[RVL97a] RVL. Book review: [untitled]. Journal of the American Statistical Association, 92(437):389, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291494.

RVL:1997:BRBb

[RVL97b] RVL. Book review: [untitled]. Journal of the American Statistical Association, 92(437):389, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291495.

RVL:1997:BRBc

[RVL97c] RVL. Book review: [untitled]. Journal of the American Statistical Association, 92(438):799, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965743.

RVL:1997:BRBd

[RVL97d] RVL. Book review: [untitled]. Journal of the American Statistical Association, 92(438):800, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965754.

RVL:1997:BRBe

[RVL97e] RVL. Book review: [untitled]. Journal of the American Statistical Association, 92(438):800, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965756.

RVL:1997:BRBf

[RVL97f] RVL. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1656, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965462.

RVL:1999:BRBa

[RVL99a] RVL. Book review: [untitled]. Journal of the American Statistical Association, 94(445):350, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669727.

RVL:1999:BRBb

[RVL99b] RVL. Book review: [untitled]. Journal of the American Statistical Association, 94(446):657–658, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670206.

RVL:1999:BRBc

[RVL99c] RVL. Book review: [untitled]. Journal of the American Statistical Association, 94(446):658, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670208.

Rousseeuw:1990:UMOb

[RvZ90] Peter J. Rousseeuw and Bert C. van Zomeren. Unmasking multivariate outliers and leverage points: Rejoinder. Journal of the American Statistical Association, 85(411):648–651, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289999.

Rocke:1996:IOM

[RW96] David M. Rocke and David L. Woodruff. Identification of outliers in multivariate data. Journal of the American Statistical Association, 91(435):1047–1061, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291724.

Roeder:1997:PBD

[RW97] Kathryn Roeder and Larry Wasserman. Practical Bayesian density estimation using mixtures of normals. Journal of the American Statistical Association, 92(439):894–902, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2965553.

Rogan:1998:CCO

[RW98a] Walter J. Rogan and James H. Ware. Comment: Correcting for omitted-variables and measurement-error bias in regression with an application to the effect of lead on IQ. Journal of the American Statistical Association, 93(442):513–??, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rogan:1998:COV

[RW98b] Walter J. Rogan and James H. Ware. Correcting for omittedvariables and measurement-error bias in regression with an application to the effect of lead on IQ: Comment. Journal of the American Statistical Association, 93(442):513–514, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670097.

Ryu:1993:EMO

[Ryu93] Keunkwan Ryu. An extension of Marshall and Olkin’s bivariate exponential distribution. Journal of the American Statistical Association, 88(424):1458–1465, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291293.

Ryu:1994:GDA

[Ryu94] Keunkwan Ryu. Group duration analysis of the proportional hazard model: Minimum chi-squared estimators and specification tests. Journal of the American Statistical Association, 89(428): 1386–1397, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291001.

Raftery:1998:EBW

[RZ98] Adrian E. Raftery and Judith E. Zeh. Estimating bowhead whale population size and rate of increase from the 1993 census. Journal of the American Statistical Association, 93(442): 451–463, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/june/RAFTERY.HTM;
http://www.jstor.org/stable/2670092.

Sobel:1992:MHF

[SA92] Michael E. Sobel and Gerhard Arminger. Modeling household fertility decisions: a nonlinear simultaneous probit model. Journal of the American Statistical Association, 87(417):38–47, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290449.

Shively:1990:FED

[SAK90] Thomas S. Shively, Craig F. Ansley, and Robert Kohn. Fast evaluation of the distribution of the Durbin-Watson and other invariant test statistics in time series regression. Journal of the American Statistical Association, 85(411):676–685, September 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290002.

Sallas:1993:BRB

[Sal93] William M. Sallas. Book review: [untitled]. Journal of the American Statistical Association, 88(421):376–377, March 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290737.

Samaniego:1990:BRB

[Sam90a] Francisco J. Samaniego. Book review: [untitled]. Journal of the American Statistical Association, 85(412):1166–1167, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289618.

Sampson:1990:MCM

[Sam90b] Michael Sampson. A Markov chain model for unskilled workers and the highly mobile. Journal of the American Statistical Association, 85(409):177–180, March 1990. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289541.

Samarov:1993:ERS

[Sam93a] Alexander M. Samarov. Exploring regression structure using nonparametric functional estimation. Journal of the American Statistical Association, 88(423):836–847, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290772.

Samuels:1993:SPR

[Sam93b] Myra L. Samuels. Simpson’s paradox and related phenomena. Journal of the American Statistical Association, 88(421): 81–88, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290700.

Samuels:1997:BRB

[Sam97] Stephen M. Samuels. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1223–1224, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2965604.

Santner:1990:BRB

[San90] Thomas Santner. Book review: [untitled]. Journal of the American Statistical Association, 85(410):596–597, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289813.

Sandefur:1991:BRB

[San91] Gary D. Sandefur. Book review: [untitled]. Journal of the American Statistical Association, 86(413):241–242, March 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289737.

Santos:1998:ACS

[San98a] Robert L. Santos. Applications and case studies — comment: Not asked and not answered: Multiple imputation for multiple surveys. Journal of the American Statistical Association, 93(443): 865–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Santos:1998:AAM

[San98b] Robert L. Santos. Not asked and not answered: Multiple imputation for multiple surveys: Comment. Journal of the American Statistical Association, 93(443):865–866, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669823.

Saphire:1991:BRB

[Sap91] Diane G. Saphire. Book review: [untitled]. Journal of the American Statistical Association, 86(415):829–830, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290428.

Sarkar:1991:BRB

[Sar91a] Sanat K. Sarkar. Book review: [untitled]. Journal of the American Statistical Association, 86(416):1144–1145, December 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290544.

Sarle:1991:BRB

[Sar91b] Warren S. Sarle. Book review: [untitled]. Journal of the American Statistical Association, 86(415):830–832, September 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290430.

Sarkar:1994:BRB

[Sar94] Sanat K. Sarkar. Book review: [untitled]. Journal of the American Statistical Association, 89(425):362–363, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291247.

Sarndal:1996:EES

[Sär96] Carl-Erik Särndal. Efficient estimators with simple variance in unequal probability sampling. Journal of the American Statistical Association, 91(435):1289–1300, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291747.

Sasieni:1993:MWP

[Sas93] Peter Sasieni. Maximum weighted partial likelihood estimators for the Cox model. Journal of the American Statistical Association, 88(421):144–152, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290707.

Sastry:1997:NFM

[Sas97] Narayan Sastry. A nested frailty model for survival data, with an application to the study of child survival in northeast Brazil. Journal of the American Statistical Association, 92(438):426–435, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2965690.

Satorra:1990:BRB

[Sat90] Albert Satorra. Book review: [untitled]. Journal of the American Statistical Association, 85(411):910–911, September 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290050.

Satten:1995:ULB

[Sat95] Glen A. Satten. Upper and lower bound distributions that give simultaneous confidence intervals for quantiles. Journal of the American Statistical Association, 90(430):747–752, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291087.

Stasny:1990:DRC

[SB90] Elizabeth A. Stasny and Harold R. Bauer. Deviations from randomness in Children’s early speech. Journal of the American Statistical Association, 85(409):46–57, March 1990. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289524.

Stefanski:1995:IVE

[SB95] L. A. Stefanski and J. S. Buzas. Instrumental variable estimation in binary regression measurement error models. Journal of the American Statistical Association, 90(430):541–550, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291065.

Sengupta:1997:ROC

[SB97] Debasis Sengupta and Pochiraju Bhimasankaram. On the roles of observations in collinearity in the linear model. Journal of the American Statistical Association, 92(439):1024– 1032, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/sept/sengupta.htm;
 http://www.jstor.org/stable/2965567.

Sheiner:1997:ANCa

[SBD97a] Lewis B. Sheiner, Stuart L. Beal, and Adrian Dunne. Analysis of nonrandomly censored ordered categorical longitudinal data from analgesic trials. Journal of the American Statistical Association, 92(440):1235–1244, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/december/SHEINER.HTM;   http://www.jstor.
org/stable/2965391.

Sheiner:1997:ANCb

[SBD97b] Lewis B. Sheiner, Stuart L. Beal, and Adrian Dunne. Analysis of nonrandomly censored ordered categorical longitudinal data from analgesic trials: Rejoinder. Journal of the American Statistical Association, 92(440):1253–1255, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965395.

Sheiner:1997:RAN

[SBD97c] Lewis B. Sheiner, Stuart L. Beal, and Adrian Dunne. Rejoinder: Analysis of nonrandomly censored ordered categorical longitudinal data from analgesic trials. Journal of the American Statistical Association, 92(440):1253–??, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sain:1994:CVM

[SBS94] Stephan R. Sain, Keith A. Baggerly, and David W. Scott. Crossvalidation of multivariate densities. Journal of the American Statistical Association, 89(427):807–817, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290906.

Schlax:1992:FDD

[SC92a] Michael G. Schlax and Dudley B. Chelton. Frequency domain diagnostics for linear smoothers. Journal of the American Statistical Association, 87(420):1070–1081, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290644.

StLaurent:1992:LSN

[SC92b] Roy T. St. Laurent and R. Dennis Cook. Leverage and superleverage in nonlinear regression. Journal of the American Statistical Association, 87(420):985–990, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290635.

Sherman:1994:NEM

[SC94] Michael Sherman and Edward Carlstein. Nonparametric estimation of the moments of a general statistic computed from spatial data. Journal of the American Statistical Association, 89(426): 496–500, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290851.

Stefanski:1995:SEM

[SC95] L. A. Stefanski and J. R. Cook. Simulation-extrapolation: The measurement error jackknife. Journal of the American Statistical Association, 90(432):1247–1256, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291515.

Samuel-Cahn:1996:OSR

[SC96a] Ester Samuel-Cahn. Optimal stopping with random horizon with application to the full-information best-choice problem with random freeze. Journal of the American Statistical Association, 91 (433):357–364, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291415.

Sherman:1996:RH

[SC96b] Michael Sherman and Edward Carlstein. Replicate histograms. Journal of the American Statistical Association, 91(434):566–576, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291653.

Sakkar:1997:SMM

[SC97a] Sanat K. Sakkar and Chung-Kuei Chang. The Simes method for multiple hypothesis testing with positively dependent test statistics. Journal of the American Statistical Association, 92(440): 1601–1608, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/SARKAR.HTM.

Sarkar:1997:SMM

[SC97b] Sanat K. Sarkar and Chung-Kuei Chang. The Simes method for multiple hypothesis testing with positively dependent test statistics. Journal of the American Statistical Association, 92(440): 1601–1608, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/SARKAR.HTM;
http://www.jstor.org/stable/2965431.

SullivanPepe:1997:MPC

[SC97c] Margaret Sullivan Pepe and David Couper. Modeling partly conditional means with longitudinal data. Journal of the American Statistical Association, 92(439):991–??, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sherman:1999:CPS

[SC99a] Michael Sherman and Edward Carlstein. Comment: Prediction of spatial cumulative distribution functions using subsampling. Journal of the American Statistical Association, 94(445):102–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sherman:1999:PSC

[SC99b] Michael Sherman and Edward Carlstein. Prediction of spatial cumulative distribution functions using subsampling: Comment. Journal of the American Statistical Association, 94(445):102– 105, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2669684.

Scarsini:1992:BRB

[Sca92] Marco Scarsini. Book review: [untitled]. Journal of the American Statistical Association, 87(417):250–251, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290479.

Shao:1998:BRR

[SCC98] Jun Shao, Ying Chen, and Yinzhong Chen. Balanced repeated replication for stratified multistage survey data under imputation. Journal of the American Statistical Association, 93 (442):819–831, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/june/SHAO.HTM;   http:/
/www.jstor.org/stable/2670131.

Symons:1993:BNC

[SCF93] Michael J. Symons, Chu-Chih Chen, and Michael R. Flynn. Bayesian nonparametrics for compliance to exposure standards. Journal of the American Statistical Association, 88(424):1237– 1241, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291262.

Schegloff:1990:ITF

[Sch90a] Emanuel A. Schegloff. Interactional troubles in face-to-face survey interviews: Comment. Journal of the American Statistical Association, 85(409):248–250, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289554.

Schervish:1990:BRB

[Sch90b] Mark J. Schervish. Book review: [untitled]. Journal of the American Statistical Association, 85(412):1169–1170, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289623.

Schmidely:1990:BRB

[Sch90c] A. Dianne Schmidely. Book review: [untitled]. Journal of the American Statistical Association, 85(410):603–604, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289825.

Schaefer:1991:BRB

[Sch91a] Robert L. Schaefer. Book review: [untitled]. Journal of the American Statistical Association, 86(415):830, September 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290429.

Schirm:1991:ECU

[Sch91b] Allen L. Schirm. The effects of census undercount adjustment on Congressional apportionment. Journal of the American Statistical Association, 86(414):526–541, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290604.

Schott:1991:TSP

[Sch91c] James R. Schott. A test for a specific principal component of a correlation matrix. Journal of the American Statistical Association, 86(415):747–751, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290407.

Schultz:1991:BRB

[Sch91d] T. Paul Schultz. Book review: [untitled]. Journal of the American Statistical Association, 86(416):1145–1146, December 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290545.

Schafer:1993:ABB

[Sch93a] Joseph L. Schafer. Assessing between-block heterogeneity within the post-strata of the 1990 post-enumeration survey: Comment. Journal of the American Statistical Association, 88(423):1125– 1127, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290807.

Schenker:1993:UCS

[Sch93b] Nathaniel Schenker. Undercount in the 1990 census: Special section. Journal of the American Statistical Association, 88(423): 1044–1046, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290797.

Scheaffer:1994:BRB

[Sch94a] Richard L. Scheaffer. Book review: [untitled]. Journal of the American Statistical Association, 89(426):711–712, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290878.

Schick:1994:BRB

[Sch94b] Anton Schick. Book review: [untitled]. Journal of the American Statistical Association, 89(428):1565–1566, December 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291028.

Schmidt:1994:BRB

[Sch94c] Steven R. Schmidt. Book review: [untitled]. Journal of the American Statistical Association, 89(425):360–361, March 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291244.

Schmoyer:1994:PTC

[Sch94d] Richard L. Schmoyer. Permutation tests for correlation in regression errors. Journal of the American Statistical Association, 89 (428):1507–1516, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291013.

Schott:1994:DDS

[Sch94e] James R. Schott. Determining the dimensionality in sliced inverse regression. Journal of the American Statistical Association, 89 (425):141–148, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291210.

Scheuren:1995:BRB

[Sch95a] Fritz Scheuren. Book review: [untitled]. Journal of the American Statistical Association, 90(429):386–388, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291170.

Schucany:1995:ABC

[Sch95b] William R. Schucany. Adaptive bandwidth choice for kernel regression. Journal of the American Statistical Association, 90(430): 535–540, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291064.

Schweder:1995:CID

[Sch95c] Tore Schweder. Comment: Inference from a deterministic population dynamics model for bowhead whales. Journal of the American Statistical Association, 90(430):420–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schweder:1995:IDP

[Sch95d] Tore Schweder. Inference from a deterministic population dynamics model for bowhead whales: Comment. Journal of the American Statistical Association, 90(430):420–423, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291050.

Schwenke:1995:BRB

[Sch95e] James Schwenke. Book review: [untitled]. Journal of the American Statistical Association, 90(429):390, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291172.

Schachermayer:1996:BRB

[Sch96a] W. Schachermayer. Book review: [untitled]. Journal of the American Statistical Association, 91(434):912–913, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291694.

Schervish:1996:BRB

[Sch96b] Mark J. Schervish. Book review: [untitled]. Journal of the American Statistical Association, 91(435):1376–1377, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291759.

Schmid:1996:EAF

[Sch96c] Christopher H. Schmid. An EM algorithm fitting first-order conditional autoregressive models to longitudinal data. Journal of the American Statistical Association, 91(435):1322–1330, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291750.

Schottl:1997:BRB

[Sch97] A. Schottl. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1650–1651, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965445.

Schmock:1998:BRB

[Sch98a] Uwe Schmock. Book review: [untitled]. Journal of the American Statistical Association, 93(442):840–842, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670142.

Schott:1998:BRB

[Sch98b] James R. Schott. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1236–1237, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669871.

Schell:1999:BDE

[Sch99a] Michael J. Schell. Bridging different eras in sports: Comment. Journal of the American Statistical Association, 94(447):681–684, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2669976.

Schmertmann:1999:EMT

[Sch99b] Carl P. Schmertmann. Estimating multistate transition hazards from last-move data. Journal of the American Statistical Association, 94(445):53–63, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669677.

Schott:1999:BRB

[Sch99c] James R. Schott. Book review: [untitled]. Journal of the American Statistical Association, 94(448):1388, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669960.

Samuels:1991:IBRb

[SCM91a] Myra L. Samuels, George Casella, and George P. McCabe. Interpreting blocks and random factors: Rejoinder. Journal of the American Statistical Association, 86(415):818–821, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290420.

Samuels:1991:IBRa

[SCM91b] Myra L. Samuels, George Casella, and George P. McCabe. Interpreting blocks and random factors. With comments and a rejoinder by the authors. Journal of the American Statistical Association, 86(415):798–808, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290415.

Scott:1990:BRB

[Sco90] David W. Scott. Book review: [untitled]. Journal of the American Statistical Association, 85(410):598, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289816.

Scott:1991:TDE

[Sco91] David W. Scott. Transformations in density estimation: Comment. Journal of the American Statistical Association, 86(414): 359, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290574.

Sconing:1999:BRB

[Sco99] James Sconing. Book review: [untitled]. Journal of the American Statistical Association, 94(445):335–337, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669708.

Strawderman:1996:PSS

[SCW96] Robert L. Strawderman, George Casella, and Martin T. Wells. Practical small-sample asymptotics for regression problems. Journal of the American Statistical Association, 91(434):643–654, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291660. See

correction [SCW97].

Strawderman:1997:CPS

[SCW97] Robert L. Strawderman, George Casella, and Martin T. Wells. Correction to “Practical Small-Sample Asymptotics for Regression Problems”. Journal of the American Statistical Association, 92(440):1657–??, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [SCW96].

Sinha:1997:SBA

[SD97] Debajyoti Sinha and Dipak K. Dey. Semiparametric Bayesian analysis of survival data. Journal of the American Statistical Association, 92(439):1195–1212, September 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_97/
sept/sinha.htm;   http://www.jstor.org/stable/2965586.

Satten:1998:IBI

[SDW98] Glen A. Satten, Somnath Datta, and John M. Williamson. Inference based on imputed failure times for the proportional hazards model with interval censored data. Journal of the American Statistical Association, 93(441):318–327, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
SATTEN.HTM;   http://www.jstor.org/stable/2669628.

Seaman:1998:BRB

[Sea98] John W. Seaman, Jr. Book review: [untitled]. Journal of the American Statistical Association, 93(442):838–839, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670138.

Seely:1994:BRB

[See94] Justus Seely. Book review: [untitled]. Journal of the American Statistical Association, 89(426):714, June 1994. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2290882.

Seeber:1995:BRB

[See95] G. U. H. Seeber. Book review: [untitled]. Journal of the American Statistical Association, 90(429):385–386, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291168.

Segal:1992:TSM

[Seg92] Mark Robert Segal. Tree-structured methods for longitudinal data. Journal of the American Statistical Association, 87(418): 407–418, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290271.

Soofi:1995:IDA

[SEH95] Ehsan S. Soofi, Nader Ebrahimi, and Mohamed Habibullah. Information distinguishability with application to analysis of failure data. Journal of the American Statistical Association, 90(430): 657–668, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291079.

Sekkappan:1990:LE

[Sek90] R. M. Sekkappan. Letters to the editor. Journal of the American Statistical Association, 85(410):608, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289831.

Selvin:1995:BRB

[Sel95] Steve Selvin. Book review: [untitled]. Journal of the American Statistical Association, 90(430):798–799, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291094.

Selvin:1996:BRB

[Sel96] Steve Selvin. Book review: [untitled]. Journal of the American Statistical Association, 91(433):433, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291427.

Sen:1992:BRB

[Sen92] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 87(419):909–910, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290249.

Sen:1993:BRB

[Sen93] A. R. Sen. Book review: [untitled]. Journal of the American Statistical Association, 88(424):1471–1472, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291302.

Sen:1994:BRB

[Sen94] Pranab K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 89(425):358–359, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291240.

Serlin:1991:BRB

[Ser91] Ronald C. Serlin. Book review: [untitled]. Journal of the American Statistical Association, 86(413):247–248, March 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289744.

Serfozo:1993:BRB

[Ser93] Richard F. Serfozo. Book review: [untitled]. Journal of the American Statistical Association, 88(423):1168, September 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290817.

Sethuraman:1995:BRB

[Set95] Jayaram Sethuraman. Book review: [untitled]. Journal of the American Statistical Association, 90(429):384, March 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291166.

Severini:1991:CML

[Sev91] Thomas A. Severini. A comparison of the maximum likelihood estimator and the posterior mean in the single-parameter case. Journal of the American Statistical Association, 86(416):997– 1000, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290516.

Severini:1994:ABI

[Sev94] Thomas A. Severini. Approximately Bayesian inference. Journal of the American Statistical Association, 89(425):242–249, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291220.

Severini:1995:ICI

[Sev95] Thomas A. Severini. Information and conditional inference. Journal of the American Statistical Association, 90(432):1341–1346, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291524.

Severini:1996:MSR

[Sev96] Thomas A. Severini. Measures of the sensitivity of regression estimates to the choice of estimator. Journal of the American Statistical Association, 91(436):1651–1658, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291592.

Seymour:1997:BRB

[Sey97] Lynne Seymour. Book review: [untitled]. Journal of the American Statistical Association, 92(440):1647, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965440.

Seymour:1999:BRB

[Sey99] Lynne Seymour. Book review: [untitled]. Journal of the American Statistical Association, 94(448):1389, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669962.

Stein:1997:COE

[SF97a] Michael L. Stein and Dongping Fang. Comment: Ozone exposure and population density in Harris County, Texas. Journal of the American Statistical Association, 92(438):405–??, June 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CCG^{+}97a].

Stein:1997:OEP

[SF97b] Michael L. Stein and Dongping Fang. Ozone exposure and population density in Harris County, Texas: Comment. Journal of the American Statistical Association, 92(438):408–411, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2965686. See

[CCG^{+}97a].

Sampson:1992:NEN

[SG92] Paul D. Sampson and Peter Guttorp. Nonparametric estimation of nonstationary spatial covariance structure. Journal of the American Statistical Association, 87(417):108–119, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290458.

Sison:1995:CSC

[SG95a] Christina P. Sison and Joseph Glaz. Correction: “Simultaneous confidence intervals and sample size determination for multinomial proportions” [J. Amer. Statist. Assoc. 90 (1995), no. 429, 366–369, MR 96e:62039]. Journal of the American Statistical Association, 90(432):1494, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [SG95b].

Sison:1995:SCI

[SG95b] Cristina P. Sison and Joseph Glaz. Simultaneous confidence intervals and sample size determination for multinomial proportions. Journal of the American Statistical Association, 90(429): 366–369, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2291162. See correction [SG95a].

Seifert:1996:FSV

[SG96a] Burkhardt Seifert and Theo Gasser. Finite-sample variance of local polynomials: Analysis and solutions. Journal of the American Statistical Association, 91(433):267–275, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291404.

Sinclair:1996:PEE

[SG96b] Michael D. Sinclair and Joseph L. Gastwirth. On procedures for evaluating the effectiveness of reinterview survey methods: Application to labor force data. Journal of the American Statistical Association, 91(435):961–969, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291715.

Swanson:1997:IRF

[SG97] Norman R. Swanson and Clive W. J. Granger. Impulse response functions based on a causal approach to residual orthogonalization in vector autoregressions. Journal of the American Statistical Association, 92(437):357–367, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/march/
swanson.htm;   http://www.jstor.org/stable/2291481.

Schwartz:1995:TMC

[SGC95] Pamela F. Schwartz, Chris Gennings, and Vernon M. Chinchilli. Threshold models for combination data from reproductive and developmental experiments. Journal of the American Statistical Association, 90(431):862–870, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291320.

Stute:1998:BAM

[SGP98] W. Stute, W. González Manteiga, and M. Presedo Quindimil. Bootstrap approximations in model checks for regression. Journal of the American Statistical Association, 93(441):141– 149, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/STUTE.HTM;   http:/
/www.jstor.org/stable/2669611.

Sleeper:1990:RSC

[SH90] Lynn A. Sleeper and David P. Harrington. Regression splines in the Cox model with application to covariate effects in liver disease. Journal of the American Statistical Association, 85(412): 941–949, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289591.

Sanchez:1992:OSR

[SH92] Susan M. Sanchez and Julia L. Higle. Observational studies of rare events: a subset selection approach. Journal of the American Statistical Association, 87(419):878–883, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290227.

Shaffer:1991:PIA

[Sha91] Juliet Popper Shaffer. On the problem of interactions in the analysis of variance: Comment. Journal of the American Statistical Association, 86(414):367–369, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290577. See [Fab91b].

Shafer:1992:BRB

[Sha92a] Glenn Shafer. Book review: [untitled]. Journal of the American Statistical Association, 87(418):585–586, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290300.

Shanmugam:1992:BRB

[Sha92b] Ramalingam Shanmugam. Book review: [untitled]. Journal of the American Statistical Association, 87(417):253–254, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290483.

Shanmugam:1993:BRB

[Sha93a] Ramalingam Shanmugam. Book review: [untitled]. Journal of the American Statistical Association, 88(423):1172, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290825.

Shao:1993:LMS

[Sha93b] Jun Shao. Linear model selection by cross-validation. Journal of the American Statistical Association, 88(422):486–494, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290328.

Shao:1996:BMS

[Sha96] Jun Shao. Bootstrap model selection. Journal of the American Statistical Association, 91(434):655–665, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291661.

Shaked:1998:BRB

[Sha98] Moshe Shaked. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1237–1238, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669872.

Shafer:1999:BRB

[Sha99] Glenn Shafer. Book review: [untitled]. Journal of the American Statistical Association, 94(446):648–649, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670186.

Spiegelhalter:1994:EEP

[SHBF94] David J. Spiegelhalter, Nomi L. Harris, Kate Bull, and Rodney C. G. Franklin. Empirical evaluation of prior beliefs about frequencies: Methodology and a case study in congenital heart disease. Journal of the American Statistical Association, 89(426): 435–443, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290843.

Shephard:1993:MLE

[She93] Neil Shephard. Maximum likelihood estimation of regression models with stochastic trend components. Journal of the American Statistical Association, 88(422):590–595, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290340.

Sherman:1995:BRB

[She95] Malcolm J. Sherman. Book review: [untitled]. Journal of the American Statistical Association, 90(432):1492, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291550.

Sherman:1997:SMR

[She97] Michael Sherman. Subseries methods in regression. Journal of the American Statistical Association, 92(439):1041– 1048, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/sept/sherman.htm;
http://www.jstor.org/stable/2965569.

Sun:1996:BMV

[SHGL96] Li Sun, John S. J. Hsu, Irwin Guttman, and Tom Leonard. Bayesian methods for variance component models. Journal of the American Statistical Association, 91(434):743–752, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291669.

Shi:1991:THO

[Shi91] Ning-Zhong Shi. A test of homogeneity of odds ratios against order restrictions. Journal of the American Statistical Association, 86(413):154–158, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289725.

Shlay:1990:BRB

[Shl90] Anne B. Shlay. Book review: [untitled]. Journal of the American Statistical Association, 85(410):604–605, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289826.

Simpson:1998:CEP

[SHL98a] Douglas G. Simpson, Xuming He, and Yao-Tsorng Liu. Comment: Edge preserving smoothers for image processing. Journal of the American Statistical Association, 93(442):544–??, June 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Simpson:1998:EPS

[SHL98b] Douglas G. Simpson, Xuming He, and Yao-Tsorng Liu. Edgepreserving smoothers for image processing: Comment. Journal of the American Statistical Association, 93(442):544–548, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670102.

Short:1992:BRB

[Sho92] Thomas H. Short. Book review: [untitled]. Journal of the American Statistical Association, 87(419):901–902, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290238.

Short:1993:ADM

[Sho93] Thomas H. Short. An algorithm for the detection and measurement of rail surface defects. Journal of the American Statistical Association, 88(422):436–440, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290322.

Shumway:1990:BRB

[Shu90] Robert H. Shumway. Book review: [untitled]. Journal of the American Statistical Association, 85(412):1168, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289621.

Shumway:1993:BRB

[Shu93] Robert H. Shumway. Book review: [untitled]. Journal of the American Statistical Association, 88(422):705–706, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290369.

Shun:1997:ALS

[Shu97] Zhenming Shun. Another look at the salamander mating data: a modified Laplace approximation approach. Journal of the American Statistical Association, 92(437):341–349, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_97/march/shun.htm;   http://www.jstor.
org/stable/2291479.

Strauss:1990:PES

[SI90] David Strauss and Michael Ikeda. Pseudolikelihood estimation for social networks. Journal of the American Statistical Association, 85(409):204–212, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289546.

Siegel:1993:SCI

[Sie93] Andrew F. Siegel. A surprising covariance involving the minimum of multivariate normal variables. Journal of the American Statistical Association, 88(421):77–80, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290699.

Silvapulle:1992:RWT

[Sil92] Mervyn J. Silvapulle. Robust Wald-type tests of one-sided hypotheses in the linear model. Journal of the American Statistical Association, 87(417):156–161, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290464.

Silvapulle:1996:TPN

[Sil96] Mervyn J. Silvapulle. A test in the presence of nuisance parameters. Journal of the American Statistical Association, 91(436): 1690–1693, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291597.

Silliman:1997:HSM

[Sil97a] Nancy Paul Silliman. Hierarchical selection models with applications in meta- analysis. Journal of the American Statistical Association, 92(439):926–936, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/sept/
silliman.htm;   http://www.jstor.org/stable/2965556.

Silvapulle:1997:CTP

[Sil97b] Mervyn J. Silvapulle. Correction: a test in the presence of nuisance parameters. Journal of the American Statistical Association, 92(438):801, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2965759.

Simonoff:1993:IMO

[Sim93a] Jeffrey S. Simonoff. The identification of multiple outliers: Comment. Journal of the American Statistical Association, 88(423): 793–794, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290764.

Simpson:1993:BRB

[Sim93b] Douglas G. Simpson. Book review: [untitled]. Journal of the American Statistical Association, 88(423):1172, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290824.

Simonoff:1994:BRB

[Sim94] Jeffrey S. Simonoff. Book review: [untitled]. Journal of the American Statistical Association, 89(428):1559–1560, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291020.

Sinha:1993:SBA

[Sin93] Debajyoti Sinha. Semiparametric Bayesian analysis of multiple event time data. Journal of the American Statistical Association, 88(423):979–983, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290789.

Singh:1997:BRB

[Sin97] Kesar Singh. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1214, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965588.

Sitter:1992:RPC

[Sit92] R. R. Sitter. A resampling procedure for complex survey data. Journal of the American Statistical Association, 87(419):755–765, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290213.

Sitter:1997:VER

[Sit97] R. R. Sitter. Variance estimation for the regression estimator in two-phase sampling. Journal of the American Statistical Association, 92(438):780–787, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2965727.

Suchman:1990:ITFa

[SJ90a] Lucy Suchman and Brigitte Jordan. Interactional troubles in faceto-face survey interviews. Journal of the American Statistical Association, 85(409):232–241, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289550.

Suchman:1990:ITFb

[SJ90b] Lucy Suchman and Brigitte Jordan. Interactional troubles in faceto-face survey interviews: Rejoinder. Journal of the American Statistical Association, 85(409):252–253, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289556.

Shiboski:1992:SAT

[SJ92] Stephen C. Shiboski and Nicholas P. Jewell. Statistical analysis of the time dependence of HIV infectivity based on partner study data. Journal of the American Statistical Association, 87(418): 360–372, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290266.

Storer:1990:EPS

[SK90] Barry E. Storer and Choongrak Kim. Exact properties of some exact test statistics for comparing two binomial proportions. Journal of the American Statistical Association, 85(409):146–155, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289537.

Shumway:1991:APM

[SK91] Robert H. Shumway and Myron J. Katzoff. Adjustment of provisional mortality series: The dynamic linear model with structured measurement errors. Journal of the American Statistical Association, 86(415):611–617, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290389.

Satten:1993:IAE

[SK93a] Glen A. Satten and Lawrence L. Kupper. Inferences about exposure–disease associations using probability-of-exposure information. Journal of the American Statistical Association, 88(421): 200–208, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290714.

Sun:1993:ACS

[SK93b] Jianguo Sun and John D. Kalbfleisch. The analysis of current status data on point processes. Journal of the American Statistical Association, 88(424):1449–1454, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291291.

Smith:1997:BAN

[SK97] Michael Smith and Robert Kohn. A Bayesian approach to nonparametric bivariate regression. Journal of the American Statistical Association, 92(440):1522–1535, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/december/SMITH_M.HTM;   http://www.jstor.
org/stable/2965423.

Sepanski:1994:SCA

[SKC94] J. H. Sepanski, R. Knickerbocker, and R. J. Carroll. A semiparametric correction for attenuation. Journal of the American Statistical Association, 89(428):1366–1373, December 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290998.

Stasny:1998:FDP

[SKF98] Elizabeth A. Stasny, Joseph B. Kadane, and Kathleen S. Fritsch. On the fairness of death-penalty jurors: a comparison of Bayesian models with different levels of hierarchy and various missingdata mechanisms. Journal of the American Statistical Association, 93(442):464–477, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.amstat.org/publications/jasa/3r0c0005.pdf;   http:
//www.amstat.org/publications/jasa/abstracts_98/june/
STASNY.HTM;   http://www.jstor.org/stable/2670093.

Skinner:1991:ERR

[Ski91] C. J. Skinner. On the efficiency of raking ratio estimation for multiple frame surveys. Journal of the American Statistical Association, 86(415):779–784, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290412.

Skinner:1996:BRB

[Ski96] Chris Skinner. Book review: [untitled]. Journal of the American Statistical Association, 91(434):911–912, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291692.

Shively:1999:VSFa

[SKW99a] Thomas S. Shively, Robert Kohn, and Sally Wood. Variable selection and function estimation in additive nonparametric regression using a data-based prior. Journal of the American Statistical Association, 94(447):777–794, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669990.

Shively:1999:VSFb

[SKW99b] Thomas S. Shively, Robert Kohn, and Sally Wood. Variable selection and function estimation in additive nonparametric regression using a data-based prior: Rejoinder. Journal of the American Statistical Association, 94(447):804–806, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669995.

Saikkonen:1993:TMA

[SL93a] Pentti Saikkonen and Ritva Luukkonen. Testing for a moving average unit root in autoregressive integrated moving average models. Journal of the American Statistical Association, 88(422): 596–601, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290341.

Su:1993:LCM

[SL93b] John Q. Su and Jun S. Liu. Linear combinations of multiple diagnostic markers. Journal of the American Statistical Association, 88(424):1350–1355, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291276.

Spiekerman:1998:MRM

[SL98a] C. F. Spiekerman and D. Y. Lin. Marginal regression models for multivariate failure time data. Journal of the American Statistical Association, 93(443):1164–1175, September 1998.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/sept/spiekerman.html;   http://www.jstor.
org/stable/2669859.

Staniswalis:1998:NRA

[SL98b] Joan G. Staniswalis and J. Jack Lee. Nonparametric regression analysis of longitudinal data. Journal of the American Statistical Association, 93(444):1403–1418, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/dec/staniswalis.html;   http://www.jstor.
org/stable/2670055.

Slate:1993:BRB

[Sla93] Elizabeth H. Slate. Book review: [untitled]. Journal of the American Statistical Association, 88(421):384–385, March 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290748.

Slate:1994:PNE

[Sla94] Elizabeth H. Slate. Parameterizations for natural exponential families with quadratic variance functions. Journal of the American Statistical Association, 89(428):1471–1482, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291009.

Shah:1997:REM

[SLS97] Amrik Shah, Nan Laird, and David Schoenfeld. A random effects model for multiple characteristics with possibly missing data. Journal of the American Statistical Association, 92(438):775–779, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2965726.

Sheather:1990:KQE

[SM90a] Simon J. Sheather and J. S. Marron. Kernel quantile estimators. Journal of the American Statistical Association, 85(410):410–416, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289777.

Spencer:1990:NDE

[SM90b] Bruce D. Spencer and Lincoln E. Moses. Needed data expenditure for an ambiguous decision problem. Journal of the American Statistical Association, 85(412):1099–1104, December 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289607.

Spall:1992:FBE

[SM92] James C. Spall and John L. Maryak. A feasible Bayesian estimator of quantiles for projectile accuracy from non-iid data. Journal of the American Statistical Association, 87(419):676–681, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290205.

Seillier-Moiseiwitsch:1993:BRB

[SM93a] Françoise Seillier-Moiseiwitsch. Book review: [untitled]. Journal of the American Statistical Association, 88(422):711–713, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290377.

Stern:1993:SAH

[SM93b] Hal S. Stern and Carl N. Morris. A statistical analysis of hitting streaks in baseball: Comment. Journal of the American Statistical Association, 88(424):1189–1194, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291256.

Seillier-Moiseiwitsch:1999:BRB

[SM99] Françoise Seillier-Moiseiwitsch. Book review: [untitled]. Journal of the American Statistical Association, 94(446):649, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670187.

Seillier-Moiseiwitsch:1993:TVS

[SMD93] F. Seillier-Moiseiwitsch and A. P. Dawid. On testing the validity of sequential probability forecasts. Journal of the American Statistical Association, 88(421):355–359, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290731.

Seo:1994:GTC

[SMF94] Takashi Seo, Shigeru Mano, and Yasunori Fujikoshi. A generalized Tukey conjecture for multiple comparisons among mean vectors. Journal of the American Statistical Association, 89(426):676–679, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290870.

Smith:1994:BRB

[Smi94] Paul J. Smith. Book review: [untitled]. Journal of the American Statistical Association, 89(425):356–357, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291237.

Smith:1996:BRBb

[Smi96a] Paul J. Smith. Book review: [untitled]. Journal of the American Statistical Association, 91(436):1752, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291610.

Smith:1996:BRBa

[Smi96b] Philip J. Smith. Book review: [untitled]. Journal of the American Statistical Association, 91(436):1751–1752, December 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291609.

Smith:1998:BRB

[Smi98] Paul J. Smith. Book review: [untitled]. Journal of the American Statistical Association, 93(441):407–408, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669651.

Senchaudhuri:1995:EEP

[SMP95a] Pralay Senchaudhuri, Cyrus R. Mehta, and Nitin R. Patel. Estimating exact p values by the method of control variates or Monte Carlo rescue. Journal of the American Statistical Association, 90 (430):640–648, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291077.

Senchaudhuri:1995:EEV

[SMP95b] Pralay Senchaudhuri, Cyrus R. Mehta, and Nitin R. Patel. Estimating exact p values by the method of control variates or Monte Carlo rescue. Journal of the American Statistical Association, 90 (430):640–648, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Spurrier:1990:SSA

[SN90] John D. Spurrier and Azhar Nizam. Sample size allocation for simultaneous inference in comparison with control experiments.

Journal of the American Statistical Association, 85(409):181– 186, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289542.

Samaniego:1996:HBB

[SN96] Francisco J. Samaniego and Andrew A. Neath. How to be a better Bayesian. Journal of the American Statistical Association, 91 (434):733–742, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291668.

Stehman:1994:CVE

[SO94] Stephen V. Stehman and W. Scott Overton. Comparison of variance estimators of the Horvitz–Thompson estimator for randomized variable probability systematic sampling. Journal of the American Statistical Association, 89(425):30–43, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291198.

Sobel:1991:BMS

[Sob91] Marc J. Sobel. Bayes and minimax sample-admissible multivariate selection procedures. Journal of the American Statistical Association, 86(416):1022–1023, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290520.

Sobel:1993:BEB

[Sob93] Marc J. Sobel. Bayes and empirical Bayes procedures for comparing parameters. Journal of the American Statistical Association, 88(422):687–693, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290352.

Sobel:1996:BRB

[Sob96] Michael E. Sobel. Book review: [untitled]. Journal of the American Statistical Association, 91(433):438, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291433.

Sobel:1998:BRB

[Sob98] Michael E. Sobel. Book review: [untitled]. Journal of the American Statistical Association, 93(442):837, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670136.

Solow:1991:EAO

[Sol91] Andrew R. Solow. An exploratory analysis of the occurrence of explosive volcanism in the Northern Hemisphere, 1851–1985. Journal of the American Statistical Association, 86(413):49–54, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289714.

Solow:1994:BRB

[Sol94] Andrew R. Solow. Book review: [untitled]. Journal of the American Statistical Association, 89(428):1564, December 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291025.

Solow:1995:EAR

[Sol95] Andrew R. Solow. An exploratory analysis of a record of el nino events, 1800- 1987. Journal of the American Statistical Association, 90(429):72–77, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291131.

Soofi:1992:GFC

[Soo92] Ehsan S. Soofi. A generalizable formulation of conditional logit with diagnostics. Journal of the American Statistical Association, 87(419):812–816, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290219.

Soofi:1994:CIC

[Soo94] Ehsan S. Soofi. Capturing the intangible concept of information. Journal of the American Statistical Association, 89(428):1243– 1254, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290988.

Spino:1991:ECP

[SP91] Cathie Spino and Marcello Pagano. Efficient calculation of the permutation distribution of trimmed means. Journal of the American Statistical Association, 86(415):729–737, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290405.

Speed:1991:IBR

[Spe91] T. P. Speed. Interpreting blocks and random factors: Comment. Journal of the American Statistical Association, 86(415):808–811, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290416.

Speckman:1993:BRB

[Spe93] Paul Speckman. Book review: [untitled]. Journal of the American Statistical Association, 88(424):1467–1468, December 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291295.

Spence:1998:BRB

[Spe98] Ian Spence. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1524, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670070.

Scheike:1999:RAR

[SPM99] Thomas H. Scheike, Jorgen H. Petersen, and Torben Martinussen. Retrospective ascertainment of recurrent events: An application to time to pregnancy. Journal of the American Statistical Association, 94(447):713–725, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669984.

Spurrier:1999:ECB

[Spu99] John D. Spurrier. Exact confidence bounds for all contrasts of three or more regression lines. Journal of the American Statistical Association, 94(446):483–488, June 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/spurrier.
pdf;   http://www.jstor.org/stable/2670168.

Stromberg:1992:BNR

[SR92] Arnold J. Stromberg and David Ruppert. Breakdown in nonlinear regression. Journal of the American Statistical Association, 87 (420):991–997, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290636.

Samaniego:1994:TRB

[SR94] Francisco J. Samaniego and Dana M. Reneau. Toward a reconciliation of the Bayesian and frequentist approaches to point estimation. Journal of the American Statistical Association, 89(427): 947–957, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290920.

Singh:1995:AGF

[SR95] A. C. Singh and J. N. K. Rao. On the adjustment of gross flow estimates for classification error with application to data from the Canadian Labour Force Survey. Journal of the American Statistical Association, 90(430):478–488, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291058.

Skinner:1996:EDF

[SR96] C. J. Skinner and J. N. K. Rao. Estimation in dual frame surveys with complex designs. Journal of the American Statistical Association, 91(433):349–356, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291414.

Simpson:1992:OSG

[SRC92] D. G. Simpson, D. Ruppert, and R. J. Carroll. On one-step GM estimates and stability of inferences in linear regression. Journal of the American Statistical Association, 87(418):439–450, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290275.

Srinivasan:1993:BRB

[Sri93] T. N. Srinivasan. Book review: [untitled]. Journal of the American Statistical Association, 88(421):377, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290738.

SRJ:1999:BRBa

[SRJ99a] SRJ. Book review: [untitled]. Journal of the American Statistical Association, 94(445):350, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669728.

SRJ:1999:BRBb

[SRJ99b] SRJ. Book review: [untitled]. Journal of the American Statistical Association, 94(446):657, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670202.

SRJ:1999:BRBc

[SRJ99c] SRJ. Book review: [untitled]. Journal of the American Statistical Association, 94(446):657, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670205.

SRJ:1999:BRBd

[SRJ99d] SRJ. Book review: [untitled]. Journal of the American Statistical Association, 94(447):990, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670029.

SRJ:1999:BRBe

[SRJ99e] SRJ. Book review: [untitled]. Journal of the American Statistical Association, 94(447):990, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2670031.

SRJ:1999:BRBf

[SRJ99f] SRJ. Book review: [untitled]. Journal of the American Statistical Association, 94(448):1390, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669969.

SRJ:1999:BRBg

[SRJ99g] SRJ. Book review: [untitled]. Journal of the American Statistical Association, 94(448):1390, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669970.

Silber:1995:CCG

[SRR95] Jeffrey H. Silber, Paul R. Rosenbaum, and Richard N. Ross. Comparing the contributions of groups of predictors: Which outcomes vary with hospital rather than patient characteristics? Journal of the American Statistical Association, 90(429):7–18, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291124.

Scharfstein:1999:ANDa

[SRR99a] Daniel O. Scharfstein, Andrea Rotnitzky, and James M. Robins. Adjusting for nonignorable drop-out using semiparametric nonresponse models. Journal of the American Statistical Association, 94(448):1096–1120, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2669923.

Scharfstein:1999:ANDb

[SRR99b] Daniel O. Scharfstein, Andrea Rotnitzky, and James M. Robins. Adjusting for nonignorable drop-out using semiparametric nonresponse models: Rejoinder. Journal of the American Statistical Association, 94(448):1135–1146, December 1999. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2669930.

Stahel:1994:REA

[SRSD94] Werner A. Stahel, Andreas F. Ruckstuhl, Peter Senn, and Kurt Dressler. Robust estimation in the analysis of complex molecular spectra. Journal of the American Statistical Association, 89(427): 788–795, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290904.

Smith:1990:RBL

[SS90a] Stanley K. Smith and Terry Sincich. The relationship between the length of the base period and population forecast errors. Journal of the American Statistical Association, 85(410):367–375, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289772.

Stine:1990:BAS

[SS90b] Robert A. Stine and Paul Shaman. Bias of autoregressive spectral estimators. Journal of the American Statistical Association, 85 (412):1091–1098, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289606.

Shumway:1991:DLM

[SS91a] R. H. Shumway and D. S. Stoffer. Dynamic linear models with switching. Journal of the American Statistical Association, 86 (415):763–769, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290410.

Staniswalis:1991:DAR

[SS91b] Joan G. Staniswalis and Thomas A. Severini. Diagnostics for assessing regression models. Journal of the American Statistical Association, 86(415):684–692, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290398.

Shumway:1992:CDL

[SS92] R. H. Shumway and D. S. Stoffer. Correction: Dynamic linear models with switching. Journal of the American Statistical Association, 87(419):913, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290252.

Severini:1994:QLE

[SS94] Thomas A. Severini and Joan G. Staniswalis. Quasi-likelihood estimation in semiparametric models. Journal of the American Statistical Association, 89(426):501–511, June 1994. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2290852. See correction [Sta94a].

Silvapulle:1995:STA

[SS95a] Mervyn J. Silvapulle and Paramsothy Silvapulle. A score test against one-sided alternatives. Journal of the American Statistical Association, 90(429):342–349, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291159.

Smith:1995:EPP

[SS95b] Stanley K. Smith and Mohammed Shahidullah. An evaluation of population projection errors for census tracts. Journal of the American Statistical Association, 90(429):64–71, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291130.

Sain:1996:LAD

[SS96a] Stephan R. Sain and David W. Scott. On locally adaptive density estimation. Journal of the American Statistical Association, 91 (436):1525–1534, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291578.

Shao:1996:BIS

[SS96b] Jun Shao and Randy R. Sitter. Bootstrap for imputed survey data. Journal of the American Statistical Association, 91(435): 1278–1288, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291746.

Solna:1996:TTE

[SS96c] Knut Sølna and Paul Switzer. Time trend estimation for a geographic region. Journal of the American Statistical Association, 91(434):577–589, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291654.

Schell:1997:RMR

[SS97] Michael J. Schell and Bahadur Singh. The reduced monotonic regression method. Journal of the American Statistical Association, 92(437):128–135, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/march/schell.htm;
http://www.jstor.org/stable/2291456.

Shao:1999:VES

[SS99] Jun Shao and Philip Steel. Variance estimation for survey data with composite imputation and nonnegligible sampling fractions. Journal of the American Statistical Association, 94(445):254– 265, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2669700.

Sackrowitz:1994:RUH

[SSC94] Harold Sackrowitz and Ester Samuel-Cahn. Rationality and unbiasedness in hypothesis testing. Journal of the American Statistical Association, 89(427):967–971, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290922.

Schervish:1990:SDU

[SSK90] Mark J. Schervish, Teddy Seidenfeld, and Joseph B. Kadane. State-dependent utilities. Journal of the American Statistical Association, 85(411):840–847, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290023.

Shen:1999:RSL

[SSW99] Xiaotong Shen, Jian Shi, and Wing Hung Wong. Random sieve likelihood and general regression models. Journal of the American Statistical Association, 94(447):835–846, September 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669998.

Soper:1992:ETS

[ST92] Keith A. Soper and Philip Troilo. Exact tests for in situ hybridization experiments. Journal of the American Statistical Association, 87(417):78–83, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290454.

Severini:1996:CEB

[ST96a] Thomas A. Severini and Martin A. Tanner. Comment: Empirical Bayes methods for combining likelihoods. Journal of the American Statistical Association, 91(434):538–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Severini:1996:EBM

[ST96b] Thomas A. Severini and Martin A. Tanner. Empirical Bayes methods for combining likelihoods: Comment. Journal of the American Statistical Association, 91(434):560–562, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291651.

Siqueira:1999:TEL

[ST99a] Arminda Lucia Siqueira and Jeremy M. G. Taylor. Treatment effects in a logistic model involving the Box–Cox transformation. Journal of the American Statistical Association, 94(445): 240–246, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2669698.

Siquiera:1999:TEL

[ST99b] Arminda Lucia Siquiera and Jeremy M. G. Taylor. Treatment effects in a logistic model involving the Box–Cox transformation. Journal of the American Statistical Association, 94(445):240–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Staniswalis:1990:CKE

[Sta90] J. G. Staniswalis. Correction: The kernel estimate of a regression function in likelihood based models. Journal of the American Statistical Association, 85(412):1182, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289640.

Stasny:1991:HMP

[Sta91] Elizabeth A. Stasny. Hierarchical models for the probabilities of a survey classification and nonresponse: An example from the National Crime Survey. Journal of the American Statistical Association, 86(414):296–303, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290561.

Stasny:1993:BRB

[Sta93] Elizabeth A. Stasny. Book review: [untitled]. Journal of the American Statistical Association, 88(421):377–378, March 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290739.

Staniswalis:1994:CQL

[Sta94a] Joan G. Staniswalis. Correction to: “Quasi-likelihood estimation in semiparametric models” [J. Amer. Statist. Assoc. 89 (1994), no. 426, 501–511 and 1294076] by T. A. Severini and Joan G. Staniswalis. Journal of the American Statistical Association, 89 (428):1572, ???? 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [SS94].

Staniswalis:1994:CSS

[Sta94b] Joan G. Staniswalis. Correction to severini and staniswalis (1994): Quasi-likelihood estimation in semiparametric models. Journal of the American Statistical Association, 89(428):1572, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291043.

Stahel:1998:BRB

[Sta98a] Werner Stahel. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1240–1241, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669876.

Starica:1998:BRB

[Sta98b] Catalin Starica. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1247–1248, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669889.

Stangl:1999:BRB

[Sta99] Dalene Stangl. Book review: [untitled]. Journal of the American Statistical Association, 94(447):985–986, September 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670020.

Steele:1990:BRB

[Ste90a] J. Michael Steele. Book review: [untitled]. Journal of the American Statistical Association, 85(410):596, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289812.

Steffey:1990:BRB

[Ste90b] Duane Steffey. Book review: [untitled]. Journal of the American Statistical Association, 85(412):1171–1172, December 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289625.

Stern:1990:MDP

[Ste90c] Hal Stern. Models for distributions on permutations. Journal of the American Statistical Association, 85(410):558–564, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289798.

Stewart:1992:BRB

[Ste92] Paul W. Stewart. Book review: [untitled]. Journal of the American Statistical Association, 87(417):255–256, March 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290486.

Stern:1994:BMM

[Ste94] Hal S. Stern. A Brownian motion model for the progress of sports scores. Journal of the American Statistical Association, 89(427): 1128–1134, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290943.

Stein:1995:FDA

[Ste95] Michael L. Stein. Fixed-domain asymptotics for spatial periodograms. Journal of the American Statistical Association, 90 (432):1277–1288, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291518.

Stefanov:1998:BRB

[Ste98] Valeri Stefanov. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1515–1516, December 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670065.

Stein:1999:CPS

[Ste99a] Michael L. Stein. Comment: Prediction of spatial cumulative distribution functions using subsampling. Journal of the American Statistical Association, 94(445):106–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stein:1999:PSC

[Ste99b] Michael L. Stein. Prediction of spatial cumulative distribution functions using subsampling: Comment. Journal of the American Statistical Association, 94(445):106–107, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669685.

Stine:1992:BRB

[Sti92] Robert A. Stine. Book review: [untitled]. Journal of the American Statistical Association, 87(419):903–904, September 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290240.

Stigler:1995:BRB

[Sti95] Stephen M. Stigler. Book review: [untitled]. Journal of the American Statistical Association, 90(430):803–804, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291101.

Stout:1990:BRB

[Sto90] William Stout. Book review: [untitled]. Journal of the American Statistical Association, 85(411):901–902, September 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290036.

Stoffer:1991:WFAa

[Sto91a] David S. Stoffer. Walsh–Fourier analysis and its statistical applications. Journal of the American Statistical Association, 86(414): 461–479, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290595.

Stoffer:1991:WFAb

[Sto91b] David S. Stoffer. Walsh–Fourier analysis and its statistical applications: Rejoinder. Journal of the American Statistical Association, 86(414):483–485, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290599.

Stoker:1993:SBD

[Sto93] Thomas M. Stoker. Smoothing bias in density derivative estimation. Journal of the American Statistical Association, 88(423): 855–863, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290774.

Stoffer:1994:BRB

[Sto94] David S. Stoffer. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1138–1139, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290946.

Stoffer:1997:BRB

[Sto97] David S. Stoffer. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1215, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965590.

Stoffer:1999:DCS

[Sto99] David S. Stoffer. Detecting common signals in multiple time series using the spectral envelope. Journal of the American Statistical Association, 94(448):1341–1356, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669947.

Stromberg:1993:CHB

[Str93a] Arnold J. Stromberg. Computation of high breakdown nonlinear regression parameters. Journal of the American Statistical Association, 88(421):237–244, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290718.

Stroup:1993:BRB

[Str93b] Walter W. Stroup. Book review: [untitled]. Journal of the American Statistical Association, 88(421):381–382, March 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290744.

Stroup:1994:BRB

[Str94] Walter W. Stroup. Book review: [untitled]. Journal of the American Statistical Association, 89(426):713–714, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290881.

Stromberg:1995:BRB

[Str95] Arnold J. Stromberg. Book review: [untitled]. Journal of the American Statistical Association, 90(429):386, March 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291169.

Scharfstein:1997:SEI

[STR97] Daniel O. Scharfstein, Anastasios A. Tsiatis, and James M. Robins. Semiparametric efficiency and its implication on the design and analysis of group- sequential studies. Journal of the American Statistical Association, 92(440):1342–1350, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_97/december/SCHARFST.HTM;   http://www.
jstor.org/stable/2965404.

Strawderman:1998:BRB

[Str98a] Robert Strawderman. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1529, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670080.

Stroup:1998:BRB

[Str98b] Donna F. Stroup. Book review: [untitled]. Journal of the American Statistical Association, 93(442):843, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670146.

Strawderman:1999:BRBa

[Str99a] Robert Strawderman. Book review: [untitled]. Journal of the American Statistical Association, 94(445):349, March 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669724.

Strawderman:1999:BRBb

[Str99b] Robert L. Strawderman. Book review: [untitled]. Journal of the American Statistical Association, 94(445):349–350, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669725.

Strawderman:1999:BRBc

[Str99c] Robert L. Strawderman. Book review: [untitled]. Journal of the American Statistical Association, 94(445):350, March 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669726.

Strawderman:1999:BRBd

[Str99d] Robert L. Strawderman. Book review: [untitled]. Journal of the American Statistical Association, 94(447):984–985, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670019.

Stuetzle:1992:BRB

[Stu92] Werner Stuetzle. Book review: [untitled]. Journal of the American Statistical Association, 87(419):897–899, September 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290233.

Stute:1999:CSM

[Stu99] Winfried Stute. Comment on stute, manteiga, and quindimil: Reply. Journal of the American Statistical Association, 94(446):660, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670213.

Sudman:1990:BRB

[Sud90] Seymour Sudman. Book review: [untitled]. Journal of the American Statistical Association, 85(410):601–602, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289823.

Sun:1999:BRB

[Sun99] Don X. Sun. Book review: [untitled]. Journal of the American Statistical Association, 94(446):650–651, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670189.

Sutton:1993:CIM

[Sut93] Clifton D. Sutton. Computer-intensive methods for tests about the mean of an asymmetrical distribution. Journal of the American Statistical Association, 88(423):802–810, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290769.

Scarsini:1993:BDN

[SV93] Marco Scarsini and Achilles Venetoulias. Bivariate distributions with nonmonotone dependence structure. Journal of the American Statistical Association, 88(421):338–344, March 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290729.

Shmueli:1990:PMC

[SW90] Uri Shmueli and George H. Weiss. Probabilistic methods in crystal structure analysis. Journal of the American Statistical Association, 85(409):6–19, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289520.

Stoffer:1991:BSS

[SW91a] David S. Stoffer and Kent D. Wall. Bootstrapping statespace models: Gaussian maximum likelihood estimation and the Kalman filter. Journal of the American Statistical Association, 86 (416):1024–1033, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290521.

Su:1991:LFT

[SW91b] John Q. Su and L. J. Wei. A lack-of-fit test for the mean function in a generalized linear model. Journal of the American Statistical Association, 86(414):420–426, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290587.

Sitter:1993:AFI

[SW93] R. R. Sitter and C. F. J. Wu. On the accuracy of Fieller intervals for binary response data. Journal of the American Statistical Association, 88(423):1021–1025, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290794.

SW:1994:BRBa

[SW94a] SW. Book review: [untitled]. Journal of the American Statistical Association, 89(428):1570, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291037.

SW:1994:BRBb

[SW94b] SW. Book review: [untitled]. Journal of the American Statistical Association, 89(428):1570, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291039.

Sakata:1995:ADF

[SW95a] Shinichi Sakata and Halbert White. An alternative definition of finite-sample breakdown point with applications to regression model estimators. Journal of the American Statistical Association, 90(431):1099–1106, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291348.

SW:1995:BRBa

[SW95b] SW. Book review: [untitled]. Journal of the American Statistical Association, 90(429):398, March 1995. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291185.

SW:1995:BRBb

[SW95c] SW. Book review: [untitled]. Journal of the American Statistical Association, 90(430):809, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291110.

SW:1995:BRBc

[SW95d] SW. Book review: [untitled]. Journal of the American Statistical Association, 90(430):810, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291118.

SW:1996:BRB

[SW96] SW. Book review: [untitled]. Journal of the American Statistical Association, 91(435):1380, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291764.

Strawderman:1997:ABC

[SW97a] Robert L. Strawderman and Martin T. Wells. Accurate bootstrap confidence limits for the cumulative hazard and survivor functions under random censoring. Journal of the American Statistical Association, 92(440):1356–1374, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_97/
december/STRAWDER.HTM;   http://www.jstor.org/stable/
2965406.

SW:1997:BRBa

[SW97b] SW. Book review: [untitled]. Journal of the American Statistical Association, 92(438):799, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965745.

SW:1997:BRBb

[SW97c] SW. Book review: [untitled]. Journal of the American Statistical Association, 92(438):799, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2965749.

Stock:1998:MUE

[SW98a] James H. Stock and Mark W. Watson. Median unbiased estimation of coefficient variance in a time varying parameter model. Journal of the American Statistical Association, 93 (441):349–358, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/STOCK.HTM;   http:/
/www.jstor.org/stable/2669631.

Strawderman:1998:AEIa

[SW98b] Robert L. Strawderman and Martin T. Wells. Approximately exact inference for the common odds ratio in several 2 × 2 tables. Journal of the American Statistical Association, 93(444): 1294–1307, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/dec/strawderman.html;
http://www.jstor.org/stable/2670044.

Strawderman:1998:AEIb

[SW98c] Robert L. Strawderman and Martin T. Wells. Approximately exact inference for the common odds ration in several 2×2 tables: Rejoinder. Journal of the American Statistical Association, 93 (444):1316–1320, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2670048.

Strawderman:1998:RAE

[SW98d] Robert L. Strawderman and Martin T. Wells. Rejoinder: Approximately exact inference for the common odds ratio in several 2×2 tables. Journal of the American Statistical Association, 93(444): 1316–??, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

SW:1998:BRBa

[SW98e] SW. Book review: [untitled]. Journal of the American Statistical Association, 93(441):412, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669665.

SW:1998:BRBb

[SW98f] SW. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1251, September 1998. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669898.

SW:1999:BRB

[SW99] SW. Book review: [untitled]. Journal of the American Statistical Association, 94(448):1390, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669968.

Swanson:1999:BRB

[Swa99] Norman R. Swanson. Book review: [untitled]. Journal of the American Statistical Association, 94(445):347, March 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669719.

Switzer:1990:BRB

[Swi90] Paul Switzer. Book review: [untitled]. Journal of the American Statistical Association, 85(411):906–907, September 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290044.

Sun:1995:RPB

[SY95] Dong Chu Sun and Keying Ye. Reference prior Bayesian analysis for normal mean products. Journal of the American Statistical Association, 90(430):589–597, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291070.

Syski:1996:BRB

[Sys96] R. Syski. Book review: [untitled]. Journal of the American Statistical Association, 91(436):1752–1753, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291611.

Stromberg:1999:BRB

[SZ99] Arnold Stromberg and Mai Zhou. Book review: [untitled]. Journal of the American Statistical Association, 94(446):653, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670193.

Thompson:1990:EIR

[TA90] G. L. Thompson and L. P. Ammann. Efficiencies of interblock rank statistics for repeated measures designs. Journal of the American Statistical Association, 85(410):519–528, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289793.

Tjostheim:1994:NINa

[TA94a] Dag Tjøstheim and Bjørn H. Auestad. Nonparametric identification of nonlinear time series: Projections. Journal of the American Statistical Association, 89(428):1398–1409, December 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291002.

Tjostheim:1994:NINb

[TA94b] Dag Tjøstheim and Bjørn H. Auestad. Nonparametric identification of nonlinear time series: Selecting significant lags. Journal of the American Statistical Association, 89(428):1410–1419, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291003.

Tableman:1990:BIR

[Tab90] Mara Tableman. Bounded-influence rank regression: a one-step estimator based on Wilcoxon scores. Journal of the American Statistical Association, 85(410):508–513, June 1990. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289791.

Tahir:1992:BRB

[Tah92] Mohamed Tahir. Book review: [untitled]. Journal of the American Statistical Association, 87(420):1249–1250, December 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290680.

Takane:1992:BRB

[Tak92] Yoshio Takane. Book review: [untitled]. Journal of the American Statistical Association, 87(418):587–588, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290303.

Tam:1995:ORS

[Tam95] S. M. Tam. Optimal and robust strategies for cluster sampling. Journal of the American Statistical Association, 90(429): 379–382, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291164.

Tanur:1990:BRB

[Tan90] Judith M. Tanur. Book review: [untitled]. Journal of the American Statistical Association, 85(410):591–592, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289805.

Tang:1993:OAB

[Tan93] Boxin Tang. Orthogonal array-based Latin hypercubes. Journal of the American Statistical Association, 88(424):1392–1397, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291282.

Tang:1994:UMP

[Tan94] Dei-In Tang. Uniformly more powerful tests in a one-sided multivariate problem. Journal of the American Statistical Association, 89(427):1006–1011, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290927.

Tarpey:1999:SCP

[Tar99] Thaddeus Tarpey. Self-consistency and principal component analysis. Journal of the American Statistical Association, 94(446): 456–467, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2670166.

Tasker:1998:BRB

[Tas98] Gary Tasker. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1233, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669866.

Taylor:1990:BRB

[Tay90] Jeremy M. G. Taylor. Book review: [untitled]. Journal of the American Statistical Association, 85(410):593, June 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289808.

Taylor:1994:SMA

[TCS94] J. M. G. Taylor, W. G. Cumberland, and J. P. Sy. A stochastic model for analysis of longitudinal AIDS data. Journal of the American Statistical Association, 89(427):727–736, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290898.

Tsiatis:1995:MRS

[TDW95] A. A. Tsiatis, Victor DeGruttola, and M. S. Wulfsohn. Modeling the relationship of survival to longitudinal data measured with error. applications to survival and CD4 counts in patients with AIDS. Journal of the American Statistical Association, 90(429): 27–37, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291126.

Teitel:1990:BRB

[Tei90] Robert F. Teitel. Book review: [untitled]. Journal of the American Statistical Association, 85(410):605–606, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289827.

Terrell:1990:MSP

[Ter90] George R. Terrell. The maximal smoothing principle in density estimation. Journal of the American Statistical Association, 85 (410):470–477, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289786.

Terasvirta:1994:SEE

[Ter94] Timo Terasvirta. Specification, estimation, and evaluation of smooth transition autoregressive models. Journal of the American Statistical Association, 89(425):208–218, March 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291217.

Tingley:1990:SSC

[TF90] Maureen Tingley and Christopher Field. Small-sample confidence intervals. Journal of the American Statistical Association, 85 (410):427–434, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289779.

Tamura:1994:CSA

[TFAH94] Roy N. Tamura, Douglas E. Faries, John S. Andersen, and John H. Heiligenstein. A case study of an adaptive clinical trial in the treatment of out-patients with depressive disorder. Journal of the American Statistical Association, 89(427):768–776, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290902.

Tarter:1990:MNS

[TFP90] Michael E. Tarter, William R. Freeman, and Lincoln Polissar. Modular nonparametric subsurvival estimation. Journal of the American Statistical Association, 85(409):29–37, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289522.

Theodossiou:1993:PSM

[The93a] Panayiotis T. Theodossiou. Predicting shifts in the mean of a multivariate time series process: An application in predicting business failures. Journal of the American Statistical Association, 88(422): 441–449, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290323.

Therneau:1993:BRB

[The93b] Terry M. Therneau. Book review: [untitled]. Journal of the American Statistical Association, 88(422):698–699, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290358.

Theberge:1999:ECE

[Thé99] Alain Théberge. Extensions of calibration estimators in survey sampling. Journal of the American Statistical Association, 94 (446):635–644, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2670183.

Thisted:1994:BRB

[Thi94] Ronald A. Thisted. Book review: [untitled]. Journal of the American Statistical Association, 89(426):719–720, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290889.

Thomas:1990:ICR

[Tho90a] William Thomas. Influence on confidence regions for regression coefficients in generalized linear models. Journal of the American Statistical Association, 85(410):393–397, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289775.

Thompson:1990:BRB

[Tho90b] Howard E. Thompson. Book review: [untitled]. Journal of the American Statistical Association, 85(411):906, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290043.

Thompson:1990:ACS

[Tho90c] Steven K. Thompson. Adaptive cluster sampling. Journal of the American Statistical Association, 85(412):1050–1059, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289601.

Thomas:1991:IDC

[Tho91a] William Thomas. Influence diagnostics for the cross-validated smoothing parameter in spline smoothing. Journal of the American Statistical Association, 86(415):693–698, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290399.

Thompson:1991:UAR

[Tho91b] G. L. Thompson. A unified approach to rank tests for multivariate and repeated measures designs. Journal of the American Statistical Association, 86(414):410–419, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290586.

Tiwari:1999:BRB

[Tiw99] Ram Tiwari. Book review: [untitled]. Journal of the American Statistical Association, 94(445):347–348, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669720.

Tuljapurkar:1995:CID

[TL95a] Shripad Tuljapurkar and Ronald Lee. Comment: Inference from a deterministic population dynamics model for bowhead whales. Journal of the American Statistical Association, 90(430):424–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tuljapurkar:1995:IDP

[TL95b] Shripad Tuljapurkar and Ronald Lee. Inference from a deterministic population dynamics model for bowhead whales: Comment. Journal of the American Statistical Association, 90(430):424–425, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291051.

Tang:1997:ALR

[TL97] Dei-In Tang and Shang P. Lin. An approximate likelihood ratio test for comparing several treatments to a control. Journal of the American Statistical Association, 92(439):1155– 1162, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/sept/tang.htm;   http:/
/www.jstor.org/stable/2965581.

Tu:1993:AEE

[TMP93] Xin Ming Tu, Xiao-Li Meng, and Marcello Pagano. The AIDS epidemic: Estimating survival after AIDS diagnosis from surveillance data. Journal of the American Statistical Association, 88 (421):26–36, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290688.

Toman:1994:BOD

[Tom94] Blaza Toman. Bayes optimal designs for two- and three-level factorial experiments. Journal of the American Statistical Association, 89(427):937–946, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290919.

Toman:1996:BED

[Tom96] Blaza Toman. Bayesian experimental design for multiple hypothesis testing. Journal of the American Statistical Association, 91 (433):185–190, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291394.

Tourangeau:1990:ITF

[Tou90] Roger Tourangeau. Interactional troubles in face-to-face survey interviews: Comment. Journal of the American Statistical Association, 85(409):250–251, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289555.

Tsou:1995:RL

[TR95] Tsung-Shan Tsou and Richard M. Royall. Robust likelihoods. Journal of the American Statistical Association, 90(429):316– 320, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291156.

Tam:1997:TSM

[TR97] Wing Kuen Tam and Gregory C. Reinsel. Tests for seasonal moving average unit root in ARIMA models. Journal of the American Statistical Association, 92(438):725–738, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965721.

Troendle:1995:SRM

[Tro95a] James F. Troendle. A stepwise resampling method of multiple hypothesis testing. Journal of the American Statistical Association, 90(429):370–378, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291163.

Trout:1995:BRB

[Tro95b] J. Richard Trout. Book review: [untitled]. Journal of the American Statistical Association, 90(429):396, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291182.

Thombs:1990:BPI

[TS90] Lori A. Thombs and William R. Schucany. Bootstrap prediction intervals for autoregression. Journal of the American Statistical Association, 85(410):486–492, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289788.

Taylor:1992:HBM

[TS92] Jeremy M. G. Taylor and Mark R. Segal. Hierarchical Bayes models for the progression of HIV infection using longitudinal CD4 T-cell numbers: Comment. Journal of the American Statistical Association, 87(419):628–631, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290196.

Treder:1996:BST

[TS96] R. P. Treder and J. Sedransk. Bayesian sequential two-phase sampling. Journal of the American Statistical Association, 91(434): 782–790, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291673.

Tsay:1990:BRB

[Tsa90] Ruey S. Tsay. Book review: [untitled]. Journal of the American Statistical Association, 85(409):258–259, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289561.

Tsay:1998:TMM

[Tsa98] Ruey S. Tsay. Testing and modeling multivariate threshold models. Journal of the American Statistical Association, 93(443): 1188–1202, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/sept/tsay.html;   http:
//www.jstor.org/stable/2669861.

Tsutakawa:1994:BRB

[Tsu94] Robert K. Tsutakawa. Book review: [untitled]. Journal of the American Statistical Association, 89(425):354, March 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291233.

Turner:1990:BRB

[Tur90] Margery Austin Turner. Book review: [untitled]. Journal of the American Statistical Association, 85(411):904–905, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290039.

Turmon:1998:BRB

[Tur98] Michael J. Turmon. Book review: [untitled]. Journal of the American Statistical Association, 93(442):833–834, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670132.

Tebaldi:1998:BINa

[TW98a] Claudia Tebaldi and Mike West. Bayesian inference on network traffic using link count data. Journal of the American Statistical Association, 93(442):557–573, June 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/june/TEBALDI.HTM;   http://www.jstor.org/

stable/2670105. With comments by Y. Vardi and Robert Mc-Culloch, and a rejoinder by the authors.

Tebaldi:1998:BINb

[TW98b] Claudia Tebaldi and Mike West. Bayesian inference on network traffic using link count data: Rejoinder. Journal of the American Statistical Association, 93(442):576, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670108.

Tebaldi:1998:RBI

[TW98c] Claudia Tebaldi and Mike West. Rejoinder: Bayesian inference on network traffic using link count data. Journal of the American Statistical Association, 93(442):576–??, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Uebersax:1993:SME

[Ueb93] John S. Uebersax. Statistical modeling of expert ratings on medical treatment appropriateness. Journal of the American Statistical Association, 88(422):421–427, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290320.

Utts:1993:BRB

[Utt93] Jessica Utts. Book review: [untitled]. Journal of the American Statistical Association, 88(422):695, June 1993. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2290354.

Utts:1994:BRB

[Utt94] Jessica Utts. Book review: [untitled]. Journal of the American Statistical Association, 89(426):715, June 1994. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2290883.

Valliant:1990:CCP

[Val90] Richard Valliant. Correction: Conditional properties of some estimators in stratified sampling. Journal of the American Statistical Association, 85(409):272, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289583.

Valliant:1993:PCV

[Val93] Richard Valliant. Poststratification and conditional variance estimation. Journal of the American Statistical Association, 88(421): 89–96, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290701.

Vardi:1996:NTE

[Var96] Y. Vardi. Network tomography: Estimating source-destination traffic intensities from link data. Journal of the American Statistical Association, 91(433):365–377, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291416.

Vardi:1998:BIN

[Var98a] Y. Vardi. Bayesian inference on network traffic using link count data: Comment. Journal of the American Statistical Association, 93(442):573–574, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2670106.

Vardi:1998:CBI

[Var98b] Y. Vardi. Comment: Bayesian inference on network traffic using link count data. Journal of the American Statistical Association, 93(442):573–??, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vaupel:1990:BRB

[Vau90] James W. Vaupel. Book review: [untitled]. Journal of the American Statistical Association, 85(411):905, September 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290041.

VanBrummelen:1997:DIH

[VB97] Glen Van Brummelen and Kenneth Butler. Determining the interdependence of historical astronomical tables. Journal of the American Statistical Association, 92(437):41–48, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/march/vanbrumm.htm;   http://www.jstor.org/
stable/2291448.

vanderLaan:1999:AND

[vdL99] Mark van der Laan. Adjusting for nonignorable drop-out using semiparametric nonresponse models: Comment. Journal of the American Statistical Association, 94(448):1125–1127, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669926.

Velilla:1995:DRE

[Vel95] Santiago Velilla. Diagnostics and robust estimation in multivariate data transformations. Journal of the American Statistical Association, 90(431):945–951, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291329.

Velilla:1998:ANL

[Vel98a] Santiago Velilla. Assessing the number of linear components in a general regression problem. Journal of the American Statistical Association, 93(443):1088–1098, September 1998. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_98/sept/
velilla.html;   http://www.jstor.org/stable/2669852.

Velleman:1998:BRB

[Vel98b] Paul F. Velleman. Book review: [untitled]. Journal of the American Statistical Association, 93(444):1526, December 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2670074.

Velleman:1999:BRB

[Vel99] Paul F. Velleman. Book review: [untitled]. Journal of the American Statistical Association, 94(447):980–981, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670013.

Verrill:1993:PSS

[Ver93] Steve Verrill. Predictor sort sampling, tight T ’s, and the analysis of covariance. Journal of the American Statistical Association, 88 (421):119–124, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290704.

Victoria-Feser:1997:REG

[VFR97] Maria-Pia Victoria-Feser and Elvezio Ronchetti. Robust estimation for grouped data. Journal of the American Statistical Association, 92(437):333–340, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/march/
victoria.htm;   http://www.jstor.org/stable/2291478.

Vollset:1991:FCE

[VHE91] Stein E. Vollset, Karim F. Hirji, and Robert M. Elashoff. Fast computation of exact confidence limits for the common odds ratio in a series of 2 × 2 tables. Journal of the American Statistical Association, 86(414):404–409, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290585.

Vidakovic:1998:NWS

[Vid98] Brani Vidakovic. Nonlinear wavelet shrinkage with Bayes rules and Bayes factors. Journal of the American Statistical Association, 93(441):173–179, March 1998. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
VIDAKOVI.HTM;   http://www.jstor.org/stable/2669614.

Villegas:1990:BIM

[Vil90] C. Villegas. Bayesian inference in models with Euclidean structures. Journal of the American Statistical Association, 85(412): 1159–1164, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289616.

Vining:1993:BRB

[Vin93] Geoff Vining. Book review: [untitled]. Journal of the American Statistical Association, 88(422):704–705, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290368.

Verdinelli:1992:BDM

[VK92] Isabella Verdinelli and Joseph B. Kadane. Bayesian designs for maximizing information and outcome. Journal of the American Statistical Association, 87(418):510–515, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290284.

Verbeke:1996:LME

[VL96] Geert Verbeke and Emmanuel Lesaffre. A linear mixed-effects model with heterogeneity in the random-effects population. Journal of the American Statistical Association, 91(433):217–221, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291398.

Vu:1996:LRT

[VM96] H. T. V. Vu and R. A. Maller. The likelihood ratio test for Poisson versus binomial distributions. Journal of the American Statistical Association, 91(434):818–824, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291676.

Vogel:1992:BRB

[Vog92] Frederic A. Vogel. Book review: [untitled]. Journal of the American Statistical Association, 87(420):1245–1246, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290676.

Vogelsang:1999:BRB

[Vog99] Timothy J. Vogelsang. Book review: [untitled]. Journal of the American Statistical Association, 94(445):344–345, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669716.

Voss:1993:CRS

[Vos93] Daniel T. Voss. Characterization of regularity for single replicate factorial designs. Journal of the American Statistical Association, 88(423):924–925, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290783.

vanderLaan:1998:LEE

[vR98] Mark J. van der Laan and James M. Robins. Locally efficient estimation with current status data and time-dependent covariates. Journal of the American Statistical Association, 93(442):693–701, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670119.

VanHouwelingen:1993:MEB

[VS93] Hans C. Van Houwelingen and Theo Stijnen. Monotone empirical Bayes estimators based on more informative samples. Journal of the American Statistical Association, 88(424):1438–1443, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291289.

Verdinelli:1995:CBF

[VW95] Isabella Verdinelli and Larry Wasserman. Computing Bayes factors using a generalization of the Savage-Dickey density ratio. Journal of the American Statistical Association, 90(430):614–618, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291073.

Wachter:1993:HLR

[Wac93] Kenneth W. Wachter. Hierarchical logistic regression models for imputation of unresolved enumeration status in undercount estimation: Comment: Ignoring nonignorable effects. Journal of the American Statistical Association, 88(423):1161–1163, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290814.

Waksberg:1991:BRB

[Wak91] Joseph Waksberg. Book review: [untitled]. Journal of the American Statistical Association, 86(414):550, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290616.

Wakefield:1996:BAP

[Wak96] Jon Wakefield. The Bayesian analysis of population pharmacokinetic models. Journal of the American Statistical Association, 91 (433):62–75, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291383.

Walke:1991:BRB

[Wal91] Roger Walke. Book review: [untitled]. Journal of the American Statistical Association, 86(413):240–241, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289735.

Wallman:1993:ESL

[Wal93] Katherine K. Wallman. Enhancing statistical literacy: Enriching our society. Journal of the American Statistical Association, 88(421):1–8, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290686.

Walker:1995:BRB

[Wal95] Esteban Walker. Book review: [untitled]. Journal of the American Statistical Association, 90(431):1127, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291356.

Walker:1996:BRB

[Wal96] Esteban Walker. Book review: [untitled]. Journal of the American Statistical Association, 91(433):434–435, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291429.

Wall:1998:BRB

[Wal98] Kent D. Wall. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1248–1249, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669890.

Wang:1991:NEC

[Wan91] Mei-Cheng Wang. Nonparametric estimation from cross-sectional survival data. Journal of the American Statistical Association, 86 (413):130–143, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289722.

Wang:1992:ARA

[Wan92] Mei-Cheng Wang. The analysis of retrospectively ascertained data in the presence of reporting delays. Journal of the American Statistical Association, 87(418):397–406, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290270.

Wang:1995:SLO

[Wan95] Julian Z. Wang. Selection of the K largest order statistics for the domain of attraction of the Gumbel distribution. Journal of the American Statistical Association, 90(431):1055–1061, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291342.

Wang:1996:LRT

[Wan96] Yazhen Wang. A likelihood ratio test against stochastic ordering in several populations. Journal of the American Statistical Association, 91(436):1676–1683, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291595.

Wang:1997:OUT

[Wan97] Weizhen Wang. Optimal unbiased tests for equivalence in intrasubject variability. Journal of the American Statistical Association, 92(439):1163–1170, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/sept/
wang_w.htm;   http://www.jstor.org/stable/2965582.

Wang:1998:CCE

[Wan98a] Yazhen Wang. Change curve estimation via wavelets. Journal of the American Statistical Association, 93(441):163–172, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_98/WANG.HTM;   http://www.jstor.org/stable/
2669613.

Wang:1998:SSMa

[Wan98b] Yuedong Wang. Smoothing spline models with correlated random errors. Journal of the American Statistical Association, 93(441):341–348, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/WANG_Y.HTM;   http:
//www.jstor.org/stable/2669630.

Wang:1999:TIB

[Wan99] Weizhen Wang. On testing of individual bioequivalence. Journal of the American Statistical Association, 94(447):880–887, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670003.

Ware:1991:BRB

[War91] James H. Ware. Book review: [untitled]. Journal of the American Statistical Association, 86(413):250, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289748.

Warner:1997:BRB

[War97] Bradley A. Warner. Book review: [untitled]. Journal of the American Statistical Association, 92(438):791–792, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965731.

Wasserman:1990:BRB

[Was90] Larry Wasserman. Book review: [untitled]. Journal of the American Statistical Association, 85(410):597, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289814.

Wasserman:1991:BRB

[Was91] Larry Wasserman. Book review: [untitled]. Journal of the American Statistical Association, 86(414):546, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290609.

Wasserman:1993:BRB

[Was93] Larry Wasserman. Book review: [untitled]. Journal of the American Statistical Association, 88(422):700–702, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290362.

Wasserman:1994:BRB

[Was94] Stanley Wasserman. Book review: [untitled]. Journal of the American Statistical Association, 89(426):715–717, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290884.

Wassell:1997:BRB

[Was97] James T. Wassell. Book review: [untitled]. Journal of the American Statistical Association, 92(439):1225, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965607.

Wassell:1998:BRBa

[Was98a] James T. Wassell. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1241, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2669877.

Wassell:1998:BRBb

[Was98b] James T. Wassell. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1241–1242, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2669878.

Wasserman:1999:ECE

[Was99a] Larry Wasserman. Estimation of the causal effect of a timevarying exposure on the marginal mean of a repeated binary outcome: Comment. Journal of the American Statistical Association, 94(447):704–706, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669981.

Wasserman:1999:BRB

[Was99b] Stanley Wasserman. Book review: [untitled]. Journal of the American Statistical Association, 94(448):1384–1385, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669954.

Waternaux:1990:BRB

[Wat90] Christine Waternaux. Book review: [untitled]. Journal of the American Statistical Association, 85(410):600, June 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289820.

Wakefield:1996:BMC

[WB96] Jon Wakefield and James Bennett. The Bayesian modeling of covariates for population pharmacokinetic models. Journal of the American Statistical Association, 91(435):917–927, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291710.

Wood:1993:SAC

[WBB93] Andrew T. A. Wood, James G. Booth, and Ronald W. Butler. Saddlepoint approximations to the CDF of some statistics with nonnormal limit distributions. Journal of the American Statistical Association, 88(422):680–686, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290351.

Wolter:1991:EPI

[WC91] Kirk M. Wolter and Beverley D. Causey. Evaluation of procedures for improving population estimates for small areas. Journal of the American Statistical Association, 86(414):278–284, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290559.

Windham:1992:IRV

[WC92] Michael P. Windham and Adele Cutler. Information ratios for validating mixture analyses. Journal of the American Statistical Association, 87(420):1188–1192, December 1992. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2290659.

Wang:1999:NER

[WC99] Mei-Cheng Wang and Shu-Hui Chang. Nonparametric estimation of a recurrent survival function. Journal of the American Statistical Association, 94(445):146–153, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669690.

Wu:1998:ACR

[WCH98] Colin O. Wu, Chin-Tsang Chiang, and Donald R. Hoover. Asymptotic confidence regions for kernel smoothing of a varyingcoefficient model with longitudinal data. Journal of the American Statistical Association, 93(444):1388–1402, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_98/dec/wu.html;   http://www.jstor.org/
stable/2670054.

Welsh:1994:FHR

[WCR94] A. H. Welsh, R. J. Carroll, and D. Ruppert. Fitting heteroscedastic regression models. Journal of the American Statistical Association, 89(425):100–116, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2291205.

Waller:1997:HST

[WCXG97] Lance A. Waller, Bradley P. Carlin, Hong Xia, and Alan E. Gelfand. Hierarchical spatio-temporal mapping of disease rates. Journal of the American Statistical Association, 92(438):607–617, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2965708.

Weerahandi:1991:TVC

[Wee91] Samaradasa Weerahandi. Testing variance components in mixed models with generalized p values. Journal of the American Statistical Association, 86(413):151–153, March 1991. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289724.

Weerahandi:1993:GCI

[Wee93] Samaradasa Weerahandi. Generalized confidence intervals. Journal of the American Statistical Association, 88(423):899–905, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2290779. See correction [Wee94].

Weerahaudi:1994:CGC

[Wee94] Samaradasa Weerahaudi. Correction: “Generalized confidence intervals” [J. Amer. Statist. Assoc. 88 (1993), no. 423, 899–905, MR 94e:62031]. Journal of the American Statistical Association, 89 (426):726, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2290895. See [Wee93].

Wegge:1990:BRB

[Weg90a] Leon L. Wegge. Book review: [untitled]. Journal of the American Statistical Association, 85(409):255–256, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289558.

Wegman:1990:HDA

[Weg90b] Edward J. Wegman. Hyperdimensional data analysis using parallel coordinates. Journal of the American Statistical Association, 85(411):664–675, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290001.

Wegman:1991:FBM

[Weg91] Edward J. Wegman. A focus on Bayesian methods: [introduction]. Journal of the American Statistical Association, 86(416): 923, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290507.

Wehrly:1994:BRB

[Weh94] Thomas E. Wehrly. Book review: [untitled]. Journal of the American Statistical Association, 89(428):1567, December 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291030.

Wei:1990:BRB

[Wei90a] William W. S. Wei. Book review: [untitled]. Journal of the American Statistical Association, 85(409):258, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289560.

Weiss:1990:BRB

[Wei90b] Robert E. Weiss. Book review: [untitled]. Journal of the American Statistical Association, 85(412):1172–1173, December 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289627.

Wei:1991:BRB

[Wei91a] William W. S. Wei. Book review: [untitled]. Journal of the American Statistical Association, 86(413):246–247, March 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289742.

Weisberg:1991:BRB

[Wei91b] Sanford Weisberg. Book review: [untitled]. Journal of the American Statistical Association, 86(416):1142, December 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290540.

Weiss:1992:EHU

[Wei92] Andrew A. Weiss. The effects of helmet use on the severity of head injuries in motorcycle accidents. Journal of the American Statistical Association, 87(417):48–56, March 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290450. See comments

[Bro92b, Bro92a].

Weiss:1994:BRB

[Wei94] Robert Weiss. Book review: [untitled]. Journal of the American Statistical Association, 89(425):359–360, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291242.

Weiss:1995:BRB

[Wei95a] Robert E. Weiss. Book review: [untitled]. Journal of the American Statistical Association, 90(432):1488–1489, December 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2291543.

Weiss:1995:IVS

[Wei95b] Robert E. Weiss. The influence of variable selection: a Bayesian diagnostic perspective. Journal of the American Statistical Association, 90(430):619–625, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2291074.

Weir:1997:FBR

[Wei97] Iain S. Weir. Fully Bayesian reconstructions from single-photon emission computed tomography data. Journal of the American Statistical Association, 92(437):49–60, March 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/march/
weir.htm;   http://www.jstor.org/stable/2291449.

Weir:1998:BRB

[Wei98] Bruce S. Weir. Book review: [untitled]. Journal of the American Statistical Association, 93(443):1250, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669893.

Welch:1990:CPT

[Wel90a] William J. Welch. Construction of permutation tests. Journal of the American Statistical Association, 85(411):693–698, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290004.

Wells:1990:REG

[Wel90b] Martin T. Wells. The relative efficiency of goodness-of-fit statistics in the simple and composite hypothesis-testing problem. Journal of the American Statistical Association, 85(410):459–463, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289784.

Welsh:1993:IMO

[Wel93] A. H. Welsh. The identification of multiple outliers: Comment. Journal of the American Statistical Association, 88(423):796–797, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290767.

Welsh:1994:BRB

[Wel94] A. H. Welsh. Book review: [untitled]. Journal of the American Statistical Association, 89(427):1136–1138, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290945.

Wells:1999:BRB

[Wel99] Martin T. Wells. Book review: [untitled]. Journal of the American Statistical Association, 94(445):348–349, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669723.

West:1995:BIC

[Wes95] Mike West. Bayesian inference in cyclical component dynamic linear models. Journal of the American Statistical Association, 90 (432):1301–1312, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291520.

West:1997:HMM

[Wes97a] Mike West. Hierarchical mixture models in neurological transmission analysis. Journal of the American Statistical Association, 92 (438):587–606, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2965707.

Westfall:1997:MTG

[Wes97b] Peter H. Westfall. Multiple testing of general contrasts’ using logical constraints and correlations. Journal of the American Statistical Association, 92(437):299–306, March 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/march/
westfall.htm;   http://www.jstor.org/stable/2291474.

Westfall:1998:AMT

[Wes98] Peter Westfall. Addendum: “Multiple testing of general contrasts using logical constraints and correlations” [J. Amer. Statist. Assoc. 92 (1997), no. 437, 299–306; MR 98c:62053]. Journal of the American Statistical Association, 93(441):413, 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Worsley:1991:LSC

[WEST91] K. J. Worsley, A. C. Evans, S. C. Strother, and J. L. Tyler. A linear spatial correlation model, with applications to positron emission tomography. Journal of the American Statistical Association, 86(413):55–67, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289715.

Wang:1990:TSI

[WH90] Jane-Ling Wang and Thomas P. Hettmansperger. Two-sample inference for median survival times based on one-sample procedures for censored survival data. Journal of the American Statistical Association, 85(410):529–536, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289794.

Whittemore:1998:TCS

[WHG98] Alice S. Whittemore, Jerry Halpern, and Gail Gong. Testing covariance structure in multivariate models: Application to family disease data. Journal of the American Statistical Association, 93(442):518–525, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/june/WHITTEMO.HTM;
 http://www.jstor.org/stable/2670099.

Wolfe:1996:WPS

[WHGA96] Barbara Wolfe, Robert Haveman, Donna Ginther, and Chong Bum An. The “window problem” in studies of Children’s attainments: a methodological exploration. Journal of the American Statistical Association, 91(435):970–982, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291716.

White:1989:SAR

[Whi89] Halbert White. Some asymptotic results for learning in single hidden layer feedforward network models. Journal of the American Statistical Association, 84(408):1003–1013, December 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290076. See comment

[Fab94]. See correction [Whi92b].

White:1992:CSA

[Whi92a] Halbert White. Correction: Some asymptotic results for learning in single hidden-layer feedforward network models. Journal of the American Statistical Association, 87(420):1252, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290683.

White:1992:SAR

[Whi92b] Halbert White. “Some asymptotic results for learning in single hidden-layer feedforward network models” [J. Amer. Statist. Assoc. 84 (1989), no. 408, 1003–1013, MR 92e:62119]. Journal of the American Statistical Association, 87(420):1252, ???? 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Whi89, Fab94].

Wiens:1998:MRD

[Wie98] Douglas P. Wiens. Minimax robust designs and weights for approximately specified regression models with heteroscedastic errors. Journal of the American Statistical Association, 93(444): 1440–1450, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/dec/wiens.html;   http:
//www.jstor.org/stable/2670058.

Williams:1991:BRB

[Wil91] William A. Williams. Book review: [untitled]. Journal of the American Statistical Association, 86(415):834, September 1991.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290435.

Wilcox:1999:CSM

[Wil99] Rand R. Wilcox. Comment on stute, manteiga, and quindimil. Journal of the American Statistical Association, 94(446):659–660, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670212.

Winship:1996:BRB

[Win96] Christopher Winship. Book review: [untitled]. Journal of the American Statistical Association, 91(435):1373–1374, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291756.

Weerakkody:1992:EWM

[WJ92] Govinda J. Weerakkody and Dallas E. Johnson. Estimation of within model parameters in regression models with a nested error structure. Journal of the American Statistical Association, 87 (419):708–713, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290208.

Wand:1993:CSP

[WJ93] M. P. Wand and M. C. Jones. Comparison of smoothing parameterizations in bivariate kernel density estimation. Journal of the American Statistical Association, 88(422):520–528, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290332.

Wasserman:1992:CBE

[WK92] Larry Wasserman and Joseph B. Kadane. Computing bounds on expectations. Journal of the American Statistical Association, 87 (418):516–522, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290285.

Wang:1994:SVC

[WK94] Morgan C. Wang and William J. Kennedy. Self-validating computations of probabilities for selected central and noncentral univariate probability functions. Journal of the American Statistical Association, 89(427):878–887, September 1994. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290912.

Wood:1998:BAR

[WK98] Sally Wood and Robert Kohn. A Bayesian approach to robust binary nonparametric regression. Journal of the American Statistical Association, 93(441):203–213, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669617.

Williamson:1995:ABO

[WKL95] John M. Williamson, KyungMann Kim, and Stuart R. Lipsitz. Analyzing bivariate ordinal data using a global odds ratio. Journal of the American Statistical Association, 90(432):1432–1437, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291535.

Waclawiw:1993:PRE

[WL93] Myron A. Waclawiw and Kung-Yee Liang. Prediction of random effects in the generalized linear model. Journal of the American Statistical Association, 88(421):171–178, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290711.

Wang:1998:BAS

[WLGC98] Naisyin Wang, Xihong Lin, Roberto G. Gutierrez, and Raymond J. Carroll. Bias analysis and SIMEX approach in generalized linear mixed measurement error models. Journal of the American Statistical Association, 93(441):249–261, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.amstat.org/publications/
jasa/abstracts_98/WANG_N.HTM;   http://www.jstor.org/
stable/2669621.

Welsh:1990:RES

[WM90] A. H. Welsh and H. L. Morrison. Robust L estimation of scale with an application in astronomy. Journal of the American Statistical Association, 85(411):729–743, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290009.

Wong:1991:CSE

[WM91] George Y. Wong and William M. Mason. Contextually specific effects and other generalizations of the hierarchical linear model for comparative analysis. Journal of the American Statistical Association, 86(414):487–503, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290601.

Wang:1998:CLR

[WM98] Yining Wang and Michael P. McDermott. Conditional likelihood ratio test for a nonnegative normal mean vector. Journal of the American Statistical Association, 93(441):380– 386, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/WANG_YIN.HTM;   http://
www.jstor.org/stable/2669634.

Wand:1991:TDEb

[WMR91a] M. P. Wand, J. S. Marron, and D. Ruppert. Transformations in density estimation: Rejoinder. Journal of the American Statistical Association, 86(414):360–361, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290575.

Wand:1991:TDEa

[WMR91b] M. P. Wand, J. S. Marron, and D. Ruppert. Transformations in density estimation. With discussion and a rejoinder by the authors. Journal of the American Statistical Association, 86(414): 343–353, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290569.

Wolf:1990:BRB

[Wol90] Frank L. Wolf. Book review: [untitled]. Journal of the American Statistical Association, 85(409):263–264, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289570.

Wolter:1991:BRB

[Wol91] Kirk M. Wolter. Book review: [untitled]. Journal of the American Statistical Association, 86(413):250–251, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289749.

Wolpin:1993:BRB

[Wol93] Kenneth I. Wolpin. Book review: [untitled]. Journal of the American Statistical Association, 88(424):1475–1476, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291309.

Wolpert:1995:CID

[Wol95a] Robert L. Wolpert. Comment: Inference from a deterministic population dynamics model for bowhead whales. Journal of the American Statistical Association, 90(430):426–??, ???? 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wolpert:1995:IDP

[Wol95b] Robert L. Wolpert. Inference from a deterministic population dynamics model for bowhead whales: Comment. Journal of the American Statistical Association, 90(430):426–427, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291052.

Wong:1993:BRB

[Won93] Weng Kee Wong. Book review: [untitled]. Journal of the American Statistical Association, 88(421):383–384, March 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290747.

Wong:1996:BRB

[Won96] Weng Kee Wong. Book review: [untitled]. Journal of the American Statistical Association, 91(434):909–910, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291688.

Wood:1990:BRB

[Woo90] Kevin Wood. Book review: [untitled]. Journal of the American Statistical Association, 85(409):265–266, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289574.

Wood:1991:BRB

[Woo91] Graham R. Wood. Book review: [untitled]. Journal of the American Statistical Association, 86(415):835–836, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290438.

Woolson:1993:BRB

[Woo93] Robert F. Woolson. Book review: [untitled]. Journal of the American Statistical Association, 88(421):382–383, March 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290745.

Wainer:1992:BRB

[WP92] Howard Wainer and Peter Pashley. Book review: [untitled]. Journal of the American Statistical Association, 87(419):899–900, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290235.

Waternaux:1998:CPU

[WPD98] Christine Waternaux, Eva Petkova, and William DuMouchel. Comment: Problems with using auxiliary information to correct for omitted variables when estimating the effect of lead on IQ. Journal of the American Statistical Association, 93(442):505–513, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2670096.

West:1999:ECEa

[WPK99a] Mike West, Raquel Prado, and Andrew D. Krystal. Evaluation and comparison of EEG traces: Latent structure in nonstationary time series. Journal of the American Statistical Association, 94(446):375–387, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/west.pdf;
 http://www.jstor.org/stable/2670154.

West:1999:ECEb

[WPK99b] Mike West, Raquel Prado, and Andrew D. Krystal. Evaluation and comparison of EEG traces: Latent structure in nonstationary time series. Journal of the American Statistical Association, 94 (448):1083–1095, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2669922.

Woodruff:1994:CRE

[WR94] David L. Woodruff and David M. Rocke. Computable robust estimation of multivariate location and shape in high dimension using compound estimators. Journal of the American Statistical Association, 89(427):888–896, September 1994. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290913.

Wang:1995:NET

[WR95] Naisyin Wang and David Ruppert. Nonparametric estimation of the transformation in the transform-both-sides regression model. Journal of the American Statistical Association, 90(430):522–534, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291063.

Wachter:1991:TEP

[WS91] Kenneth W. Wachter and Terence P. Speed. Total error in PES estimates of population: Comment. Journal of the American Statistical Association, 86(416):858–861, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290497.

Wang:1994:MRN

[WS94] Ferdinand T. Wang and David W. Scott. The L_{1} method for robust nonparametric regression. Journal of the American Statistical Association, 89(425):65–76, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291201.

Wendell:1996:EIP

[WS96] John P. Wendell and Josef Schmee. Exact inference for proportions from a stratified finite population. Journal of the American Statistical Association, 91(434):825–830, June 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291677.

Wang:1992:SDA

[WSJ92] R. Wang, J. Sedransk, and J. H. Jinn. Secondary data analysis when there are missing observations. Journal of the American Statistical Association, 87(420):952–961, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290631.

Wei:1990:SID

[WSLP90] L. J. Wei, R. T. Smythe, D. Y. Lin, and T. S. Park. Statistical inference with data-dependent treatment allocation rules. Journal of the American Statistical Association, 85(409):156–162, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289538.

Wei:1990:MCI

[WT90a] Greg C. G. Wei and Martin A. Tanner. A Monte Carlo implementation of the EM algorithm and the poor Man’s data augmentation algorithms. Journal of the American Statistical Association, 85(411):699–704, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290005.

Wei:1990:PCC

[WT90b] Greg C. G. Wei and Martin A. Tanner. Posterior computations for censored regression data. Journal of the American Statistical Association, 85(411):829–839, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290022.

Wu:1998:BRB

[Wu98] John Chih-Hang Wu. Book review: [untitled]. Journal of the American Statistical Association, 93(442):842–843, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670144.

Wuensch:1996:BRB

[Wue96] Karl L. Wuensch. Book review: [untitled]. Journal of the American Statistical Association, 91(436):1750–1751, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291607.

Wang:1991:ACA

[WW91a] J. C. Wang and C. F. J. Wu. An approach to the construction of asymmetrical orthogonal arrays. Journal of the American Statistical Association, 86(414):450–456, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290593.

West:1991:DBE

[WW91b] Mike West and Robert L. Winkler. Data base error trapping and prediction. Journal of the American Statistical Association, 86 (416):987–996, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290515.

Walden:1993:DBT

[WW93] Andrew T. Walden and Mark L. Williams. Deconvolution, bandwidth, and the trispectrum. Journal of the American Statistical Association, 88(424):1323–1329, December 1993. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2291273.

Wang:1998:CSS

[WW98a] Yuedong Wang and Grace Wahba. Comment: Smoothing spline models for the analysis of nested and crossed samples of curves. Journal of the American Statistical Association, 93(443):976–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:1998:SSMb

[WW98b] Yuedong Wang and Grace Wahba. Smoothing spline models for the analysis of nested and crossed samples of curves: Comment. Journal of the American Statistical Association, 93(443):976–980, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2669838.

Wang:1997:WSE

[WWZtO97] C. Y. Wang, Suojin Wang, Lue Ping Zhao, and Shyh tyan Ou. Weighted semiparametric estimation in regression analysis with missing covariate data. Journal of the American Statistical Association, 92(438):512–525, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2965700.

Wells:1996:DEB

[WY96] Martin T. Wells and Kwee Poo Yeo. Density estimation with bivariate censored data. Journal of the American Statistical Association, 91(436):1566–1574, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291582.

Wellner:1997:HAC

[WZ97a] Jon A. Wellner and Yihui Zhan. A hybrid algorithm for computation of the nonparametric maximum likelihood estimator from censored data. Journal of the American Statistical Association, 92(439):945–959, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/sept/wellner.htm;
http://www.jstor.org/stable/2965558.

Wiens:1997:RDB

[WZ97b] Douglas P. Wiens and Julie Zhou. Robust designs based on the infinitesimal approach. Journal of the American Statistical Association, 92(440):1503–1511, December 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_97/
december/WIENS.HTM;   http://www.jstor.org/stable/2965421.

Xiong:1995:CSC

[Xio95] Xiaoping Xiong. A class of sequential conditional probability ratio tests. Journal of the American Statistical Association, 90(432): 1463–1473, December 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291539.

Xia:1999:SIC

[XL99] Yingcun Xia and W. K. Li. On single-index coefficient regression models. Journal of the American Statistical Association, 94(448): 1275–1285, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2669941.

Xiao-Li:1991:UEO

[XLR91] Meng Xiao-Li and Donald B. Rubin. Using EM to obtain asymptotic variance-covariance matrices: The SEM algorithm. Journal of the American Statistical Association, 86(416):899–??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xie:1992:APM

[XP92] Yu Xie and Ellen Efron Pimentel. Age patterns of marital fertility: Revising the Coale–Trussell method. Journal of the American Statistical Association, 87(420):977–984, December 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290634.

Xin:1993:AEE

[XXP93] Ming Tu Xin, Meng Xiao- Li, and Marcello Pagano. The AIDS epidemic: Estimating survival after AIDS diagnosis from surveillance data. Journal of the American Statistical Association, 88 (421):26–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yamaguchi:1990:HSD

[Yam90] Kazuo Yamaguchi. Homophily and social distance in the choice of multiple friends: An analysis based on conditionally symmetric log-bilinear association model. Journal of the American Statistical Association, 85(410):356–366, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289771.

Yamaguchi:1992:AFT

[Yam92] Kazuo Yamaguchi. Accelerated failure-time regression models with a regression model of surviving fraction: An application to the analysis of “permanent employment” in Japan. Journal of the American Statistical Association, 87(418):284–292, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290258.

Yandell:1990:BRB

[Yan90] Brian S. Yandell. Book review: [untitled]. Journal of the American Statistical Association, 85(411):905–906, September 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290042.

Yang:1997:EWL

[Yan97] Song Yang. Extended weighted log-rank estimating functions in censored regression. Journal of the American Statistical Association, 92(439):977–984, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

amstat.org/publications/jasa/abstracts_97/sept/yang_s.
htm;   http://www.jstor.org/stable/2965561.

Yandell:1998:BRB

[Yan98] Brian S. Yandell. Book review: [untitled]. Journal of the American Statistical Association, 93(442):837–838, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2670137.

Yang:1999:CMR

[Yan99] Song Yang. Censored median regression using weighted empirical survival and hazard functions. Journal of the American Statistical Association, 94(445):137–145, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2669689.

Yashchin:1995:LRM

[Yas95] Emmanuel Yashchin. Likelihood ratio methods for monitoring parameters of a nested random effect model. Journal of the American Statistical Association, 90(430):729–738, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291085.

Yuan:1997:MCS

[YB97] Ke Hai Yuan and Peter M. Bentler. Mean and covariance structure analysis: Theoretical and practical improvements. Journal of the American Statistical Association, 92(438):767–774, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2965725.

Ye:1993:RPW

[Ye93] Keying Ye. Reference priors when the stopping rule depends on the parameter of interest. Journal of the American Statistical Association, 88(421):360–363, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290732.

Ye:1998:MCE

[Ye98a] Jianming Ye. On measuring and correcting the effects of data mining and model selection. Journal of the American Statistical Association, 93(441):120–131, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.amstat.org/publications/jasa/abstracts_98/YE.HTM;
 http://www.jstor.org/stable/2669609.

Ye:1998:OCL

[Ye98b] Kenny Q. Ye. Orthogonal column Latin hypercubes and their application in computer experiments. Journal of the American Statistical Association, 93(444):1430–1439, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_98/
dec/ye.html;   http://www.jstor.org/stable/2670057.

Yanagawa:1990:HTG

[YF90] Takashi Yanagawa and Yoshinori Fujii. Homogeneity test with a generalized Mantel-Haenzel estimator for L2 × K contingency tables. Journal of the American Statistical Association, 85(411): 744–748, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290010.

Yanagawa:1995:PMM

[YF95] Takashi Yanagawa and Yoshinori Fujii. Projection-method Mantel–Haenszel estimator for K2 × J tables. Journal of the American Statistical Association, 90(430):649–656, June 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2291078.

Yu:1998:LLQ

[YJ98] Keming Yu and M. C. Jones. Local linear quantile regression. Journal of the American Statistical Association, 93(441):228–237, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.amstat.org/publications/jasa/abstracts_98/YU.HTM;
 http://www.jstor.org/stable/2669619.

Ying:1995:SAM

[YJW95] Z. Ying, S. H. Jung, and L. J. Wei. Survival analysis with median regression models. Journal of the American Statistical Association, 90(429):178–184, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2291141.

Yamaguchi:1996:PES

[YK96] Kazuo Yamaguchi and Denise B. Kandel. Parametric event sequence analysis: An application to an analysis of gender and racial/ethnic differences in patterns of drug-use progression. Journal of the American Statistical Association, 91(436):1388–1399, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291565.

Yanagawa:1997:NOA

[YKB97] Takashi Yanagawa, Yasuki Kikuchi, and Kenneth G. Brown. Noobserved-adverse-effect levels in severity data. Journal of the American Statistical Association, 92(438):449–454, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2965692.

Yasui:1997:RMS

[YL97] Yutaka Yasui and Subhash Lele. A regression method for spatial disease rates: An estimating function approach. Journal of the American Statistical Association, 92(437):21–32, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/march/yasui.htm;   http://www.jstor.org/
stable/2291446.

Ylvisaker:1993:ABB

[Ylv93] Donald Ylvisaker. Assessing between-block heterogeneity within the post-strata of the 1990 post-enumeration survey: Comment. Journal of the American Statistical Association, 88(423):1127– 1128, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290808.

Yang:1999:ITK

[YM99] Lijian Yang and James S. Marron. Iterated transformation-kernel density estimation. Journal of the American Statistical Association, 94(446):580–589, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2670178.

Young:1997:BRB

[You97] David A. Young. Book review: [untitled]. Journal of the American Statistical Association, 92(438):789–791, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965730.

Yang:1999:SIP

[YP99] Song Yang and Ross L. Prentice. Semiparametric inference in the proportional odds regression model. Journal of the American Statistical Association, 94(445):125–136, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669688.

Yap:1995:ETU

[YR95] Sook Fwe Yap and Gregory C. Reinsel. Estimation and testing for unit roots in a partially nonstationary vector autoregressive moving average model. Journal of the American Statistical Association, 90(429):253–267, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2291150.

Zajac:1993:BRB

[Zaj93] Edward E. Zajac. Book review: [untitled]. Journal of the American Statistical Association, 88(423):1171–1172, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290823.

Zaslavsky:1993:CCD

[Zas93] Alan M. Zaslavsky. Combining census, dual-system, and evaluation study data to estimate population shares. Journal of the American Statistical Association, 88(423):1092–1105, September 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290804.

Zhang:1992:BCV

[ZB92] Ji Zhang and Dennis D. Boos. Bootstrap critical values for testing homogeneity of covariance matrices. Journal of the American Statistical Association, 87(418):425–429, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290273.

Zeng:1997:BAC

[ZD97] Qi Zeng and Marie Davidian. Bootstrap-adjusted calibration confidence intervals for immunoassay. Journal of the American Statistical Association, 92(437):278–290, March 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/abstracts_97/march/
zeng.htm;   http://www.jstor.org/stable/2291472.

Zackin:1996:NME

[ZDL96] Robert Zackin, Victor De Gruttola, and Nan Laird. Nonparametric mixed-effects models for repeated binary data arising in serial dilution assays: An application to estimating viral burden in AIDS. Journal of the American Statistical Association, 91(433): 52–61, March 1996. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2291382.

Zeh:1995:BRB

[Zeh95] Judy Zeh. Book review: [untitled]. Journal of the American Statistical Association, 90(430):800–801, June 1995. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2291097.

Zellner:1992:SSP

[Zel92] Arnold Zellner. Statistics, science and public policy. Journal of the American Statistical Association, 87(417):1–6, March 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290445.

Zelterman:1993:SBT

[Zel93] Daniel Zelterman. A semiparametric bootstrap technique for simulating extreme order statistics. Journal of the American Statistical Association, 88(422):477–485, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290327.

Zerbe:1996:BRB

[Zer96] Gary Zerbe. Book review: [untitled]. Journal of the American Statistical Association, 91(436):1751, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291608.

Zerner:1999:BRB

[Zer99] Martin P. W. Zerner. Book review: [untitled]. Journal of the American Statistical Association, 94(447):989, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2670025.

Zhang:1992:DPM

[Zha92] Ping Zhang. On the distributional properties of model selection criteria. Journal of the American Statistical Association, 87(419): 732–737, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290211.

Zhang:1996:NIS

[Zha96] Ping Zhang. Nonparametric importance sampling. Journal of the American Statistical Association, 91(435):1245–1253, September 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2291743.

Zhang:1998:BCM

[Zha98a] Heping Zhang. Bayesian CART model search: Comment. Journal of the American Statistical Association, 93(443):948–950, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669833.

Zhang:1998:CTM

[Zha98b] Heping Zhang. Classification trees for multiple binary responses. Journal of the American Statistical Association, 93 (441):180–193, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/ZHANG.HTM;   http:/
/www.jstor.org/stable/2669615.

Zhang:1998:CBC

[Zha98c] Heping Zhang. Comment: Bayesian CART model search. Journal of the American Statistical Association, 93(443):945–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhao:1999:IEN

[Zha99] Linda H. Zhao. Improved estimators in nonparametric regression problems. Journal of the American Statistical Association, 94 (445):164–173, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2669692.

Zieschang:1990:SWM

[Zie90] Kimberly D. Zieschang. Sample weighting methods and estimation of totals in the consumer expenditure survey. Journal of the American Statistical Association, 85(412):986–1001, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289595.

Zimmerman:1992:BRB

[Zim92] Dale Zimmerman. Book review: [untitled]. Journal of the American Statistical Association, 87(418):589–590, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290307.

Zeger:1991:GLM

[ZK91] Scott L. Zeger and M. Rezaul Karim. Generalized linear models with random effects; A Gibbs sampling approach. Journal of the American Statistical Association, 86(413):79–86, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289717.

Zhang:1999:IED

[ZKJ99] S. Zhang, R. J. Karunamuni, and M. C. Jones. An improved estimator of the density function at the boundary. Journal of the American Statistical Association, 94(448):1231–1241, December 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2669937.

Zeger:1991:CEV

[ZL91] Scott L. Zeger and Kung-Yee Liang. Compliance as an explanatory variable in clinical trials: Comment: Dose-response estimands. Journal of the American Statistical Association, 86(413): 18–19, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289708.

Zheng:1995:CVS

[ZL95] Xiaodong Zheng and Wei-Yin Loh. Consistent variable selection in linear models. Journal of the American Statistical Association, 90(429):151–156, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2291138.

Zeger:1999:ECE

[ZL99] Scott L. Zeger and Kung-Yee Liang. Estimation of the causal effect of a time-varying exposure on the marginal mean of a repeated binary outcome: Comment. Journal of the American Statistical Association, 94(447):706–707, September 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2669982.

Zhang:1998:SSM

[ZLRS98] Daowen Zhang, Xihong Lin, Jonathan Raz, and MaryFran Sowers. Semiparametric stochastic mixed models for longitudinal data. Journal of the American Statistical Association, 93 (442):710–719, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/june/ZHANG.HTM;   http:
//www.jstor.org/stable/2670121.

Zhou:1993:CIT

[ZM93] Leping Zhou and Thomas Mathew. Combining independent tests in linear models. Journal of the American Statistical Association, 88(422):650–655, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290347.

Zellner:1995:GSC

[ZM95] Arnold Zellner and Chung-Ki Min. Gibbs sampler convergence criteria. Journal of the American Statistical Association, 90(431): 921–927, September 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291326.

Zeger:1997:EMS

[ZT97] Linda M. Zeger and Neal Thomas. Efficient matrix sampling instruments for correlated latent traits: Examples from the National Assessment of Educational Progress. Journal of the American Statistical Association, 92(438):416–425, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965689.

Zucker:1998:RML

[Zuc98] David M. Zucker. Restricted mean life with covariates: Modification and extension of a useful survival analysis method. Journal of the American Statistical Association, 93(442):702–709, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/june/ZUCKER.HTM;   http://www.jstor.org/
stable/2670120.

Zhou:1996:REM

[ZWWB96] Haibo Zhou, Clarice R. Weinberg, Allen J. Wilcox, and Donna D. Baird. A random-effects model for cycle viability in fertility studies. Journal of the American Statistical Association, 91(436): 1413–1422, December 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2291567.